### v1

Abstract
We study real-valued mean-field backward stochastic differential equations (BSDEs) of the form
$Y_{t}=\xi+\int_{t}^{T}\widetilde{\mathbb{E}}\,g(s,Z_{s},\tilde{Z}_{s})\,ds-\int_{t}^{T}Z_{s}\,dW_{s},$
where $\tilde{Z}$ denotes an independent copy of $Z$ and $\widetilde{\mathbb{E}}$ the expectation with respect to $\tilde{Z}$. Under a separately quadratic growth assumption (Hq) on the generator $g$ in (Z,Z~)($\tilde{Z}$), together with a bounded terminal condition, we prove existence and uniqueness of solutions in $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$. Our approach departs from classical fixed-point arguments and instead combines Malliavin calculus with refined BMO and stability estimates: we first obtain uniform $\mathbb{S}^{\infty}$-bounds for $Z$ in a Lipschitz setting with bounded Malliavin derivative, and then pass to the quadratic case by approximation, using a stability result in $\mathbb{S}^{2}\times\mathbb{H}^{2}$. This closes the gap between the quadratic BSDE results of cheridito2017bse and hao2022mean.
In the second part of the paper, we extend the framework to generators of the form $g(t,z,\tilde{z},\bar{z})$ satisfying a fully coupled quadratic condition (Hq). In this general regime we establish existence and uniqueness under a smallness condition on the centered terminal variable. As an application, we solve a mean-field exponential utility maximization problem with a collective liability, thereby generalizing the classical utility maximization framework of hu2005utility to a fully coupled quadratic mean-field setting.
Key words: mean-field BSDEs; quadratic BSDEs; BMO martingales; Malliavin calculus; exponential utility maximization.
MSC: 60H30; 60H07; 60H10.
1 Introduction
Let $(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P})$ be a filtered probability space where $W$ is a $d$-dimensional Brownian motion and $\{\mathcal{F}_{t}\}_{t\geq 0}$ is the Brownian filtration. Consider the following mean-field BSDE:
$Y_{t}=\xi+\int_{t}^{T}\int_{\tilde{\Omega}}g(s,Z_{s},\tilde{Z}_{s})\mathop{}\!d\tilde{\mathbb{P}}ds-\int_{t}^{T}Z_{s}dW_{s},$ (1.1)
[…]
The origin of linear BSDEs can be traced back to their role as adjoint equations in stochastic control problems, first studied by bismut1973conjugate. The seminal paper by pardoux1990adapted established the existence and uniqueness of adapted solutions $(Y,Z)$. The form of the equation is as follows:
$Y_{t}=\xi+\int_{t}^{T}F(s,Y_{s},Z_{s})\,ds-\int_{t}^{T}Z_{s}dW_{s},\quad t\in[0,T],$ (1.2)
under the conditions that the generator $F$ is Lipschitz continuous and $\xi$ is square integrable. For a comprehensive presentation of the theory, see the review by el1997backward books by CohenStocal2015 or ZhangBSDEbook2017. Numerous studies have aimed to improve existence and uniqueness results by relaxing the Lipschitz condition on the coefficients or the conditions on the terminal value $\xi$. Consequently, there has been a growing interest in BSDEs with applications in various fields such as stochastic control, mathematical finance, and partial differential equations.
BSDEs with generators exhibiting quadratic growth in the variable $z$ have attracted significant attention. These equations arise, for example, in stochastic linear-quadratic control with random coefficients bismut1976linearquad, in utility optimization problems with exponential utility functions hu2005utility or in risk minimization problems for the entropic risk measure. Kobylanski:2000cy first provided the existence and uniqueness result for scalar-valued quadratic BSDEs with bounded terminal values, where $F$ is quadratic with respect to its $Z$ component. A detailed review of the literature can be found in el2008bsde. briand2013simple have continued to explore this area, demonstrating the existence of unique solutions to quadratic BSDEs with bounded terminal conditions using purely probabilistic methods. Investigations by briand2006bsde extend the previous results to unbounded unbounded terminal conditions. Some studies generalized to the multidimensional case by hu2016multi; tang2023MultiMFBSDE; luo2019triangle; tevzadze2008solvability.
Mean-field stochastic differential equations, also known as McKean-Vlasov equations, describe stochastic systems whose evolution is influenced by both the microscopic state and the macroscopic distribution of particles. These equations date back to kac1956foundations’s work on the Boltzmann and Vlasov equations in the 1950s. Mean-field game theory began with the groundbreaking contributions of lasry2007mfgame. Building on this foundation, Buckdahn2009MFlimit introduced mean-field BSDEs, derived from approximations involving weakly interacting $N$-particle systems, and investigated them using probabilistic methods. This has led to a growing interest in mean-field SDEs and BSDEs. For instance, Buckdahn2009MFBSDE studied the existence and uniqueness of solutions and a comparison theorem for mean-field BSDEs with Lipschitz continuous generators and square integrable terminal conditions. More recently, hao2022mean explored the solvability of local and global solutions for one-dimensional mean-field BSDEs with quadratic growth, while tang2023MultiMFBSDE established the existence and uniqueness of local and global solutions for multidimensional BSDEs with generators depending on the expectation of both variables. Li2018MFBSDEcontinuous examined mean-field BSDEs whose coefficients depend not only on the solution $(Y,Z)$ but also on the law of $Y$.
pardoux1990adapted interpreted the key structural feature of BSDEs as a nonlinear conditional expectation, highlighting their close link to Malliavin calculus. In fact, Buckdahn2009MFBSDE states that, in a Clark–Ocone type formula, the martingale density component of the solution pair of a BSDE with a classical globally Lipschitz generator on a Gaussian basis is the Malliavin derivative of the other component (see Proposition 5.3). The close relationship between Malliavin calculus and BSDEs stems from the use of Malliavin calculus to prove the regularity of trajectories. For example, briand2013simple provided a straightforward approach to the existence and uniqueness of quadratic BSDEs with bounded terminal conditions using Malliavin calculus to overcome difficulties, and Imkeller2010path proposed a numerical scheme for BSDEs with generators of quadratic growth as one of the applications.
Continuous time portfolio optimization under the utility-maximization framework introduced by merton1969portfolio is a foundational topic in mathematical finance. Merton analyzed the problem in a complete market where the risky asset follows the Black-Scholes dynamics and no constraints are imposed on admissible trading strategies. A major development in the early 2000s was the emergence of a connection between this class of optimization problems and the theory of BSDEs, which has since become a central approach in the field. The core idea of the BSDE methodology is to construct a stochastic process, dependent on the investment strategy, that coincides with the investor’s utility of terminal wealth. In this context, elrouge2000utility used BSDEs to compute the value function and optimal strategy for exponential utility, assuming trading strategies are restricted to a convex cone. Building on this framework, hu2005utility developed a BSDE approach to utility maximization in incomplete markets and extended the methodology to include logarithmic and power utility functions, allowing for general strategy constraints within a closed set.
Our methods are inspired by the works of tevzadze2008solvability and briand2013simple. The former studied the existence and uniqueness of solutions for multidimensional quadratic BSDEs, utilizing BMO martingale theory. The latter provided a straightforward approach to establishing the existence and uniqueness of quadratic BSDEs with bounded terminal conditions using Malliavin calculus. Building on these foundations, we bridge the gap between the approaches of cheridito2017bse and hao2022mean. Specifically, hao2022mean established local existence and uniqueness by applying a contraction mapping with the Reverse Hölder Inequality (RHI) and iterating backward in time, while cheridito2017bse used the Krasnoselskii fixed-point theorem on a small time interval, iterating backward to obtain a global solution. Our contribution extends the results of cheridito2017bse by demonstrating in Proposition 4.9 that the function $f^{1}$ satisfies the conditions set forth in hao2022mean, thereby proving the existence of a solution for quadratic mean-field BSDEs.
Our main contributions are as follows. First, we establish the existence and uniqueness of a solution when the generator exhibits separately quadratic growth in both $Z$ and the mean-field term $\tilde{Z}$ with bounded terminal condition $\xi$. Our approach builds upon the theory of BMO martingales and Malliavin calculus to effectively address the analytical challenges arising from the quadratic structure. A key novelty of this work lies in the inclusion of a quadratic mean-field component $\tilde{Z}$, which distinguishes it from existing results in the literature. Moreover, our method is notable for its simplicity, as it avoids the use of fixed-point arguments commonly employed in previous studies, such as juan2023MFBSDE and tevzadze2008solvability.
The rest of the paper is organized as follows. In Section 2, we present the mathematical framework and state the main result. Section 3 derives an almost sure upper bound for the solution pair $(Y,Z)$ of Lipschitz mean-field BSDEs, assuming the terminal condition admits a uniformly bounded Malliavin derivative, as shown in Proposition 3.1. In the main Section 4, we provide a priori estimates and establish a powerful stability result for mean-field BSDEs with separately quadratic growth. In Section 5, we utilize the bounds obtained in Section 3, which are independent of the Lipschitz constants in $z$ and $\tilde{z}$, together with the stability result and a density argument, to prove the existence and uniqueness of solutions to our quadratic mean-field BSDEs. Finally, in Section 6, we extend the existence and uniqueness results for mean-field quadratic BSDE with separately quadratic driver to a fully coupled quadratic driver with small terminal conditions, and show how the corresponding BSDE characterizes a utility maximization problem with a particular class of constraint sets.
2 Preliminaries and statement of main results
2.1 Preliminaries
[…]
By $\left|\cdot\right|$, we denote the Euclidean norm on $\mathbb{R}^{d}$, and for $p>1$, we denote by
  • $L^{p}$: all real-valued random variables $X$ s.t. $\|X\|_{L^{p}}^{p}:={\mathbb{E}}^{\mathbb{P}}\big[|X|^{p}\big]<\infty$,
  • $L^{\infty}$: all essentially bounded random variables $X$ s.t. $\|X\|_{L^{\infty}}:=\operatorname*{ess\,sup}_{\omega\in\Omega}|X(\omega)|<\infty$,
  • $\mathbb{S}^{p}$: the set of real valued $\mathcal{F}$-adapted continuous processes $(Y_{t})_{0\leq t\leq T}$ satisfying
    $\left\|Y\right\|_{\mathbb{S}^{p}}:=\mathbb{E}\left[\sup_{0\leq r\leq T}\left|Y_{r}\right|^{p}\right]^{\frac{1}{p}}<\infty,$
  • $\mathbb{S}^{\infty}$: the set of real valued $\mathcal{F}$-adapted continuous processes $(Y_{t})_{0\leq t\leq T}$ satisfying
    $\left\|Y\right\|_{\mathbb{S}^{\infty}}:=\sup_{\omega\in\Omega}\sup_{0\leq t\leq T}\left|Y_{t}(\omega)\right|<\infty,$
  • $\mathbb{H}^{p}$: the set of predictable $\mathbb{R}^{d}$-valued process $Z$ satisfying $\left\|Z\right\|_{\mathbb{H}^{p}}:=\mathbb{E}\left[\left(\int_{0}^{T}\left|Z_{r}\right|^{2}\,dr\right)^{\frac{p}{2}}\right]^{\frac{1}{p}}<\infty,$
  • $\mathbb{H}^{2}_{BMO}$: the subsets of $\mathbb{H}^{2}$ with predictable $\mathbb{R}^{d}$-valued process $Z$ satisfying $\int_{0}^{\cdot}Z_{s}\cdot\,dW_{s}$ is a Bounded in Mean Oscillation (BMO) martingale, that is, there exists a non-negative constant $C$ such that, for each stopping time $\tau\leq T$,
    $\mathbb{E}\left[\int_{\tau}^{T}\left|Z_{s}\right|^{2}\,ds|\mathcal{F}_{\tau}\right]\leq C^{2}.$
    The $\left\|\int_{0}^{\cdot}Z_{s}\cdot\,dW_{s}\right\|_{BMO}$ is defined as the smallest non-negative constant $C$ for which the above inequality is satisfied.
  • Let $C$ be a closed subset of $\mathbb{R}^{n}$ and $a\in\mathbb{R}^{n}$. The distance between $a$ and the set $C$ is given by
    $\text{dist}(a,C)=\min_{b\in C}\left|a-b\right|.$
    The set $\Pi_{C}(a)$ consists of all points $b\in C$ at which this minimum is attained, and we refer to this as the projection set.
When a different underlying measure is involved, we indicate it explicitly in the notation, for example by writing $\mathbb{E}$$^{\mathbb{Q}}$$\cdot$] and $\|\cdot\|_{BMO(\mathbb{Q})}$ for expectations and BMO norms taken with respect to a probability measure $\mathbb{Q}$. Here are some inequalities from BMO martingales theory, all the proofs can be found in kazamaki2006BMO.
Lemma 2.1 (Reverse Hölder Inequality (RHI)).
Let $M$ be a one–dimensional continuous $\mathrm{BMO}$ martingale under $\mathbb{P}$. Then the Doléans–Dade exponential $\mathcal{E}(M)$ satisfies the reverse Hölder inequality: there exists $q^{*}>1$ such that for all $q\in(1,q^{*})$ and all stopping times $\tau\leq T$,
$\mathbb{E}\!\left[\left.\left(\frac{\mathcal{E}(M)_{T}}{\mathcal{E}(M)_{\tau}}\right)^{\!q}\;\right|\mathcal{F}_{\tau}\right]\;\leq\;C_{q}^{*},$ (2.3)
where the constants $q^{*}$ and $C_{q}^{*}$ depend explicitly on the $\mathrm{BMO}$ norm $\|M\|_{\mathrm{BMO}(\mathbb{P})}$:
$q^{*}\;:=\;\phi^{-1}\!\Big(\|M\|_{\mathrm{BMO}}\Big),\qquad\phi(x)\;:=\;\Big(1+\tfrac{1}{x^{2}}\log\frac{2x-1}{2x-2}\Big)^{1/2}-1,$
and for $1<q<q^{*}$,
$C_{q}^{*}\;:=\;2\Bigg(1-\frac{2q-2}{2q-1}\,\exp\!\Big\{q^{2}\|M\|_{\mathrm{BMO}}^{2}+2\|M\|_{\mathrm{BMO}}\Big\}\Bigg)^{-1}.$
In particular, $q^{*}$ decreases and $C_{q}^{*}$ increases with respect to $\|M\|_{\mathrm{BMO}}$.
Lemma 2.2.
For $K>0$. There exist constants $c_{1}(K),c_{2}(K)>0$, depending only on $K$, such that the following holds. Let $N$ be a one–dimensional continuous $\mathrm{BMO}$ martingale under $\mathbb{P}$ with $\|N\|_{\mathrm{BMO}(\mathbb{P})}\leq K$, and define the equivalent measure $\tilde{\mathbb{P}}$ by
$\frac{d\tilde{\mathbb{P}}}{d\mathbb{P}}\;=\;\mathcal{E}(N)_{T}.$
For any continuous $\mathrm{BMO}$ martingale $M$ under $\mathbb{P}$, set
$\tilde{M}\;:=\;M-\langle M,N\rangle,$
[…]
$c_{1}(K)\,\|M\|_{\mathrm{BMO}(\mathbb{P})}\;\leq\;\|\tilde{M}\|_{\mathrm{BMO}(\tilde{\mathbb{P}})}\;\leq\;c_{2}(K)\,\|M\|_{\mathrm{BMO}(\mathbb{P})}.$ (2.4)
2.2 Statement of the main theorem
In this paper we consider the following general mean-field BSDE
$Y_{t}=\xi+\int_{t}^{T}\int_{\tilde{\Omega}}g(s,Z_{s},\tilde{Z}_{s})d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}\cdot dW_{s},\quad t\in[0,T].$ (2.5)
The main result of this paper establishes the existence and uniqueness of solutions to the mean-field BSDE under a bounded terminal condition. To this end, we introduce the following assumptions. Let $K_{0},K_{z},K_{\tilde{z}}>0$ be fixed constants.
  1. $(\rm H_{0})$}
    The random variable $\xi$ is bounded almost surely by $K_{0}>0$, and likewise, the deterministic supremum $\sup_{t\in[0,T]}|g(t,0,0)|$ is also bounded by the same constant $K_{0}>0$.
  2. $(\rm H_{q})$}
    There exist constants $K_{z}>0$ and $K_{\tilde{z}}>0$ such that:
[…]
    for any $(t,z,\tilde{z},z^{\prime},\tilde{z}^{\prime})\in[0,T]\times[\mathbb{R}^{d}\times\mathbb{R}^{d}]^{2}$.
Theorem 2.3.
Let Assumptions $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q})$ hold. Then, there exists a unique solution $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ to the BSDE (2.5).
Remark 2.4.
In the statement and proof of our main theorem we may, without loss of generality, assume that
$g(t,0,\tilde{z})=0,\qquad\forall\,(t,\tilde{z})\in[0,T]\times\mathbb{R}^{d}.$
This is not an additional structural condition on the original generator, but merely a convenient normalisation. Given any generator $g$ satisfying Assumptions $(\rm H_{0})$$(\rm H_{q})$}, we may introduce the shifted generator
[…]
and consider the mean-field BSDE driven by $\hat{g}$. For any solution $(Y,Z)$ of the original BSDE with generator $g$, the process
$U_{t}:=Y_{t}-\int_{t}^{T}\widetilde{{\mathbb{E}}}\big[g(s,0,\tilde{Z}_{s})\big]\,ds,\qquad 0\leq t\leq T,$
together with the same $Z$, solves the BSDE with generator $\hat{g}$. Conversely, given a solution $(U,Z)$ to the BSDE with generator $\hat{g}$, the process
$Y_{t}:=U_{t}+\int_{t}^{T}\widetilde{{\mathbb{E}}}\big[g(s,0,\tilde{Z}_{s})\big]\,ds$
[…]
3 Terminal conditions with bounded Malliavin derivative
This section is devoted to deriving tractable upper bounds for the solution $(Y,Z)$ of the BSDE (2.5) under the assumption that the driver $g$ is Lipschitz continuous. These bounds become particularly useful when the terminal condition $\xi$ is Malliavin differentiable. Although this may appear to be a stronger condition, it serves a crucial role in our analysis. In the proof of Lemma 5.1, we will show that the general case, where the driver exhibits quadratic growth and where $\xi$ is bounded, can be reduced to this setting through suitable approximation techniques. Therefore, we initially work under stronger regularity assumptions to facilitate this reduction. For this purpose, we suppose all along this section that $g$ satisfies the following classical Lipschitz condition.
  • $({\rm H_{lip}})$
    There exist constants $L_{z}>0$ such that:
[…]
    for any $(t,z,\tilde{z},z^{\prime},\tilde{z}^{\prime})\in[0,T]\times[\mathbb{R}^{d}\times\mathbb{R}^{d}]^{2}$.
The existence and uniqueness of a solution $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}$ to the BSDE (2.5) under assumptions $(\rm H_{0})$$({\rm H_{lip}})$ is a well-established result, as shown in Theorem 3.1 of Buckdahn2009MFBSDE. In the following proposition, we demonstrate that the process $Y$, which lies in $\mathbb{S}^{\infty}$, is bounded by a constant that is independent of the Lipschitz constant $L_{z}$. More remarkably, we show that the same independence holds for the process $Z$, provided the terminal condition $\xi$ admits uniformly bounded Malliavin derivatives.
Proposition 3.1.
If $(\rm H_{0})$} and $({\rm H_{lip}})$ hold, then the unique solution $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}$ of the Lipschitz mean-field BSDE satisfies the following properties.
  1. 1.
    $Y\in\mathbb{S}^{\infty}$ with the upper bound:
    $\|Y\|_{\mathbb{S}^{\infty}}\;\leq\;\left\|\xi\right\|_{L^{\infty}}.$ (3.6)
  2. 2.
    If $\xi$ is Malliavin differentiable and satisfies $\left\|D\xi\right\|_{\mathbb{S}^{\infty}}=\sup_{0\leq t\leq T}\left\|D_{t}\xi\right\|_{L^{\infty}}<\infty$, then $Z\in\mathbb{S}^{\infty}$, and its norm is also bounded independently of $L_{z}$:
    $\left\|Z\right\|_{\mathbb{S}^{\infty}}\leq\left\|D\xi\right\|_{\mathbb{S}^{\infty}},$ (3.7)
Proof.
Unlike standard Lipschitz BSDE, the comparison theorem cannot be applied here (see Theorem 3.2 in Buckdahn2009MFBSDE). Observe that (3.6) can be derived following the line of proof of Proposition 4.4. In this Lipschitz context, the proof is simpler since no BMO martingale arguments are necessary.
[…]
$D_{\theta}Y_{t}=D_{\theta}\xi+\int_{t}^{T}\widetilde{{\mathbb{E}}}\!\left[\partial_{z}g(s,Z_{s},\widetilde{Z}_{s})\right]\!\cdot D_{\theta}Z_{s}\,ds-\int_{t}^{T}D_{\theta}Z_{s}\cdot dW_{s}.$
Under assumption $({\rm H_{lip}})$, we have $\left|\partial_{z}g\right|\leq L_{z}$. Therefore, the Novikov condition is satisfied, and we may apply the Girsanov transform to obtain the following representation with new Brownian motion $W^{\mathbb{Q}}$ under the new probability measure $\mathbb{Q}$:
$D_{\theta}Y_{t}=D_{\theta}\xi-\int_{t}^{T}D_{\theta}Z_{s}\cdot dW_{s}^{\mathbb{Q}},\qquad 0\leq\theta\leq t\leq T.$
Taking the conditional expectation under $\mathbb{Q}$, we deduce
$\left|D_{\theta}Y_{t}\right|\leq\left\|D\xi\right\|_{\mathbb{S}^{\infty}}.$
Finally, as $(Z_{t})_{0\leq t\leq T}$ is a version of $(D_{\theta}Y_{t})_{0\leq t\leq T}$ and due to the arbitrariness of $t$, we establish 3.7. ∎
4 A priori estimates and stability for BSDEs
In this section, we establish the stability of the mean-field BSDE (2.5), as stated in Theorem 4.6. The strategy of the proof is as follows. We first derive a priori upper bounds on the $\mathbb{S}^{\infty}$ norm of $Y$ and the BMO norm of $\int_{0}^{\cdot}Z_{s}\cdot dW_{s}$ for any solution $(Y,Z)$ to the BSDE (2.5). These estimates, together with fundamental properties of BMO martingales, lead to the stability result in Theorem 4.6. We begin with a useful estimate that will play an important role in the analysis.
Remark 4.1.
Under assumptions $(\rm H_{0})$} and $(\rm H_{q})$}, we have the following estimate:
[…]
for any $(t,z,\tilde{z})\in[0,T]\times\mathbb{R}^{d}\times\mathbb{R}^{d}$. Taking expectations with respect to $\tilde{\mathbb{P}}$ yields:
$\tilde{\mathbb{E}}\left|g(t,Z_{t},\tilde{Z}_{t})\right|\leq K_{0}+\frac{1}{2}K_{z}+\frac{1}{2}K_{\tilde{z}}+\frac{3K_{z}}{2}\left|Z_{t}\right|^{2}+\frac{3K_{\tilde{z}}}{2}\tilde{\mathbb{E}}\left|\tilde{Z}_{t}\right|^{2}.$ (4.8)
We begin by noting that if the solution $Y\in\mathbb{S}^{\infty}$, then for any solution pair $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}$ to the quadratic mean-field BSDE, the stochastic integral $\int_{0}^{\cdot}Z_{s}\cdot dW_{s}$ defines a BMO martingale. Furthermore, the BMO norm of this martingale can be explicitly bounded in terms of the uniform norm $\left\|Y\right\|_{\mathbb{S}^{\infty}}$ and certain constants in the assumptions.
Proposition 4.2.
Suppose that (H0) and (Hq) hold. Let $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}$ be a bounded solution to the mean-field BSDE (2.5). Then the stochastic integral process $\left(\int_{0}^{t}Z_{s}\cdot dW_{s}\right)_{0\leq t\leq T}$ defines a BMO martingale. Moreover, its BMO norm satisfies the following bound:
$\left\|\int_{0}^{\cdot}Z_{s}\cdot dW_{s}\right\|_{\mathrm{BMO}}^{2}\leq(1\vee T)\big(4K_{0}+K_{z}+K_{\tilde{z}}\big)\,\frac{e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}}{3K_{z}}\left(\frac{K_{\tilde{z}}}{K_{z}}e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}+1\right).$ (4.9)
Proof.
The argument follows the two steps scheme of (briand2008stobsdes, p. 831) with modified $C^{2}$ function
$\varphi(x):=\frac{e^{3K_{z}|x|}-1-6cK_{z}|x|}{(3K_{z})^{2}},\qquad x\in\mathbb{R}.$
In Step 1 we express the $\mathrm{BMO}$ norm of $Z$ in terms of its $\mathbb{H}^{2}$ norm; in Step 2 we derive an explicit $\mathbb{H}^{2}$ bound in terms of the model coefficients and $\|Y\|_{\mathbb{S}^{\infty}}$.
Step 1 (Estimate of $\|Z\|_{\mathrm{BMO}}$ in terms of $\|Z\|_{\mathbb{H}^{2}}$). Fix $c=\tfrac{1}{2}$ then $\varphi(0)=0$ and, for all $x\in\mathbb{R}$,
$\varphi^{\prime}(x)=\frac{e^{3K_{z}|x|}-1}{3K_{z}}\operatorname{sgn}(x),\qquad\varphi^{\prime\prime}(x)=e^{3K_{z}|x|},\qquad\varphi^{\prime\prime}(x)-3K_{z}\,|\varphi^{\prime}(x)|\equiv 1.$
Moreover, for $|x|\leq\|Y\|_{\mathbb{S}^{\infty}}$ we have the deterministic bound
$0\leq|\varphi^{\prime}(x)|\leq C_{Y}:=\frac{1}{3K_{z}}\,e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}.$ (4.10)
Let $\tau$ be any stopping time and apply Itô’s formula to $\varphi(Y)$ and using the growth estimate 4.8 for $g$ together with the symmetry $\widetilde{\mathbb{E}}|\tilde{Z}_{s}|^{2}=\mathbb{E}|Z_{s}|^{2}$ yields
$\varphi(Y_{\tau})$ $=\varphi(Y_{T})+\!\int_{\tau}^{T}\!\Big(-\varphi^{\prime}(Y_{s})\,\widetilde{\mathbb{E}}[g(s,Z_{s},\tilde{Z}_{s})]-\tfrac{1}{2}\,\varphi^{\prime\prime}(Y_{s})|Z_{s}|^{2}\Big)ds-\!\int_{\tau}^{T}\!\varphi^{\prime}(Y_{s})Z_{s}\cdot dW_{s}$
$\leq\varphi(Y_{T})+\!\int_{\tau}^{T}\!\Big(\tfrac{3}{2}|\varphi^{\prime}(Y_{s})|-\tfrac{1}{2}\varphi^{\prime\prime}(Y_{s})\Big)|Z_{s}|^{2}ds+\!\int_{\tau}^{T}\!|\varphi^{\prime}(Y_{s})|\Big(K+\tfrac{3}{2}K_{\tilde{z}}\,\mathbb{E}|Z_{s}|^{2}\Big)ds$
$\leq\varphi(Y_{T})-\tfrac{1}{2}\!\int_{\tau}^{T}\!|Z_{s}|^{2}ds+\!\int_{\tau}^{T}\!|\varphi^{\prime}(Y_{s})|\Big(K+\tfrac{3}{2}K_{\tilde{z}}\,\mathbb{E}|Z_{s}|^{2}\Big)ds-\!\int_{\tau}^{T}\!\varphi^{\prime}(Y_{s})Z_{s}\cdot dW_{s},$
where $K:=K_{0}+\frac{1}{2}K_{z}+\frac{1}{2}K_{\tilde{z}}$. Taking $\mathbb{E}[\cdot\mid\mathcal{F}_{\tau}]$ and using (4.10) together with $|\varphi(x)|\leq C_{Y}|x|$ (since $\varphi(0)=0$ and $\varphi^{\prime}$ is bounded on the range of $Y$), we get
$\tfrac{1}{2}\,\mathbb{E}\!\Big[\!\int_{\tau}^{T}\!|Z_{s}|^{2}ds\,\Big|\,\mathcal{F}_{\tau}\Big]$ $\leq\mathbb{E}\!\big[\varphi(Y_{T})-\varphi(Y_{\tau})\,\big|\,\mathcal{F}_{\tau}\big]+C_{Y}K(T-\tau)+\tfrac{3}{2}K_{\tilde{z}}C_{Y}\!\!\int_{0}^{T}\!\mathbb{E}|Z_{s}|^{2}ds$
$\leq C_{Y}\,\mathbb{E}\!\big[|\xi|\,\big|\,\mathcal{F}_{\tau}\big]+C_{Y}KT+\tfrac{3}{2}K_{\tilde{z}}C_{Y}\!\!\int_{0}^{T}\!\mathbb{E}|Z_{s}|^{2}ds.$
we thus obtain
$\tfrac{1}{2}\,{\mathbb{E}}\!\Big[\!\int_{\tau}^{T}\!|Z_{s}|^{2}ds\,\Big|\,\mathcal{F}_{\tau}\Big]\leq(3K_{z})^{-1}e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}\left(\|\xi\|_{\infty}+K\,T+\tfrac{3}{2}K_{\tilde{z}}\!\!\int_{0}^{T}\!\!{\mathbb{E}}|Z_{s}|^{2}ds\right).$ (4.11)
Since $\tau$ is arbitrary, (4.11) shows that $\big\|\int_{0}^{\cdot}Z_{s}\cdot dW_{s}\big\|_{\mathrm{BMO}}^{2}$ is controlled by $\|Z\|_{\mathbb{H}^{2}}^{2}$ and model parameters.
Step 2 (An explicit $\mathbb{H}^{2}$ bound). Write $\alpha:=K_{\tilde{z}}/K_{z}>0$. We may, without loss of generality (enlarge $K_{z}$ or $K_{\tilde{z}}$) assume $\alpha$ satisfies
$0<\alpha\;<\;\min\!\Big\{\tfrac{1}{2},\ \frac{e^{3K_{z}}-1}{e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}}\Big\}.$
Define positive constants
$c_{1}:=\frac{K_{\tilde{z}}}{2K_{z}}e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}+\frac{1}{2},\qquad c_{2}:=\tfrac{1}{2}e^{3K_{z}}.$
Note $c_{1}$ can also be expressed as $c_{1}=\tfrac{1}{2}\big(\alpha e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}+1\big)=\frac{3K_{\tilde{z}}}{2}C_{Y}+\tfrac{1}{2}$. By the choice of $\alpha$, it is easy to verify $c_{2}>c_{1}$. By density of rational numbers, pick any $c\in(c_{1},c_{2})$ and consider the same $C^{2}$ function at in Step 1, so that $\varphi(0)=0$ and, for all $x$,
$|\varphi^{\prime}(x)|=\frac{e^{3K_{z}|x|}-2c}{3K_{z}}>\frac{e^{3K_{z}|x|}-2c_{2}}{3K_{z}}=0.$
So that we have identity $\varphi^{\prime\prime}(\cdot)-3K_{z}\left|\varphi^{\prime}(\cdot)\right|=2c$. Applying Itô to $\varphi(Y)$ on $[0,T]$ and arguing as in Step 1 gives
$\varphi(Y_{0})$ $\leq\varphi(\xi)-\int_{0}^{T}c\,|Z_{s}|^{2}ds+\frac{3K_{\tilde{z}}}{2}\int_{0}^{T}|\varphi^{\prime}(Y_{s})|\,\mathbb{E}|Z_{s}|^{2}ds+K\int_{0}^{T}|\varphi^{\prime}(Y_{s})|ds-\int_{0}^{T}\varphi^{\prime}(Y_{s})Z_{s}\cdot dW_{s}.$
Taking expectations and using $|\varphi^{\prime}(x)|=\frac{e^{3K_{z}x}-2c}{3K_{z}}\leq\frac{e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}}{3K_{z}}=C_{Y}$ for $|x|\leq\|Y\|_{\mathbb{S}^{\infty}}$,
$\varphi(Y_{0})\leq\mathbb{E}\!\left[\varphi(\xi)\right]+KC_{Y}T-c\,\mathbb{E}\!\left[\int_{0}^{T}|Z_{s}|^{2}ds\right]+\frac{3K_{\tilde{z}}}{2}\,C_{Y}\,\mathbb{E}\!\left[\int_{0}^{T}|Z_{s}|^{2}ds\right].$
By the definition of $c_{1}$ and $c>c_{1}$ we have $\frac{3K_{\tilde{z}}}{2}C_{Y}\leq c-\tfrac{1}{2}$. Hence
$\frac{1}{2}\,\mathbb{E}\!\left[\int_{0}^{T}|Z_{s}|^{2}ds\right]\ \leq\ \mathbb{E}\!\left[\varphi(\xi)\right]+KC_{Y}T\ \leq\ \frac{e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}}{3K_{z}}\,\big(\left\|\xi\right\|_{L^{\infty}}+KT\big).$ (4.12)
Finally, plugging (4.12) into (4.11) and taking the essential supremum over $\tau$ yields an explicit $\mathrm{BMO}$ bound for $\int_{0}^{\cdot}Z_{s}\cdot dW_{s}$ in terms of $\|Y\|_{\mathbb{S}^{\infty}}$ and the model parameters. This completes the proof. ∎
Next, suppose the solution $Z\in\mathbb{H}^{2}_{\mathrm{BMO}}$. Then, for any solution pair $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ to the quadratic mean-field BSDE, we derive the following proposition showing that $Y\in\mathbb{S}^{\infty}$ which is independent of the BMO norm $\left\|\int_{0}^{\cdot}Z_{s}\cdot dW_{s}\right\|_{\mathrm{BMO}}$. We introduce a simple but useful lemma which will be used later on when we work under equivalent probability measures.
Lemma 4.3.
Let $\mathbb{P}$ and $\mathbb{Q}$ be equivalent probability measure on $(\Omega,\mathcal{F})$. For any random variable $X$,
$\|X\|_{L^{\infty}(\mathbb{P})}=\|X\|_{L^{\infty}(\mathbb{Q})}.$
In particular, $X$ is bounded $\mathbb{P}$-a.s. iff it is bounded $\mathbb{Q}$-a.s., with the same bound.
Proof.
Set $K:=\|X\|_{L^{\infty}(\mathbb{P})}$. For every $\varepsilon>0$, $\mathbb{P}(|X|>K+\varepsilon)=0$. Since $\mathbb{Q}\ll\mathbb{P}$, also $\mathbb{Q}(|X|>K+\varepsilon)=0$, so $\|X\|_{L^{\infty}(\mathbb{Q})}\leq K$. By symmetry ($\mathbb{P}\ll\mathbb{Q}$), the reverse inequality holds. ∎
Proposition 4.4.
Suppose Assumptions $(\rm H_{0})$} and $(\rm H_{q})$} hold, and that we work with the shifted generator satisfying
$g(t,0,\tilde{z})=0,\qquad\forall\,(t,\tilde{z})\in[0,T]\times\mathbb{R}^{d}$
(cf. Remark 2.4). Then for any solution $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ of BSDE (2.5), we have $Y\in\mathbb{S}^{\infty}$ and the following estimate holds:
$\left\|Y\right\|_{\mathbb{S}^{\infty}}\leq\left\|\xi\right\|_{L^{\infty}}.$ (4.13)
Proof.
The proof relies on a linearization argument combined with the BMO property. Under the standing normalisation $g(t,0,\tilde{z})=0$, the BSDE (2.5) can be rewritten as
$Y_{t}=\xi+\int_{t}^{T}\Big(\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big)\cdot Z_{s}\,ds-\int_{t}^{T}Z_{s}\cdot dW_{s},\qquad t\in[0,T],$ (4.14)
where the process $(L_{s})_{0\leq s\leq T}$ is defined by
$L_{s}:=\frac{g(s,Z_{s},\tilde{Z}_{s})-g(s,0,\tilde{Z}_{s})}{\left|Z_{s}\right|^{2}}\,Z_{s}\,\mathbf{1}_{\{\left|Z_{s}\right|>0\}},\qquad 0\leq s\leq T.$
Since $g(s,0,\tilde{z})=0$ by construction, the numerator is simply $g(s,Z_{s},\tilde{Z}_{s})$. Moreover, Assumption $(\rm H_{q})$} yields
$\Big|\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big|\leq\int_{\tilde{\Omega}}\left|L_{s}\right|\,d\tilde{\mathbb{P}}\leq K_{z}\big(1+\left|Z_{s}\right|\big),\qquad 0\leq s\leq T.$
Since $Z\in\mathbb{H}^{2}_{\mathrm{BMO}}$ by assumption, $\big(\int_{0}^{t}\Big(\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big)\cdot dW_{s}\big)_{0\leq t\leq T}$ is a BMO martingale, and its stochastic exponential $\mathcal{E}_{t}$ satisfies Kazamaki’s condition. Define the probability measure $\mathbb{P}^{L}$ by $d\mathbb{P}^{L}:=\mathcal{E}_{T}\,d\mathbb{P}$. By the above BMO argument, $(\mathcal{E}_{t})$ is a uniformly integrable martingale, so $\mathbb{P}^{L}$ is equivalent to $\mathbb{P}$. Girsanov’s theorem then implies that the process
$W_{t}^{L}:=W_{t}+\int_{0}^{t}\Big(\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big)\,ds,\qquad t\in[0,T],$
is a dd–dimensional Brownian motion under $\mathbb{P}^{L}$, thus we obtain
$Y_{t}=\xi-\int_{t}^{T}Z_{s}\cdot dW_{s}^{L}.$
Hence,
$Y_{t}={\mathbb{E}}^{\mathbb{P}^{L}}\big[\xi\,\big|\,\mathcal{F}_{t}\big],\qquad 0\leq t\leq T.$
Since $\mathbb{P}^{L}$ is equivalent to $\mathbb{P}$, the same bound holds $\mathbb{P}$–almost surely (Lemma 4.3), and taking the essential supremum over $t\in[0,T]$ completes the proof. ∎
Then we have a direct result from previous two propositions.
Corollary 4.5.
Let $(\rm H_{0})$} and $(\rm H_{q})$} hold. Then, for any solution pair $(Y,Z)$ to the quadratic BSDE (2.5), if either
$(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}\quad\text{or}\quad(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}_{\mathrm{BMO}},$
it follows that $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ and
$\left\|Y\right\|_{\mathbb{S}^{\infty}}+\Big\|\textstyle\int_{0}^{\cdot}Z_{s}\cdot dW_{s}\Big\|_{\mathrm{BMO}}\leq C,$ (4.15)
where $C$ is a constant depending only on $\left\|\xi\right\|_{L^{\infty}}$, $T$, and the constants appearing in Assumptions $(\rm H_{0})$} and $(\rm H_{q})$}.
Proof.
We only sketch how the two estimates interact, as the details follow directly from Propositions 4.2 and 4.4. If $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}$, then Proposition 4.2 yields
$Z\in\mathbb{H}^{2}_{\mathrm{BMO}}\quad\text{and}\quad\Big\|\textstyle\int_{0}^{\cdot}Z_{s}\cdot dW_{s}\Big\|_{\mathrm{BMO}}\ \text{is bounded in terms of }\|Y\|_{\mathbb{S}^{\infty}}.$
Since we are now in the setting of Proposition 4.4, we obtain an $\mathbb{S}^{\infty}$ estimate for $Y$ which depends only on $\left\|\xi\right\|_{L^{\infty}}$. Combining these two facts shows that both $\|Y\|_{\mathbb{S}^{\infty}}$ and the BMO norm of $\int Z\,dW$ are bounded by some constant $C$ depending only on the model parameters ($K_{0},K_{z},K_{\tilde{z}},T$). The other case $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ is the same. ∎
Theorem 4.6.
For $i=1,2$, let $(Y^{i},Z^{i})\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ solve
$Y_{t}^{i}=\xi^{i}+\int_{t}^{T}\!\int_{\tilde{\Omega}}g(s,Z_{s}^{i},\tilde{Z}_{s}^{i})\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}^{i}\cdot dW_{s},\qquad t\in[0,T],$
with $\xi^{1},\xi^{2},g$ satisfying $(\mathrm{H}_{0})$ and $(\mathrm{H}_{g})$. Then
$\|Y^{1}-Y^{2}\|_{\mathbb{S}^{2}}^{2}+\|Z^{1}-Z^{2}\|_{\mathbb{H}^{2}}^{2}\ \leq\ C_{\mathrm{global}}\;\|\xi^{1}-\xi^{2}\|_{L^{4}},$
where $C_{\mathrm{global}}$ depends only on $T$, $K_{z}$, $K_{\tilde{z}}$ and $K_{0}$.
Proof.
Let $\Delta Y:=Y^{1}-Y^{2}$, $\Delta Z:=Z^{1}-Z^{2}$, $\Delta\xi:=\xi^{1}-\xi^{2}$ and
$\Delta g(s,Z_{s},\tilde{Z}_{s}):=g(s,Z_{s}^{1},\tilde{Z}_{s}^{1})-g(s,Z_{s}^{2},\tilde{Z}_{s}^{2}),\qquad s\in[0,T].$
Then
$\Delta Y_{t}=\Delta\xi+\int_{t}^{T}\!\int_{\tilde{\Omega}}\!\Delta g(s,Z_{s},\tilde{Z}_{s})\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}\Delta Z_{s}\cdot dW_{s},\qquad t\in[0,T].$ (4.16)
Preliminary step (linearization and RHI). As in the proof of Proposition 4.4, we use a linearization argument, there exists processes $L_{s}$ and $\tilde{L}_{s}$ live on the product space such that, for all $s\in[0,T]$,
$|L_{s}|\leq K_{z}\big(1+|Z_{s}^{1}|+|Z_{s}^{2}|\big),\qquad|\tilde{L}_{s}|\leq K_{\tilde{z}}\big(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|\big),$ (4.17)
and $\Delta g(s,Z_{s},\tilde{Z}_{s})=L_{s}\Delta Z_{s}+\tilde{L}_{s}\,\Delta\tilde{Z}_{s}$. Fix a small $\delta>0$ and write $I:=[T-\delta,T]$. According to Proposition 4.2, we have $Z^{1},Z^{2}\in\mathbb{H}^{2}_{\mathrm{BMO}}$, so in particular $\big(\int_{0}^{t}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\big)_{0\leq t\leq T}$ is a $\mathrm{BMO}$ martingale by (4.17). Moreover, by Corollary 4.5, there exists a constant $C>0$, depending only on $(T,K_{z},K_{\tilde{z}},K_{0})$, such that
$\Big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\Big\|_{\mathrm{BMO}}+\Big\|\int_{0}^{\cdot}Z^{1}_{s}\cdot dW_{s}\Big\|_{\mathrm{BMO}}+\Big\|\int_{0}^{\cdot}Z^{2}_{s}\cdot dW_{s}\Big\|_{\mathrm{BMO}}\;\leq\;C.$ (4.18)
Hence $\Lambda:=\mathcal{E}(\int_{0}^{\cdot}\widetilde{\mathbb{E}}[L_{s}]\cdot dW_{s})$ defines an equivalent probability measure $\mathbb{P}^{L}$ by Girsanov’s theorem with Brownian motion $W^{L}_{t}:=W_{t}+\int_{0}^{t}\widetilde{\mathbb{E}}[L_{s}]\,ds$. In addition, by Lemma 2.2, the same processes admit $\mathrm{BMO}$ norms under the equivalent measure $\mathbb{P}^{L}$, up to multiplicative constants depending on the same constant. We then obtain,
$\Delta Y_{u}=\Delta\xi+\int_{u}^{T}\widetilde{\mathbb{E}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds-\int_{u}^{T}\Delta Z_{s}\cdot dW^{L}_{s},\qquad u\in I.$ (4.19)
The Doléans–Dade exponential $\Lambda=\mathcal{E}\!\left(\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\right)$ satisfies RHI from the Lemma 2.1,
$\bar{{\mathbb{E}}}\!\left[\left(\Lambda_{T}\right)^{q}\,\big|\,\bar{\mathcal{F}}_{t}\right]\ \leq\ C_{q}^{*}\,\left(\Lambda_{t}\right)^{q},\qquad 0\leq t\leq T,\quad 1\leq q\leq q^{*},$
where
$q^{*}$ $:=\phi\!\left(\Big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\Big\|_{\mathrm{BMO}}\right),\qquad\phi(x):=\left(1+\frac{1}{x^{2}}\log\frac{2x-1}{2x-2}\right)^{\!1/2}-1,$
$C_{q}^{*}$ $:=2\left(1-\frac{2q-1}{2q-2}\,\exp\!\left\{q^{2}\Big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2}+2\Big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\Big\|_{\mathrm{BMO}}\right\}\right)^{-1}\!.$
By the sliceability of BMO martingales and the function $\phi$ is strictly decreasing with respect to BMO norm, we can restrict to a short interval $I=[T-\delta,T]$ so that the local norm $\big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\big\|_{\mathrm{BMO}(I)}$ is small enough to ensure $q^{*}>2$. Then we have the RHI for every nonnegative $X\in\mathcal{F}_{T}$, and we pick $q=2$:
$\mathbb{E}\!\left[X\ \middle|\ \mathcal{F}_{T-\delta}\right]\ \leq\ C_{2,-}(I)^{1/2}\,\Big(\mathbb{E}^{L}\!\left[X^{2}\,\middle|\,\mathcal{F}_{T-\delta}\right]\Big)^{1/2},$ (4.20)
where the explicit constant $C_{2,-}(I)\ :=\ C_{q}^{*},$ with $q=2$ on interval $I$.
Step 1 (local stability on $I$). Applying Itô’s Lemma to $|\Delta Y|^{2}$ and taking $\mathbb{E}^{L}$ gives
$\mathbb{E}^{L}|\Delta Y_{u}|^{2}+\mathbb{E}^{L}\!\int_{u}^{T}|\Delta Z_{s}|^{2}\,ds\ \leq\ \mathbb{E}^{L}|\Delta\xi|^{2}+2\mathbb{E}^{L}\int_{u}^{T}\!\big[\,|\Delta Y_{s}|\,\widetilde{\mathbb{E}}|\tilde{L}_{s}\,\Delta\tilde{Z}_{s}|\,\big]ds.$ (4.21)
By Corollary 4.5 we have $\|\Delta Y\|_{\mathbb{S}^{\infty}}<\infty$ and the bound depends only on model parameters; hence, for all $u\in I$,
$\mathbb{E}^{L}|\Delta Y_{u}|^{2}+\mathbb{E}^{L}\!\int_{u}^{T}|\Delta Z_{s}|^{2}ds\ \leq\ \mathbb{E}^{L}|\Delta\xi|^{2}+2\|\Delta Y\|_{\mathbb{S}^{\infty}}\,\mathbb{E}^{L}[M_{\delta}],\qquad M_{\delta}:=\int_{T-\delta}^{T}\widetilde{\mathbb{E}}|\tilde{L}_{s}\,\Delta\tilde{Z}_{s}|\,ds.$ (4.22)
Evaluating (4.22) at $u=T-\delta$,
$\mathbb{E}^{L}\!\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}ds\ \leq\ \mathbb{E}^{L}|\Delta\xi|^{2}+2\|\Delta Y\|_{\mathbb{S}^{\infty}}\,M_{\delta}.$ (4.23)
By Cauchy–Schwarz inequality on the copy space and (4.17), for $s\in I$,
$\widetilde{{\mathbb{E}}}\,|\tilde{L}_{s}\,\Delta\tilde{Z}_{s}|\;\leq\;\big(\widetilde{{\mathbb{E}}}|\tilde{L}_{s}|^{2}\big)^{1/2}\big(\widetilde{{\mathbb{E}}}|\Delta\tilde{Z}_{s}|^{2}\big)^{1/2}\;\leq\;\sqrt{3}\,K_{\tilde{z}}\,\Big(1+{\mathbb{E}}|Z_{s}^{1}|^{2}+{\mathbb{E}}|Z_{s}^{2}|^{2}\Big)^{1/2}\big({\mathbb{E}}|\Delta Z_{s}|^{2}\big)^{1/2}.$
Cauchy–Schwarz in time yields
$M_{\delta}$ $\leq\sqrt{3}\,K_{\tilde{z}}\int_{T-\delta}^{T}\Big(1+{\mathbb{E}}|Z_{s}^{1}|^{2}+{\mathbb{E}}|Z_{s}^{2}|^{2}\Big)^{1/2}\big({\mathbb{E}}|\Delta Z_{s}|^{2}\big)^{1/2}\,ds$ (4.24)
$\leq\sqrt{3}\,K_{\tilde{z}}\,\Bigg(\int_{T-\delta}^{T}\Big(1+{\mathbb{E}}|Z_{s}^{1}|^{2}+{\mathbb{E}}|Z_{s}^{2}|^{2}\Big)\,ds\Bigg)^{\!1/2}\Bigg(\int_{T-\delta}^{T}{\mathbb{E}}|\Delta Z_{s}|^{2}\,ds\Bigg)^{\!1/2}.$
For the first term we simply use BMO energy inequality (kazamaki2006BMO, Sec. 2.1)
$\int_{T-\delta}^{T}\Big(1+{\mathbb{E}}|Z_{s}^{1}|^{2}+{\mathbb{E}}|Z_{s}^{2}|^{2}\Big)\,ds\;$ $\leq\;\delta+\Big\|\int_{T-\delta}^{\cdot}Z_{s}^{1}\,\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2}+\Big\|\int_{T-\delta}^{\cdot}Z_{s}^{2}\,\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2}=:\;C_{\mathrm{energy}},$
where $C_{\mathrm{energy}}<\infty$ and increases with respect to the BMO norm. Hence,
$M_{\delta}\;\leq\;\sqrt{3}\,K_{\tilde{z}}\,\sqrt{C_{\mathrm{energy}}}\,\Bigg({\mathbb{E}}\!\int_{T-\delta}^{T}\!|\Delta Z_{s}|^{2}\,ds\Bigg)^{\!1/2}.$ (4.25)
Since the right-hand side is deterministic (depends only on $\mathbb{E}$$\cdot$]), the outer expectation ${\mathbb{E}}^{L}$ drops out; thus ${\mathbb{E}}^{L}[M_{\delta}]$ has the same bound. Next, bridge $\mathbb{E}$} and ${\mathbb{E}}^{L}$ by the RHI (4.20) with $X=\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}\,ds$:
${\mathbb{E}}\!\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}\,ds\;\leq\;\sqrt{C_{2,-}(I)}\,\Bigg({\mathbb{E}}^{L}\!\Big[\Big(\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}\,ds\Big)^{2}\Big]\Bigg)^{\!1/2}.$ (4.26)
Combining (4.25) and (4.26) yields
$M_{\delta}\;\leq\;\sqrt{3}\,K_{\tilde{z}}\,\sqrt{C_{\mathrm{energy}}}\,C_{2,-}(I)^{\frac{1}{4}}\;\Bigg({\mathbb{E}}^{L}\!\Big[\Big(\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}\,ds\Big)^{2}\Big]\Bigg)^{\!1/4}.$ (4.27)
To bound the right hand side, pick $u\in I$ and set $N_{u}:=\int_{T-\delta}^{u}\Delta Z_{s}\cdot dW^{L}_{s}$. Then $N]_{T}=\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}ds$ and the Burkholder–Davis–Gundy (BDG) lower bound at $p=4$ gives $c_{4}^{\downarrow}\,{\mathbb{E}}^{L}\!\big[[N]_{T}^{2}\big]\leq{\mathbb{E}}^{L}|N_{T}|^{4}$. From (4.19),
$N_{T}=X-{\mathbb{E}}^{L}[X|\mathcal{F}_{T-\delta}],\qquad X:=\Delta\xi+\int_{T-\delta}^{T}\widetilde{{\mathbb{E}}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds\in\mathcal{F}_{T}.$
Thus, Jensen’s inequality and $(|a|+|b|)^{4}\leq 8(|a|^{4}+|b|^{4})$ yield
${\mathbb{E}}^{L}|N_{T}|^{4}\leq 16\,{\mathbb{E}}^{L}|X|^{4}\;\leq\;128\Big({\mathbb{E}}^{L}|\Delta\xi|^{4}+{\mathbb{E}}^{L}\big|\textstyle\int_{T-\delta}^{T}\widetilde{{\mathbb{E}}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds\big|^{4}\Big).$
Hence, using (4.27),
${\mathbb{E}}^{L}\big|\textstyle\int_{T-\delta}^{T}\widetilde{{\mathbb{E}}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds\big|^{4}\;\leq\;M_{\delta}^{4}\;\leq\;9\,K_{\tilde{z}}^{4}\,C_{\mathrm{energy}}^{2}\,C_{2,-}(I)\;{\mathbb{E}}^{L}\!\Big[\Big(\int_{T-\delta}^{T}\!|\Delta Z_{s}|^{2}\,ds\Big)^{\!2}\Big].$
Collecting the displays, there exists $C_{\mathrm{BDG}}>0$ (absorbing numeric constants and $c_{4}^{\downarrow}$ only) such that
${\mathbb{E}}^{L}\!\Big[\Big(\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}\,ds\Big)^{2}\Big]\;\leq\;C_{\mathrm{BDG}}\Big({\mathbb{E}}^{L}|\Delta\xi|^{4}\;+\;K_{\tilde{z}}^{4}\,C_{\mathrm{energy}}^{2}\,C_{2,-}(I)\;{\mathbb{E}}^{L}\!\Big[\Big(\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}\,ds\Big)^{2}\Big]\Big).$
By choosing $\delta>0$ small, we can make the local BMO norms of $Z^{1},Z^{2}$ and $\widetilde{{\mathbb{E}}}[L]$ on the slice $I=[T-\delta,T]$ arbitrarily small. Consequently, both the energy constant $C_{\mathrm{energy}}$ (from the BMO energy inequality) and the reverse Hölder constant $C_{2,-}(I)$ are controlled by these norms and $\delta$, hence decrease simultaneously to zero as $\delta\downarrow 0$. In particular, the coefficient in front of ${\mathbb{E}}^{L}\big[(\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}ds)^{2}\big]$ can be made strictly less than $\tfrac{1}{4}$, so we absorb it to the left-hand side and conclude
${\mathbb{E}}^{L}\!\Big[\Big(\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}\,ds\Big)^{2}\Big]\;\leq\;4\,C_{\mathrm{BDG}}\;{\mathbb{E}}^{L}|\Delta\xi|^{4}.$
Finally, plug this into (4.27) to get
$M_{\delta}\;\leq\;\,\sqrt{6}\,K_{\tilde{z}}\,\sqrt{C_{\mathrm{energy}}}\,C_{2,-}(I)^{\frac{1}{4}}\;C_{\mathrm{BDG}}^{\frac{1}{4}}\;\big({\mathbb{E}}^{L}|\Delta\xi|^{4}\big)^{1/4}.$ (4.28)
$\mathbb{H}^{2}$–control of $\Delta Z$ on $I$. Plugging (4.28) into (4.23) gives
$\mathbb{E}^{L}\!\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}ds\$$\leq\ \mathbb{E}^{L}|\Delta\xi|^{2}+2\|\Delta Y\|_{\mathbb{S}^{\infty}}\,\sqrt{6}\,K_{\tilde{z}}\,\sqrt{C_{\mathrm{energy}}}\,C_{2,-}(I)^{\frac{1}{4}}\;C_{\mathrm{BDG}}^{\frac{1}{4}}\,(\mathbb{E}^{L}|\Delta\xi|^{4})^{1/4}$
$\leq\ \left(2K_{0}+2\|\Delta Y\|_{\mathbb{S}^{\infty}}\,\sqrt{6}\,K_{\tilde{z}}\,\sqrt{C_{\mathrm{energy}}}\,C_{2,-}(I)^{\frac{1}{4}}\;C_{\mathrm{BDG}}^{\frac{1}{4}}\right)(\mathbb{E}^{L}|\Delta\xi|^{4})^{1/4},$
the last inequality using Cauchy–Schwarz and $\Delta\xi\in L^{\infty}(\mathbb{P})$ and thus $\Delta\xi\in L^{\infty}(\mathbb{P}^{L})$ with the same bound $K_{0}$ by Lemma 4.3, and therefore, $\mathbb{E}^{L}|\Delta\xi|^{2}\leq(\mathbb{E}^{L}|\Delta\xi|^{4})^{1/2}\leq 2K_{0}\,(\mathbb{E}^{L}|\Delta\xi|^{4})^{1/4}$.
$\mathbb{S}^{2}$–control of $\Delta Y$ on $I$. Recall $X=\Delta\xi+\int_{T-\delta}^{T}\widetilde{{\mathbb{E}}}\!\big[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}\big]\,ds\in\mathcal{F}_{T}$ from (4.19); for $u\in I$,
$\Delta Y_{u}={\mathbb{E}}^{L}\!\big[X\,\big|\,\mathcal{F}_{u}\big]-\int_{T-\delta}^{u}\widetilde{{\mathbb{E}}}\!\big[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}\big]\,ds.$
Hence
$\sup_{u\in I}|\Delta Y_{u}|\ \leq\ \sup_{u\in I}\Big|{\mathbb{E}}^{L}\!\big[X\,\big|\,\mathcal{F}_{u}\big]\Big|\ +\ \int_{T-\delta}^{T}\widetilde{{\mathbb{E}}}\big|\tilde{L}_{s}\,\Delta\tilde{Z}_{s}\big|\,ds\ =\ \sup_{u\in I}\Big|{\mathbb{E}}^{L}\!\big[X\,\big|\,\mathcal{F}_{u}\big]\Big|\ +\ M_{\delta}.$
By $(a+b)^{2}\leq 2a^{2}+2b^{2}$ and Doob’s maximal inequality,
${\mathbb{E}}^{L}\!\Big[\sup_{u\in I}|\Delta Y_{u}|^{2}\Big]\ \leq\ 8\,{\mathbb{E}}^{L}|X|^{2}\;+2\;{\mathbb{E}}^{L}[M_{\delta}^{2}].$
Using again $(a+b)^{2}\leq 2a^{2}+2b^{2}$,
${\mathbb{E}}^{L}|X|^{2}={\mathbb{E}}^{L}\Big|\Delta\xi+\int_{T-\delta}^{T}\widetilde{{\mathbb{E}}}\!\big[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}\big]\,ds\Big|^{2}\ \leq\ 2\,{\mathbb{E}}^{L}|\Delta\xi|^{2}\;+2\;{\mathbb{E}}^{L}[M_{\delta}^{2}],$
we obtain
${\mathbb{E}}^{L}\!\Big[\sup_{u\in I}|\Delta Y_{u}|^{2}\Big]\ \leq\ 16\,{\mathbb{E}}^{L}|\Delta\xi|^{2}\;+\;18\,{\mathbb{E}}^{L}[M_{\delta}^{2}].$
Invoking the bound (4.28) and $\mathbb{E}^{L}|\Delta\xi|^{2}\leq(\mathbb{E}^{L}|\Delta\xi|^{4})^{1/2}$, we conclude
${\mathbb{E}}^{L}\!\Big[\sup_{u\in I}|\Delta Y_{u}|^{2}\Big]\ \leq\ \left(16\;+\;108\,K_{\tilde{z}}^{2}\,C_{\mathrm{energy}}\,\sqrt{C_{2,-}(I)}\,\sqrt{C_{\mathrm{BDG}}}\right)\big({\mathbb{E}}^{L}|\Delta\xi|^{4}\big)^{1/2}.$ (4.29)
We arrive at the local estimate $({\mathbb{E}}^{L}|\Delta\xi|^{4})^{1/2}\leq 2K_{0}({\mathbb{E}}^{L}|\Delta\xi|^{4})^{1/4}$:
${\mathbb{E}}^{L}\!\Big[\sup_{u\in I}|\Delta Y_{u}|^{2}\Big]\;+\;{\mathbb{E}}^{L}\!\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}ds\ \leq\ C_{\mathrm{loc}}(I)\,({\mathbb{E}}^{L}|\Delta\xi|^{4})^{1/4},$ (4.30)
with the constant $C_{\mathrm{loc}}$ depending on $\delta$, $K_{0}$, $K_{z}$, $K_{\tilde{z}}$ and on the BMO/RHI/BDG constants.
Step 2 (global stability). Let $0=t_{0}<t_{1}<\cdots<t_{N}=T$ with $t_{k+1}-t_{k}=\delta$ (small enough so the local estimates hold on each slice $I_{k}:=[t_{k},t_{k+1}]$). By Proposition 4.4 we have $\Delta Y_{t_{k+1}}\in L^{\infty}(\mathbb{P})$. Since essential suprema are invariant under equivalent changes of measure (Lemma 4.3), in particular, $\Delta\xi$ and $\Delta Y_{t_{k+1}}$ belong to $L^{\infty}$ under both $\mathbb{P}$ and $\mathbb{P}^{L}$. Hence (4.30) applies on every $I_{k}$ with terminal datum $\Delta Y_{t_{k+1}}\in L^{\infty}(\mathbb{P}^{L})$:
$\mathbb{E}^{L}\!\Big[\sup_{u\in I_{k}}|\Delta Y_{u}|^{2}\Big]\;+\;\mathbb{E}^{L}\!\int_{I_{k}}|\Delta Z_{s}|^{2}\,ds\ \leq\ C_{\mathrm{loc}}(I_{k})\;\Big(\mathbb{E}^{L}|\Delta Y_{t_{k+1}}|^{4}\Big)^{1/4}.$ (4.31)
Repeating Step 1 on $I_{k}$ with terminal $\Delta Y_{t_{k+1}}$ under $\mathbb{P}^{L}$ read
$\Delta Y_{u}=\Delta Y_{t_{k+1}}+\int_{u}^{t_{k+1}}\widetilde{{\mathbb{E}}}\!\big[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}\big]\,ds-\int_{u}^{t_{k+1}}\Delta Z_{s}\cdot dW_{s}^{L},\qquad u\in I_{k}.$
Evaluating at $u=t_{k}$ and setting
$X_{k}:=\Delta Y_{t_{k+1}}+\int_{t_{k}}^{t_{k+1}}\widetilde{{\mathbb{E}}}\!\big[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}\big]\,ds,\qquad M_{\delta,k}:=\int_{t_{k}}^{t_{k+1}}\widetilde{{\mathbb{E}}}\!\big|\tilde{L}_{s}\,\Delta\tilde{Z}_{s}\big|\,ds,$
we get
$\Delta Y_{t_{k}}={\mathbb{E}}^{L}\!\big[X_{k}\,\big|\,\mathcal{F}_{t_{k}}\big]-\int_{t_{k}}^{t_{k+1}}\widetilde{{\mathbb{E}}}\!\big[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}\big]\,ds.$
Using $(a+b)^{4}\leq 8(a^{4}+b^{4})$ and Jensen’s inequality,
${\mathbb{E}}^{L}\!|\Delta Y_{t_{k}}|^{4}$ $\leq 8\,{\mathbb{E}}^{L}|X_{k}|^{4}+8\,{\mathbb{E}}^{L}\Big|\textstyle\int_{t_{k}}^{t_{k+1}}\widetilde{{\mathbb{E}}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds\Big|^{4}\leq 8\,{\mathbb{E}}^{L}|X_{k}|^{4}+8\,M_{\delta,k}^{4}$
$\leq 8(8\,{\mathbb{E}}^{L}|\Delta Y_{t_{k+1}}|^{4}+8\,M_{\delta,k}^{4})+8\,M_{\delta,k}^{4}$
$=64\,{\mathbb{E}}^{L}\!|\Delta Y_{t_{k+1}}|^{4}\;+\;72\,M_{\delta,k}^{4}$
By repeating the bound for $M_{\delta}$ derived on a single slice (with $\Delta\xi$ replaced by the terminal datum $\Delta Y_{t_{k+1}}$), there exists a constant $C_{\delta}(I_{k})$ (depending only on $K_{z},K_{\tilde{z}},K_{0}$ and the BMO energy/RHI/BDG constants) such that
$M_{\delta,k}\;\leq\;C_{\delta}(I_{k})\,\Big({\mathbb{E}}^{L}\big[|\Delta Y_{t_{k+1}}|^{4}\big]\Big)^{1/4}.$
Therefore
$\Big({\mathbb{E}}^{L}|\Delta Y_{t_{k}}|^{4}\Big)^{1/4}\;\leq\;\alpha_{\delta}(I_{k})\;\Big({\mathbb{E}}^{L}\!\big[|\Delta Y_{t_{k+1}}|^{4}\big]\Big)^{1/4},\qquad\alpha_{\delta}(I_{k}):=\big(64+72\,C_{\delta}(I_{k})^{4}\big)^{1/4}.$ (4.32)
Iterating gives
$\Big(\mathbb{E}^{L}|\Delta Y_{t_{k+1}}|^{4}\Big)^{1/4}\ \leq\ \prod\limits_{i=k}^{N-k-1}\alpha_{\delta}(I_{i+1})\;\Big(\mathbb{E}^{L}|\Delta\xi|^{4}\Big)^{1/4}.$ (4.33)
Summing (4.31) over $k$ and using (4.33),
$\mathbb{E}^{L}\!\Big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{2}\Big]+\mathbb{E}^{L}\!\int_{0}^{T}|\Delta Z_{s}|^{2}\,ds$ $\leq\sum_{k=0}^{N-1}\left(\mathbb{E}^{L}\!\Big[\sup_{u\in I_{k}}|\Delta Y_{u}|^{2}\Big]+\mathbb{E}^{L}\!\int_{I_{k}}|\Delta Z_{s}|^{2}\,ds\right)$
$\leq\left(\sum_{k=0}^{N-1}C_{\mathrm{loc}}(I_{k})\,\prod\limits_{i=k}^{N-k-1}\alpha_{\delta}(I_{i+1})\right)\Big(\mathbb{E}^{L}|\Delta\xi|^{4}\Big)^{1/4}$
$=:C_{\mathrm{global}}\Big(\mathbb{E}^{L}|\Delta\xi|^{4}\Big)^{1/4}.$
Finally, the Girsanov kernel is a BMO martingale, the RHI transfers this global $\mathbb{P}^{L}$ bound back to $\mathbb{P}$, up to a multiplicative factor. This completes the proof. ∎
Remark 4.7.
Theorem 4.6 provides a stability estimate in $\mathbb{S}^{2}\times\mathbb{H}^{2}$. The proof is based on a global linearisation argument and can be generalized to obtain analogous $L^{p}$ stability estimates (for $p>2$) by similar argument. We do not pursue this level of generality here, as the $\mathbb{S}^{2}\times\mathbb{H}^{2}$ control is already sufficient for our purposes in the sense that the Banach space $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ is continuously embedded in $\mathbb{S}^{2}\times\mathbb{H}^{2}$.
In particular, in the existence proof for the main result below, we construct approximating solutions and show that the sequence is Cauchy in $\mathbb{S}^{2}\times\mathbb{H}^{2}$ by applying Theorem 4.6. The limit in $\mathbb{S}^{2}\times\mathbb{H}^{2}$ then automatically belongs to $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$.
5 Proof: existence and uniqueness of solutions
In this section, we establish the main result for the mean-field BSDE (2.5) under Assumptions $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q})$. The proof proceeds in two main steps. First, we prove the existence of a solution when the terminal condition $\xi$ is Malliavin differentiable with bounded derivative. This is achieved in Lemma 5.1 via a truncation and regularization argument. With the help of the estimates for Lipschitz BSDEs derived in Proposition 3.1, this allows us to apply the classical result from Buckdahn2009MFBSDE.
In the second step, we extend this existence result to the case of general bounded terminal conditions by approximating $\xi$ with a sequence of Malliavin differentiable random variables. For each approximating BSDE, existence is guaranteed by Lemma 5.1, and we invoke the stability result from Theorem 4.6 to show that the sequence of solutions is Cauchy in suitable norms. The limit of this sequence is then shown to solve the original mean-field BSDE (2.5). Uniqueness follows directly from the same stability result. The full argument is completed in the proof of the main theorem below.
Lemma 5.1.
Let $(\rm H_{0})$} and $(\rm H_{q})$} holds and suppose $\xi$ is Malliavin differentiable with bounded derivative $\left\|D\xi\right\|_{\mathbb{S}^{\infty}}=\sup_{0\leq t\leq T}\left\|D_{t}\xi\right\|_{L^{\infty}}<\infty\text{ almost surely}$, then the mean-field BSDE admits a solution in $\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}.$
Proof.
We define a sequence of driver functions $(\bar{g}_{k})_{k\in\mathbb{N}}$ as follows:
$\bar{g}_{k}(t,z,\tilde{z}):=g\left(t,\frac{\left|z\right|\wedge k}{\left|z\right|}z,\frac{\left|\tilde{z}\right|\wedge k}{\left|\tilde{z}\right|}\tilde{z}\right),\quad k\in\mathbb{N}.$
Since $g$ satisfies Assumption $(\rm H_{q})$}, it is straightforward to verify that the new driver $\bar{g}_{k}$ is Lipschitz, with
[…]
for all $k\in\mathbb{N}$ and $(t,z,z^{\prime},\tilde{z},\tilde{z}^{\prime})\in[0,T]\times(\mathbb{R}^{d})^{4}$. Thus, the corresponding BSDE with driver $\bar{g}_{k}$ satisfies $(\rm H_{0})$} and $(H_{\mathrm{lip}})$. Then, the mean-field BSDE
$Y_{t}^{k}=\xi+\int_{t}^{T}\int_{\tilde{\Omega}}\bar{g}_{k}(s,Z_{s}^{k},\tilde{Z}_{s}^{k})\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}^{k}\cdot dW_{s},\quad 0\leq t\leq T,$ (5.34)
admits a unique solution $(Y^{k},Z^{k})\in\mathbb{S}^{2}\times\mathbb{H}^{2}$ (see Theorem 3.1 in Buckdahn2009MFBSDE). Moreover, Proposition 3.1 ensures that for each $k\in\mathbb{N}$, the pair $(Y^{k},Z^{k})$ belongs to $\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}$, and the process $Z^{k}$ satisfies a uniform bound independent of the Lipschitz constant of the driver with respect to $z$ and $\tilde{z}$. In particular, there exists a constant $C_{Z}>0$ such that $\left\|Z^{k}\right\|_{\mathbb{S}^{\infty}}\leq C_{Z}$ for all $k\in\mathbb{N}$. Since $(\tilde{Y}^{k},\tilde{Z}^{k})$ are independent copies of $(Y^{k},Z^{k})$, the same uniform bound holds, we can choose $k^{*}\in\mathbb{N}$ large enough so that
$k^{*}>\sup_{k\in\mathbb{N}}\left\|Z^{k}\right\|_{\mathbb{S}^{\infty}}.$
For such $k^{*}$, the truncation is inactive and
$\bar{g}_{k^{*}}(t,Z_{s}^{k^{*}},\tilde{Z}_{s}^{k^{*}})=g(t,Z_{s}^{k^{*}},\tilde{Z}_{s}^{k^{*}}),\quad 0\leq s\leq T.$
Hence, the pair $(Y^{k^{*}},Z^{k^{*}})$ indeed solves the mean-field BSDE (2.5), which completes the proof. ∎
We are now in a position to state the main result of this section. For the reader’s convenience, we restate the theorem below. To complete the proof of following theorem, it remains to extend the existence result from Malliavin differentiable terminal conditions to the general bounded case. Uniqueness then follows naturally from the stability result established in Theorem 4.6.
Theorem 5.2.
Let Assumptions $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q})$ hold. Then, there exists a unique solution $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ to the BSDE (2.5).
Proof.
We begin by approximating the bounded terminal condition $\xi$ with a sequence of Malliavin differentiable random variables of the form $\xi^{n}=\Phi^{n}(W_{t_{1}},\dots,W_{t_{n}})$, where each $\Phi^{n}$ is a bounded continuous function. Since $\xi$ is bounded, the sequence $(\xi^{n})_{n\in\mathbb{N}}$ can be chosen such that $\xi^{n}$ converges to $\xi$ in in any $L^{p}$ space for $p\geq 1$.
To ensure that the approximating BSDEs satisfy Assumption $(\mathrm{H}_{0})$, we truncate each $\Phi^{n}$ by setting $\Phi^{n}:=\Phi^{n}\wedge\left\|\xi\right\|_{L^{\infty}}$. Hence, each $\xi^{n}$ remains bounded, and $(\xi^{n},g)$ satisfies $(\mathrm{H}_{0})$ for all $n\in\mathbb{N}$. By Lemma 5.1, each approximating BSDE with terminal condition $\xi^{n}$ admits a solution $(Y^{n},Z^{n})\in\mathbb{S}^{2}\times\mathbb{S}^{\infty}$, satisfying
$Y_{t}^{n}=\xi^{n}+\int_{t}^{T}\int_{\tilde{\Omega}}g(s,Z_{s}^{n},\tilde{Z}_{s}^{n})\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}^{n}\cdot dW_{s}.$ (5.35)
Furthermore, the stability result in Theorem 4.6 implies that (pick $p=4$), we have
$\|Y^{n}-Y^{m}\|_{\mathbb{S}^{2}}^{2}+\|Z^{n}-Z^{m}\|_{\mathbb{H}^{2}}^{2}\ \leq\ C_{\mathrm{global}}\;\|\xi^{1}-\xi^{2}\|_{L^{4}},$
where $C_{\mathrm{global}}$ depends only on $T$, $K_{z}$, $K_{\tilde{z}}$, and $K_{0}$, thus are independent of $n$ and $m$.
Since $(\xi^{n})_{n\in\mathbb{N}}$ is Cauchy in $L^{4}$, it follows that $(Y^{n},Z^{n})_{n\in\mathbb{N}}$ is also Cauchy in $\mathbb{S}^{2}\times\mathbb{H}^{2}$. As each $(Y^{n},Z^{n})$ lies in the Banach space $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$, we conclude the existence of a limit $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$. Passing to the limit in (5.35), we verify that the pair $(Y,Z)$ indeed solves the original mean-field BSDE (2.5), establishing existence. Uniqueness then follows directly from the stability result in Theorem 4.6, which concludes the proof. ∎
6 Application: Utility Maximization
In this section, we provide an extension of the foundational work of hu2005utility to the mean-field case. Specifically, we explore how a trader’s decision-making process evolves in a financial market influenced by collective behavior. Unlike previous studies that assume fixed constraints, we consider a dynamic setting where a trader’s strategy is shaped not only by individual preferences but also by the actions of others. This interaction leads to a more flexible and realistic model, where part of the trading constraints are given by market conditions rather than being imposed arbitrarily.
This section is structured as follows: In Section 6.1, we introduce the utility maximization problem and the associated mathematical framework. We then extend the given constraint set to incorporate a dependence on the stochastic process. It is important to note that, at this stage, we do not make any assumptions about this process beyond considering it as a general stochastic process. Next, we establish the connection between the utility maximization problem and the BSDE (6.37) with a thoughtfully chosen driver. We will subsequently demonstrate that the process defined earlier for the constraint set is precisely the solution $Z$ in the BSDE (6.37). With the complete framework in place, we offer a financial interpretation of the components, which are discussed in Remarks 6.4.
To connect this application to our earlier analysis about BSDE, we generalize the separately quadratic driver from BSDE (2.5) to fully coupled quadratic driver from BSDE (6.39) and establish the existence and uniqueness of solutions in Proposition 6.7. To do this we adopte a localization method in 6.6 then we extend it to the global case. Then this is the proof part for BSDE, then we connect this back to the utility maximization problem by presenting the final result in Theorem 6.9, we first derive a key estimate in Lemma 6.3. This estimate is then used in Theorem 6.9 to verify that the driver of BSDE (6.37) satisfies conditions $(\rm H_{0})$} and $(\rm H_{q}^{\prime})$}, ensuring the well-posedness of BSDE (6.39). Leveraging these lemmas, we conclude with the main result in Theorem 6.9, which is supported by the proof of Theorem 7 in hu2005utility.
6.1 Financial market model setup
Consider a financial market composed of $d$ risky stocks. In this market, there are countably infinite number of homogeneous investors who trades dynamically between the risk-free bond and the risky assets. An investor, a representative of homogeneous investors, uses the following model to maximize the utility. The price process of the $i$-th stock, denoted by $S^{i}$, evolves according to the following stochastic differential equation:
$\frac{\mathrm{d}S_{t}^{i}}{S_{t}^{i}}=b_{t}^{i}\,\mathrm{d}t+\sigma_{t}^{i}\,\mathrm{d}W_{t},\quad i=1,\dots,d,$
where $W$ is $d$-dimensional Brownian motion, $b^{i}$ is a real-valued predictable, bounded stochastic process, representing the drift, and $\sigma^{i}$ is an $\mathbb{R}^{d}$-valued predictable bounded stochastic process, representing the volatility. The volatility matrix $\sigma_{t}=(\sigma_{t}^{1},\dots,\sigma_{t}^{d})^{\top}$ is assumed to have full rank $d$.
The market price of risk $\theta_{t}$, also referred to as the risk premium, is defined by:
$\theta_{t}:=\sigma_{t}^{\top}(\sigma_{t}\sigma_{t}^{\top})^{-1}b_{t},\quad t\in[0,T],$
where $\theta_{t}$ is an $\mathbb{R}^{n}$-valued bounded predictable process.
For $1\leq i\leq d$, let $\pi^{i}_{t}$ represent the amount of money invested in stock $i$ at time $t$, meaning the number of shares held in stock $i$ is $\frac{\pi^{i}_{t}}{S^{i}_{t}}$. A predictable process $\pi=(\pi_{t}^{1},\dots,\pi_{t}^{d})_{0\leq t\leq T}\in\mathbb{R}^{d}$, or simply $(\pi_{t})_{0\leq t\leq T}$, is called a trading strategy if $\int_{0}^{\cdot}\pi_{u}\frac{\mathop{}\!dS_{u}}{S_{u}}$ is well-defined, which holds if $\int_{0}^{T}\|\pi_{t}\sigma_{t}\|^{2}\mathop{}\!dt<\infty$ almost surely. The corresponding wealth process $X^{\pi}$, representing the investor’s wealth over time until $T$, is given by:
$X^{\pi}_{t}=x+\sum_{i=1}^{d}\int_{0}^{t}\frac{\pi^{i}_{u}}{S^{i}_{u}}\mathop{}\!dS^{i}_{u}=x+\int_{0}^{t}\pi_{u}\sigma_{u}(\mathop{}\!dW_{u}+\theta_{u}\mathop{}\!du),$
where $x\in\mathbb{R}$ is the initial capital.
At maturity $T$, the investor’s total wealth is determined by both their terminal wealth $X^{\pi}_{T}$ and an $\mathcal{F}_{T}$-measurable random liability $F$, where $F\geq 0$ represents a payment, and $F\leq 0$ represents income. The total wealth at maturity is thus $X^{\pi}_{T}-F$. The investor’s preferences are modeled by an exponential utility function applied to their terminal wealth $X^{\pi}_{T}-F$. The exponential utility function is given by:
$U(x)=-\exp(-\alpha x),\quad x\in\mathbb{R}.$
For the investors other than the representative, we assume the risk aversion parameter $\alpha>0$ are identical. The model for the risky assets are of the same type but with $b,\sigma,W$ being independent and identically distributed to the representative’s model. In addition, we assume the terminal liabilities are independent and has distribution identical to $F$.
6.2 Constraint set and aggregated market signal
To ensure that no arbitrage opportunities exist, the definition of admissible trading strategies imposes constraints on the set of allowable positions. Below, we outline the constraint set in Definition 6.1, which accounts for market-wide influences. To incorporate for the influence of market-wide behavior, we denote $\tilde{Z}$ for the trading strategy adjusted by $\theta/\alpha$, which is an $\mathbb{R}^{1\times d}$-valued stochastic process that satisfies the necessary integrability conditions. More precisely, $\tilde{Z}_{t}\sigma_{t}+\frac{\theta_{t}}{\alpha}$ yields the optimal strategy at time $t\in[0,T]$. In addition, for $t\in[0,T]$ and $\omega\in\Omega$, we define
$Z_{t}(\omega):=\tilde{Z_{t}}(\omega)\sigma_{t}(\omega),$
where $Z_{t}(\omega)$ is an element of $\mathbb{R}^{1\times n}$. These two stochastic processes are essentially equivalent because the matrix $\sigma$ has full rank. Later, we will show this $Z$ will precisely corresponds to the martingale density process of BSDE (6.37). We now present the following expressions for the constraint set under different settings. Specifically, we denote the constraint without market influence by $\tilde{D}\subseteq\mathbb{R}^{d}$, where $\tilde{D}$ is a closed set. When market influence depending on $\mathbb{E}[\tilde{Z}$] is present, the strategy $\pi_{t}(\omega)$ is constrained to $\tilde{C}(\mathbb{E}[\tilde{Z}_{t}])$, where $\tilde{C}(\mathbb{E}[\tilde{Z}_{t}])\subseteq\mathbb{R}^{d}$ is also a closed set.
The expectation of $\tilde{Z}$ is referred to as the aggregated market signal, representing the expected collective adjusted trading strategy of all market participants. This signal plays a crucial role in defining the feasible trading region (constraint set), as described below.
Definition 6.1 (Mean-field adaptive constraint set).
Let $\tilde{D}$ be a closed set in $\mathbb{R}^{d}$, and define the mean-field adaptive constraint set as:
$\tilde{C}(\mathbb{E}[\tilde{Z}_{t}]):=\tilde{D}\odot L(\mathbb{E}[\tilde{Z}_{t}]),\quad t\in[0,T],$ (6.36)
where $\odot$ denotes the Hadamard product (component-wise multiplication) and $L(\mathbb{E}[\tilde{Z}_{t}])$ is a linear mapping depending on the value of $\mathbb{E}[\tilde{Z}_{t}]$ for fixed $t\in[0,T]$, representing how the trader adjusts decisions based on the market’s behavior, defined by:
$L(\mathbb{E}[\tilde{Z}_{t}]):=(\gamma\odot\mathbb{E}[\tilde{Z}_{t}])+\eta,$
where $\gamma:=(\gamma_{1},\dots,\gamma_{d})$ is the market sensitivity vector and $\eta\in\mathbb{R}^{1\times d}$ is the market shift vector, with both being constant vectors.
With the help of the previous definitions, for a fixed $t\in[0,T]$, we start with a set $\tilde{D}\in\mathbb{R}^{d}$, which represents certain constraints, such as risk limits or position constraints. Upon incorporating information from the market, we could apply scaling and shifting to this set, which now reflects both the original constraints and the market’s influence. This results in a new set that is much more flexible and realistic, capturing the dynamic behavior of the market. Therefore, we denote this updated set as $\tilde{C}(\mathbb{E}[\tilde{Z}_{t}])$, emphasizing its dependence on the aggregated market signal $\mathbb{E}[\tilde{Z_{t}}$] at time $t\in[0,T]$.
Definition 6.2 (Admissible strategies with constraints $\tilde{\mathcal{A}}$).
Let $\tilde{C}(\mathbb{E}[\tilde{Z}_{t}])$ be a closed set in $\mathbb{R}^{d}$ for fixed $t\in[0,T]$. The set of admissible trading strategies $\tilde{\mathcal{A}}$ consists of all $\mathbb{R}^{d}$-valued predictable processes $\pi$ satisfying $\pi_{t}\sigma_{t}\in L^{2}[0,T]$, $\pi_{t}\in\tilde{C}(\mathbb{E}[\tilde{Z}_{t}])$ a.s. $\forall t\in[0,T]$, and
[…]
forms a uniformly integrable family.
The investor’s objective is to choose an admissible self-financing trading strategy $\pi^{*}$ to maximize the expected utility of the net wealth at maturity $T$:
$V(0,x):=\sup_{\pi\in\tilde{\mathcal{A}}}\mathbb{E}\left[-\exp\left(-\alpha\left(x+\int_{0}^{T}\pi_{u}\frac{dS_{u}}{S_{u}}-F\right)\right)\right],$
where $V(0,\cdot)$ is called the value function at initial time 0. Solving this optimization problem requires selecting an admissible set from which we select the optimal trading strategy $\pi^{*}$.
To make this notation more compact, we introduce the following notations. For $t\in[0,T],\omega\in\Omega$, define the set $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)\subseteq\mathbb{R}^{n}$ and (volatility-scaled) trading strategy $p_{t}$ such that
$C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right):=\tilde{C}(\mathbb{E}[\tilde{Z}_{t}])\,\sigma_{t}(\omega)\quad\text{and}\quad p_{t}:=\pi_{t}\sigma_{t}.$
Lemma 6.3.
For every $(\omega,t)$, the set $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)$ is closed, and we have the following estimate:
$\max\{|c|:c\in C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)\}\leq K_{1}\left(1+\mathbb{E}\left|Z_{t}\right|\right)\quad\text{a.s., for all }t\in[0,T],$
with some constant $K_{1}\geq 0$.
Proof.
Since the market sensitivity vector $\gamma$ and the market shift vector $\eta$ are both constant, the process $\sigma$ is uniformly bounded, and $\tilde{D}$ is closed, the result follows. ∎
Note that the admissible trading strategy set $\mathcal{A}$, which is equivalent to $\tilde{\mathcal{A}}$ (since $\sigma$ has full rank), consists of all $\mathbb{R}^{n}$-valued predictable processes $p$ that satisfy $p\in L^{2}[0,T]$ and $p_{t}(\omega)\in C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)$ almost surely for all $t\in[0,T]$. Additionally, the family
[…]
is uniformly integrable. Finally, the dependency of the set $\tilde{C}(\mathbb{E}[\tilde{Z}_{t}])$ on the process $(\tilde{Z}_{t})_{0\leq t\leq T}$ is encoded by the dependency of the set $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)$ on the process $(Z_{t})_{0\leq t\leq T}$. We keep the notation $X^{p}$ for the wealth process corresponding to investing strategy $p$.
With the help of these notations, we can rewrite the problem as
$\sup_{p\in\mathcal{A}}\mathbb{E}\left[-\exp\left(-\alpha\left(x+\int_{0}^{T}p_{t}(\mathop{}\!dW_{t}+\theta_{t}\mathop{}\!dt)-F\right)\right)\right]=\mathbb{E}\left[-\exp(-\alpha(X_{T}^{p^{*}}-F))\right]$
for the optimal control $p^{*}\in\mathcal{A}$. Let us introduce a family of processes $\left\{V^{(p)}(\cdot,x):p\in\mathcal{A}\right\}$ as follows: let
$V^{(p)}(t,x):=-\exp(-\alpha(X_{t}^{p}-Y_{t})),\quad t\in[0,T],\,p\in\mathcal{A},$
where $(Y,Z)$ is a solution of the BSDE
$Y_{t}=F+\int_{t}^{T}f(s,Z_{s},\mathbb{E}[Z_{s}])\,ds-\int_{t}^{T}Z_{s}\cdot dW_{s},\quad t\in[0,T],.$ (6.37)
Here, the driver $f$ to be chosen such that $V^{(p)}(\cdot,x)$ is a supermartingale for all $p\in\mathcal{A}$, and there exists $p^{*}\in\mathcal{A}$ such that $V^{(p^{*})}(\cdot,x)$ is a true martingale. If we are able to choose such $f$, which will be shown later, we have
$\mathbb{E}V^{(p)}(T,x)\leq V^{(p)}(0,x)=V^{(p^{*})}(0,x)=\mathbb{E}V^{(p^{*})}(T,x).$
for all $p\in\mathcal{A}$ and the optimal control $p^{*}\in\mathcal{A}$.
Earlier, we described the process $(Z_{t})_{0\leq t\leq T}$ simply as a stochastic process. However, in this context, we now define it as the martingale density process $(Z_{t})_{0\leq t\leq T}$ for the BSDE in equation (6.37). Consequently, as the constraint set depends on $\mathbb{E}[Z_{t}]$, the driver has dependence on $\mathbb{E}[Z_{t}]$: see (6.38).
Let us specify the driver $f$ for the construction of $\left\{V^{(p)}(\cdot,x):p\in\mathcal{A}\right\}$. By Itô’s formula,
$V^{(p)}(t,x)=V^{(p)}(0,x)\tilde{A}^{p}\mathcal{E}_{t}\left(\int_{0}^{\cdot}\alpha(Z_{s}-p_{s})\cdot\mathop{}\!dW_{s}\right),$
where
$\tilde{A}^{p}=-\exp\left(\int_{0}^{t}\alpha v(s,p_{s},Z_{s})\mathop{}\!ds\right)$
and
$v(t,p_{t},Z_{t})=f(t,Z_{t},\mathbb{E}[Z_{t}])-p_{t}\theta_{t}+\frac{\alpha}{2}\left|p_{t}-Z_{t}\right|^{2}.$
Note that
$\mathcal{E}_{t}\left(\int_{0}^{\cdot}\alpha(Z_{s}-p_{s})\cdot\mathop{}\!dW_{s}\right)$
is a true martingale as $\alpha(Z-p)$ is square integrable. We require $\tilde{A}^{p}$ to be a non-increasing process such that $v(t,p_{t},Z_{t})\geq 0$ for all $p\in\mathcal{A}$, and there exists a $p^{*}\in\mathcal{A}$ such that $v(t,p^{*}_{t},Z_{t})=0$. Therefore, if we choose
$f(t,Z_{t},\mathbb{E}[Z_{t}])=-\frac{\alpha}{2}\text{dist}^{2}\left(Z_{t}+\frac{\theta_{t}}{\alpha},C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)\right)+Z_{t}\theta_{t}+\frac{\left|\theta_{t}\right|^{2}}{2\alpha},$ (6.38)
then we have
$v(t,p_{t},Z_{t})$ $=f(t,Z_{t},\mathbb{E}[Z_{t}])+\frac{\alpha}{2}\left|p_{t}\right|^{2}+\frac{\alpha}{2}\left|Z_{t}\right|^{2}-\alpha p_{t}\left(Z_{t}+\frac{\theta_{t}}{\alpha}\right)$
$=f(t,Z_{t},\mathbb{E}[Z_{t}])+\frac{\alpha}{2}\left|p_{t}-\left(Z_{t}+\frac{\theta_{t}}{\alpha}\right)\right|^{2}-Z_{t}\theta_{t}-\frac{\left|\theta_{t}\right|^{2}}{2\alpha}$
$=\frac{\alpha}{2}\left|p_{t}-\left(Z_{t}+\frac{\theta_{t}}{\alpha}\right)\right|^{2}-\frac{\alpha}{2}\text{dist}^{2}\left(Z_{t}+\frac{\theta_{t}}{\alpha},C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)\right)\geq 0.$
Thus, a family of processes $\left\{V^{(p)}(\cdot,x):p\in\mathcal{A}\right\}$ becomes a colletion of supermartingales where, for $p_{t}^{*}\in\Pi_{C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)}\left(Z_{t}+\frac{\theta_{t}}{\alpha}\right)$, we have $v(\cdot,p^{*},Z)=0$ for every $t\in[0,T]$. Therefore, $V^{(p^{*})}(\cdot,x)$ becomes a true martingale.
We have now completed the construction, and it is clear how the utility maximization problem connects to some types of BSDEs. In hu2005utility, the optimization problem was solved under the assumption that the random endowment $F$ is bounded and that the trading strategy $p$ takes values in the admissible set $\mathcal{A}$. However, this set differs from ours because, in their case, the constraint set for trading strategies is simply closed in $\mathbb{R}^{d}$. In contrast, we introduce a particular constraint set where the trader’s behavior depends on the market (i.e., the behavior of other traders) in a mean-field sense. Specifically, we expand the driver to include the expectation of the control process $Z$. Thus, we define a constraint set that depends on time, reflecting the market’s influence at each moment. This approach introduces greater flexibility and a more realistic model. To establish the connection between the control process $Z$ in the BSDE (6.37) and the trading strategy $p\in\mathcal{A}$, we provide the following remarks, which bridge the mathematical framework with financial interpretations.
Remark 6.4.
If we take $\tilde{D}$ to be the entire space, i.e., $\tilde{D}=\mathbb{R}^{d}$, then the resulting constraint set $\tilde{C}(\mathbb{E}[\tilde{Z}_{t}])$ is simply $\mathbb{R}^{d}$ as well. Consequently, we obtain $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)=\mathbb{R}^{n}$, implying the absence of constraints. In this case, the optimal trading strategy is given by
$p_{t}^{*}=Z_{t}+\frac{\theta_{t}}{\alpha},\quad\text{a.s., for all }t\in[0,T].$
Thus, the process $Z_{t}=p_{t}^{*}-\frac{\theta_{t}}{\alpha}$ can be interpreted as the unconstrained trading strategy adjusted by $\frac{\theta_{t}}{\alpha}$ at time $t\in[0,T]$. In general, the constraint set $C$ depends on the average of the adjusted trading strategies among the market participants. The convergence is guaranteed by the homogeneity and exchangeability of market participants.
6.3 Existence and uniqueness for BSDEs with aggregated market signal
We are now ready to present the final result. We first borrow some ideas from tevzadze2008solvability, under small terminal value and lipschitz coefficient, the mean-field BSDE 6.39 admits a unique adapted solution. Next, in Proposition 6.7 we demonstrate that the driver in BSDE 6.37 guarantees the existence and uniqueness of solution. Finally, we complete this section by presenting the maximum expected utility value and corresponding trading strategy in in Theorem 6.9.
For the purpose of the final result, we introduce the BSDE (6.39) with the same formulation and the same $(\rm H_{0})$} in the main result but with a fully coupled quadratic growth condition and assume the following BSDE with driver $g$ and terminal condition $\xi$, satisfying the conditions specified below:
$Y_{t}=\xi+\int_{t}^{T}\int_{\tilde{\Omega}}g(s,Z_{s},\tilde{Z}_{s})\mathop{}\!d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}\cdot\mathop{}\!dW_{s},\quad t\in[0,T].$ (6.39)
  • $(\rm H_{q}^{\prime})$}
    There exist constants $L_{z}>0$, and $K_{z}>0$ such that:
[…]
    for any $(t,z,\tilde{z},z^{\prime},\tilde{z}^{\prime})\in[0,T]\times[\mathbb{R}^{d}\times\mathbb{R}^{d}]^{2}$.
This is the fully coupled quadratic case, in contrast with the separately quadratic $(\rm H_{q})$} of the previous sections. We provide an explicit example here.
Example 6.5.
For simplicity, consider the one-dimensional case $d=1$ and define
$g(t,z,\tilde{z}):=z\,\tilde{z},\qquad(t,z,\tilde{z})\in[0,T]\times\mathbb{R}^{2}.$
This $g$ satisfies the fully coupled quadratic condition $(H_{q}^{\prime})$, but does not satisfy the separately quadratic condition $(\rm H_{q})$}.
We begin by establishing existence and uniqueness of solutions on a sufficiently small time interval in Lemma 6.6, where the required smallness conditions for the terminal value and time horizon are satisfied. By concatenating these local solutions, we extend the result to the global interval $[0,T]$ in Proposition 6.7. First, we note that the terminal condition in BSDE 6.37 is $F$, representing a random liability. It is not realistic to assume $F$ is close to zero. Instead, we relax this requirement and require that $F$ is close to its mean. Accordingly, we rewrite the general BSDE 6.39 in terms of the centered random variable $\eta$:
$U_{t}=\eta+\int_{t}^{T}\int_{\tilde{\Omega}}\big[g(s,Z_{s},\tilde{Z}_{s})-g(s,0,0)\big]\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}\cdot dW_{s},\qquad 0\leq t\leq T,$ (6.40)
where
$U_{t}:=Y_{t}-g(t,0,0)-\mathbb{E}[\xi],\qquad\eta:=\xi-\mathbb{E}[\xi].$
Lemma 6.6.
Let $T<(64L_{z}^{2})^{-1}$ and assume that the driver $g$ satisfies  $(\mathrm{H}_{q}^{\prime})$ and terminal condition $\eta$ be a bounded terminal condition such that
$\left\|\eta\right\|_{L^{\infty}}\leq\rho,$
where
$\rho^{2}:=\frac{1-64L_{z}^{2}T}{65K_{z}}.$
Then the BSDE 6.39 admits a unique solution $(U,Z)\in\mathcal{B}_{\rho}$, where
$\mathcal{B}_{\rho}:=\left\{(U,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}:\left\|U\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|Z\right\|_{BMO}^{2}\leq\rho^{2}\right\}.$
Proof.
Step 1. We first establish existence and uniqueness for the auxiliary BSDE:
$U_{t}=\eta+\int_{t}^{T}\int_{\tilde{\Omega}}g(s,Z_{s},\tilde{Z}_{s})-g(s,0,0)\,d\tilde{\mathbb{P}}ds-\int_{t}^{T}Z_{s}\cdot dW_{s},\qquad 0\leq t\leq T.$
For each $(u,z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$, with $\tilde{z}$ a copy of $z$, define $(U,Z)=\Gamma(u,z)$ to be the solution to
$U_{t}=\eta+\int_{t}^{T}\int_{\tilde{\Omega}}g(s,z_{s},\tilde{z}_{s})-g(s,0,0)\,d\tilde{\mathbb{P}}ds-\int_{t}^{T}Z_{s}\cdot dW_{s}.$
Applying Itô’s formula to $|U_{t}|^{2}$ on $[t,T]$, we obtain
$|U_{t}|^{2}+\int_{t}^{T}|Z_{s}|^{2}ds=|\eta|^{2}+\int_{t}^{T}2U_{s}\int_{\tilde{\Omega}}g(s,z_{s},\tilde{z}_{s})-g(s,0,0)\,d\tilde{\mathbb{P}}ds-2\int_{t}^{T}U_{s}\cdot Z_{s}dW_{s}.$
Taking the conditional expectation and applying $2ab\leq\frac{1}{2}a^{2}+2b^{2}$, we deduce
$|U_{t}|^{2}+\mathbb{E}\left[\int_{t}^{T}|Z_{s}|^{2}ds\mid\mathcal{F}_{t}\right]$ $\leq\left\|\eta\right\|_{L^{\infty}}^{2}+2\left\|U\right\|_{\mathbb{S}^{\infty}}\mathbb{E}\left[\int_{t}^{T}\Big|\int_{\tilde{\Omega}}g(s,z_{s},\tilde{z}_{s})-g(s,0,0)d\tilde{\mathbb{P}}\Big|ds\mid\mathcal{F}_{t}\right]$
$\leq\left\|\eta\right\|_{L^{\infty}}^{2}+\frac{1}{2}\left\|U\right\|_{\mathbb{S}^{\infty}}^{2}+2\mathbb{E}\left[\int_{t}^{T}\int_{\tilde{\Omega}}|g(s,z_{s},\tilde{z}_{s})-g(s,0,0)|^{2}d\tilde{\mathbb{P}}ds\Big|\mathcal{F}_{t}\right].$ (6.41)
Together with the $(\rm H_{q}^{\prime})$} condition and Jensen’s inequality, we have
$\left|\int_{\tilde{\Omega}}g(s,z_{s},\tilde{z}_{s})-g(s,0,0)d\tilde{\mathbb{P}}\right|\leq L_{z}\big(|z_{s}|+\mathbb{E}|z_{s}|\big)+2K_{z}\big(|z_{s}|^{2}+\mathbb{E}|z_{s}|^{2}\big).$
Thus,
$\mathbb{E}\left[\int_{t}^{T}\left|\int_{\tilde{\Omega}}g(s,z_{s},\tilde{z}_{s})-g(s,0,0)d\tilde{\mathbb{P}}\right|ds\mid\mathcal{F}_{t}\right]$ (6.42)
$\leq\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{z}|z_{s}|^{2}ds\mid\mathcal{F}_{t}\right]+\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{z}|z_{s}|^{2}ds\right].$
Combining (6.41) and (6.42) gives
$\frac{1}{2}\left\|U\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|Z\right\|_{BMO}^{2}$
$\leq\left\|\eta\right\|_{L^{\infty}}^{2}+2\operatorname*{ess\,sup}_{(t,\omega)\in[0,T]\times\Omega}\left[\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{z}|z_{s}|^{2}ds\mid\mathcal{F}_{t}\right]+\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{z}|z_{s}|^{2}ds\right]\right]^{2}$
$\leq\left\|\eta\right\|_{L^{\infty}}^{2}+4\operatorname*{ess\,sup}_{(t,\omega)\in[0,T]\times\Omega}\left(\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{z}|z_{s}|^{2}ds\mid\mathcal{F}_{t}\right]^{2}+\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{z}|z_{s}|^{2}ds\right]^{2}\right)$
$\leq\left\|\eta\right\|_{L^{\infty}}^{2}+16\left(L_{z}^{2}T\left\|z\right\|_{BMO}^{2}+4K_{z}^{2}\left\|z\right\|_{BMO}^{4}\right).$
The last inequality uses Jensen’s inequality:
$\left(\int_{t}^{T}L_{z}|z_{s}|\,ds\right)^{2}\leq(T-t)\left(\int_{t}^{T}L_{z}^{2}|z_{s}|^{2}\,ds\right).$
Now, using $a^{2}+b^{2}\leq(|a|+|b|)^{2}$, it follows that
$\left\|U\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|Z\right\|_{BMO}^{2}$ $\leq 2\left\|U\right\|_{\mathbb{S}^{\infty}}^{2}+4\left\|Z\right\|_{BMO}^{2}$
$\leq 4\left\|\eta\right\|_{L^{\infty}}^{2}+\alpha^{2}(\left\|u\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|z\right\|_{BMO}^{2})+\beta^{2}(\left\|u\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|z\right\|_{BMO}^{2})^{2},$
where $\alpha:=8L_{z}\sqrt{T}$, $\beta:=16K_{z}$. To obtain a uniform bound, set $R>0$ such that
$4\left\|\eta\right\|_{L^{\infty}}^{2}+\alpha^{2}R^{2}+\beta^{2}R^{4}\leq R^{2},$
which holds provided $\alpha^{2}-1<0$ (that is, $T<(64L_{z}^{2})^{-1}$), and we require
$\left\|\eta\right\|_{L^{\infty}}\leq\frac{1-\alpha^{2}}{4\beta}.$ (6.43)
In this case, we may take
$R^{2}=\frac{8}{(1-\alpha^{2})}\left\|\eta\right\|_{L^{\infty}}^{2}.$ (6.44)
The ball
$\mathcal{B}_{R}:=\left\{(U,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{BMO}:\left\|U\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|Z\right\|_{BMO}^{2}\leq R^{2}\right\}$
is then such that $\Gamma(\mathcal{B}_{R})\subset\mathcal{B}_{R}$.
Step 2. Next, we prove that $\Gamma$ is a contraction on $\mathcal{B}_{R}$. Let $(u^{i},z^{i})\in\mathcal{B}_{R}$ ($i=1,2$), with $(U^{i},Z^{i})=\Gamma(u^{i},z^{i})$, and set
$\Delta u=u^{1}-u^{2},\quad\Delta z=z^{1}-z^{2},\quad\Delta U=U^{1}-U^{2},\quad\Delta Z=Z^{1}-Z^{2}.$
A similar computation as above yields
$\frac{1}{2}\left\|\Delta U\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|\Delta Z\right\|_{BMO}^{2}$
$\leq 2\,\operatorname*{ess\,sup}_{(t,\omega)\in[0,T]\times\Omega}\left[\mathbb{E}\left(\int_{t}^{T}|g(s,z^{1}_{s},\tilde{z}^{1}_{s})-g(s,z^{2}_{s},\tilde{z}^{2}_{s})|ds\mid\mathcal{F}_{t}\right)\right]^{2}$
$\leq 4L_{z}^{2}T\,\operatorname*{ess\,sup}_{(t,\omega)}\mathbb{E}\left[\int_{t}^{T}|\Delta z_{s}|^{2}+\mathbb{E}|\Delta z_{s}|^{2}ds\mid\mathcal{F}_{t}\right]$
$\quad+16K_{z}^{2}\,\operatorname*{ess\,sup}_{(t,\omega)}\mathbb{E}\left[\int_{t}^{T}(|z^{1}_{s}|^{2}+|z^{2}_{s}|^{2}+\mathbb{E}|z^{1}_{s}|^{2}+\mathbb{E}|z^{2}_{s}|^{2})ds\mid\mathcal{F}_{t}\right]\mathbb{E}\left[\int_{t}^{T}(|\Delta z_{s}|+\mathbb{E}|\Delta z_{s}|)^{2}ds\mid\mathcal{F}_{t}\right]$
$\leq\left(8L_{z}^{2}T+32K_{z}^{2}(\left\|z^{1}\right\|_{BMO}^{2}+\left\|z^{2}\right\|_{BMO}^{2})\right)\left(\left\|\Delta u\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|\Delta z\right\|_{BMO}^{2}\right).$
Since $\left\|z^{i}\right\|_{BMO}^{2}\leq R^{2}$ for $i=1,2$, we obtain
$\left\|\Delta U\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|\Delta Z\right\|_{BMO}^{2}\leq(16L_{z}^{2}T+128K_{z}^{2}R^{2})\left(\left\|\Delta u\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|\Delta z\right\|_{BMO}^{2}\right).$
Now, recall from the previous step that $R^{2}=\frac{8}{(1-\alpha^{2})}\left\|\eta\right\|_{L^{\infty}}^{2}$. In order for $R^{2}$ to be positive, we require $T<(64L_{z}^{2})^{-1}$, which is more restrictive than the condition here. Furthermore, if
$\left\|\eta\right\|_{L^{\infty}}^{2}<\frac{(4-\alpha^{2})(1-\alpha^{2})}{16\beta^{2}},$ (6.45)
then it follows that
$16L_{z}^{2}T+128K_{z}^{2}R^{2}<1,$
and thus the mapping $\Gamma$ is a contraction and admits a unique fixed point in $\mathcal{B}_{R}$.
Step 3. Finally, together with the local condition $T<(64L_{z}^{2})^{-1}$, the smallness condition (6.43) is in fact more stringent than (6.45). Consequently, if we define $\rho>0$ by
$\rho^{2}:=\frac{1-64L_{z}^{2}T}{65K_{z}},$
then, provided that $\left\|\eta\right\|_{L^{\infty}}\leq\rho$, the general BSDE admits a unique adapted solution $(U,Z)\in\mathcal{B}_{\rho}$ on this local time interval, under the imposed smallness conditions on both the terminal value and the time horizon. This completes the proof of local existence and uniqueness. ∎
Proposition 6.7.
Let $T>0$, and assume that $g$ satisfies $(\mathrm{H}_{q}^{\prime})$. Suppose there exists $\rho>0$ such that
$\left\|\eta\right\|_{L^{\infty}}\leq\rho,$
where
$\rho^{2}=\frac{1}{65^{2}K_{z}}.$
Then the BSDE (6.39) admits a unique solution $(U,Z)\in\mathcal{B}_{\rho}$ on $[0,T]$, where
$\mathcal{B}_{\rho}:=\Bigl\{(U,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}:\ \|U\|_{\mathbb{S}^{\infty}}^{2}+\|Z\|_{BMO}^{2}\leq\rho^{2}\Bigr\}.$
Proof.
Let $\delta:=(65L_{z}^{2})^{-1}<(64L_{z}^{2})^{-1}$, and divide $[0,T]$ into $N$ subintervals of length at most $\delta$: $0=t_{0}<t_{1}<\cdots<t_{N}=T$, where $t_{k+1}-t_{k}\leq\delta$ for all $k$. By construction, $\rho^{2}=\frac{1-64L_{z}^{2}\delta}{65K_{z}}=\frac{1}{65^{2}K_{z}}$. We construct the solution recursively, starting from the last interval $t_{N-1},t_{N}]$. On this interval, the terminal value is $\eta$, and the smallness condition is satisfied by assumption. Thus, by applying the local existence and uniqueness lemma, we obtain a unique solution $(U,Z)$ on $t_{N-1},t_{N}]$ with a priori bound in $\mathcal{B}_{\rho}$. Then we proceed recursively: at each interval $t_{k-1},t_{k}]$, we solve the BSDE using $U_{t_{k}}$ as the terminal value. The a priori bound $\|U\|_{\mathbb{S}^{\infty}}\leq\rho$ ensures that the smallness condition $\|U_{t_{k}}\|_{L^{\infty}}\leq\rho$ continues to hold at each step. Finally, By concatenating these local solutions, we obtain a unique adapted solution $(U,Z)$ on $[0,T]$. ∎
We remark that, given the unique solution $(U,Z)\in\mathcal{B}_{\rho}$ constructed above for the auxiliary BSDE 6.40, the solution to the original BSDE 6.39 is obtained by the explicit transformation
$Y_{t}=U_{t}+g(t,0,0)+\mathbb{E}[\xi],\qquad 0\leq t\leq T.$
Furthermore, we emphasize that the smallness condition $\left\|\eta\right\|_{L^{\infty}}\leq\rho$, with $\eta:=\xi-\mathbb{E}[\xi]$, requires the terminal condition $\xi$ to be close to its mean. This is a slightly more general requirement than simply demanding that $\xi$ itself be small, as it allows for arbitrary (but fixed) expectations and only controls the oscillation of $\xi$ around its mean.
Remark 6.8.
Proposition 6.7 was established for drivers of the form $\int_{\tilde{\Omega}}g(s,Z_{s},\tilde{Z}_{s})\,d\tilde{\mathbb{P}}$, which naturally extends to the case where the driver depends on the conditional expectation, i.e., $g(s,Z_{s},\mathbb{E}[Z_{s}])$. In particular, the existence and uniqueness result applies to BSDEs of the type
$Y_{t}=\xi+\int_{t}^{T}g(s,Z_{s},\mathbb{E}[Z_{s}])\,ds-\int_{t}^{T}Z_{s}\cdot dW_{s},\quad t\in[0,T],$
which will be useful for analyzing the driver (6.38) in this section.
While the above analysis focuses on drivers with $Z$ and mean-field dependence, it is very easy to extend existence and uniqueness results to more general settings where the driver $g$ may also depend on the process $Y$ and its expectation $\mathbb{E}[Y]$ as in hibon2017MFquadratic. In the present section, however, the application of interest involves drivers depending solely on $Z$ and its expectation, so the above results apply directly. We now present the main result of this section.
Theorem 6.9.
Suppose the liability $F$, the terminal condition of BSDE (6.37) satisfies the condition such that $\left\|F-\mathbb{E}[F]\right\|_{L^{\infty}}\leq\rho$ where
$\rho^{2}=\frac{1}{65^{2}K_{z}}.$
Let $(Y,Z)$ denote the unique solution to BSDE (6.37) with driver $f$ (6.38). Then the value function $V^{(p)}(t,x)$ associated with the utility maximization problem over the admissible set $\mathcal{A}$ and constraint set $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)$ is given by
$V^{(p)}(0,x)=-\exp(-\alpha(x-Y_{0})),$
and there exists an optimal trading strategy $p^{*}\in\mathcal{A}$ such that, for all $t\in[0,T]$ and $\omega\in\Omega$,
$p^{*}_{t}(\omega)\in\Pi_{C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)}\left(Z_{t}+\frac{\theta_{t}}{\alpha}\right).$
Proof.
First we will show the driver $f$ (6.38) satisfies the condition $(\mathrm{H}_{q}^{\prime})$. To verify $(\mathrm{H}_{q}^{\prime})$, we apply Lemma 6.3. For any $\omega\in\Omega$ and $t\in[0,T]$, since $Z_{t}(\omega)\in\mathbb{R}^{d}$, we have
$\text{dist}^{2}\left(Z_{t}+\frac{\theta_{t}}{\alpha},\,C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)\right)$ $\leq 2|Z_{t}|^{2}+2\left(\frac{|\theta_{t}|}{\alpha}+K_{1}\mathbb{E}|Z_{t}|\right)^{2}$
$\leq 2\frac{|\theta_{t}|^{2}}{\alpha^{2}}+\frac{K_{1}|\theta_{t}|}{\alpha}+2|Z_{t}|^{2}+2\left(K_{1}^{2}+\frac{K_{1}|\theta_{t}|}{\alpha}\right)\mathbb{E}|Z_{t}|^{2}.$
Now, for $t\in[0,T]$ and $Z_{t}^{1}(\omega),Z_{t}^{2}(\omega)\in\mathbb{R}^{d}$, we observe
$f(t,Z_{t}^{1},\mathbb{E}[Z_{t}^{1}])-f(t,Z_{t}^{2},\mathbb{E}[Z_{t}^{2}])$
$\qquad=-\frac{\alpha}{2}\left[\text{dist}^{2}\left(Z_{t}^{1}+\frac{\theta_{t}}{\alpha},\,C_{t}(\omega,\mathbb{E}[Z_{t}^{1}])\right)-\text{dist}^{2}\left(Z_{t}^{2}+\frac{\theta_{t}}{\alpha},\,C_{t}(\omega,\mathbb{E}[Z_{t}^{2}])\right)\right]+(Z_{t}^{1}-Z_{t}^{2})\theta_{t}.$
By the Lipschitz property of the distance function from a closed set, it follows that, for all $t\in[0,T]$
$\left|f\left(t,Z_{t}^{1},\mathbb{E}[Z_{t}^{1}]\right)-f\left(t,Z_{t}^{2},\mathbb{E}[Z_{t}^{2}]\right)\right|$
$\quad\leq\frac{\alpha}{2}\left(|Z_{t}^{1}|+|Z_{t}^{2}|+2\frac{|\theta_{t}|}{\alpha}+K_{1}\big(\mathbb{E}|Z_{t}^{1}|+\mathbb{E}|Z_{t}^{2}|\big)\right)\left(|Z_{t}^{1}-Z_{t}^{2}|+K_{1}\mathbb{E}|Z_{t}^{1}-Z_{t}^{2}|\right)+|\theta_{t}|\,|Z_{t}^{1}-Z_{t}^{2}|$
$\quad\leq 2|\theta_{t}|\,|Z_{t}^{1}-Z_{t}^{2}|+\frac{\alpha}{2}(1\vee K_{1})^{2}\left(|Z_{t}^{1}|+|Z_{t}^{2}|+\mathbb{E}|Z_{t}^{1}|+\mathbb{E}|Z_{t}^{2}|\right)\left(|Z_{t}^{1}-Z_{t}^{2}|+\mathbb{E}|Z_{t}^{1}-Z_{t}^{2}|\right),\quad\text{a.s.}$
Together with the boundedness of $\theta$, we obtain, for some constants $c_{1},c_{2}>0$, for all $t\in[0,T]$
$\left|f\left(t,Z_{t}^{1},\mathbb{E}[Z_{t}^{1}]\right)-f\left(t,Z_{t}^{2},\mathbb{E}[Z_{t}^{2}]\right)\right|$
$\quad\leq c_{1}\,|Z_{t}^{1}-Z_{t}^{2}|+c_{2}\left(|Z_{t}^{1}|+|Z_{t}^{2}|+\mathbb{E}|Z_{t}^{1}|+\mathbb{E}|Z_{t}^{2}|\right)\left(|Z_{t}^{1}-Z_{t}^{2}|+\mathbb{E}|Z_{t}^{1}-Z_{t}^{2}|\right),\quad\text{a.s.}$
which is a sufficient condition for $(\rm H_{q}^{\prime})$}. This ensures the existence and uniqueness of a solution $(Y,Z)$ to BSDE (6.37) by Proposition 6.7. It remains to show that for any $p\in\mathcal{A}$, the process
$V^{(p)}(t,x):=-\exp\left(-\alpha(X_{t}^{p}-Y_{t})\right),\quad t\in[0,T],$
is a supermartingale, and that it becomes a martingale when $p=p^{*}$ as defined above. This verification follows directly from arguments in Theorem 7 of hu2005utility, and we omit the details here. ∎
7 Conclusion
In this paper we have developed a general framework for quadratic mean-field BSDEs whose generators exhibit quadratic growth in the control and its law. Working under the separately quadratic structure (Hq) with bounded terminal condition, we proved global existence and uniqueness of solutions in the Banach space $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$, without imposing any smallness condition. The key ingredients were a Malliavin calculus argument for Lipschitz mean-field BSDEs, which yields uniform bounds on the control process, together with BMO estimates and a global stability result that allows us to pass from Lipschitz to quadratic generators.
We then extended the analysis to the fully coupled quadratic regime (Hq), in which the generator admit genuine cross-interaction terms. In this more general setting we established existence and uniqueness under a natural smallness condition on the centred terminal condition. As an application, we showed how the fully coupled theory can be used to formulate and solve a mean-field exponential utility maximization problem, thereby generalizing the classical framework of hu2005utility.

### v2

@@
 Abstract
-We study real-valued mean-field backward stochastic differential equations (BSDEs) of the form
-$Y_{t}=\xi+\int_{t}^{T}\widetilde{\mathbb{E}}\,g(s,Z_{s},\tilde{Z}_{s})\,ds-\int_{t}^{T}Z_{s}\,dW_{s},$
-where $\tilde{Z}$ denotes an independent copy of $Z$ and $\widetilde{\mathbb{E}}$ the expectation with respect to $\tilde{Z}$.
-Under a separately quadratic growth assumption (Hq) on the generator $g$ in (Z,Z~)($\tilde{Z}$), together with a bounded terminal condition, we prove existence and uniqueness of solutions in $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$.
-Our approach departs from classical fixed-point arguments and instead combines Malliavin calculus with refined BMO and stability estimates: we first obtain uniform $\mathbb{S}^{\infty}$-bounds for $Z$ in a Lipschitz setting with bounded Malliavin derivative, and then pass to the quadratic case by approximation, using a stability result in $\mathbb{S}^{2}\times\mathbb{H}^{2}$.
-This closes the gap between the quadratic BSDE results of cheridito2017bse and hao2022mean.
-In the second part of the paper, we extend the framework to generators of the form $g(t,z,\tilde{z},\bar{z})$ satisfying a fully coupled quadratic condition (Hq).
-In this general regime we establish existence and uniqueness under a smallness condition on the centered terminal variable.
-As an application, we solve a mean-field exponential utility maximization problem with a collective liability, thereby generalizing the classical utility maximization framework of hu2005utility to a fully coupled quadratic mean-field setting.
+In this paper, we study a class of real-valued mean-field backward stochastic differential equations (BSDEs) with generator of quadratic growth in the control variable and the mean-field term.
+Under this assumption, together with a bounded terminal condition, we establish existence and uniqueness of solutions.
+Our approach departs from classical fixed-point arguments and instead combines Malliavin calculus with refined BMO and stability estimates.
+The result bridges the gap between the quadratic BSDE results of cheridito2017bse and hao2022mean.
+Moreover, motivated by the structure of the mean-field exponential utility maximization problem introduced in our paper, we extend our framework to generators satisfying a weaker quadratic condition on the generator.
+This relaxation is designed to accommodate the additional mean-field terms that arise in our utility maximization setting and that fall outside the scope of previous quadratic assumptions.
+Within this more general regime, we establish existence and uniqueness of solutions under a smallness condition on the terminal random variable.
+We then apply this extended theory to solve a mean-field exponential utility maximization problem, thereby generalizing the classical framework of hu2005utility to a fully coupled quadratic mean-field setting.
 Key words: mean-field BSDEs; quadratic BSDEs; BMO martingales; Malliavin calculus; exponential utility maximization.
@@
 Consider the following mean-field BSDE:
-$Y_{t}=\xi+\int_{t}^{T}\int_{\tilde{\Omega}}g(s,Z_{s},\tilde{Z}_{s})\mathop{}\!d\tilde{\mathbb{P}}ds-\int_{t}^{T}Z_{s}dW_{s},$ (1.1)
+$Y_{t}=\xi+\int_{t}^{T}\int_{\tilde{\Omega}}g(s,Z_{s},\tilde{Z}_{s})\mathop{}\!d\tilde{\mathbb{P}}ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},$ (1.1)
 […]
 The origin of linear BSDEs can be traced back to their role as adjoint equations in stochastic control problems, first studied by bismut1973conjugate.
-The seminal paper by pardoux1990adapted established the existence and uniqueness of adapted solutions $(Y,Z)$.
-The form of the equation is as follows:
-$Y_{t}=\xi+\int_{t}^{T}F(s,Y_{s},Z_{s})\,ds-\int_{t}^{T}Z_{s}dW_{s},\quad t\in[0,T],$ (1.2)
-under the conditions that the generator $F$ is Lipschitz continuous and $\xi$ is square integrable.
-For a comprehensive presentation of the theory, see the review by el1997backward books by CohenStocal2015 or ZhangBSDEbook2017.
-Numerous studies have aimed to improve existence and uniqueness results by relaxing the Lipschitz condition on the coefficients or the conditions on the terminal value $\xi$.
-Consequently, there has been a growing interest in BSDEs with applications in various fields such as stochastic control, mathematical finance, and partial differential equations.
+The seminal paper by pardoux1990adapted established the existence and uniqueness of adapted solutions $(Y,Z)$ for
+$Y_{t}=\xi+\int_{t}^{T}F(s,Y_{s},Z_{s})\,ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},\quad t\in[0,T],$ (1.2)
+under the conditions that the generator $F$ is Lipschitz continuous and the terminal condition $\xi$ is square integrable.
+Since then, there has been a growing interest in BSDEs with applications in various fields such as stochastic control, mathematical finance, and partial differential equations.
+For a comprehensive presentation of the theory, see el1997backward, CohenStocal2015 or ZhangBSDEbook2017.
+Numerous studies have aimed to improve existence and uniqueness results by relaxing the Lipschitz condition on the generator $F$ or the conditions on the terminal condition $\xi$.
 BSDEs with generators exhibiting quadratic growth in the variable $z$ have attracted significant attention.
-These equations arise, for example, in stochastic linear-quadratic control with random coefficients bismut1976linearquad, in utility optimization problems with exponential utility functions hu2005utility or in risk minimization problems for the entropic risk measure.
-Kobylanski:2000cy first provided the existence and uniqueness result for scalar-valued quadratic BSDEs with bounded terminal values, where $F$ is quadratic with respect to its $Z$ component.
-A detailed review of the literature can be found in el2008bsde.
+These equations arise, for example, in stochastic linear-quadratic control with random coefficients bismut1976linearquad, in utility optimization problems in an incomplete market hu2005utility.
+Kobylanski:2000cy provided the seminal existence and uniqueness result for real-valued quadratic BSDEs with bounded terminal values, where $F$ has quadratic growth with respect to its $Z$ component.
 briand2013simple have continued to explore this area, demonstrating the existence of unique solutions to quadratic BSDEs with bounded terminal conditions using purely probabilistic methods.
-Investigations by briand2006bsde extend the previous results to unbounded unbounded terminal conditions.
-Some studies generalized to the multidimensional case by hu2016multi; tang2023MultiMFBSDE; luo2019triangle; tevzadze2008solvability.
+Investigations by briand2006bsde extend the previous results to unbounded terminal conditions.
+Some studies the multidimensional cases: see hu2016multi; tang2023MultiMFBSDE; luo2019triangle; tevzadze2008solvability.
 Mean-field stochastic differential equations, also known as McKean-Vlasov equations, describe stochastic systems whose evolution is influenced by both the microscopic state and the macroscopic distribution of particles.
@@
 Mean-field game theory began with the groundbreaking contributions of lasry2007mfgame.
-Building on this foundation, Buckdahn2009MFlimit introduced mean-field BSDEs, derived from approximations involving weakly interacting $N$-particle systems, and investigated them using probabilistic methods.
-This has led to a growing interest in mean-field SDEs and BSDEs.
-For instance, Buckdahn2009MFBSDE studied the existence and uniqueness of solutions and a comparison theorem for mean-field BSDEs with Lipschitz continuous generators and square integrable terminal conditions.
-More recently, hao2022mean explored the solvability of local and global solutions for one-dimensional mean-field BSDEs with quadratic growth, while tang2023MultiMFBSDE established the existence and uniqueness of local and global solutions for multidimensional BSDEs with generators depending on the expectation of both variables.
+Building on this foundation, Buckdahn2009MFlimit introduced mean-field BSDEs, derived from approximations involving weakly interacting $N$ particle systems, and investigated them using probabilistic methods.
+This has led to a growing interest in mean-field BSDEs and FBSDEs.
+For instance, Buckdahn2009MFBSDE studied the existence and uniqueness of solutions with Lipschitz continuous generators and square integrable terminal conditions.
+More recently, hao2022mean explored the solvability of solutions for one-dimensional mean-field BSDEs with quadratic growth, while tang2023MultiMFBSDE established the existence and uniqueness of solutions for multidimensional BSDEs with generators depending on the expectation of both variables.
 Li2018MFBSDEcontinuous examined mean-field BSDEs whose coefficients depend not only on the solution $(Y,Z)$ but also on the law of $Y$.
 pardoux1990adapted interpreted the key structural feature of BSDEs as a nonlinear conditional expectation, highlighting their close link to Malliavin calculus.
-In fact, Buckdahn2009MFBSDE states that, in a Clark–Ocone type formula, the martingale density component of the solution pair of a BSDE with a classical globally Lipschitz generator on a Gaussian basis is the Malliavin derivative of the other component (see Proposition 5.3).
-The close relationship between Malliavin calculus and BSDEs stems from the use of Malliavin calculus to prove the regularity of trajectories.
-For example, briand2013simple provided a straightforward approach to the existence and uniqueness of quadratic BSDEs with bounded terminal conditions using Malliavin calculus to overcome difficulties, and Imkeller2010path proposed a numerical scheme for BSDEs with generators of quadratic growth as one of the applications.
+In fact, el1997backward and Buckdahn2009MFBSDE state that, in a Clark–Ocone type formula, the martingale density component $Z$ of the solution pair of a BSDE with a classical globally Lipschitz generator on a Gaussian basis is the Malliavin derivative of the other component $Y$ (see Proposition 5.3).
+The result has been extending our understanding of quadratic BSDEs: for example, briand2013simple provided a straightforward approach to the existence and uniqueness of quadratic BSDEs with bounded terminal conditions using Malliavin calculus to overcome difficulties with quadratic generators, and Imkeller2010path proposed a numerical scheme for BSDEs with generators of quadratic growth.
 Continuous time portfolio optimization under the utility-maximization framework introduced by merton1969portfolio is a foundational topic in mathematical finance.
@@
 Building on these foundations, we bridge the gap between the approaches of cheridito2017bse and hao2022mean.
-Specifically, hao2022mean established local existence and uniqueness by applying a contraction mapping with the Reverse Hölder Inequality (RHI) and iterating backward in time, while cheridito2017bse used the Krasnoselskii fixed-point theorem on a small time interval, iterating backward to obtain a global solution.
-Our contribution extends the results of cheridito2017bse by demonstrating in Proposition 4.9 that the function $f^{1}$ satisfies the conditions set forth in hao2022mean, thereby proving the existence of a solution for quadratic mean-field BSDEs.
+Specifically, hao2022mean established the existence and uniqueness by applying a contraction mapping with the Reverse Hölder Inequality (RHI) and iterating backward in time when the generator is quadratic in $Z$ component, while cheridito2017bse used the Krasnoselskii fixed-point theorem on a small time interval, iterating backward to obtain a global solution.
+Our contribution extends the results of cheridito2017bse by demonstrating in Proposition 4.9 that the function $f^{1}$ satisfies the conditions outlined in hao2022mean, thereby proving the existence and uniqueness of a solution for quadratic mean-field BSDEs, which is quadratic both in $Z$ and the corresponding mean-field component $\tilde{Z}$.
 Our main contributions are as follows.
@@
 A key novelty of this work lies in the inclusion of a quadratic mean-field component $\tilde{Z}$, which distinguishes it from existing results in the literature.
-Moreover, our method is notable for its simplicity, as it avoids the use of fixed-point arguments commonly employed in previous studies, such as juan2023MFBSDE and tevzadze2008solvability.
+Moreover, our method is notable for its simplicity, as it avoids the use of fixed-point arguments commonly employed in previous studies, such as juan2024MFBSDE and tevzadze2008solvability.
 The rest of the paper is organized as follows.
@@
 […]
-By $\left|\cdot\right|$, we denote the Euclidean norm on $\mathbb{R}^{d}$, and for $p>1$, we denote by
-  • $L^{p}$: all real-valued random variables $X$ s.t. $\|X\|_{L^{p}}^{p}:={\mathbb{E}}^{\mathbb{P}}\big[|X|^{p}\big]<\infty$,
+  • $L^{p}$: all real-valued random variables $X$ s.t. $\|X\|_{L^{p}}:=\left({\mathbb{E}}^{\mathbb{P}}\big[|X|^{p}\big]\right)^{\frac{1}{p}}<\infty$,
   • $L^{\infty}$: all essentially bounded random variables $X$ s.t. $\|X\|_{L^{\infty}}:=\operatorname*{ess\,sup}_{\omega\in\Omega}|X(\omega)|<\infty$,
   • $\mathbb{S}^{p}$: the set of real valued $\mathcal{F}$-adapted continuous processes $(Y_{t})_{0\leq t\leq T}$ satisfying
-    $\left\|Y\right\|_{\mathbb{S}^{p}}:=\mathbb{E}\left[\sup_{0\leq r\leq T}\left|Y_{r}\right|^{p}\right]^{\frac{1}{p}}<\infty,$
+    $\left\|Y\right\|_{\mathbb{S}^{p}}:=\left(\mathbb{E}\left[\sup_{0\leq r\leq T}\left|Y_{r}\right|^{p}\right]\right)^{\frac{1}{p}}<\infty,$
   • $\mathbb{S}^{\infty}$: the set of real valued $\mathcal{F}$-adapted continuous processes $(Y_{t})_{0\leq t\leq T}$ satisfying
-    $\left\|Y\right\|_{\mathbb{S}^{\infty}}:=\sup_{\omega\in\Omega}\sup_{0\leq t\leq T}\left|Y_{t}(\omega)\right|<\infty,$
-  • $\mathbb{H}^{p}$: the set of predictable $\mathbb{R}^{d}$-valued process $Z$ satisfying $\left\|Z\right\|_{\mathbb{H}^{p}}:=\mathbb{E}\left[\left(\int_{0}^{T}\left|Z_{r}\right|^{2}\,dr\right)^{\frac{p}{2}}\right]^{\frac{1}{p}}<\infty,$
-  • $\mathbb{H}^{2}_{BMO}$: the subsets of $\mathbb{H}^{2}$ with predictable $\mathbb{R}^{d}$-valued process $Z$ satisfying $\int_{0}^{\cdot}Z_{s}\cdot\,dW_{s}$ is a Bounded in Mean Oscillation (BMO) martingale, that is, there exists a non-negative constant $C$ such that, for each stopping time $\tau\leq T$,
-    $\mathbb{E}\left[\int_{\tau}^{T}\left|Z_{s}\right|^{2}\,ds|\mathcal{F}_{\tau}\right]\leq C^{2}.$
-    The $\left\|\int_{0}^{\cdot}Z_{s}\cdot\,dW_{s}\right\|_{BMO}$ is defined as the smallest non-negative constant $C$ for which the above inequality is satisfied.
-  • Let $C$ be a closed subset of $\mathbb{R}^{n}$ and $a\in\mathbb{R}^{n}$. The distance between $a$ and the set $C$ is given by
+    $\left\|Y\right\|_{\mathbb{S}^{\infty}}:=\sup_{0\leq r\leq T}\left\|Y_{r}\right\|_{L^{\infty}}=\sup_{\omega\in\Omega}\sup_{0\leq r\leq T}\left|Y_{r}(\omega)\right|<\infty,$
+  • $\mathbb{H}^{p}$: the set of predictable $\mathbb{R}^{d}$-valued process $Z$ satisfying $\left\|Z\right\|_{\mathbb{H}^{p}}:=\left(\mathbb{E}\left[\int_{0}^{T}\left|Z_{r}\right|^{2}\,dr\right]^{\frac{p}{2}}\right)^{\frac{1}{p}}<\infty,$
+  • $\mathbb{H}^{2}_{BMO}$: a subset of $\mathbb{H}^{2}$ with $Z$ satisfying $Z\cdot W:=\int_{0}^{\cdot}Z_{s}\!\cdot dW_{s}$ is a Bounded Mean Oscillation (BMO) martingale, that is, there exists a non-negative constant $C$ such that, for each stopping time $\tau\leq T$,
+    $\mathbb{E}\left[\int_{\tau}^{T}\left|Z_{r}\right|^{2}\,dr\Big|\mathcal{F}_{\tau}\right]\leq C^{2}.$
+    The $\left\|Z\right\|_{\mathbb{H}^{2}_{BMO}}:=\left\|\int_{0}^{\cdot}Z_{s}\cdot dW_{s}\right\|_{BMO}$ is defined as the smallest non-negative constant $C$ for which the above inequality is satisfied.
+  • For a closed subset $C$ of $\mathbb{R}^{n}$ and $a\in\mathbb{R}^{n}$, the distance between $a$ and the set $C$ is defined by
     $\text{dist}(a,C)=\min_{b\in C}\left|a-b\right|.$
     The set $\Pi_{C}(a)$ consists of all points $b\in C$ at which this minimum is attained, and we refer to this as the projection set.
-When a different underlying measure is involved, we indicate it explicitly in the notation, for example by writing $\mathbb{E}$$^{\mathbb{Q}}$$\cdot$] and $\|\cdot\|_{BMO(\mathbb{Q})}$ for expectations and BMO norms taken with respect to a probability measure $\mathbb{Q}$.
+When a different underlying measure other than $\mathbb{P}$ is involved, we indicate it explicitly in the notation, for example by writing $\mathbb{E}$$^{\mathbb{Q}}$$\cdot$] and $\|\cdot\|_{BMO(\mathbb{Q})}$ for expectations and BMO norms taken with respect to a probability measure $\mathbb{Q}$.
 Here are some inequalities from BMO martingales theory, all the proofs can be found in kazamaki2006BMO.
@@
 There exist constants $c_{1}(K),c_{2}(K)>0$, depending only on $K$, such that the following holds.
-Let $N$ be a one–dimensional continuous $\mathrm{BMO}$ martingale under $\mathbb{P}$ with $\|N\|_{\mathrm{BMO}(\mathbb{P})}\leq K$, and define the equivalent measure $\tilde{\mathbb{P}}$ by
+Let $N$ be a one-dimensional continuous $\mathrm{BMO}$ martingale under $\mathbb{P}$ with $\|N\|_{\mathrm{BMO}(\mathbb{P})}\leq K$, and define the equivalent measure $\tilde{\mathbb{P}}$ by
 $\frac{d\tilde{\mathbb{P}}}{d\mathbb{P}}\;=\;\mathcal{E}(N)_{T}.$
@@
 $c_{1}(K)\,\|M\|_{\mathrm{BMO}(\mathbb{P})}\;\leq\;\|\tilde{M}\|_{\mathrm{BMO}(\tilde{\mathbb{P}})}\;\leq\;c_{2}(K)\,\|M\|_{\mathrm{BMO}(\mathbb{P})}.$ (2.4)
+Definition 2.3 (sliceability).
+Let $M$ be a one–dimensional continuous $\mathrm{BMO}$ martingale under $\mathbb{P}$ on $[0,T]$.
+A random partition of $[0,T]$ is a finite sequence of stopping times $0=\tau_{0}\leq\tau_{1}\leq\cdots\leq\tau_{m}=T$.
+For $\delta>0$, the index of sliceability of $M$ is the function
+$N_{M}:(0,\infty)\longrightarrow\mathbb{N}\cup\{\infty\}$
+defined as follows: $N_{M}(\delta)$ is the smallest integer $m$ such that there exists a random partition $(t_{k})_{k=0}^{m}$ of $[0,T]$ with
+$\big\|M^{(k)}\big\|_{\mathrm{BMO}}\leq\delta,\qquad 1\leq k\leq m,$ (2.5)
+where the kk–th increment $M^{(k)}$ is the continuous martingale
+$M^{(k)}_{t}:=M_{t\wedge\tau_{k}}-M_{t\wedge\tau_{k-1}},\qquad 0\leq t\leq T.$
+If no such $m$ exists, we set $N_{M}(\delta)=\infty$. We say that $M$ is $\delta$–sliceable if $N_{M}(\delta)<\infty$, and simply sliceable if $M$ is $\delta$–sliceable for every $\delta>0$.
 2.2 Statement of the main theorem
-In this paper we consider the following general mean-field BSDE
+In this paper, we consider the following general mean-field BSDE
 $Y_{t}=\xi+\int_{t}^{T}\int_{\tilde{\Omega}}g(s,Z_{s},\tilde{Z}_{s})d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},\quad t\in[0,T].$ (2.6)
@@
     for any $(t,z,\tilde{z},z^{\prime},\tilde{z}^{\prime})\in[0,T]\times[\mathbb{R}^{d}\times\mathbb{R}^{d}]^{2}$.
+Example 2.4.
+The following $g$ satisfies the condition $(\mathrm{H}_{q})$:
+[…]
+where $l$ is Lipschitz in (z,z~)($\tilde{z}$), and $q_{1}$ and $q_{2}$ are quadratic polynomials in $z$ and $\tilde{z}$, respectively.
+Note that this example does not satisfy the existence and uniqueness assumption in previous literature, such as hao2022mean.
 Theorem 2.5.
@@
 Remark 2.6.
-In the statement and proof of our main theorem we may, without loss of generality, assume that
+In the statement and proof of our main theorem we may, without loss of generality, we can assume that
 $g(t,0,\tilde{z})=0,\qquad\forall\,(t,\tilde{z})\in[0,T]\times\mathbb{R}^{d}.$
@@
 $Y_{t}:=U_{t}+\int_{t}^{T}\widetilde{{\mathbb{E}}}\big[g(s,0,\tilde{Z}_{s})\big]\,ds$
-[…]
+solves the original BSDE.
+Thus there is a one-to-one correspondence in terms of existence and uniqueness.
+From this point on, we therefore work directly with the shifted generator, still denoted by $g$, and silently assume $g(t,0,\tilde{z})=0$ for all $(t,\tilde{z})\in[0,T]\times\mathbb{R}^{d}$.
 3 Terminal conditions with bounded Malliavin derivative
-This section is devoted to deriving tractable upper bounds for the solution $(Y,Z)$ of the BSDE (2.5) under the assumption that the driver $g$ is Lipschitz continuous.
-These bounds become particularly useful when the terminal condition $\xi$ is Malliavin differentiable.
-Although this may appear to be a stronger condition, it serves a crucial role in our analysis.
-In the proof of Lemma 5.1, we will show that the general case, where the driver exhibits quadratic growth and where $\xi$ is bounded, can be reduced to this setting through suitable approximation techniques.
+This section is devoted to deriving tractable upper bounds for the solution $(Y,Z)$ of the BSDE (2.6) under the assumption that the driver $g$ is Lipschitz continuous and the terminal condition $\xi$ is Malliavin differentiable.
+These bounds under the stronger condition become particularly useful in the proof of Lemma 5.1, we will show that the general case, where the driver exhibits quadratic growth and where $\xi$ is bounded, can be reduced to this setting through suitable approximation techniques.
 Therefore, we initially work under stronger regularity assumptions to facilitate this reduction.
-For this purpose, we suppose all along this section that $g$ satisfies the following classical Lipschitz condition.
+For this purpose, we suppose all along this section that $g$ satisfies the following classical Lipschitz condition and condition on $\xi$ will be introduced in Propisiton 3.1 below.
   • $({\rm H_{lip}})$
-    There exist constants $L_{z}>0$ such that:
+    There exist a constant $L_{z}>0$ such that:
 […]
@@
 Under assumption $({\rm H_{lip}})$, we have $\left|\partial_{z}g\right|\leq L_{z}$.
-Therefore, the Novikov condition is satisfied, and we may apply the Girsanov transform to obtain the following representation with new Brownian motion $W^{\mathbb{Q}}$ under the new probability measure $\mathbb{Q}$:
+Therefore, the Novikov condition is satisfied and $\widetilde{{\mathbb{E}}}\!\left[\partial_{z}g(s,Z_{s},\widetilde{Z}_{s})\right]$ is measurable to $mathcal{F_{s}}$, and we may apply the Girsanov transform to obtain the following representation with new Brownian motion $W^{\mathbb{Q}}$ under the new probability measure $\mathbb{Q}$:
 $D_{\theta}Y_{t}=D_{\theta}\xi-\int_{t}^{T}D_{\theta}Z_{s}\!\cdot dW_{s}^{\mathbb{Q}},\qquad 0\leq\theta\leq t\leq T.$
@@
 $\left|D_{\theta}Y_{t}\right|\leq\left\|D\xi\right\|_{\mathbb{S}^{\infty}}.$
-Finally, as $(Z_{t})_{0\leq t\leq T}$ is a version of $(D_{\theta}Y_{t})_{0\leq t\leq T}$ and due to the arbitrariness of $t$, we establish 3.7. ∎
+Finally, as $(Z_{t})_{0\leq t\leq T}$ is a version of $(D_{\theta}Y_{t})_{0\leq t\leq T}$ and due to the arbitrariness of $t$, we establish 3.8. ∎
 4 A priori estimates and stability for BSDEs
-In this section, we establish the stability of the mean-field BSDE (2.5), as stated in Theorem 4.6.
+In this section, we will work under the assumptions $(\rm H_{0})$} and $(\rm H_{q})$} to establish the stability of the mean-field BSDE (2.6), as stated in Theorem 4.6.
 The strategy of the proof is as follows.
 We first derive a priori upper bounds on the $\mathbb{S}^{\infty}$ norm of $Y$ and the BMO norm of $\int_{0}^{\cdot}Z_{s}\!\cdot dW_{s}$ for any solution $(Y,Z)$ to the BSDE (2.6).
-These estimates, together with fundamental properties of BMO martingales, lead to the stability result in Theorem 4.6.
+These estimates, together with fundamental properties of BMO martingales, lead to the stability result.
 We begin with a useful estimate that will play an important role in the analysis.
@@
 Suppose that (H0) and (Hq) hold.
-Let $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}$ be a bounded solution to the mean-field BSDE (2.5).
-Then the stochastic integral process $\left(\int_{0}^{t}Z_{s}\cdot dW_{s}\right)_{0\leq t\leq T}$ defines a BMO martingale.
+Let $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}$ be a solution to the mean-field BSDE (2.6).
+Then the stochastic process $\left(\int_{0}^{t}Z_{s}\!\cdot dW_{s}\right)_{0\leq t\leq T}$ defines a BMO martingale.
 Moreover, its BMO norm satisfies the following bound:
@@
 Proof.
-The argument follows the two steps scheme of (briand2008stobsdes, p. 831) with modified $C^{2}$ function
-$\varphi(x):=\frac{e^{3K_{z}|x|}-1-6cK_{z}|x|}{(3K_{z})^{2}},\qquad x\in\mathbb{R}.$
+The argument follows from arguments in Proposition 2.1 of briand2008stobsdes with modified $C^{2}$ function
+$\varphi(x):=\frac{e^{3K_{z}|x|}-1-6cK_{z}|x|}{(3K_{z})^{2}},\qquad c\in\mathbb{R},x\in\mathbb{R}.$
 In Step 1 we express the $\mathrm{BMO}$ norm of $Z$ in terms of its $\mathbb{H}^{2}$ norm; in Step 2 we derive an explicit $\mathbb{H}^{2}$ bound in terms of the model coefficients and $\|Y\|_{\mathbb{S}^{\infty}}$.
-Step 1 (Estimate of $\|Z\|_{\mathrm{BMO}}$ in terms of $\|Z\|_{\mathbb{H}^{2}}$). Fix $c=\tfrac{1}{2}$ then $\varphi(0)=0$ and, for all $x\in\mathbb{R}$,
-$\varphi^{\prime}(x)=\frac{e^{3K_{z}|x|}-1}{3K_{z}}\operatorname{sgn}(x),\qquad\varphi^{\prime\prime}(x)=e^{3K_{z}|x|},\qquad\varphi^{\prime\prime}(x)-3K_{z}\,|\varphi^{\prime}(x)|\equiv 1.$
+Step 1 (Estimate of $\|Z\|_{\mathbb{H}^{2}_{\mathrm{BMO}}}$ in terms of $\|Z\|_{\mathbb{H}^{2}}$). Fix $c=\tfrac{1}{2}$ then $\varphi(0)=0$ and, for all $x\in\mathbb{R}$,
+$\varphi(x)\geq 0,\qquad\varphi^{\prime}(x)=\frac{e^{3K_{z}|x|}-1}{3K_{z}}\operatorname{sgn}(x),\qquad\varphi^{\prime\prime}(x)=e^{3K_{z}|x|},\qquad\varphi^{\prime\prime}(x)-3K_{z}\,|\varphi^{\prime}(x)|\equiv 1.$
 Moreover, for $|x|\leq\|Y\|_{\mathbb{S}^{\infty}}$ we have the deterministic bound
 $0\leq|\varphi^{\prime}(x)|\leq C_{Y}:=\frac{1}{3K_{z}}\,e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}.$ (4.11)
-Let $\tau$ be any stopping time and apply Itô’s formula to $\varphi(Y)$ and using the growth estimate 4.8 for $g$ together with the symmetry $\widetilde{\mathbb{E}}|\tilde{Z}_{s}|^{2}=\mathbb{E}|Z_{s}|^{2}$ yields
-$\varphi(Y_{\tau})$ $=\varphi(Y_{T})+\!\int_{\tau}^{T}\!\Big(-\varphi^{\prime}(Y_{s})\,\widetilde{\mathbb{E}}[g(s,Z_{s},\tilde{Z}_{s})]-\tfrac{1}{2}\,\varphi^{\prime\prime}(Y_{s})|Z_{s}|^{2}\Big)ds-\!\int_{\tau}^{T}\!\varphi^{\prime}(Y_{s})Z_{s}\cdot dW_{s}$
-$\leq\varphi(Y_{T})+\!\int_{\tau}^{T}\!\Big(\tfrac{3}{2}|\varphi^{\prime}(Y_{s})|-\tfrac{1}{2}\varphi^{\prime\prime}(Y_{s})\Big)|Z_{s}|^{2}ds+\!\int_{\tau}^{T}\!|\varphi^{\prime}(Y_{s})|\Big(K+\tfrac{3}{2}K_{\tilde{z}}\,\mathbb{E}|Z_{s}|^{2}\Big)ds$
-$\leq\varphi(Y_{T})-\tfrac{1}{2}\!\int_{\tau}^{T}\!|Z_{s}|^{2}ds+\!\int_{\tau}^{T}\!|\varphi^{\prime}(Y_{s})|\Big(K+\tfrac{3}{2}K_{\tilde{z}}\,\mathbb{E}|Z_{s}|^{2}\Big)ds-\!\int_{\tau}^{T}\!\varphi^{\prime}(Y_{s})Z_{s}\cdot dW_{s},$
+Let $\tau$ be any stopping time and apply Itô’s formula to $\varphi(Y)$ and using the growth estimate 4.9 for $g$ together with the symmetry $\widetilde{\mathbb{E}}|\tilde{Z}_{s}|^{2}=\mathbb{E}|Z_{s}|^{2}$ yields
+$\varphi(Y_{\tau})$ $=\varphi(\xi)+\!\int_{\tau}^{T}\!\Big(-\varphi^{\prime}(Y_{s})\,\widetilde{\mathbb{E}}[g(s,Z_{s},\tilde{Z}_{s})]-\tfrac{1}{2}\,\varphi^{\prime\prime}(Y_{s})|Z_{s}|^{2}\Big)ds-\!\int_{\tau}^{T}\!\varphi^{\prime}(Y_{s})Z_{s}\cdot dW_{s}$
+$\leq\varphi(\xi)+\!\int_{\tau}^{T}\!\Big(\frac{3K_{z}}{2}|\varphi^{\prime}(Y_{s})|-\tfrac{1}{2}\varphi^{\prime\prime}(Y_{s})\Big)|Z_{s}|^{2}ds+\!\int_{\tau}^{T}\!|\varphi^{\prime}(Y_{s})|\Big(K+\tfrac{3}{2}K_{\tilde{z}}\,\mathbb{E}|Z_{s}|^{2}\Big)ds$
+$\leq\varphi(\xi)-\tfrac{1}{2}\!\int_{\tau}^{T}\!|Z_{s}|^{2}ds+\!\int_{\tau}^{T}\!|\varphi^{\prime}(Y_{s})|\Big(K+\tfrac{3}{2}K_{\tilde{z}}\,\mathbb{E}|Z_{s}|^{2}\Big)ds-\!\int_{\tau}^{T}\!\varphi^{\prime}(Y_{s})Z_{s}\cdot dW_{s},$
 where $K:=K_{0}+\frac{1}{2}K_{z}+\frac{1}{2}K_{\tilde{z}}$.
 Taking $\mathbb{E}[\cdot\mid\mathcal{F}_{\tau}]$ and using (4.11) together with $|\varphi(x)|\leq C_{Y}|x|$ (since $\varphi(0)=0$ and $\varphi^{\prime}$ is bounded on the range of $Y$), we get
-$\tfrac{1}{2}\,\mathbb{E}\!\Big[\!\int_{\tau}^{T}\!|Z_{s}|^{2}ds\,\Big|\,\mathcal{F}_{\tau}\Big]$ $\leq\mathbb{E}\!\big[\varphi(Y_{T})-\varphi(Y_{\tau})\,\big|\,\mathcal{F}_{\tau}\big]+C_{Y}K(T-\tau)+\tfrac{3}{2}K_{\tilde{z}}C_{Y}\!\!\int_{0}^{T}\!\mathbb{E}|Z_{s}|^{2}ds$
+$\tfrac{1}{2}\,\mathbb{E}\!\Big[\!\int_{\tau}^{T}\!|Z_{s}|^{2}ds\,\Big|\,\mathcal{F}_{\tau}\Big]$ $\leq\mathbb{E}\!\big[\varphi(\xi)-\varphi(Y_{\tau})\,\big|\,\mathcal{F}_{\tau}\big]+C_{Y}K(T-\tau)+\tfrac{3}{2}K_{\tilde{z}}C_{Y}\!\!\int_{0}^{T}\!\mathbb{E}|Z_{s}|^{2}ds$
 $\leq C_{Y}\,\mathbb{E}\!\big[|\xi|\,\big|\,\mathcal{F}_{\tau}\big]+C_{Y}KT+\tfrac{3}{2}K_{\tilde{z}}C_{Y}\!\!\int_{0}^{T}\!\mathbb{E}|Z_{s}|^{2}ds.$
 we thus obtain
-$\tfrac{1}{2}\,{\mathbb{E}}\!\Big[\!\int_{\tau}^{T}\!|Z_{s}|^{2}ds\,\Big|\,\mathcal{F}_{\tau}\Big]\leq(3K_{z})^{-1}e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}\left(\|\xi\|_{\infty}+K\,T+\tfrac{3}{2}K_{\tilde{z}}\!\!\int_{0}^{T}\!\!{\mathbb{E}}|Z_{s}|^{2}ds\right).$ (4.11)
+$\tfrac{1}{2}\,{\mathbb{E}}\!\Big[\!\int_{\tau}^{T}\!|Z_{s}|^{2}ds\,\Big|\,\mathcal{F}_{\tau}\Big]\leq C_{Y}\left(\|\xi\|_{L^{\infty}}+K\,T+\tfrac{3}{2}K_{\tilde{z}}\!\!\int_{0}^{T}\!\!{\mathbb{E}}|Z_{s}|^{2}ds\right).$ (4.12)
 Since $\tau$ is arbitrary, (4.12) shows that $\big\|\int_{0}^{\cdot}Z_{s}\cdot dW_{s}\big\|_{\mathrm{BMO}}^{2}$ is controlled by $\|Z\|_{\mathbb{H}^{2}}^{2}$ and model parameters.
-Step 2 (An explicit $\mathbb{H}^{2}$ bound). Write $\alpha:=K_{\tilde{z}}/K_{z}>0$. We may, without loss of generality (enlarge $K_{z}$ or $K_{\tilde{z}}$) assume $\alpha$ satisfies
-$0<\alpha\;<\;\min\!\Big\{\tfrac{1}{2},\ \frac{e^{3K_{z}}-1}{e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}}\Big\}.$
-Define positive constants
+Step 2 (An explicit $\mathbb{H}^{2}$ bound). Write $\alpha:=K_{\tilde{z}}/K_{z}>0$. We may, without loss of generality, assume $\alpha$ satisfies
+$0<\alpha\;<\;\min\!\Big\{\tfrac{1}{2},\ \frac{e^{3K_{z}}-2}{e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}}\Big\}.$
+by enlarging $K_{z}$ or $K_{\tilde{z}}$. Define positive constants
 $c_{1}:=\frac{K_{\tilde{z}}}{2K_{z}}e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}+\frac{1}{2},\qquad c_{2}:=\tfrac{1}{2}e^{3K_{z}}.$
-Note $c_{1}$ can also be expressed as $c_{1}=\tfrac{1}{2}\big(\alpha e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}+1\big)=\frac{3K_{\tilde{z}}}{2}C_{Y}+\tfrac{1}{2}$.
-By the choice of $\alpha$, it is easy to verify $c_{2}>c_{1}$.
-By density of rational numbers, pick any $c\in(c_{1},c_{2})$ and consider the same $C^{2}$ function at in Step 1, so that $\varphi(0)=0$ and, for all $x$,
-$|\varphi^{\prime}(x)|=\frac{e^{3K_{z}|x|}-2c}{3K_{z}}>\frac{e^{3K_{z}|x|}-2c_{2}}{3K_{z}}=0.$
-So that we have identity $\varphi^{\prime\prime}(\cdot)-3K_{z}\left|\varphi^{\prime}(\cdot)\right|=2c$. Applying Itô to $\varphi(Y)$ on $[0,T]$ and arguing as in Step 1 gives
+Note that, by definition of $\alpha$, $c_{1}$ can also be expressed as $c_{1}=\tfrac{1}{2}\big(\alpha e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}+1\big)$.
+By the choice of $\alpha$, it is easy to verify $c_{1}<c_{2}$.
+Pick any $c\in(c_{1},c_{2})$ and consider the same $\varphi$ at in Step 1 with a different $c$ value, so that $\varphi(0)=0$ and, for all $x$,
+$|\varphi^{\prime}(x)|=\frac{e^{3K_{z}|x|}-2c}{3K_{z}}>\frac{e^{3K_{z}|x|}-2c_{2}}{3K_{z}}\geq 0.$
+We also have identity $\varphi^{\prime\prime}(\cdot)-3K_{z}\left|\varphi^{\prime}(\cdot)\right|=2c$. Applying Itô to $\varphi(Y)$ on $[0,T]$ and arguing as in Step 1 with $\tau=0$ gives
 $\varphi(Y_{0})$ $\leq\varphi(\xi)-\int_{0}^{T}c\,|Z_{s}|^{2}ds+\frac{3K_{\tilde{z}}}{2}\int_{0}^{T}|\varphi^{\prime}(Y_{s})|\,\mathbb{E}|Z_{s}|^{2}ds+K\int_{0}^{T}|\varphi^{\prime}(Y_{s})|ds-\int_{0}^{T}\varphi^{\prime}(Y_{s})Z_{s}\cdot dW_{s}.$
@@
 $\varphi(Y_{0})\leq\mathbb{E}\!\left[\varphi(\xi)\right]+KC_{Y}T-c\,\mathbb{E}\!\left[\int_{0}^{T}|Z_{s}|^{2}ds\right]+\frac{3K_{\tilde{z}}}{2}\,C_{Y}\,\mathbb{E}\!\left[\int_{0}^{T}|Z_{s}|^{2}ds\right].$
-By the definition of $c_{1}$ and $c>c_{1}$ we have $\frac{3K_{\tilde{z}}}{2}C_{Y}\leq c-\tfrac{1}{2}$. Hence
+By the definition of $c_{1}$ and the fact that $c>c_{1}=\frac{3K_{\tilde{z}}}{2}C_{Y}+\tfrac{1}{2}$ we have $\frac{3K_{\tilde{z}}}{2}C_{Y}<c-\tfrac{1}{2}$. Hence
 $\frac{1}{2}\,\mathbb{E}\!\left[\int_{0}^{T}|Z_{s}|^{2}ds\right]\ \leq\ \mathbb{E}\!\left[\varphi(\xi)\right]+KC_{Y}T\ \leq\ \frac{e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}}{3K_{z}}\,\big(\left\|\xi\right\|_{L^{\infty}}+KT\big).$ (4.13)
@@
 Proposition 4.4.
-Suppose Assumptions $(\rm H_{0})$} and $(\rm H_{q})$} hold, and that we work with the shifted generator satisfying
-$g(t,0,\tilde{z})=0,\qquad\forall\,(t,\tilde{z})\in[0,T]\times\mathbb{R}^{d}$
-(cf. Remark 2.4). Then for any solution $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ of BSDE (2.5), we have $Y\in\mathbb{S}^{\infty}$ and the following estimate holds:
+Suppose Assumptions $(\rm H_{0})$} and $(\rm H_{q})$} hold.
+Then for any solution $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ of BSDE (2.6), we have $Y\in\mathbb{S}^{\infty}$ and the following estimate holds:
 $\left\|Y\right\|_{\mathbb{S}^{\infty}}\leq\left\|\xi\right\|_{L^{\infty}}.$ (4.14)
 Proof.
-The proof relies on a linearization argument combined with the BMO property. Under the standing normalisation $g(t,0,\tilde{z})=0$, the BSDE (2.5) can be rewritten as
+The proof relies on a linearization argument combined with the BMO property. Using Remark 2.6, without the loss of generality, we assume
+$g(t,0,\tilde{z})=0,\qquad\forall\,(t,\tilde{z})\in[0,T]\times\mathbb{R}^{d}.$
+Then, the BSDE (2.6) can be rewritten as
 $Y_{t}=\xi+\int_{t}^{T}\Big(\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big)\cdot Z_{s}\,ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},\qquad t\in[0,T],$ (4.15)
 where the process $(L_{s})_{0\leq s\leq T}$ is defined by
-$L_{s}:=\frac{g(s,Z_{s},\tilde{Z}_{s})-g(s,0,\tilde{Z}_{s})}{\left|Z_{s}\right|^{2}}\,Z_{s}\,\mathbf{1}_{\{\left|Z_{s}\right|>0\}},\qquad 0\leq s\leq T.$
+$L_{s}:=\frac{g(s,Z_{s},\tilde{Z}_{s})-g(s,0,\tilde{Z}_{s})}{|\Delta Z_{s}|^{2}}\,\Delta Z_{s}\,\mathbf{1}_{\{\Delta Z_{s}\neq 0\}}\qquad 0\leq s\leq T,$
 Since $g(s,0,\tilde{z})=0$ by construction, the numerator is simply $g(s,Z_{s},\tilde{Z}_{s})$. Moreover, Assumption $(\rm H_{q})$} yields
 $\Big|\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big|\leq\int_{\tilde{\Omega}}\left|L_{s}\right|\,d\tilde{\mathbb{P}}\leq K_{z}\big(1+\left|Z_{s}\right|\big),\qquad 0\leq s\leq T.$
-Since $Z\in\mathbb{H}^{2}_{\mathrm{BMO}}$ by assumption, $\big(\int_{0}^{t}\Big(\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big)\cdot dW_{s}\big)_{0\leq t\leq T}$ is a BMO martingale, and its stochastic exponential $\mathcal{E}_{t}$ satisfies Kazamaki’s condition.
+Since $Z\in\mathbb{H}^{2}_{\mathrm{BMO}}$ by assumption, $\big(\int_{0}^{t}\Big(\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big)\cdot dW_{s}\big)_{t\in[0,T]}$ is a BMO martingale, and its stochastic exponential $\mathcal{E}_{t}$ satisfies Kazamaki’s condition.
 Define the probability measure $\mathbb{P}^{L}$ by $d\mathbb{P}^{L}:=\mathcal{E}_{T}\,d\mathbb{P}$.
@@
 $Y_{t}={\mathbb{E}}^{\mathbb{P}^{L}}\big[\xi\,\big|\,\mathcal{F}_{t}\big],\qquad 0\leq t\leq T.$
-Since $\mathbb{P}^{L}$ is equivalent to $\mathbb{P}$, the same bound holds $\mathbb{P}$–almost surely (Lemma 4.3), and taking the essential supremum over $t\in[0,T]$ completes the proof. ∎
-Then we have a direct result from previous two propositions.
+Since $\mathbb{P}^{L}$ is equivalent to $\mathbb{P}$, the same bound holds $\mathbb{P}$–almost surely by Lemma 4.3, and taking the essential supremum over $t\in[0,T]$ completes the proof. ∎
+Then we have a direct result from the previous two propositions.
 Corollary 4.5.
@@
 $\left\|Y\right\|_{\mathbb{S}^{\infty}}+\Big\|\textstyle\int_{0}^{\cdot}Z_{s}\cdot dW_{s}\Big\|_{\mathrm{BMO}}\leq C,$ (4.16)
-where $C$ is a constant depending only on $\left\|\xi\right\|_{L^{\infty}}$, $T$, and the constants appearing in Assumptions $(\rm H_{0})$} and $(\rm H_{q})$}.
+where $C$ is a constant depending only on $(T,K_{z},K_{\tilde{z}},K_{0})$.
 Proof.
@@
 Theorem 4.6.
-For $i=1,2$, let $(Y^{i},Z^{i})\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ solve
+For $i=1,2$, let $(Y^{i},Z^{i})\in\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}$ solve
 $Y_{t}^{i}=\xi^{i}+\int_{t}^{T}\!\int_{\tilde{\Omega}}g(s,Z_{s}^{i},\tilde{Z}_{s}^{i})\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}^{i}\cdot dW_{s},\qquad t\in[0,T],$
-with $\xi^{1},\xi^{2},g$ satisfying $(\mathrm{H}_{0})$ and $(\mathrm{H}_{g})$. Then
-$\|Y^{1}-Y^{2}\|_{\mathbb{S}^{2}}^{2}+\|Z^{1}-Z^{2}\|_{\mathbb{H}^{2}}^{2}\ \leq\ C_{\mathrm{global}}\;\|\xi^{1}-\xi^{2}\|_{L^{4}},$
-where $C_{\mathrm{global}}$ depends only on $T$, $K_{z}$, $K_{\tilde{z}}$ and $K_{0}$.
+with $\xi^{1},\xi^{2},g$ satisfying $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q})$. Then there exist $p^{*}>1$ and, for every $p>p^{*}$, a constant $C_{p}>0$ such that
+$\|Y^{1}-Y^{2}\|_{\mathbb{S}^{2p}}+\|Z^{1}-Z^{2}\|_{\mathbb{H}^{p}}\ \leq\ C_{p}\;\|\xi^{1}-\xi^{2}\|_{L^{2p}},$
+where $p^{*}>1$ and $C_{p}$ depend only on $T$, $K_{z}$, $K_{\tilde{z}}$, $K_{0}$, and do not depend on $\|Z^{i}\|_{\mathbb{S}^{\infty}}$.
 Proof.
@@
 $\Delta g(s,Z_{s},\tilde{Z}_{s}):=g(s,Z_{s}^{1},\tilde{Z}_{s}^{1})-g(s,Z_{s}^{2},\tilde{Z}_{s}^{2}),\qquad s\in[0,T].$
-Then
+Then we obtain
 $\Delta Y_{t}=\Delta\xi+\int_{t}^{T}\!\int_{\tilde{\Omega}}\!\Delta g(s,Z_{s},\tilde{Z}_{s})\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}\Delta Z_{s}\cdot dW_{s},\qquad t\in[0,T].$ (4.17)
-Preliminary step (linearization and RHI).
-As in the proof of Proposition 4.4, we use a linearization argument, there exists processes $L_{s}$ and $\tilde{L}_{s}$ live on the product space such that, for all $s\in[0,T]$,
+Preliminary step (linearization, RHI and sliceability).
+As in the proof of Proposition 4.4, we use a linearization argument: there exist processes $L_{s}$ and $\tilde{L}_{s}$ on the product space such that, for all $s\in[0,T]$,
+$L_{s}:=\frac{g(s,Z^{1}_{s},\tilde{Z}^{1}_{s})-g(s,Z^{2}_{s},\tilde{Z}^{1}_{s})}{|\Delta Z_{s}|^{2}}\,\Delta Z_{s}\,\mathbf{1}_{\{\Delta Z_{s}\neq 0\}},\qquad\tilde{L}_{s}:=\frac{g(s,Z^{2}_{s},\tilde{Z}^{1}_{s})-g(s,Z^{2}_{s},\tilde{Z}^{2}_{s})}{|\Delta\tilde{Z}_{s}|^{2}}\,\Delta\tilde{Z}_{s}\,\mathbf{1}_{\{\Delta\tilde{Z}_{s}\neq 0\}}$
+with estimates
 $|L_{s}|\leq K_{z}\big(1+|Z_{s}^{1}|+|Z_{s}^{2}|\big),\qquad|\tilde{L}_{s}|\leq K_{\tilde{z}}\big(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|\big),$ (4.18)
 and $\Delta g(s,Z_{s},\tilde{Z}_{s})=L_{s}\Delta Z_{s}+\tilde{L}_{s}\,\Delta\tilde{Z}_{s}$.
-Fix a small $\delta>0$ and write $I:=[T-\delta,T]$.
-According to Proposition 4.2, we have $Z^{1},Z^{2}\in\mathbb{H}^{2}_{\mathrm{BMO}}$, so in particular $\big(\int_{0}^{t}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\big)_{0\leq t\leq T}$ is a $\mathrm{BMO}$ martingale by (4.17).
+According to Proposition 4.2, we have $Z^{1},Z^{2}\in\mathbb{H}^{2}_{\mathrm{BMO}}$, so in particular $\big(\int_{0}^{t}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\big)_{t\in[0,T]}$ is a $\mathrm{BMO}$ martingale by (4.18) and symmetry.
 Moreover, by Corollary 4.5, there exists a constant $C>0$, depending only on $(T,K_{z},K_{\tilde{z}},K_{0})$, such that
@@
 Hence $\Lambda:=\mathcal{E}(\int_{0}^{\cdot}\widetilde{\mathbb{E}}[L_{s}]\cdot dW_{s})$ defines an equivalent probability measure $\mathbb{P}^{L}$ by Girsanov’s theorem with Brownian motion $W^{L}_{t}:=W_{t}+\int_{0}^{t}\widetilde{\mathbb{E}}[L_{s}]\,ds$.
-In addition, by Lemma 2.2, the same processes admit $\mathrm{BMO}$ norms under the equivalent measure $\mathbb{P}^{L}$, up to multiplicative constants depending on the same constant.
-We then obtain,
-$\Delta Y_{u}=\Delta\xi+\int_{u}^{T}\widetilde{\mathbb{E}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds-\int_{u}^{T}\Delta Z_{s}\cdot dW^{L}_{s},\qquad u\in I.$ (4.19)
-The Doléans–Dade exponential $\Lambda=\mathcal{E}\!\left(\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\right)$ satisfies RHI from the Lemma 2.1,
-$\bar{{\mathbb{E}}}\!\left[\left(\Lambda_{T}\right)^{q}\,\big|\,\bar{\mathcal{F}}_{t}\right]\ \leq\ C_{q}^{*}\,\left(\Lambda_{t}\right)^{q},\qquad 0\leq t\leq T,\quad 1\leq q\leq q^{*},$
+In addition, by Lemma 2.2, the same processes admit $\mathrm{BMO}$ norms under the equivalent measure $\mathbb{P}^{L}$, up to multiplicative constants depending only on $C$.
+Under $\mathbb{P}^{L}$ we can rewrite (4.17) as
+$\Delta Y_{u}=\Delta\xi+\int_{u}^{T}\widetilde{\mathbb{E}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds-\int_{u}^{T}\Delta Z_{s}\cdot dW^{L}_{s},\qquad u\in[0,T].$ (4.20)
+The Doléans–Dade exponential $\Lambda=\mathcal{E}\!\left(\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\right)$ satisfies a reverse Hölder inequality by Lemma 2.1: there exists $q^{*}>1$ such that, for all $1\leq q\leq q^{*}$,
+$\bar{{\mathbb{E}}}\!\left[\left(\Lambda_{T}\right)^{q}\,\big|\,\bar{\mathcal{F}}_{t}\right]\ \leq\ C_{q}^{*}\,\left(\Lambda_{t}\right)^{q},\qquad 0\leq t\leq T,$
 where
@@
 $C_{q}^{*}$ $:=2\left(1-\frac{2q-1}{2q-2}\,\exp\!\left\{q^{2}\Big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2}+2\Big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\Big\|_{\mathrm{BMO}}\right\}\right)^{-1}\!.$
-By the sliceability of BMO martingales and the function $\phi$ is strictly decreasing with respect to BMO norm, we can restrict to a short interval $I=[T-\delta,T]$ so that the local norm $\big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\big\|_{\mathrm{BMO}(I)}$ is small enough to ensure $q^{*}>2$.
-Then we have the RHI for every nonnegative $X\in\mathcal{F}_{T}$, and we pick $q=2$:
-$\mathbb{E}\!\left[X\ \middle|\ \mathcal{F}_{T-\delta}\right]\ \leq\ C_{2,-}(I)^{1/2}\,\Big(\mathbb{E}^{L}\!\left[X^{2}\,\middle|\,\mathcal{F}_{T-\delta}\right]\Big)^{1/2},$ (4.20)
-where the explicit constant $C_{2,-}(I)\ :=\ C_{q}^{*},$ with $q=2$ on interval $I$.
-Step 1 (local stability on $I$). Applying Itô’s Lemma to $|\Delta Y|^{2}$ and taking $\mathbb{E}^{L}$ gives
-$\mathbb{E}^{L}|\Delta Y_{u}|^{2}+\mathbb{E}^{L}\!\int_{u}^{T}|\Delta Z_{s}|^{2}\,ds\ \leq\ \mathbb{E}^{L}|\Delta\xi|^{2}+2\mathbb{E}^{L}\int_{u}^{T}\!\big[\,|\Delta Y_{s}|\,\widetilde{\mathbb{E}}|\tilde{L}_{s}\,\Delta\tilde{Z}_{s}|\,\big]ds.$ (4.21)
-By Corollary 4.5 we have $\|\Delta Y\|_{\mathbb{S}^{\infty}}<\infty$ and the bound depends only on model parameters; hence, for all $u\in I$,
-$\mathbb{E}^{L}|\Delta Y_{u}|^{2}+\mathbb{E}^{L}\!\int_{u}^{T}|\Delta Z_{s}|^{2}ds\ \leq\ \mathbb{E}^{L}|\Delta\xi|^{2}+2\|\Delta Y\|_{\mathbb{S}^{\infty}}\,\mathbb{E}^{L}[M_{\delta}],\qquad M_{\delta}:=\int_{T-\delta}^{T}\widetilde{\mathbb{E}}|\tilde{L}_{s}\,\Delta\tilde{Z}_{s}|\,ds.$ (4.22)
-Evaluating (4.22) at $u=T-\delta$,
-$\mathbb{E}^{L}\!\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}ds\ \leq\ \mathbb{E}^{L}|\Delta\xi|^{2}+2\|\Delta Y\|_{\mathbb{S}^{\infty}}\,M_{\delta}.$ (4.23)
-By Cauchy–Schwarz inequality on the copy space and (4.17), for $s\in I$,
-$\widetilde{{\mathbb{E}}}\,|\tilde{L}_{s}\,\Delta\tilde{Z}_{s}|\;\leq\;\big(\widetilde{{\mathbb{E}}}|\tilde{L}_{s}|^{2}\big)^{1/2}\big(\widetilde{{\mathbb{E}}}|\Delta\tilde{Z}_{s}|^{2}\big)^{1/2}\;\leq\;\sqrt{3}\,K_{\tilde{z}}\,\Big(1+{\mathbb{E}}|Z_{s}^{1}|^{2}+{\mathbb{E}}|Z_{s}^{2}|^{2}\Big)^{1/2}\big({\mathbb{E}}|\Delta Z_{s}|^{2}\big)^{1/2}.$
-Cauchy–Schwarz in time yields
-$M_{\delta}$ $\leq\sqrt{3}\,K_{\tilde{z}}\int_{T-\delta}^{T}\Big(1+{\mathbb{E}}|Z_{s}^{1}|^{2}+{\mathbb{E}}|Z_{s}^{2}|^{2}\Big)^{1/2}\big({\mathbb{E}}|\Delta Z_{s}|^{2}\big)^{1/2}\,ds$ (4.24)
-$\leq\sqrt{3}\,K_{\tilde{z}}\,\Bigg(\int_{T-\delta}^{T}\Big(1+{\mathbb{E}}|Z_{s}^{1}|^{2}+{\mathbb{E}}|Z_{s}^{2}|^{2}\Big)\,ds\Bigg)^{\!1/2}\Bigg(\int_{T-\delta}^{T}{\mathbb{E}}|\Delta Z_{s}|^{2}\,ds\Bigg)^{\!1/2}.$
-For the first term we simply use BMO energy inequality (kazamaki2006BMO, Sec. 2.1)
-$\int_{T-\delta}^{T}\Big(1+{\mathbb{E}}|Z_{s}^{1}|^{2}+{\mathbb{E}}|Z_{s}^{2}|^{2}\Big)\,ds\;$ $\leq\;\delta+\Big\|\int_{T-\delta}^{\cdot}Z_{s}^{1}\,\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2}+\Big\|\int_{T-\delta}^{\cdot}Z_{s}^{2}\,\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2}=:\;C_{\mathrm{energy}},$
-where $C_{\mathrm{energy}}<\infty$ and increases with respect to the BMO norm. Hence,
-$M_{\delta}\;\leq\;\sqrt{3}\,K_{\tilde{z}}\,\sqrt{C_{\mathrm{energy}}}\,\Bigg({\mathbb{E}}\!\int_{T-\delta}^{T}\!|\Delta Z_{s}|^{2}\,ds\Bigg)^{\!1/2}.$ (4.25)
-Since the right-hand side is deterministic (depends only on $\mathbb{E}$$\cdot$]), the outer expectation ${\mathbb{E}}^{L}$ drops out; thus ${\mathbb{E}}^{L}[M_{\delta}]$ has the same bound.
-Next, bridge $\mathbb{E}$} and ${\mathbb{E}}^{L}$ by the RHI (4.20) with $X=\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}\,ds$:
-${\mathbb{E}}\!\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}\,ds\;\leq\;\sqrt{C_{2,-}(I)}\,\Bigg({\mathbb{E}}^{L}\!\Big[\Big(\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}\,ds\Big)^{2}\Big]\Bigg)^{\!1/2}.$ (4.26)
-Combining (4.25) and (4.26) yields
-$M_{\delta}\;\leq\;\sqrt{3}\,K_{\tilde{z}}\,\sqrt{C_{\mathrm{energy}}}\,C_{2,-}(I)^{\frac{1}{4}}\;\Bigg({\mathbb{E}}^{L}\!\Big[\Big(\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}\,ds\Big)^{2}\Big]\Bigg)^{\!1/4}.$ (4.27)
-To bound the right hand side, pick $u\in I$ and set $N_{u}:=\int_{T-\delta}^{u}\Delta Z_{s}\cdot dW^{L}_{s}$.
-Then $N]_{T}=\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}ds$ and the Burkholder–Davis–Gundy (BDG) lower bound at $p=4$ gives $c_{4}^{\downarrow}\,{\mathbb{E}}^{L}\!\big[[N]_{T}^{2}\big]\leq{\mathbb{E}}^{L}|N_{T}|^{4}$.
-From (4.19),
-$N_{T}=X-{\mathbb{E}}^{L}[X|\mathcal{F}_{T-\delta}],\qquad X:=\Delta\xi+\int_{T-\delta}^{T}\widetilde{{\mathbb{E}}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds\in\mathcal{F}_{T}.$
-Thus, Jensen’s inequality and $(|a|+|b|)^{4}\leq 8(|a|^{4}+|b|^{4})$ yield
-${\mathbb{E}}^{L}|N_{T}|^{4}\leq 16\,{\mathbb{E}}^{L}|X|^{4}\;\leq\;128\Big({\mathbb{E}}^{L}|\Delta\xi|^{4}+{\mathbb{E}}^{L}\big|\textstyle\int_{T-\delta}^{T}\widetilde{{\mathbb{E}}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds\big|^{4}\Big).$
-Hence, using (4.27),
-${\mathbb{E}}^{L}\big|\textstyle\int_{T-\delta}^{T}\widetilde{{\mathbb{E}}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds\big|^{4}\;\leq\;M_{\delta}^{4}\;\leq\;9\,K_{\tilde{z}}^{4}\,C_{\mathrm{energy}}^{2}\,C_{2,-}(I)\;{\mathbb{E}}^{L}\!\Big[\Big(\int_{T-\delta}^{T}\!|\Delta Z_{s}|^{2}\,ds\Big)^{\!2}\Big].$
-Collecting the displays, there exists $C_{\mathrm{BDG}}>0$ (absorbing numeric constants and $c_{4}^{\downarrow}$ only) such that
-${\mathbb{E}}^{L}\!\Big[\Big(\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}\,ds\Big)^{2}\Big]\;\leq\;C_{\mathrm{BDG}}\Big({\mathbb{E}}^{L}|\Delta\xi|^{4}\;+\;K_{\tilde{z}}^{4}\,C_{\mathrm{energy}}^{2}\,C_{2,-}(I)\;{\mathbb{E}}^{L}\!\Big[\Big(\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}\,ds\Big)^{2}\Big]\Big).$
-By choosing $\delta>0$ small, we can make the local BMO norms of $Z^{1},Z^{2}$ and $\widetilde{{\mathbb{E}}}[L]$ on the slice $I=[T-\delta,T]$ arbitrarily small.
-Consequently, both the energy constant $C_{\mathrm{energy}}$ (from the BMO energy inequality) and the reverse Hölder constant $C_{2,-}(I)$ are controlled by these norms and $\delta$, hence decrease simultaneously to zero as $\delta\downarrow 0$.
-In particular, the coefficient in front of ${\mathbb{E}}^{L}\big[(\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}ds)^{2}\big]$ can be made strictly less than $\tfrac{1}{4}$, so we absorb it to the left-hand side and conclude
-${\mathbb{E}}^{L}\!\Big[\Big(\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}\,ds\Big)^{2}\Big]\;\leq\;4\,C_{\mathrm{BDG}}\;{\mathbb{E}}^{L}|\Delta\xi|^{4}.$
-Finally, plug this into (4.27) to get
-$M_{\delta}\;\leq\;\,\sqrt{6}\,K_{\tilde{z}}\,\sqrt{C_{\mathrm{energy}}}\,C_{2,-}(I)^{\frac{1}{4}}\;C_{\mathrm{BDG}}^{\frac{1}{4}}\;\big({\mathbb{E}}^{L}|\Delta\xi|^{4}\big)^{1/4}.$ (4.28)
-$\mathbb{H}^{2}$–control of $\Delta Z$ on $I$. Plugging (4.28) into (4.23) gives
-$\mathbb{E}^{L}\!\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}ds\$$\leq\ \mathbb{E}^{L}|\Delta\xi|^{2}+2\|\Delta Y\|_{\mathbb{S}^{\infty}}\,\sqrt{6}\,K_{\tilde{z}}\,\sqrt{C_{\mathrm{energy}}}\,C_{2,-}(I)^{\frac{1}{4}}\;C_{\mathrm{BDG}}^{\frac{1}{4}}\,(\mathbb{E}^{L}|\Delta\xi|^{4})^{1/4}$
-$\leq\ \left(2K_{0}+2\|\Delta Y\|_{\mathbb{S}^{\infty}}\,\sqrt{6}\,K_{\tilde{z}}\,\sqrt{C_{\mathrm{energy}}}\,C_{2,-}(I)^{\frac{1}{4}}\;C_{\mathrm{BDG}}^{\frac{1}{4}}\right)(\mathbb{E}^{L}|\Delta\xi|^{4})^{1/4},$
-the last inequality using Cauchy–Schwarz and $\Delta\xi\in L^{\infty}(\mathbb{P})$ and thus $\Delta\xi\in L^{\infty}(\mathbb{P}^{L})$ with the same bound $K_{0}$ by Lemma 4.3, and therefore, $\mathbb{E}^{L}|\Delta\xi|^{2}\leq(\mathbb{E}^{L}|\Delta\xi|^{4})^{1/2}\leq 2K_{0}\,(\mathbb{E}^{L}|\Delta\xi|^{4})^{1/4}$.
-$\mathbb{S}^{2}$–control of $\Delta Y$ on $I$.
-Recall $X=\Delta\xi+\int_{T-\delta}^{T}\widetilde{{\mathbb{E}}}\!\big[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}\big]\,ds\in\mathcal{F}_{T}$ from (4.19); for $u\in I$,
-$\Delta Y_{u}={\mathbb{E}}^{L}\!\big[X\,\big|\,\mathcal{F}_{u}\big]-\int_{T-\delta}^{u}\widetilde{{\mathbb{E}}}\!\big[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}\big]\,ds.$
-Hence
-$\sup_{u\in I}|\Delta Y_{u}|\ \leq\ \sup_{u\in I}\Big|{\mathbb{E}}^{L}\!\big[X\,\big|\,\mathcal{F}_{u}\big]\Big|\ +\ \int_{T-\delta}^{T}\widetilde{{\mathbb{E}}}\big|\tilde{L}_{s}\,\Delta\tilde{Z}_{s}\big|\,ds\ =\ \sup_{u\in I}\Big|{\mathbb{E}}^{L}\!\big[X\,\big|\,\mathcal{F}_{u}\big]\Big|\ +\ M_{\delta}.$
-By $(a+b)^{2}\leq 2a^{2}+2b^{2}$ and Doob’s maximal inequality,
-${\mathbb{E}}^{L}\!\Big[\sup_{u\in I}|\Delta Y_{u}|^{2}\Big]\ \leq\ 8\,{\mathbb{E}}^{L}|X|^{2}\;+2\;{\mathbb{E}}^{L}[M_{\delta}^{2}].$
-Using again $(a+b)^{2}\leq 2a^{2}+2b^{2}$,
-${\mathbb{E}}^{L}|X|^{2}={\mathbb{E}}^{L}\Big|\Delta\xi+\int_{T-\delta}^{T}\widetilde{{\mathbb{E}}}\!\big[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}\big]\,ds\Big|^{2}\ \leq\ 2\,{\mathbb{E}}^{L}|\Delta\xi|^{2}\;+2\;{\mathbb{E}}^{L}[M_{\delta}^{2}],$
-we obtain
-${\mathbb{E}}^{L}\!\Big[\sup_{u\in I}|\Delta Y_{u}|^{2}\Big]\ \leq\ 16\,{\mathbb{E}}^{L}|\Delta\xi|^{2}\;+\;18\,{\mathbb{E}}^{L}[M_{\delta}^{2}].$
-Invoking the bound (4.28) and $\mathbb{E}^{L}|\Delta\xi|^{2}\leq(\mathbb{E}^{L}|\Delta\xi|^{4})^{1/2}$, we conclude
-${\mathbb{E}}^{L}\!\Big[\sup_{u\in I}|\Delta Y_{u}|^{2}\Big]\ \leq\ \left(16\;+\;108\,K_{\tilde{z}}^{2}\,C_{\mathrm{energy}}\,\sqrt{C_{2,-}(I)}\,\sqrt{C_{\mathrm{BDG}}}\right)\big({\mathbb{E}}^{L}|\Delta\xi|^{4}\big)^{1/2}.$ (4.29)
-We arrive at the local estimate $({\mathbb{E}}^{L}|\Delta\xi|^{4})^{1/2}\leq 2K_{0}({\mathbb{E}}^{L}|\Delta\xi|^{4})^{1/4}$:
-${\mathbb{E}}^{L}\!\Big[\sup_{u\in I}|\Delta Y_{u}|^{2}\Big]\;+\;{\mathbb{E}}^{L}\!\int_{T-\delta}^{T}|\Delta Z_{s}|^{2}ds\ \leq\ C_{\mathrm{loc}}(I)\,({\mathbb{E}}^{L}|\Delta\xi|^{4})^{1/4},$ (4.30)
-with the constant $C_{\mathrm{loc}}$ depending on $\delta$, $K_{0}$, $K_{z}$, $K_{\tilde{z}}$ and on the BMO/RHI/BDG constants.
-Step 2 (global stability).
-Let $0=t_{0}<t_{1}<\cdots<t_{N}=T$ with $t_{k+1}-t_{k}=\delta$ (small enough so the local estimates hold on each slice $I_{k}:=[t_{k},t_{k+1}]$).
-By Proposition 4.4 we have $\Delta Y_{t_{k+1}}\in L^{\infty}(\mathbb{P})$.
-Since essential suprema are invariant under equivalent changes of measure (Lemma 4.3), in particular, $\Delta\xi$ and $\Delta Y_{t_{k+1}}$ belong to $L^{\infty}$ under both $\mathbb{P}$ and $\mathbb{P}^{L}$.
-Hence (4.30) applies on every $I_{k}$ with terminal datum $\Delta Y_{t_{k+1}}\in L^{\infty}(\mathbb{P}^{L})$:
-$\mathbb{E}^{L}\!\Big[\sup_{u\in I_{k}}|\Delta Y_{u}|^{2}\Big]\;+\;\mathbb{E}^{L}\!\int_{I_{k}}|\Delta Z_{s}|^{2}\,ds\ \leq\ C_{\mathrm{loc}}(I_{k})\;\Big(\mathbb{E}^{L}|\Delta Y_{t_{k+1}}|^{4}\Big)^{1/4}.$ (4.31)
-Repeating Step 1 on $I_{k}$ with terminal $\Delta Y_{t_{k+1}}$ under $\mathbb{P}^{L}$ read
-$\Delta Y_{u}=\Delta Y_{t_{k+1}}+\int_{u}^{t_{k+1}}\widetilde{{\mathbb{E}}}\!\big[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}\big]\,ds-\int_{u}^{t_{k+1}}\Delta Z_{s}\cdot dW_{s}^{L},\qquad u\in I_{k}.$
-Evaluating at $u=t_{k}$ and setting
-$X_{k}:=\Delta Y_{t_{k+1}}+\int_{t_{k}}^{t_{k+1}}\widetilde{{\mathbb{E}}}\!\big[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}\big]\,ds,\qquad M_{\delta,k}:=\int_{t_{k}}^{t_{k+1}}\widetilde{{\mathbb{E}}}\!\big|\tilde{L}_{s}\,\Delta\tilde{Z}_{s}\big|\,ds,$
-we get
-$\Delta Y_{t_{k}}={\mathbb{E}}^{L}\!\big[X_{k}\,\big|\,\mathcal{F}_{t_{k}}\big]-\int_{t_{k}}^{t_{k+1}}\widetilde{{\mathbb{E}}}\!\big[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}\big]\,ds.$
-Using $(a+b)^{4}\leq 8(a^{4}+b^{4})$ and Jensen’s inequality,
-${\mathbb{E}}^{L}\!|\Delta Y_{t_{k}}|^{4}$ $\leq 8\,{\mathbb{E}}^{L}|X_{k}|^{4}+8\,{\mathbb{E}}^{L}\Big|\textstyle\int_{t_{k}}^{t_{k+1}}\widetilde{{\mathbb{E}}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds\Big|^{4}\leq 8\,{\mathbb{E}}^{L}|X_{k}|^{4}+8\,M_{\delta,k}^{4}$
-$\leq 8(8\,{\mathbb{E}}^{L}|\Delta Y_{t_{k+1}}|^{4}+8\,M_{\delta,k}^{4})+8\,M_{\delta,k}^{4}$
-$=64\,{\mathbb{E}}^{L}\!|\Delta Y_{t_{k+1}}|^{4}\;+\;72\,M_{\delta,k}^{4}$
-By repeating the bound for $M_{\delta}$ derived on a single slice (with $\Delta\xi$ replaced by the terminal datum $\Delta Y_{t_{k+1}}$), there exists a constant $C_{\delta}(I_{k})$ (depending only on $K_{z},K_{\tilde{z}},K_{0}$ and the BMO energy/RHI/BDG constants) such that
-$M_{\delta,k}\;\leq\;C_{\delta}(I_{k})\,\Big({\mathbb{E}}^{L}\big[|\Delta Y_{t_{k+1}}|^{4}\big]\Big)^{1/4}.$
-Therefore
-$\Big({\mathbb{E}}^{L}|\Delta Y_{t_{k}}|^{4}\Big)^{1/4}\;\leq\;\alpha_{\delta}(I_{k})\;\Big({\mathbb{E}}^{L}\!\big[|\Delta Y_{t_{k+1}}|^{4}\big]\Big)^{1/4},\qquad\alpha_{\delta}(I_{k}):=\big(64+72\,C_{\delta}(I_{k})^{4}\big)^{1/4}.$ (4.32)
-Iterating gives
-$\Big(\mathbb{E}^{L}|\Delta Y_{t_{k+1}}|^{4}\Big)^{1/4}\ \leq\ \prod\limits_{i=k}^{N-k-1}\alpha_{\delta}(I_{i+1})\;\Big(\mathbb{E}^{L}|\Delta\xi|^{4}\Big)^{1/4}.$ (4.33)
-Summing (4.31) over $k$ and using (4.33),
-$\mathbb{E}^{L}\!\Big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{2}\Big]+\mathbb{E}^{L}\!\int_{0}^{T}|\Delta Z_{s}|^{2}\,ds$ $\leq\sum_{k=0}^{N-1}\left(\mathbb{E}^{L}\!\Big[\sup_{u\in I_{k}}|\Delta Y_{u}|^{2}\Big]+\mathbb{E}^{L}\!\int_{I_{k}}|\Delta Z_{s}|^{2}\,ds\right)$
-$\leq\left(\sum_{k=0}^{N-1}C_{\mathrm{loc}}(I_{k})\,\prod\limits_{i=k}^{N-k-1}\alpha_{\delta}(I_{i+1})\right)\Big(\mathbb{E}^{L}|\Delta\xi|^{4}\Big)^{1/4}$
-$=:C_{\mathrm{global}}\Big(\mathbb{E}^{L}|\Delta\xi|^{4}\Big)^{1/4}.$
-Finally, the Girsanov kernel is a BMO martingale, the RHI transfers this global $\mathbb{P}^{L}$ bound back to $\mathbb{P}$, up to a multiplicative factor. This completes the proof. ∎
-Remark 4.7.
-Theorem 4.6 provides a stability estimate in $\mathbb{S}^{2}\times\mathbb{H}^{2}$.
-The proof is based on a global linearisation argument and can be generalized to obtain analogous $L^{p}$ stability estimates (for $p>2$) by similar argument.
-We do not pursue this level of generality here, as the $\mathbb{S}^{2}\times\mathbb{H}^{2}$ control is already sufficient for our purposes in the sense that the Banach space $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ is continuously embedded in $\mathbb{S}^{2}\times\mathbb{H}^{2}$.
-In particular, in the existence proof for the main result below, we construct approximating solutions and show that the sequence is Cauchy in $\mathbb{S}^{2}\times\mathbb{H}^{2}$ by applying Theorem 4.6.
-The limit in $\mathbb{S}^{2}\times\mathbb{H}^{2}$ then automatically belongs to $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$.
-5 Proof: existence and uniqueness of solutions
+By boundedness of $\Big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\Big\|_{\mathrm{BMO}}$, the constants $q^{*},C_{q}^{*}$ depend only on $(T,K_{z},K_{\tilde{z}},K_{0})$. Finally, set
+$M_{t}:=\int_{0}^{t}(\tilde{\mathbb{E}}|K_{\tilde{z}}\big(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|\big)|^{2})^{1/2}dW_{s},\qquad t\in[0,T].$
+Then $M\in\mathcal{H}^{\infty}(\mathbb{P})$, and by schachermayer1996institut it is sliceable.
+Furthermore, since $Z^{1},Z^{2}\in\mathbb{S}^{\infty}$, we can (by (jackson2022triangle, Remark 2.8)) construct a deterministic partition $0=t_{0}<t_{1}<\dots<t_{N}=T$ with arbitrarily small mesh size bounded by a prescribed $\varepsilon>0$.
+$\mathbb{S}^{p}$–control of $\Delta Y$.
+Fix $\varepsilon$, $k\in\{0,\dots,N-1\}$ and write $I_{k}=[t_{k},t_{k+1}]$.
+By subtracting (4.20) at $u=t_{k+1}$ from that at $u\in I_{k}$, we obtain the local BSDE
+$\Delta Y_{u}=\Delta Y_{t_{k+1}}+\int_{u}^{t_{k+1}}\widetilde{\mathbb{E}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds-\int_{u}^{t_{k+1}}\Delta Z_{s}\cdot dW^{L}_{s},\qquad u\in I_{k}.$
+Using Cauchy-Schwarz on the copy space, the growth bound (4.18), and then Fefferman’s inequality (Theorem 2.5 in kazamaki2006BMO) on $I_{k}$, we obtain
+$\int_{u}^{t_{k+1}}\tilde{\mathbb{E}}|\tilde{L}_{s}||\Delta\tilde{Z}_{s}|ds$ $\leq\mathbb{E}\int_{0}^{T}(\tilde{\mathbb{E}}\big|K_{\tilde{z}}\big(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|\big)\big|^{2})^{1/2}(\tilde{\mathbb{E}}|\Delta\tilde{Z}_{s}|^{2})^{1/2}ds$ (4.21)
+$\leq\sqrt{2}\left\|\int_{0}^{\cdot}(\tilde{\mathbb{E}}|K_{\tilde{z}}\big(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|\big)|^{2})^{1/2}dW_{s}\right\|_{BMO}\int_{u}^{t_{k+1}}(\mathbb{E}|\Delta Z_{s}|^{2})^{1/2}ds$
+$\leq\varepsilon\int_{u}^{t_{k+1}}(\mathbb{E}|\Delta Z_{s}|^{2})^{1/2}ds\qquad u\in I_{k}.$
+Taking $\mathbb{E}^{L}[\cdot\,|\,\mathcal{F}_{u}]$ in the local BSDE yields
+$|\Delta Y_{u}|$ $\leq\mathbb{E}^{L}[|\Delta Y_{t_{k+1}}|\,|\,\mathcal{F}_{u}]+\varepsilon\int_{u}^{t_{k+1}}\big(\mathbb{E}|\Delta Z_{s}|^{2}\big)^{1/2}ds,\qquad u\in I_{k}.$
+Iterating this estimate over the intervals $I_{k+1},I_{k+2},\dots,I_{N-1}$ and using $\Delta Y_{T}=\Delta\xi$, we obtain
+$|\Delta Y_{u}|\leq\mathbb{E}^{L}[|\Delta\xi|\,|\,\mathcal{F}_{u}]+\varepsilon\sqrt{T}\,\left(\int_{u}^{T}\mathbb{E}|\Delta Z_{s}|^{2}\,ds\right)^{1/2},\qquad u\in[0,T].$
+By BMO energy inequality (kazamaki2006BMO, Sec. 2.1) and Corollary 4.5, the last term is bounded by a constant depending only on $(T,K_{z},K_{\tilde{z}},K_{0})$ since:
+$\int_{u}^{T}\Big({\mathbb{E}}|Z_{s}^{1}|^{2}+{\mathbb{E}}|Z_{s}^{2}|^{2}\Big)\,ds\;$ $\leq\;\Big\|\int_{0}^{\cdot}Z_{s}^{1}\,\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2}+\Big\|\int_{0}^{\cdot}Z_{s}^{2}\,\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2}.$
+Therefore, letting $\varepsilon\to 0$ yields
+$|\Delta Y_{u}|\leq\mathbb{E}^{L}[|\Delta\xi|\,|\,\mathcal{F}_{u}]=(\Lambda_{u})^{-1}\,\mathbb{E}[\Lambda_{T}|\Delta\xi|\,|\,\mathcal{F}_{u}],\qquad u\in[0,T].$ (4.22)
+Now fix $q\in(1,q^{*})$ and let $p$ be its conjugate exponent. By the reverse Hölder inequality and conditional Hölder’s inequality,
+$|\Delta Y_{u}|^{p}$ $\leq(\Lambda_{u})^{-p}\,\mathbb{E}[(\Lambda_{T})^{q}\,|\,\mathcal{F}_{u}]^{p/q}\,\mathbb{E}[|\Delta\xi|^{p}\,|\,\mathcal{F}_{u}]$
+$\leq(C_{q}^{*})^{p/q}\,\mathbb{E}[|\Delta\xi|^{p}\,|\,\mathcal{F}_{u}],\qquad u\in[0,T].$
+By Doob’s maximal inequality,
+$\mathbb{E}\Big(\sup_{u\in[0,T]}|\Delta Y_{u}|^{p}\Big)\ \leq\ C_{Y}(p)\,\mathbb{E}|\Delta\xi|^{p},\qquad p\geq p^{*},$ (4.23)
+where $p^{*}$ is the conjugate exponent of $q^{*}$ and $C_{Y}(p):=\Big(\frac{p}{p-1}\Big)^{p}(C_{q}^{*})^{p/q}$.
+They both depending only on $(T,K_{z},K_{\tilde{z}},K_{0})$ as we discussed in the Preliminary step.
+$\mathbb{H}^{p}$–control of $\Delta Z$. Fix $p>p^{*}$, and applying Itô’s formula to $|\Delta Y|^{2}$ on $[0,T]$. we obtain
+$|\Delta Y_{0}|^{2}+\!\int_{0}^{T}|\Delta Z_{s}|^{2}\,ds=|\Delta\xi|^{2}+2\int_{0}^{T}\Delta Y_{s}\,\widetilde{\mathbb{E}}[\Delta g(s,Z_{s},\tilde{Z}_{s})]\,ds-2\int_{0}^{T}\Delta Y_{s}\,Z_{s}\cdot dW_{s}$
+Using the same Fefferman/Cauchy–Schwarz argument as above and $\|\Delta Y\|_{\mathbb{S}^{\infty}}<\infty$, we obtain
+$2\int_{0}^{T}\Delta Y_{s}\,\widetilde{{\mathbb{E}}}\!\big[\Delta g(s,Z_{s},\tilde{Z}_{s})\big]\,ds$ $\leq 2K_{z}\int_{0}^{T}(1+|Z_{s}^{1}|+|Z_{s}^{2}|)\,|\Delta Y_{s}|\,|\Delta Z_{s}|\,ds$
+$\quad+2K_{\tilde{z}}\int_{0}^{T}|\Delta Y_{s}|\,\widetilde{{\mathbb{E}}}\!\Big[(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|)\,|\Delta\tilde{Z}_{s}|\Big]\,ds$
+$\leq\sup_{t\in[0,T]}|\Delta Y_{t}|^{2}\Big(2K_{z}^{2}\int_{0}^{T}(1+|Z_{s}^{1}|+|Z_{s}^{2}|)^{2}\,ds\Big)+\frac{1}{2}\int_{0}^{T}|\Delta Z_{s}|^{2}\,ds$
+$\quad+2K_{\tilde{z}}\|\Delta Y\|_{\mathbb{S}^{\infty}}\int_{0}^{T}\widetilde{{\mathbb{E}}}\!\Big[(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|)\,|\Delta\tilde{Z}_{s}|\Big]\,ds.$
+The last term vanishes in the limit $\varepsilon\to 0$ by the same slicing argument as in (4.21). Next, let $K:=\int_{0}^{T}(1+|Z_{s}^{1}|+|Z_{s}^{2}|)^{2}\,ds$. Then
+${\mathbb{E}}\Big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{p}\,K^{p/2}\Big]\leq\Big({\mathbb{E}}\big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{2p}\big]\Big)^{1/2}\Big({\mathbb{E}}\big[K^{p}\big]\Big)^{1/2}.$ (4.24)
+If $\int_{0}^{\cdot}Z_{s}^{i}\cdot dW_{s}$ is a BMO martingale ($i=1,2$), the BMO energy inequality (see (kazamaki2006BMO, Sec. 2.1)) gives
+${\mathbb{E}}\Big[\Big(\int_{0}^{T}|Z_{s}^{i}|^{2}ds\Big)^{p}\Big]\leq p!\,\Big\|\int_{0}^{\cdot}Z_{s}^{i}\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2p}.$
+Therefore,
+${\mathbb{E}}[K^{p}]\leq 3^{p-1}\Big((3T)^{p}+3^{p}p!\,\Big\|\int_{0}^{t}Z_{s}^{1}\,\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2p}+3^{p}p!\,\Big\|\int_{0}^{t}Z_{s}^{2}\,\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2p}\Big)=:C_{\mathrm{energy}}(p).$ (4.25)
+Lastly, with the help of BDG inequality and Cauchy–Schwarz, we have
+${\mathbb{E}}\Big|\int_{0}^{T}(\Delta Y_{s})\Delta Z_{s}\cdot dW_{s}\Big|^{p/2}$ $\leq C_{\mathrm{BDG}}\,{\mathbb{E}}\Big[\Big(\int_{0}^{T}|\Delta Y_{s}|^{2}|\Delta Z_{s}|^{2}\,ds\Big)^{p/4}\Big]$
+$\leq C_{\mathrm{BDG}}\,\Big({\mathbb{E}}\big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{p}\big]\Big)^{1/2}\Big({\mathbb{E}}\big[(\int_{0}^{T}|\Delta Z_{s}|^{2}\,ds)^{p/2}\big]\Big)^{1/2}.$ (4.26)
+Putting (4.24), (4.25) and (4) together, we obtain
+${\mathbb{E}}\Big[\Big(\int_{0}^{T}|\Delta Z_{s}|^{2}ds\Big)^{p/2}\Big]$ $\leq 2^{\frac{3p}{2}-1}{\mathbb{E}}[|\Delta\xi|^{p}]+2^{\frac{3p}{2}-1}(2K_{z}^{2})^{p/2}\sqrt{C_{\mathrm{energy}}(p)}\Big({\mathbb{E}}[\sup_{t\in[0,T]}|\Delta Y_{t}|^{2p}]\Big)^{1/2}$
+$\quad+2^{4p-4}(C_{\mathrm{BDG}})^{2}\,{\mathbb{E}}[\sup_{t\in[0,T]}|\Delta Y_{t}|^{p}],$
+From previous step (4.23), there exist constants $C_{Y}(p),C_{Y}(2p)$ such that
+$\mathbb{E}\Big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{p}\Big]\leq C_{Y}(p)\,\mathbb{E}\big[|\Delta\xi|^{p}\big],\qquad\Big(\mathbb{E}\Big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{2p}\Big]\Big)^{1/2}\leq\sqrt{C_{Y}(2p)}\,\Big(\mathbb{E}\big[|\Delta\xi|^{2p}\big]\Big)^{1/2}.$
+Substituting these bounds into the previous inequality, we obtain for all $p\geq p^{*}$,
+$\mathbb{E}\Big[\Big(\int_{0}^{T}|\Delta Z_{s}|^{2}ds\Big)^{p/2}\Big]$ $\leq 2^{\frac{3p}{2}-1}\mathbb{E}\big[|\Delta\xi|^{p}\big]+2^{4p-4}(C_{\mathrm{BDG}})^{2}\,C_{Y}(p)\,\mathbb{E}\big[|\Delta\xi|^{p}\big]$
+$\quad+2^{\frac{3p}{2}-1}(2K_{z}^{2})^{p/2}\sqrt{C_{\mathrm{energy}}(p)}\,\sqrt{C_{Y}(2p)}\,\Big(\mathbb{E}\big[|\Delta\xi|^{2p}\big]\Big)^{1/2}.$
+Finally, since $\mathbb{E}|\Delta\xi|^{p}\leq\big(\mathbb{E}|\Delta\xi|^{2p}\big)^{1/2}$, we may absorb the first two terms into the last one and conclude that
+$\mathbb{E}\Big[\Big(\int_{0}^{T}|\Delta Z_{s}|^{2}ds\Big)^{p/2}\Big]$ $\leq C_{Z}(p)\,\Big(\mathbb{E}\big[|\Delta\xi|^{2p}\big]\Big)^{1/2},$
+where
+$C_{Z}(p):=2^{\frac{3p}{2}-1}+2^{4p-4}(C_{\mathrm{BDG}})^{2}\,C_{Y}(p)+2^{\frac{3p}{2}-1}(2K_{z}^{2})^{p/2}\sqrt{C_{\mathrm{energy}}(p)}\,\sqrt{C_{Y}(2p)}.$
+Combing the stability estimate (4.23) completes the proof. ∎
+5 Proof: existence and uniqueness
 In this section, we establish the main result for the mean-field BSDE (2.6) under Assumptions $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q})$.
@@
 Lemma 5.1.
-Let $(\rm H_{0})$} and $(\rm H_{q})$} holds and suppose $\xi$ is Malliavin differentiable with bounded derivative $\left\|D\xi\right\|_{\mathbb{S}^{\infty}}=\sup_{0\leq t\leq T}\left\|D_{t}\xi\right\|_{L^{\infty}}<\infty\text{ almost surely}$, then the mean-field BSDE admits a solution in $\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}.$
+Let $(\rm H_{0})$} and $(\rm H_{q})$} holds and suppose $\xi$ is Malliavin differentiable with bounded derivative $\left\|D\xi\right\|_{\mathbb{S}^{\infty}}<\infty$, then the mean-field BSDE admits a solution in $\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}.$
 Proof.
@@
 […]
-for all $k\in\mathbb{N}$ and $(t,z,z^{\prime},\tilde{z},\tilde{z}^{\prime})\in[0,T]\times(\mathbb{R}^{d})^{4}$.
+for all $k\in\mathbb{N}$ and $(t,z,z^{\prime},\tilde{z},\tilde{z}^{\prime})\in[0,T]\times[\mathbb{R}^{d}\times\mathbb{R}^{d}]^{2}$.
 Thus, the corresponding BSDE with driver $\bar{g}_{k}$ satisfies $(\rm H_{0})$} and $(H_{\mathrm{lip}})$.
@@
 We are now in a position to state the main result of this section.
-For the reader’s convenience, we restate the theorem below.
+For the reader’s convenience, we restate the main theorem below.
 To complete the proof of following theorem, it remains to extend the existence result from Malliavin differentiable terminal conditions to the general bounded case.
-Uniqueness then follows naturally from the stability result established in Theorem 4.6.
+Uniqueness then follows directly from the stability result established in Theorem 4.6.
 Theorem 5.2.
@@
 Hence, each $\xi^{n}$ remains bounded, and $(\xi^{n},g)$ satisfies $(\mathrm{H}_{0})$ for all $n\in\mathbb{N}$.
-By Lemma 5.1, each approximating BSDE with terminal condition $\xi^{n}$ admits a solution $(Y^{n},Z^{n})\in\mathbb{S}^{2}\times\mathbb{S}^{\infty}$, satisfying
+By Lemma 5.1, each approximating BSDE with terminal condition $\xi^{n}$ admits a solution $(Y^{n},Z^{n})\in\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}$, satisfying
 $Y_{t}^{n}=\xi^{n}+\int_{t}^{T}\int_{\tilde{\Omega}}g(s,Z_{s}^{n},\tilde{Z}_{s}^{n})\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}^{n}\cdot dW_{s}.$ (5.28)
-Furthermore, the stability result in Theorem 4.6 implies that (pick $p=4$), we have
-$\|Y^{n}-Y^{m}\|_{\mathbb{S}^{2}}^{2}+\|Z^{n}-Z^{m}\|_{\mathbb{H}^{2}}^{2}\ \leq\ C_{\mathrm{global}}\;\|\xi^{1}-\xi^{2}\|_{L^{4}},$
-where $C_{\mathrm{global}}$ depends only on $T$, $K_{z}$, $K_{\tilde{z}}$, and $K_{0}$, thus are independent of $n$ and $m$.
-Since $(\xi^{n})_{n\in\mathbb{N}}$ is Cauchy in $L^{4}$, it follows that $(Y^{n},Z^{n})_{n\in\mathbb{N}}$ is also Cauchy in $\mathbb{S}^{2}\times\mathbb{H}^{2}$.
+Furthermore, the stability result in Theorem 4.6 implies that
+$\|Y^{n}-Y^{m}\|_{\mathbb{S}^{2p}}+\|Z^{n}-Z^{m}\|_{\mathbb{H}^{p}}\ \leq\ C_{p}\;\|\xi^{n}-\xi^{m}\|_{L^{2p}}$
+where $C_{p}$ depends only on $(T,K_{z},K_{\tilde{z}},K_{0})$, thus is independent of $n$ and $m$.
+Pick $(\xi^{n})_{n\in\mathbb{N}}$ is Cauchy in $L^{2p}$, it follows that $(Y^{n},Z^{n})_{n\in\mathbb{N}}$ is also Cauchy in $\mathbb{S}^{2p}\times\mathbb{H}^{p}$.
 As each $(Y^{n},Z^{n})$ lies in the Banach space $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$, we conclude the existence of a limit $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$.
@@
 We will subsequently demonstrate that the process defined earlier for the constraint set is precisely the solution $Z$ in the BSDE (6.30).
-With the complete framework in place, we offer a financial interpretation of the components, which are discussed in Remarks 6.4.
-To connect this application to our earlier analysis about BSDE, we generalize the separately quadratic driver from BSDE (2.5) to fully coupled quadratic driver from BSDE (6.39) and establish the existence and uniqueness of solutions in Proposition 6.7.
-To do this we adopte a localization method in 6.6 then we extend it to the global case.
-Then this is the proof part for BSDE, then we connect this back to the utility maximization problem by presenting the final result in Theorem 6.9, we first derive a key estimate in Lemma 6.3.
+With the complete framework in place, we offer a financial interpretation of the components, which is discussed in Remarks 6.4.
+To connect this application to our earlier analysis about BSDE, we generalize the separately quadratic driver from BSDE (2.6) to a fully coupled quadratic driver in BSDE (6.32) and establish the existence and uniqueness of solutions in Proposition 6.7 when the terminal condition is small enough.
+To do this, we adopt a localization method in 6.6, then we extend it to the global case.
+Then, after we established the existence and uniqueness for BSDE (6.32), we then connect this back to the utility maximization problem by presenting the final result in Theorem 6.9.
+We first derive a key estimate in Lemma 6.3.
 This estimate is then used in Theorem 6.9 to verify that the driver of BSDE (6.30) satisfies conditions $(\rm H_{0})$} and $(\rm H_{q}^{\prime})$}, ensuring the well-posedness of BSDE (6.32).
@@
 Consider a financial market composed of $d$ risky stocks.
-In this market, there are countably infinite number of homogeneous investors who trades dynamically between the risk-free bond and the risky assets.
+In this market, there are countably infinite number of homogeneous investors who trade dynamically between the risk-free bond and the risky assets.
 An investor, a representative of homogeneous investors, uses the following model to maximize the utility.
@@
 $\frac{\mathrm{d}S_{t}^{i}}{S_{t}^{i}}=b_{t}^{i}\,\mathrm{d}t+\sigma_{t}^{i}\,\mathrm{d}W_{t},\quad i=1,\dots,d,$
-where $W$ is $d$-dimensional Brownian motion, $b^{i}$ is a real-valued predictable, bounded stochastic process, representing the drift, and $\sigma^{i}$ is an $\mathbb{R}^{d}$-valued predictable bounded stochastic process, representing the volatility.
+where $W$ is $n$-dimensional Brownian motion, $b^{i}$ is a real-valued predictable, bounded stochastic process, representing the drift, and $\sigma^{i}$ is an $\mathbb{R}^{n}$-valued predictable bounded stochastic process, representing the volatility.
 The volatility matrix $\sigma_{t}=(\sigma_{t}^{1},\dots,\sigma_{t}^{d})^{\top}$ is assumed to have full rank $d$.
-The market price of risk $\theta_{t}$, also referred to as the risk premium, is defined by:
+And we assume that $\sigma_{t}\sigma_{t}^{\top}$ is uniformly elliptic.
+Thus market price of risk $\theta_{t}$, also referred to as the risk premium, is defined by:
 $\theta_{t}:=\sigma_{t}^{\top}(\sigma_{t}\sigma_{t}^{\top})^{-1}b_{t},\quad t\in[0,T],$
-where $\theta_{t}$ is an $\mathbb{R}^{n}$-valued bounded predictable process.
+where $\theta_{t}$ is an $\mathbb{R}^{n}$-valued uniformly bounded predictable process.
 For $1\leq i\leq d$, let $\pi^{i}_{t}$ represent the amount of money invested in stock $i$ at time $t$, meaning the number of shares held in stock $i$ is $\frac{\pi^{i}_{t}}{S^{i}_{t}}$.
@@
 $U(x)=-\exp(-\alpha x),\quad x\in\mathbb{R}.$
-For the investors other than the representative, we assume the risk aversion parameter $\alpha>0$ are identical.
-The model for the risky assets are of the same type but with $b,\sigma,W$ being independent and identically distributed to the representative’s model.
-In addition, we assume the terminal liabilities are independent and has distribution identical to $F$.
+For the investors other than the representative, we assume the risk aversion parameter $\alpha>0$ is identical.
+The model for the risky assets is of the same type but with $b,\sigma,W$ being independent and identically distributed to the representative’s model.
+In addition, we assume the terminal liabilities are independent and have distribution $F$.
 6.2 Constraint set and aggregated market signal
@@
 Below, we outline the constraint set in Definition 6.1, which accounts for market-wide influences.
-To incorporate for the influence of market-wide behavior, we denote $\tilde{Z}$ for the trading strategy adjusted by $\theta/\alpha$, which is an $\mathbb{R}^{1\times d}$-valued stochastic process that satisfies the necessary integrability conditions.
-More precisely, $\tilde{Z}_{t}\sigma_{t}+\frac{\theta_{t}}{\alpha}$ yields the optimal strategy at time $t\in[0,T]$.
-In addition, for $t\in[0,T]$ and $\omega\in\Omega$, we define
-$Z_{t}(\omega):=\tilde{Z_{t}}(\omega)\sigma_{t}(\omega),$
-where $Z_{t}(\omega)$ is an element of $\mathbb{R}^{1\times n}$.
+To incorporate the influence of market-wide behavior, for $t\in[0,T]$ and $\omega\in\Omega$, we define two stochastic processes that satisfy the necessary integrability conditions such that
+$Z_{t}(\omega):=\hat{Z_{t}}(\omega)\sigma_{t}(\omega),$
+where $Z_{t}(\omega)\in\mathbb{R}^{1\times n}$ and $\hat{Z_{t}}(\omega)\in\mathbb{R}^{1\times d}$.
 These two stochastic processes are essentially equivalent because the matrix $\sigma$ has full rank.
-Later, we will show this $Z$ will precisely corresponds to the martingale density process of BSDE (6.37).
+Later on, in Remark 6.4, we will see $\hat{Z}_{t}\sigma_{t}+\frac{\theta_{t}}{\alpha}=Z_{t}+\frac{\theta_{t}}{\alpha}$ yields the optimal strategy at time $t\in[0,T]$ if no constants are imposed.
+Thus $Z$ can be interpreted as the trading strategy adjusted by $\theta/\alpha$.
+Later, we will show that this $Z$ will precisely correspond to the martingale density process of BSDE (6.30) as well.
 We now present the following expressions for the constraint set under different settings.
-Specifically, we denote the constraint without market influence by $\tilde{D}\subseteq\mathbb{R}^{d}$, where $\tilde{D}$ is a closed set.
-When market influence depending on $\mathbb{E}[\tilde{Z}$] is present, the strategy $\pi_{t}(\omega)$ is constrained to $\tilde{C}(\mathbb{E}[\tilde{Z}_{t}])$, where $\tilde{C}(\mathbb{E}[\tilde{Z}_{t}])\subseteq\mathbb{R}^{d}$ is also a closed set.
-The expectation of $\tilde{Z}$ is referred to as the aggregated market signal, representing the expected collective adjusted trading strategy of all market participants.
+Specifically, we denote the constraint without market influence by $\hat{D}\subseteq\mathbb{R}^{d}$, where $\hat{D}$ is a closed set.
+When market influence is present, the strategy $\pi_{t}(\omega)$ is constrained to a new closed set.
+The expectation of $\hat{Z}$ is referred to as the aggregated market signal, representing the expected collective adjusted trading strategy of all market participants.
 This signal plays a crucial role in defining the feasible trading region (constraint set), as described below.
 Definition 6.1 (Mean-field adaptive constraint set).
-Let $\tilde{D}$ be a closed set in $\mathbb{R}^{d}$, and define the mean-field adaptive constraint set as:
-$\tilde{C}(\mathbb{E}[\tilde{Z}_{t}]):=\tilde{D}\odot L(\mathbb{E}[\tilde{Z}_{t}]),\quad t\in[0,T],$ (6.36)
-where $\odot$ denotes the Hadamard product (component-wise multiplication) and $L(\mathbb{E}[\tilde{Z}_{t}])$ is a linear mapping depending on the value of $\mathbb{E}[\tilde{Z}_{t}]$ for fixed $t\in[0,T]$, representing how the trader adjusts decisions based on the market’s behavior, defined by:
-$L(\mathbb{E}[\tilde{Z}_{t}]):=(\gamma\odot\mathbb{E}[\tilde{Z}_{t}])+\eta,$
+Let $\hat{D}$ be a closed set in $\mathbb{R}^{d}$, and define the mean-field adaptive constraint set as:
+$\hat{C}(\mathbb{E}[\hat{Z}_{t}]):=\hat{D}\odot L(\mathbb{E}[\hat{Z}_{t}]),\quad t\in[0,T],$ (6.29)
+where $\odot$ denotes the Hadamard product (component-wise multiplication) and $L(\mathbb{E}[\hat{Z}_{t}])$ is a linear mapping depending on the value of $\mathbb{E}[\hat{Z}_{t}]$ for fixed $t\in[0,T]$, representing how the trader adjusts decisions based on the market’s behavior, defined by:
+$L(\mathbb{E}[\hat{Z}_{t}]):=(\gamma\odot\mathbb{E}[\hat{Z}_{t}])+\eta,$
 where $\gamma:=(\gamma_{1},\dots,\gamma_{d})$ is the market sensitivity vector and $\eta\in\mathbb{R}^{1\times d}$ is the market shift vector, with both being constant vectors.
-With the help of the previous definitions, for a fixed $t\in[0,T]$, we start with a set $\tilde{D}\in\mathbb{R}^{d}$, which represents certain constraints, such as risk limits or position constraints.
+With the help of the previous definitions, for a fixed $t\in[0,T]$, we start with a set $\hat{D}\subset\mathbb{R}^{d}$, which represents certain constraints, such as risk limits or position constraints.
 Upon incorporating information from the market, we could apply scaling and shifting to this set, which now reflects both the original constraints and the market’s influence.
 This results in a new set that is much more flexible and realistic, capturing the dynamic behavior of the market.
-Therefore, we denote this updated set as $\tilde{C}(\mathbb{E}[\tilde{Z}_{t}])$, emphasizing its dependence on the aggregated market signal $\mathbb{E}[\tilde{Z_{t}}$] at time $t\in[0,T]$.
-Definition 6.2 (Admissible strategies with constraints $\tilde{\mathcal{A}}$).
-Let $\tilde{C}(\mathbb{E}[\tilde{Z}_{t}])$ be a closed set in $\mathbb{R}^{d}$ for fixed $t\in[0,T]$.
-The set of admissible trading strategies $\tilde{\mathcal{A}}$ consists of all $\mathbb{R}^{d}$-valued predictable processes $\pi$ satisfying $\pi_{t}\sigma_{t}\in L^{2}[0,T]$, $\pi_{t}\in\tilde{C}(\mathbb{E}[\tilde{Z}_{t}])$ a.s.
+Therefore, we denote this updated set as $\hat{C}(\mathbb{E}[\hat{Z}_{t}])$, emphasizing its dependence on the aggregated market signal $\mathbb{E}[\hat{Z_{t}}$] at time $t\in[0,T]$.
+Definition 6.2 (Admissible strategies with constraints $\hat{\mathcal{A}}$).
+Let $\hat{C}(\mathbb{E}[\hat{Z}_{t}])$ be a closed set in $\mathbb{R}^{d}$ for fixed $t\in[0,T]$.
+The set of admissible trading strategies $\hat{\mathcal{A}}$ consists of all $\mathbb{R}^{d}$-valued predictable processes $\pi$ satisfying $\pi_{t}\sigma_{t}\in L^{2}[0,T]$, $\pi_{t}\in\hat{C}(\mathbb{E}[\hat{Z}_{t}])$ a.s.
 $\forall t\in[0,T]$, and
@@
 The investor’s objective is to choose an admissible self-financing trading strategy $\pi^{*}$ to maximize the expected utility of the net wealth at maturity $T$:
-$V(0,x):=\sup_{\pi\in\tilde{\mathcal{A}}}\mathbb{E}\left[-\exp\left(-\alpha\left(x+\int_{0}^{T}\pi_{u}\frac{dS_{u}}{S_{u}}-F\right)\right)\right],$
-where $V(0,\cdot)$ is called the value function at initial time 0. Solving this optimization problem requires selecting an admissible set from which we select the optimal trading strategy $\pi^{*}$.
+$V(0,x):=\sup_{\pi\in\hat{\mathcal{A}}}\mathbb{E}\left[-\exp\left(-\alpha\left(x+\int_{0}^{T}\pi_{u}\frac{dS_{u}}{S_{u}}-F\right)\right)\right],$
+where $V(0,\cdot)$ is called the value function at initial time 0.
+Solving this optimization problem requires selecting an admissible set from which we select the optimal trading strategy $\pi^{*}$.
 To make this notation more compact, we introduce the following notations.
 For $t\in[0,T],\omega\in\Omega$, define the set $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)\subseteq\mathbb{R}^{n}$ and (volatility-scaled) trading strategy $p_{t}$ such that
-$C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right):=\tilde{C}(\mathbb{E}[\tilde{Z}_{t}])\,\sigma_{t}(\omega)\quad\text{and}\quad p_{t}:=\pi_{t}\sigma_{t}.$
+$C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right):=\hat{C}(\mathbb{E}[\hat{Z}_{t}])\,\sigma_{t}(\omega)\quad\text{and}\quad p_{t}:=\pi_{t}\sigma_{t}.$
+Note $Z_{t}=\hat{Z_{t}}\sigma_{t}$ and $\mathbb{E}[Z_{t}]=\mathbb{E}[\hat{Z_{t}}]\sigma_{t}$.
 Lemma 6.3.
@@
 Proof.
-Since the market sensitivity vector $\gamma$ and the market shift vector $\eta$ are both constant, the process $\sigma$ is uniformly bounded, and $\tilde{D}$ is closed, the result follows. ∎
-Note that the admissible trading strategy set $\mathcal{A}$, which is equivalent to $\tilde{\mathcal{A}}$ (since $\sigma$ has full rank), consists of all $\mathbb{R}^{n}$-valued predictable processes $p$ that satisfy $p\in L^{2}[0,T]$ and $p_{t}(\omega)\in C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)$ almost surely for all $t\in[0,T]$.
+Since the market sensitivity vector $\gamma$ and the market shift vector $\eta$ are both constant, the process $\sigma$ is uniformly bounded, and $\hat{D}$ is closed, the result follows. ∎
+Note that the admissible trading strategy set $\mathcal{A}$, which is equivalent to $\hat{\mathcal{A}}$ ($\sigma$ has full rank), consists of all $\mathbb{R}^{n}$-valued predictable processes $p$ that satisfy $p\in L^{2}[0,T]$ and $p_{t}(\omega)\in C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)$ almost surely for all $t\in[0,T]$.
 Additionally, the family
@@
 is uniformly integrable.
-Finally, the dependency of the set $\tilde{C}(\mathbb{E}[\tilde{Z}_{t}])$ on the process $(\tilde{Z}_{t})_{0\leq t\leq T}$ is encoded by the dependency of the set $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)$ on the process $(Z_{t})_{0\leq t\leq T}$.
+Finally, the dependency of the set $\hat{C}(\mathbb{E}[\hat{Z}_{t}])$ on the process $(\hat{Z}_{t})_{0\leq t\leq T}$ is encoded by the dependency of the set $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)$ on the process $(Z_{t})_{0\leq t\leq T}$.
 We keep the notation $X^{p}$ for the wealth process corresponding to investing strategy $p$.
@@
 However, in this context, we now define it as the martingale density process $(Z_{t})_{0\leq t\leq T}$ for the BSDE in equation (6.30).
-Consequently, as the constraint set depends on $\mathbb{E}[Z_{t}]$, the driver has dependence on $\mathbb{E}[Z_{t}]$: see (6.38).
-Let us specify the driver $f$ for the construction of $\left\{V^{(p)}(\cdot,x):p\in\mathcal{A}\right\}$. By Itô’s formula,
+Consequently, as the constraint set depends on $\mathbb{E}[Z_{t}]$, the driver should has dependence on $\mathbb{E}[Z_{t}]$ as well: see (6.31).
+To verify this, let us specify the driver $f$ for the construction of $\left\{V^{(p)}(\cdot,x):p\in\mathcal{A}\right\}$.
+By Itô’s formula,
 $V^{(p)}(t,x)=V^{(p)}(0,x)\tilde{A}^{p}\mathcal{E}_{t}\left(\int_{0}^{\cdot}\alpha(Z_{s}-p_{s})\cdot\mathop{}\!dW_{s}\right),$
@@
 Therefore, $V^{(p^{*})}(\cdot,x)$ becomes a true martingale.
-We have now completed the construction, and it is clear how the utility maximization problem connects to some types of BSDEs.
+We have now completed the construction, and it is clear how the utility maximization problem in our setting connects to a type of mean-field type BSDEs as in hibon2017MFquadratic.
 In hu2005utility, the optimization problem was solved under the assumption that the random endowment $F$ is bounded and that the trading strategy $p$ takes values in the admissible set $\mathcal{A}$.
@@
 Remark 6.4.
-If we take $\tilde{D}$ to be the entire space, i.e., $\tilde{D}=\mathbb{R}^{d}$, then the resulting constraint set $\tilde{C}(\mathbb{E}[\tilde{Z}_{t}])$ is simply $\mathbb{R}^{d}$ as well.
+If we take $\hat{D}$ to be the entire space, i.e., $\hat{D}=\mathbb{R}^{d}$, then the resulting constraint set $\hat{C}(\mathbb{E}[\hat{Z}_{t}])$ is again $\mathbb{R}^{d}$ as well.
 Consequently, we obtain $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)=\mathbb{R}^{n}$, implying the absence of constraints.
@@
 Thus, the process $Z_{t}=p_{t}^{*}-\frac{\theta_{t}}{\alpha}$ can be interpreted as the unconstrained trading strategy adjusted by $\frac{\theta_{t}}{\alpha}$ at time $t\in[0,T]$.
-In general, the constraint set $C$ depends on the average of the adjusted trading strategies among the market participants.
+In general situation, the constraint set $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)$ depends on the average of the adjusted trading strategies among the market participants.
 The convergence is guaranteed by the homogeneity and exchangeability of market participants.
@@
 We are now ready to present the final result.
-We first borrow some ideas from tevzadze2008solvability, under small terminal value and lipschitz coefficient, the mean-field BSDE 6.39 admits a unique adapted solution.
-Next, in Proposition 6.7 we demonstrate that the driver in BSDE 6.37 guarantees the existence and uniqueness of solution.
+We first borrow some ideas from tevzadze2008solvability, under small terminal value and lipschitz coefficient, the mean-field BSDE 6.32 admits a unique adapted solution.
+Next, in Proposition 6.7 we demonstrate that the driver in BSDE 6.30 guarantees the existence and uniqueness of solution.
 Finally, we complete this section by presenting the maximum expected utility value and corresponding trading strategy in in Theorem 6.9.
-For the purpose of the final result, we introduce the BSDE (6.39) with the same formulation and the same $(\rm H_{0})$} in the main result but with a fully coupled quadratic growth condition and assume the following BSDE with driver $g$ and terminal condition $\xi$, satisfying the conditions specified below:
+For the purpose of the final result, we introduce the BSDE (6.32) with the same formulation and the same $(\rm H_{0})$} condition in the main result but with a fully coupled quadratic growth condition and assume the following BSDE with driver $g$ and terminal condition $\xi$, satisfying the conditions specified below:
 $Y_{t}=\xi+\int_{t}^{T}\int_{\tilde{\Omega}}g(s,Z_{s},\tilde{Z}_{s})\mathop{}\!d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}\cdot\mathop{}\!dW_{s},\quad t\in[0,T].$ (6.32)
@@
 $g(t,z,\tilde{z}):=z\,\tilde{z},\qquad(t,z,\tilde{z})\in[0,T]\times\mathbb{R}^{2}.$
-This $g$ satisfies the fully coupled quadratic condition $(H_{q}^{\prime})$, but does not satisfy the separately quadratic condition $(\rm H_{q})$}.
+This $g$ satisfies the fully coupled quadratic condition $(\rm H_{q}^{\prime})$}, but does not satisfy the separately quadratic condition $(\rm H_{q})$}.
 We begin by establishing existence and uniqueness of solutions on a sufficiently small time interval in Lemma 6.6, where the required smallness conditions for the terminal value and time horizon are satisfied.
 By concatenating these local solutions, we extend the result to the global interval $[0,T]$ in Proposition 6.7.
-First, we note that the terminal condition in BSDE 6.37 is $F$, representing a random liability.
+First, we note that the terminal condition in BSDE 6.30 is $F$, representing a random liability.
 It is not realistic to assume $F$ is close to zero.
 Instead, we relax this requirement and require that $F$ is close to its mean.
-Accordingly, we rewrite the general BSDE 6.39 in terms of the centered random variable $\eta$:
+Accordingly, we rewrite the general BSDE 6.32 in terms of the centered random variable $\eta$:
 $U_{t}=\eta+\int_{t}^{T}\int_{\tilde{\Omega}}\big[g(s,Z_{s},\tilde{Z}_{s})-g(s,0,0)\big]\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},\qquad 0\leq t\leq T,$ (6.33)
 where
-$U_{t}:=Y_{t}-g(t,0,0)-\mathbb{E}[\xi],\qquad\eta:=\xi-\mathbb{E}[\xi].$
+$U_{t}:=Y_{t}-g(t,0,0)-\mathbb{E}[\xi],\qquad\eta:=\xi-\mathbb{E}[\xi]$
+are both well-defined because of $(\rm H_{0})$}.
 Lemma 6.6.
-Let $T<(64L_{z}^{2})^{-1}$ and assume that the driver $g$ satisfies  $(\mathrm{H}_{q}^{\prime})$ and terminal condition $\eta$ be a bounded terminal condition such that
+Let $T<(64L_{z}^{2})^{-1}$ and assume that the driver $g$ satisfies $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q}^{\prime})$ and terminal condition $\eta$ be a bounded terminal condition such that
 $\left\|\eta\right\|_{L^{\infty}}\leq\rho,$
@@
 $\rho^{2}:=\frac{1-64L_{z}^{2}T}{65K_{z}}.$
-Then the BSDE 6.39 admits a unique solution $(U,Z)\in\mathcal{B}_{\rho}$, where
+Then the BSDE 6.33 admits a unique solution $(U,Z)\in\mathcal{B}_{\rho}$, where
 $\mathcal{B}_{\rho}:=\left\{(U,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}:\left\|U\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|Z\right\|_{BMO}^{2}\leq\rho^{2}\right\}.$
@@
 Applying Itô’s formula to $|U_{t}|^{2}$ on $[t,T]$, we obtain
-$|U_{t}|^{2}+\int_{t}^{T}|Z_{s}|^{2}ds=|\eta|^{2}+\int_{t}^{T}2U_{s}\int_{\tilde{\Omega}}g(s,z_{s},\tilde{z}_{s})-g(s,0,0)\,d\tilde{\mathbb{P}}ds-2\int_{t}^{T}U_{s}\cdot Z_{s}dW_{s}.$
+$|U_{t}|^{2}+\int_{t}^{T}|Z_{s}|^{2}ds=|\eta|^{2}+\int_{t}^{T}2U_{s}\int_{\tilde{\Omega}}g(s,z_{s},\tilde{z}_{s})-g(s,0,0)\,d\tilde{\mathbb{P}}ds-2\int_{t}^{T}U_{s}\cdot Z_{s}\!\cdot dW_{s}.$
 Taking the conditional expectation and applying $2ab\leq\frac{1}{2}a^{2}+2b^{2}$, we deduce
-$|U_{t}|^{2}+\mathbb{E}\left[\int_{t}^{T}|Z_{s}|^{2}ds\mid\mathcal{F}_{t}\right]$ $\leq\left\|\eta\right\|_{L^{\infty}}^{2}+2\left\|U\right\|_{\mathbb{S}^{\infty}}\mathbb{E}\left[\int_{t}^{T}\Big|\int_{\tilde{\Omega}}g(s,z_{s},\tilde{z}_{s})-g(s,0,0)d\tilde{\mathbb{P}}\Big|ds\mid\mathcal{F}_{t}\right]$
-$\leq\left\|\eta\right\|_{L^{\infty}}^{2}+\frac{1}{2}\left\|U\right\|_{\mathbb{S}^{\infty}}^{2}+2\mathbb{E}\left[\int_{t}^{T}\int_{\tilde{\Omega}}|g(s,z_{s},\tilde{z}_{s})-g(s,0,0)|^{2}d\tilde{\mathbb{P}}ds\Big|\mathcal{F}_{t}\right].$ (6.41)
+$|U_{t}|^{2}+\mathbb{E}\left[\int_{t}^{T}|Z_{s}|^{2}ds\mid\mathcal{F}_{t}\right]$ $\leq\left\|\eta\right\|_{L^{\infty}}^{2}+\frac{1}{2}\left\|U\right\|_{\mathbb{S}^{\infty}}^{2}+2\mathbb{E}\left[\int_{t}^{T}\int_{\tilde{\Omega}}|g(s,z_{s},\tilde{z}_{s})-g(s,0,0)|d\tilde{\mathbb{P}}ds\Big|\mathcal{F}_{t}\right]^{2}.$ (6.34)
 Together with the $(\rm H_{q}^{\prime})$} condition and Jensen’s inequality, we have
@@
 and thus the mapping $\Gamma$ is a contraction and admits a unique fixed point in $\mathcal{B}_{R}$.
-Step 3. Finally, together with the local condition $T<(64L_{z}^{2})^{-1}$, the smallness condition (6.43) is in fact more stringent than (6.45). Consequently, if we define $\rho>0$ by
+Step 3.
+Finally, together with the local condition $T<(64L_{z}^{2})^{-1}$, the smallness condition (6.36) is in fact more stringent than the smallness condition (6.38).
+Consequently, if we define $\rho>0$ by
 $\rho^{2}:=\frac{1-64L_{z}^{2}T}{65K_{z}},$
@@
 Proposition 6.7.
-Let $T>0$, and assume that $g$ satisfies $(\mathrm{H}_{q}^{\prime})$. Suppose there exists $\rho>0$ such that
+Let $T>0$, and assume that $g$ satisfies $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q}^{\prime})$. Suppose there exists $\rho>0$ such that
 $\left\|\eta\right\|_{L^{\infty}}\leq\rho,$
@@
 ∎
-We remark that, given the unique solution $(U,Z)\in\mathcal{B}_{\rho}$ constructed above for the auxiliary BSDE 6.40, the solution to the original BSDE 6.39 is obtained by the explicit transformation
+We remark that, given the unique solution $(U,Z)\in\mathcal{B}_{\rho}$ constructed above for the auxiliary BSDE 6.33, the solution to the original BSDE 6.32 is obtained by the explicit transformation
 $Y_{t}=U_{t}+g(t,0,0)+\mathbb{E}[\xi],\qquad 0\leq t\leq T.$
@@
 Remark 6.8.
-Proposition 6.7 was established for drivers of the form $\int_{\tilde{\Omega}}g(s,Z_{s},\tilde{Z}_{s})\,d\tilde{\mathbb{P}}$, which naturally extends to the case where the driver depends on the conditional expectation, i.e., $g(s,Z_{s},\mathbb{E}[Z_{s}])$.
+Proposition 6.7 was established for drivers of the form $\int_{\tilde{\Omega}}g(s,Z_{s},\tilde{Z}_{s})\,d\tilde{\mathbb{P}}$, which naturally extends to the case where the driver depends on the expectation $\mathbb{E}[Z]$ instead.
 In particular, the existence and uniqueness result applies to BSDEs of the type
 $Y_{t}=\xi+\int_{t}^{T}g(s,Z_{s},\mathbb{E}[Z_{s}])\,ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},\quad t\in[0,T],$
-which will be useful for analyzing the driver (6.38) in this section.
-While the above analysis focuses on drivers with $Z$ and mean-field dependence, it is very easy to extend existence and uniqueness results to more general settings where the driver $g$ may also depend on the process $Y$ and its expectation $\mathbb{E}[Y]$ as in hibon2017MFquadratic.
+which will be useful for analyzing the driver (6.31) we picked earlier in the context of utility maximization in this section.
+While the above analysis focuses on drivers with $Z$ and mean-field dependence, it is very easy to extend existence and uniqueness results to more general settings where the driver $g$ may also depend on the process $Y$ and its expectation $\mathbb{E}[Y]$.
 In the present section, however, the application of interest involves drivers depending solely on $Z$ and its expectation, so the above results apply directly.
@@
 Proof.
-First we will show the driver $f$ (6.38) satisfies the condition $(\mathrm{H}_{q}^{\prime})$.
+First we will show the driver $f$ (6.31) satisfies the condition $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q}^{\prime})$.
+The verification of the first condition is straightforward due to the boundedness of $\theta$.
 To verify $(\mathrm{H}_{q}^{\prime})$, we apply Lemma 6.3.
@@
 $\quad\leq 2|\theta_{t}|\,|Z_{t}^{1}-Z_{t}^{2}|+\frac{\alpha}{2}(1\vee K_{1})^{2}\left(|Z_{t}^{1}|+|Z_{t}^{2}|+\mathbb{E}|Z_{t}^{1}|+\mathbb{E}|Z_{t}^{2}|\right)\left(|Z_{t}^{1}-Z_{t}^{2}|+\mathbb{E}|Z_{t}^{1}-Z_{t}^{2}|\right),\quad\text{a.s.}$
-Together with the boundedness of $\theta$, we obtain, for some constants $c_{1},c_{2}>0$, for all $t\in[0,T]$
-$\left|f\left(t,Z_{t}^{1},\mathbb{E}[Z_{t}^{1}]\right)-f\left(t,Z_{t}^{2},\mathbb{E}[Z_{t}^{2}]\right)\right|$
-$\quad\leq c_{1}\,|Z_{t}^{1}-Z_{t}^{2}|+c_{2}\left(|Z_{t}^{1}|+|Z_{t}^{2}|+\mathbb{E}|Z_{t}^{1}|+\mathbb{E}|Z_{t}^{2}|\right)\left(|Z_{t}^{1}-Z_{t}^{2}|+\mathbb{E}|Z_{t}^{1}-Z_{t}^{2}|\right),\quad\text{a.s.}$
-which is a sufficient condition for $(\rm H_{q}^{\prime})$}.
+Together with the boundedness of $\theta$, we obtain a sufficient condition for $(\rm H_{q}^{\prime})$}.
 This ensures the existence and uniqueness of a solution $(Y,Z)$ to BSDE (6.30) by Proposition 6.7.
@@
 7 Conclusion
-In this paper we have developed a general framework for quadratic mean-field BSDEs whose generators exhibit quadratic growth in the control and its law.
+In this paper we have developed a general framework for mean-field BSDEs whose generators have quadratic growth in the control and its law.
 Working under the separately quadratic structure (Hq) with bounded terminal condition, we proved global existence and uniqueness of solutions in the Banach space $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$, without imposing any smallness condition.

### v3

@@
 Abstract
-In this paper, we study a class of real-valued mean-field backward stochastic differential equations (BSDEs) with generator of quadratic growth in the control variable and the mean-field term.
-Under this assumption, together with a bounded terminal condition, we establish existence and uniqueness of solutions.
+In this paper, we study a class of real-valued mean-field backward stochastic differential equations (BSDEs) with generators of quadratic growth in the control variable and the mean-field term.
+Under this assumption, together with a bounded terminal condition, we establish the existence and uniqueness of solutions.
 Our approach departs from classical fixed-point arguments and instead combines Malliavin calculus with refined BMO and stability estimates.
-The result bridges the gap between the quadratic BSDE results of cheridito2017bse and hao2022mean.
-Moreover, motivated by the structure of the mean-field exponential utility maximization problem introduced in our paper, we extend our framework to generators satisfying a weaker quadratic condition on the generator.
-This relaxation is designed to accommodate the additional mean-field terms that arise in our utility maximization setting and that fall outside the scope of previous quadratic assumptions.
-Within this more general regime, we establish existence and uniqueness of solutions under a smallness condition on the terminal random variable.
-We then apply this extended theory to solve a mean-field exponential utility maximization problem, thereby generalizing the classical framework of hu2005utility to a fully coupled quadratic mean-field setting.
-Key words: mean-field BSDEs; quadratic BSDEs; BMO martingales; Malliavin calculus; exponential utility maximization.
+The result bridges the gap between the quadratic BSDE results of [Ann.
+Probab.
+45 (2017), pp.
+3795–3828] and Hao et al.
+[Ann.
+Appl.
+Probab.
+35 (2025), pp.
+2128–2174].
+Moreover, motivated by the structure of the mean-field exponential utility maximization problem introduced in our paper, we extend our framework to terminal conditions without continuity or the Markovian assumption.
+We establish the existence and uniqueness of solutions under a smallness terminla value on the terminal conditions.
+We then apply this extended theory to solve a mean-field exponential utility maximization problem, which developing the classical framework of Hu et al.
+[Ann.
+Appl.
+Probab.
+15 (2005), pp.
+1691–1712] to a fully coupled quadratic mean-field setting.
+Key words: Mean-field BSDEs; quadratic BSDEs; BMO martingales; Malliavin calculus; exponential utility; maximal utility.
 MSC: 60H30; 60H07; 60H10.
@@
 […]
-The origin of linear BSDEs can be traced back to their role as adjoint equations in stochastic control problems, first studied by bismut1973conjugate.
-The seminal paper by pardoux1990adapted established the existence and uniqueness of adapted solutions $(Y,Z)$ for
+The origin of linear BSDEs can be traced back to their role as adjoint equations in stochastic control problems, first studied by Bismut (1973).
+The seminal paper by Pardoux and Peng (1990) established the existence and uniqueness of adapted solutions $(Y,Z)$ for
 $Y_{t}=\xi+\int_{t}^{T}F(s,Y_{s},Z_{s})\,ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},\quad t\in[0,T],$ (1.2)
 under the conditions that the generator $F$ is Lipschitz continuous and the terminal condition $\xi$ is square integrable.
-Since then, there has been a growing interest in BSDEs with applications in various fields such as stochastic control, mathematical finance, and partial differential equations.
-For a comprehensive presentation of the theory, see el1997backward, CohenStocal2015 or ZhangBSDEbook2017.
+Since then, there has been growing interest in BSDEs, with applications in stochastic control, mathematical finance, and partial differential equations.
+For a comprehensive presentation of the theory, see El Karoui et al.
+(1997), Cohen and Elliott (2015), or Zhang (2017).
 Numerous studies have aimed to improve existence and uniqueness results by relaxing the Lipschitz condition on the generator $F$ or the conditions on the terminal condition $\xi$.
-BSDEs with generators exhibiting quadratic growth in the variable $z$ have attracted significant attention.
-These equations arise, for example, in stochastic linear-quadratic control with random coefficients bismut1976linearquad, in utility optimization problems in an incomplete market hu2005utility.
-Kobylanski:2000cy provided the seminal existence and uniqueness result for real-valued quadratic BSDEs with bounded terminal values, where $F$ has quadratic growth with respect to its $Z$ component.
-briand2013simple have continued to explore this area, demonstrating the existence of unique solutions to quadratic BSDEs with bounded terminal conditions using purely probabilistic methods.
-Investigations by briand2006bsde extend the previous results to unbounded terminal conditions.
-Some studies the multidimensional cases: see hu2016multi; tang2023MultiMFBSDE; luo2019triangle; tevzadze2008solvability.
+BSDEs with generators of quadratic growth in the variable $z$ have attracted significant attention.
+These equations arise, for example, in stochastic linear-quadratic control with random coefficients Bismut (1976), in utility optimization problems in an incomplete market Hu et al.
+(2005).
+Kobylanski (2000) provided the seminal existence and uniqueness result for real-valued quadratic BSDEs with bounded terminal values, where $F$ has quadratic growth with respect to its $Z$ component.
+Briand and Elie (2013) have continued to explore this area, demonstrating the existence of unique solutions to quadratic BSDEs with bounded terminal conditions using purely probabilistic methods.
+Investigations by Briand and Hu (2006) extend the previous results to unbounded terminal conditions.
+Some studies the multidimensional cases: see Hu and Tang (2016); Tang and Yang (2023); Luo (2020); Tevzadze (2008).
 Mean-field stochastic differential equations, also known as McKean-Vlasov equations, describe stochastic systems whose evolution is influenced by both the microscopic state and the macroscopic distribution of particles.
-These equations date back to kac1956foundations’s work on the Boltzmann and Vlasov equations in the 1950s.
-Mean-field game theory began with the groundbreaking contributions of lasry2007mfgame.
-Building on this foundation, Buckdahn2009MFlimit introduced mean-field BSDEs, derived from approximations involving weakly interacting $N$ particle systems, and investigated them using probabilistic methods.
+These equations date back to Kac (1956)’s work on the Boltzmann and Vlasov equations in the 1950s.
+Mean-field game theory began with the groundbreaking contributions of Lasry and Lions (2007).
+Building on this foundation, Buckdahn et al.
+(2009a) introduced mean-field BSDEs, derived from approximations involving weakly interacting $N$ particle systems, and investigated them using probabilistic methods.
 This has led to a growing interest in mean-field BSDEs and FBSDEs.
-For instance, Buckdahn2009MFBSDE studied the existence and uniqueness of solutions with Lipschitz continuous generators and square integrable terminal conditions.
-More recently, hao2022mean explored the solvability of solutions for one-dimensional mean-field BSDEs with quadratic growth, while tang2023MultiMFBSDE established the existence and uniqueness of solutions for multidimensional BSDEs with generators depending on the expectation of both variables.
-Li2018MFBSDEcontinuous examined mean-field BSDEs whose coefficients depend not only on the solution $(Y,Z)$ but also on the law of $Y$.
-pardoux1990adapted interpreted the key structural feature of BSDEs as a nonlinear conditional expectation, highlighting their close link to Malliavin calculus.
-In fact, el1997backward and Buckdahn2009MFBSDE state that, in a Clark–Ocone type formula, the martingale density component $Z$ of the solution pair of a BSDE with a classical globally Lipschitz generator on a Gaussian basis is the Malliavin derivative of the other component $Y$ (see Proposition 5.3).
-The result has been extending our understanding of quadratic BSDEs: for example, briand2013simple provided a straightforward approach to the existence and uniqueness of quadratic BSDEs with bounded terminal conditions using Malliavin calculus to overcome difficulties with quadratic generators, and Imkeller2010path proposed a numerical scheme for BSDEs with generators of quadratic growth.
-Continuous time portfolio optimization under the utility-maximization framework introduced by merton1969portfolio is a foundational topic in mathematical finance.
+For instance, Buckdahn et al.
+(2009b) studied the existence and uniqueness of solutions with Lipschitz continuous generators and square integrable terminal conditions.
+More recently, Hao et al.
+(2025) explored the solvability of solutions for one-dimensional mean-field BSDEs with quadratic growth.
+In contrast, Tang and Yang (2023) established the existence and uniqueness of solutions for multidimensional BSDEs with generators depending on the expectation of both variables.
+Li et al.
+(2018) examined mean-field BSDEs whose coefficients depend not only on the solution $(Y,Z)$ but also on the law of $Y$.
+Buckdahn et al.
+(2026) derived a global stochastic maximum principle for mean-field forward–backward stochastic control problems with quadratic BSDE generators via BMO-based analysis.
+El Karoui et al.
+(1997) and Buckdahn et al.
+(2009b) state that, in a Clark–Ocone type formula, the martingale density component $Z$ of the solution pair of a BSDE with a classical globally Lipschitz generator on a Gaussian basis is the Malliavin derivative of the other component $Y$ (see Proposition 5.3).
+The result has been extending our understanding of quadratic BSDEs: for example, Briand and Elie (2013) provided a straightforward approach to the existence and uniqueness of quadratic BSDEs with bounded terminal conditions using Malliavin calculus to overcome difficulties with quadratic generators, and Imkeller and Dos Reis (2010) proposed a numerical scheme for BSDEs with generators of quadratic growth.
+Continuous time portfolio optimization under the utility-maximization framework introduced by Merton (1969) is a foundational topic in mathematical finance.
 Merton analyzed the problem in a complete market where the risky asset follows the Black-Scholes dynamics and no constraints are imposed on admissible trading strategies.
 A major development in the early 2000s was the emergence of a connection between this class of optimization problems and the theory of BSDEs, which has since become a central approach in the field.
-The core idea of the BSDE methodology is to construct a stochastic process, dependent on the investment strategy, that coincides with the investor’s utility of terminal wealth.
-In this context, elrouge2000utility used BSDEs to compute the value function and optimal strategy for exponential utility, assuming trading strategies are restricted to a convex cone.
-Building on this framework, hu2005utility developed a BSDE approach to utility maximization in incomplete markets and extended the methodology to include logarithmic and power utility functions, allowing for general strategy constraints within a closed set.
-Our methods are inspired by the works of tevzadze2008solvability and briand2013simple.
-The former studied the existence and uniqueness of solutions for multidimensional quadratic BSDEs, utilizing BMO martingale theory.
+The core idea of the BSDE methodology is to construct a stochastic process, depending on the investment strategy, that matches the investor’s utility of terminal wealth.
+In this context, Rouge and El Karoui (2000) used BSDEs to compute the value function and optimal strategy for exponential utility, assuming trading strategies are restricted to a convex cone.
+Building on this framework, Hu et al.
+(2005) developed a BSDE approach to utility maximization in incomplete markets and extended the methodology to include logarithmic and power utility functions, allowing for general strategy constraints within a closed set.
+Our methods are inspired by the works of Tevzadze (2008) and Briand and Elie (2013).
+The former studied the existence and uniqueness of solutions to multidimensional quadratic BSDEs using BMO martingale theory.
 The latter provided a straightforward approach to establishing the existence and uniqueness of quadratic BSDEs with bounded terminal conditions using Malliavin calculus.
-Building on these foundations, we bridge the gap between the approaches of cheridito2017bse and hao2022mean.
-Specifically, hao2022mean established the existence and uniqueness by applying a contraction mapping with the Reverse Hölder Inequality (RHI) and iterating backward in time when the generator is quadratic in $Z$ component, while cheridito2017bse used the Krasnoselskii fixed-point theorem on a small time interval, iterating backward to obtain a global solution.
-Our contribution extends the results of cheridito2017bse by demonstrating in Proposition 4.9 that the function $f^{1}$ satisfies the conditions outlined in hao2022mean, thereby proving the existence and uniqueness of a solution for quadratic mean-field BSDEs, which is quadratic both in $Z$ and the corresponding mean-field component $\tilde{Z}$.
+Building on these foundations, we bridge the gap between the approaches of Cheridito and Nam (2017) and Hao et al.
+(2025).
+Specifically, Hao et al.
+(2025) established the existence and uniqueness by applying a contraction mapping with the Reverse Hölder Inequality (RHI) and iterating backward in time when the generator is quadratic in $Z$ component, while Cheridito and Nam (2017) used the Krasnoselskii fixed-point theorem on a small time interval, iterating backward to obtain a global solution.
+Our contribution extends the results of Cheridito and Nam (2017) by demonstrating in Proposition 4.9 that the function $f^{1}$ satisfies the conditions outlined in Hao et al.
+(2025), thereby proving the existence and uniqueness of a solution for quadratic mean-field BSDEs, which is quadratic both in $Z$ and the mean-field component $\tilde{Z}$.
 Our main contributions are as follows.
-First, we establish the existence and uniqueness of a solution when the generator exhibits separately quadratic growth in both $Z$ and the mean-field term $\tilde{Z}$ with bounded terminal condition $\xi$.
-Our approach builds upon the theory of BMO martingales and Malliavin calculus to effectively address the analytical challenges arising from the quadratic structure.
-A key novelty of this work lies in the inclusion of a quadratic mean-field component $\tilde{Z}$, which distinguishes it from existing results in the literature.
-Moreover, our method is notable for its simplicity, as it avoids the use of fixed-point arguments commonly employed in previous studies, such as juan2024MFBSDE and tevzadze2008solvability.
+First, we prove existence and uniqueness for quadratic mean-field BSDEs when the generator satisfies the unified quadratic condition, that is, quadratic both in $Z$ and its mean-field component $\tilde{Z}$.
+In this fully coupled regime, the generator may contain genuine interactions between the control variable $Z$ and the mean-field component $\tilde{Z}$.
+To the best of our knowledge, this is the first paper that studies such a case in a general framework.
+Under this condition, we prove well-posedness for bounded terminal conditions of the form $\xi=\phi(X_{t_{1}},X_{t_{2}},\cdots,X_{t_{n}})$, where $\phi$ is $\alpha$-Hölder continuous for $\alpha\in(0,1]$, $X$ is a solution of a forward stochastic differential equation, and $t_{i}\in[0,T]$.
+Second, under the separable quadratic condition (i.e.
+the growth of the generator with respect to $\tilde{Z}$ can be bounded by a quadratic function of $\tilde{Z}$), we obtain existence and uniqueness of the form $\xi=\phi(W_{[0,T]})$, where $\phi:C([0,T];\mathbb{R}^{d})\to\mathbb{R}$ is uniformly continuous in the sup norm.
+Finally, in the application part of the paper, we consider the class of generators of the same form as unified quadratic condition up to constants, and we also prove existence and uniqueness under a smallness condition on the terminal value $\xi$, without imposing additional structural regularity.
+Our approach builds on BMO martingale theory and Malliavin calculus to address analytical difficulties arising from the quadratic structure, and it avoids fixed-point arguments commonly used in previous studies, such as Jiang et al.
+(2024) and Tevzadze (2008).
+A key novelty throughout is the inclusion of a quadratic mean-field component $\tilde{Z}$, which substantially changes the structure compared to classical quadratic (MF)BSDEs.
 The rest of the paper is organized as follows.
-In Section 2, we present the mathematical framework and state the main result.
+In Section 2, we present the framework, introduce the standing assumptions (including $(\mathrm{H}_{q}^{\mathrm{unif}})$ and $(\mathrm{H}_{q}^{\mathrm{sep}})$), and state our main results.
 Section 3 derives an almost sure upper bound for the solution pair $(Y,Z)$ of Lipschitz mean-field BSDEs, assuming the terminal condition admits a uniformly bounded Malliavin derivative, as shown in Proposition 3.1.
-In the main Section 4, we provide a priori estimates and establish a powerful stability result for mean-field BSDEs with separately quadratic growth.
-In Section 5, we utilize the bounds obtained in Section 3, which are independent of the Lipschitz constants in $z$ and $\tilde{z}$, together with the stability result and a density argument, to prove the existence and uniqueness of solutions to our quadratic mean-field BSDEs.
-Finally, in Section 6, we extend the existence and uniqueness results for mean-field quadratic BSDE with separately quadratic driver to a fully coupled quadratic driver with small terminal conditions, and show how the corresponding BSDE characterizes a utility maximization problem with a particular class of constraint sets.
+Importantly, the resulting bound is independent of the Lipschitz constants in $z$ and $\tilde{z}$.
+In the main Section 4, we provide a priori estimates and establish a stability result tailored to mean-field BSDEs with quadratic growth, which will be used to pass from Lipschitz truncations to quadratic generators.
+In Section 5, we combine the uniform Lipschitz bounds obtained in Section 3 with the stability result and a density argument to prove existence and uniqueness under $(\mathrm{H}_{q}^{\mathrm{unif}})$ and $(\mathrm{H}_{q}^{\mathrm{sep}})$.
+Finally, in Section 6, we study the class of generators of the form $(\mathrm{H}_{q}^{\mathrm{LK}})$, which is equivalent to $(\mathrm{H}_{q}^{\mathrm{unif}})$ and show how the corresponding mean-field BSDE characterizes an exponential utility maximization problem under a suitable class of strategy constraints.
 2 Preliminaries and statement of main results
@@
     $\left\|Y\right\|_{\mathbb{S}^{\infty}}:=\sup_{0\leq r\leq T}\left\|Y_{r}\right\|_{L^{\infty}}=\sup_{\omega\in\Omega}\sup_{0\leq r\leq T}\left|Y_{r}(\omega)\right|<\infty,$
-  • $\mathbb{H}^{p}$: the set of predictable $\mathbb{R}^{d}$-valued process $Z$ satisfying $\left\|Z\right\|_{\mathbb{H}^{p}}:=\left(\mathbb{E}\left[\int_{0}^{T}\left|Z_{r}\right|^{2}\,dr\right]^{\frac{p}{2}}\right)^{\frac{1}{p}}<\infty,$
+  • $\mathbb{H}^{p}$: the set of predictable $\mathbb{R}^{d}$ valued process $Z$ satisfying $\left\|Z\right\|_{\mathbb{H}^{p}}:=\left(\mathbb{E}\left[\int_{0}^{T}\left|Z_{r}\right|^{2}\,dr\right]^{\frac{p}{2}}\right)^{\frac{1}{p}}<\infty,$
   • $\mathbb{H}^{2}_{BMO}$: a subset of $\mathbb{H}^{2}$ with $Z$ satisfying $Z\cdot W:=\int_{0}^{\cdot}Z_{s}\!\cdot dW_{s}$ is a Bounded Mean Oscillation (BMO) martingale, that is, there exists a non-negative constant $C$ such that, for each stopping time $\tau\leq T$,
-    $\mathbb{E}\left[\int_{\tau}^{T}\left|Z_{r}\right|^{2}\,dr\Big|\mathcal{F}_{\tau}\right]\leq C^{2}.$
+    $\mathbb{E}_{\tau}\left[\int_{\tau}^{T}\left|Z_{r}\right|^{2}\,dr\right]\leq C^{2},$
+    where $\mathbb{E}_{\tau}$ is the conditional expectation given $\mathcal{F}_{\tau}$.
 The $\left\|Z\right\|_{\mathbb{H}^{2}_{BMO}}:=\left\|\int_{0}^{\cdot}Z_{s}\cdot dW_{s}\right\|_{BMO}$ is defined as the smallest non-negative constant $C$ for which the above inequality is satisfied.
@@
 When a different underlying measure other than $\mathbb{P}$ is involved, we indicate it explicitly in the notation, for example by writing $\mathbb{E}$$^{\mathbb{Q}}$$\cdot$] and $\|\cdot\|_{BMO(\mathbb{Q})}$ for expectations and BMO norms taken with respect to a probability measure $\mathbb{Q}$.
-Here are some inequalities from BMO martingales theory, all the proofs can be found in kazamaki2006BMO.
+Here are some inequalities from BMO martingales theory, all the proofs can be found in Kazamaki (2006).
 Lemma 2.1 (Reverse Hölder Inequality (RHI)).
@@
 Lemma 2.2.
-For $K>0$.
-There exist constants $c_{1}(K),c_{2}(K)>0$, depending only on $K$, such that the following holds.
+For $K>0$, there exist constants $c_{1}(K),c_{2}(K)>0$, depending only on $K$, such that the following holds.
 Let $N$ be a one-dimensional continuous $\mathrm{BMO}$ martingale under $\mathbb{P}$ with $\|N\|_{\mathrm{BMO}(\mathbb{P})}\leq K$, and define the equivalent measure $\tilde{\mathbb{P}}$ by
 $\frac{d\tilde{\mathbb{P}}}{d\mathbb{P}}\;=\;\mathcal{E}(N)_{T}.$
-For any continuous $\mathrm{BMO}$ martingale $M$ under $\mathbb{P}$, set
-$\tilde{M}\;:=\;M-\langle M,N\rangle,$
 […]
@@
 Let $K_{0},K_{z},K_{\tilde{z}}>0$ be fixed constants.
-  1. $(\rm H_{0})$}
+  1. $(\mathrm{H}_{0})$
     The random variable $\xi$ is bounded almost surely by $K_{0}>0$, and likewise, the deterministic supremum $\sup_{t\in[0,T]}|g(t,0,0)|$ is also bounded by the same constant $K_{0}>0$.
-  2. $(\rm H_{q})$}
+  2. $(\mathrm{H}_{q}^{\mathrm{unif}})$
+    There exist constants $K_{z}>0$ and $K_{\tilde{z}}>0$ such that:
+[…]
+[…]
+    for any $(t,z,\tilde{z},z^{\prime},\tilde{z}^{\prime})\in[0,T]\times[\mathbb{R}^{d}\times\mathbb{R}^{d}]^{2}$.
+  3. $(\mathrm{H}_{q}^{\mathrm{sep}})$
     There exist constants $K_{z}>0$ and $K_{\tilde{z}}>0$ such that:
@@
     for any $(t,z,\tilde{z},z^{\prime},\tilde{z}^{\prime})\in[0,T]\times[\mathbb{R}^{d}\times\mathbb{R}^{d}]^{2}$.
+Remark 2.4.
+The unified quadratic condition $(\mathrm{H}_{q}^{\mathrm{unif}})$} is the fully coupled case, in contrast with the separately quadratic $(\mathrm{H}_{q}^{\mathrm{sep}})$}.
+It is immediate that $(\mathrm{H}_{q}^{\mathrm{unif}})$} implies $(\mathrm{H}_{q}^{\mathrm{sep}})$}, any statement proved under $(\mathrm{H}_{q}^{\mathrm{unif}})$} also holds under $(\mathrm{H}_{q}^{\mathrm{sep}})$}, while the converse is not true.
+We provide an explicit example here.
+For simplicity, consider the one-dimensional case and define
+$g(t,z,\tilde{z}):=z\,\tilde{z},\qquad(t,z,\tilde{z})\in[0,T]\times\mathbb{R}^{2}.$
+This $g$ satisfies $(\mathrm{H}_{q}^{\mathrm{unif}})$}, but does not satisfy $(\mathrm{H}_{q}^{\mathrm{sep}})$}.
 Example 2.5.
-The following $g$ satisfies the condition $(\mathrm{H}_{q})$:
+The following $g$ satisfies the condition $(\mathrm{H}_{q}^{\mathrm{sep}})$}:
 […]
-where $l$ is Lipschitz in (z,z~)($\tilde{z}$), and $q_{1}$ and $q_{2}$ are quadratic polynomials in $z$ and $\tilde{z}$, respectively.
-Note that this example does not satisfy the existence and uniqueness assumption in previous literature, such as hao2022mean.
+where $l$ is locally Lipschitz in (z,z~)($\tilde{z}$) with quadratic growth in $z$ and linear growth in $\tilde{z}$, and $q$ is quadratic polynomials in $\tilde{z}$.
+Note that this example does not satisfy the existence and uniqueness assumption in previous literature, such as Hao et al.
+(2025).
+Let us state our main theorems.
+The first main theorem assumes a Markovian structure with $(\mathrm{H}_{q}^{\mathrm{unif}})$} while the second one assumes non-Markovian structure with $(\mathrm{H}_{q}^{\mathrm{sep}})$}.
 Theorem 2.6.
-Let Assumptions $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q})$ hold. Then, there exists a unique solution $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ to the BSDE (2.6).
+Assume that there is a constant $C$ such that $b:[0,T]\times\mathbb{R}^{k}\to\mathbb{R}^{k}$ and $\sigma:[0,T]\to\mathbb{R}^{k\times n}$ are measurable functions with
+  • $|b(t,x)-b(t,x^{\prime})|\leq C|x-x^{\prime}|$, $|b(t,0)|\leq C$ for every $(t,x)\in[0,T]\times\mathbb{R}^{k}$, and
+  • $C^{-1}|v|^{2}\leq v^{\intercal}(\sigma\sigma^{\intercal})(t)v\leq C|v|^{2}$ for all $v\in\mathbb{R}^{k}$ and $t\in[0,T]$.
+Then, let $X$ be the strong solution of
+$dX_{t}=b(t,X_{t})dt+\sigma(t)dW_{t};\qquad X_{0}=x_{0}.$
+Let us assume $(\mathrm{H}_{0})$} and $(\mathrm{H}_{q}^{\mathrm{unif}})$}. In addition, assume that
+$\xi=\phi(X_{t_{1}},X_{t_{2}},\cdots,X_{t_{n}}),$
+for a bounded function $\phi:(\mathbb{R}^{d})^{n}\to\mathbb{R}$ which is $\alpha$-Hölder continuous, that is, there exists $K>0$ and $\alpha\in(0,1]$ such that
+$|\phi(x)-\phi(x^{\prime})|\leq K\max_{i=1,2,\cdots,n}|x_{i}-x^{\prime}_{i}|^{\alpha},$
+for any $x,x^{\prime}\in(\mathbb{R}^{d})^{n}$. Then, there exists a unique solution $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ to the BSDE (2.6).
 Remark 2.7.
-In the statement and proof of our main theorem we may, without loss of generality, we can assume that
+Note that our terminal condition is not necessarily Malliavin differentiable if $\alpha<1$.
+Theorem 2.8.
+Let Assumptions $(\mathrm{H}_{0})$} and $(\mathrm{H}_{q}^{\mathrm{sep}})$} hold.
+In addition, assume that there exists a function $\phi:C([0,T];\mathbb{R}^{d})\to\mathbb{R}$ such that $\xi=\phi(W_{[0,T]})$ and uniformly continuous in sup norm.
+Then, there exists a unique solution $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ to the BSDE (2.6).
+Remark 2.9.
+In the statement and proof of our main theorem, we may, without loss of generality, assume that
 $g(t,0,\tilde{z})=0,\qquad\forall\,(t,\tilde{z})\in[0,T]\times\mathbb{R}^{d}.$
-This is not an additional structural condition on the original generator, but merely a convenient normalisation.
-Given any generator $g$ satisfying Assumptions $(\rm H_{0})$$(\rm H_{q})$}, we may introduce the shifted generator
+Given any generator $g$ satisfying Assumptions $(\mathrm{H}_{0})$$(\mathrm{H}_{q}^{\mathrm{unif}})$}, we may introduce the shifted generator
 […]
@@
 together with the same $Z$, solves the BSDE with generator $\hat{g}$. Conversely, given a solution $(U,Z)$ to the BSDE with generator $\hat{g}$, the process
-$Y_{t}:=U_{t}+\int_{t}^{T}\widetilde{{\mathbb{E}}}\big[g(s,0,\tilde{Z}_{s})\big]\,ds$
+$Y_{t}=U_{t}+\int_{t}^{T}\widetilde{{\mathbb{E}}}\big[g(s,0,\tilde{Z}_{s})\big]\,ds$
 solves the original BSDE.
@@
 This section is devoted to deriving tractable upper bounds for the solution $(Y,Z)$ of the BSDE (2.6) under the assumption that the driver $g$ is Lipschitz continuous and the terminal condition $\xi$ is Malliavin differentiable.
-These bounds under the stronger condition become particularly useful in the proof of Lemma 5.1, we will show that the general case, where the driver exhibits quadratic growth and where $\xi$ is bounded, can be reduced to this setting through suitable approximation techniques.
+These bounds under the stronger condition become particularly useful in the proof of Lemma 5.1, we will show that the general case, where the driver of quadratic growth and where $\xi$ is bounded, can be reduced to this setting through suitable approximation techniques.
 Therefore, we initially work under stronger regularity assumptions to facilitate this reduction.
 For this purpose, we suppose all along this section that $g$ satisfies the following classical Lipschitz condition and condition on $\xi$ will be introduced in Propisiton 3.1 below.
-  • $({\rm H_{lip}})$
+  • $(\mathrm{H}_{\mathrm{lip}})$
     There exist a constant $L_{z}>0$ such that:
@@
     for any $(t,z,\tilde{z},z^{\prime},\tilde{z}^{\prime})\in[0,T]\times[\mathbb{R}^{d}\times\mathbb{R}^{d}]^{2}$.
-The existence and uniqueness of a solution $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}$ to the BSDE (2.6) under assumptions $(\rm H_{0})$$({\rm H_{lip}})$ is a well-established result, as shown in Theorem 3.1 of Buckdahn2009MFBSDE.
+The existence and uniqueness of a solution $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}$ to the BSDE (2.6) under assumptions $(\mathrm{H}_{0})$$(\mathrm{H}_{\mathrm{lip}})$ is a well-established result, as shown in Theorem 3.1 of Buckdahn et al.
+(2009b).
 In the following proposition, we demonstrate that the process $Y$, which lies in $\mathbb{S}^{\infty}$, is bounded by a constant that is independent of the Lipschitz constant $L_{z}$.
@@
 Proposition 3.1.
-If $(\rm H_{0})$} and $({\rm H_{lip}})$ hold, then the unique solution $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}$ of the Lipschitz mean-field BSDE satisfies the following properties.
+If $(\mathrm{H}_{0})$ and $(\mathrm{H}_{\mathrm{lip}})$ hold, then the unique solution $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}$ of the Lipschitz mean-field BSDE satisfies the following properties.
   1. 1.
@@
     If $\xi$ is Malliavin differentiable and satisfies $\left\|D\xi\right\|_{\mathbb{S}^{\infty}}=\sup_{0\leq t\leq T}\left\|D_{t}\xi\right\|_{L^{\infty}}<\infty$, then $Z\in\mathbb{S}^{\infty}$, and its norm is also bounded independently of $L_{z}$:
-    $\left\|Z\right\|_{\mathbb{S}^{\infty}}\leq\left\|D\xi\right\|_{\mathbb{S}^{\infty}},$ (3.8)
+    $\left\|Z\right\|_{\mathbb{S}^{\infty}}\leq\left\|D\xi\right\|_{\mathbb{S}^{\infty}}.$ (3.8)
 Proof.
-Unlike standard Lipschitz BSDE, the comparison theorem cannot be applied here (see Theorem 3.2 in Buckdahn2009MFBSDE).
+Unlike standard Lipschitz BSDE, the comparison theorem cannot be applied here (see Theorem 3.2 in Buckdahn et al.
+(2009b)).
 Observe that (3.7) can be derived following the line of proof of Proposition 4.3.
@@
 $D_{\theta}Y_{t}=D_{\theta}\xi+\int_{t}^{T}\widetilde{{\mathbb{E}}}\!\left[\partial_{z}g(s,Z_{s},\widetilde{Z}_{s})\right]\!\cdot D_{\theta}Z_{s}\,ds-\int_{t}^{T}D_{\theta}Z_{s}\!\cdot dW_{s}.$
-Under assumption $({\rm H_{lip}})$, we have $\left|\partial_{z}g\right|\leq L_{z}$.
-Therefore, the Novikov condition is satisfied and $\widetilde{{\mathbb{E}}}\!\left[\partial_{z}g(s,Z_{s},\widetilde{Z}_{s})\right]$ is measurable to $mathcal{F_{s}}$, and we may apply the Girsanov transform to obtain the following representation with new Brownian motion $W^{\mathbb{Q}}$ under the new probability measure $\mathbb{Q}$:
+Under assumption $(\mathrm{H}_{\mathrm{lip}})$, we have $\left|\partial_{z}g\right|\leq L_{z}$.
+Therefore, the Novikov condition is satisfied and $\widetilde{{\mathbb{E}}}\!\left[\partial_{z}g(s,Z_{s},\widetilde{Z}_{s})\right]\in\mathcal{F}_{s}$, and we may apply the Girsanov transform to obtain the following representation with new Brownian motion $W^{\mathbb{Q}}$ under the new measure $\mathbb{Q}$:
 $D_{\theta}Y_{t}=D_{\theta}\xi-\int_{t}^{T}D_{\theta}Z_{s}\!\cdot dW_{s}^{\mathbb{Q}},\qquad 0\leq\theta\leq t\leq T.$
@@
 $\left|D_{\theta}Y_{t}\right|\leq\left\|D\xi\right\|_{\mathbb{S}^{\infty}}.$
-Finally, as $(Z_{t})_{0\leq t\leq T}$ is a version of $(D_{\theta}Y_{t})_{0\leq t\leq T}$ and due to the arbitrariness of $t$, we establish 3.8. ∎
+Finally, as $(Z_{t})_{0\leq t\leq T}$ is a version of $(D_{\theta}Y_{t})_{0\leq t\leq T}$ and due to the arbitrariness of $t$, we establish (3.8). ∎
 4 A priori estimates and stability for BSDEs
-In this section, we will work under the assumptions $(\rm H_{0})$} and $(\rm H_{q})$} to establish the stability of the mean-field BSDE (2.6), as stated in Theorem 4.6.
-The strategy of the proof is as follows.
-We first derive a priori upper bounds on the $\mathbb{S}^{\infty}$ norm of $Y$ and the BMO norm of $\int_{0}^{\cdot}Z_{s}\!\cdot dW_{s}$ for any solution $(Y,Z)$ to the BSDE (2.6).
-These estimates, together with fundamental properties of BMO martingales, lead to the stability result.
-We begin with a useful estimate that will play an important role in the analysis.
+In this section, we establish stability estimates for the mean-field BSDE (2.6) under two different quadratic structures of the generator.
+More precisely, we prove Theorem 4.6 under the unified quadratic condition $(\mathrm{H}_{q}^{\mathrm{unif}})$ (together with the additional terminal condition assumption $(\mathrm{H}_{\xi})$), and Theorem 4.7 under the separable quadratic condition $(\mathrm{H}_{q}^{\mathrm{sep}})$.
+Both results yield the same stability estimate in the Banach space $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$.
+The overall strategy in both regimes is as follows.
+We first derive a priori estimates for any solution $(Y,Z)$ to (2.6): an $\mathbb{S}^{\infty}$ bound for $Y$ and a BMO bound for $\int_{0}^{\cdot}Z_{s}\cdot dW_{s}$.
+These estimates, together with fundamental properties of BMO martingales, lead to the stability results.
+We begin with a growth bound for the generator, which will be used repeatedly throughout the section.
 Remark 4.1.
-Under assumptions $(\rm H_{0})$} and $(\rm H_{q})$}, we have the following estimate:
+Under assumptions $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q}^{\mathrm{unif}})$}, we have the following estimate:
 […]
-[…]
-for any $(t,z,\tilde{z})\in[0,T]\times\mathbb{R}^{d}\times\mathbb{R}^{d}$. Taking expectations with respect to $\tilde{\mathbb{P}}$ yields:
-$\tilde{\mathbb{E}}\left|g(t,Z_{t},\tilde{Z}_{t})\right|\leq K_{0}+\frac{1}{2}K_{z}+\frac{1}{2}K_{\tilde{z}}+\frac{3K_{z}}{2}\left|Z_{t}\right|^{2}+\frac{3K_{\tilde{z}}}{2}\tilde{\mathbb{E}}\left|\tilde{Z}_{t}\right|^{2}.$ (4.9)
+for any $(t,z,\tilde{z})\in[0,T]\times\mathbb{R}^{d}\times\mathbb{R}^{d}$, where $K_{0}:=\sup_{t\in[0,T]}|g(t,0,0)|$. Taking expectations with respect to $\tilde{\mathbb{P}}$ yields:
+$\tilde{\mathbb{E}}\,|g(t,Z_{t},\tilde{Z}_{t})|\leq K_{0}+\frac{1}{2}K_{z}+\frac{1}{2}K_{\tilde{z}}+\Big(2K_{z}+\frac{1}{2}K_{\tilde{z}}\Big)|Z_{t}|^{2}+\Big(2K_{\tilde{z}}+\frac{1}{2}K_{z}\Big)\tilde{\mathbb{E}}|\tilde{Z}_{t}|^{2},$ (4.9)
+where we used symmetry $\tilde{\mathbb{E}}|\tilde{Z}_{t}|={\mathbb{E}}|Z_{t}|$, together with $|Z_{t}|\,{\mathbb{E}}|Z_{t}|\leq\frac{1}{2}|Z_{t}|^{2}+\frac{1}{2}{\mathbb{E}}|Z_{t}|^{2}$.
 We begin by noting that if the solution $Y\in\mathbb{S}^{\infty}$, then for any solution pair $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}$ to the quadratic mean-field BSDE, the stochastic integral $\int_{0}^{\cdot}Z_{s}\!\cdot dW_{s}$ defines a BMO martingale.
@@
 Proposition 4.2.
-Suppose that (H0) and (Hq) hold.
+Suppose that $(\mathrm{H}_{0})$} and $(\mathrm{H}_{q}^{\mathrm{unif}})$} hold.
 Let $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}$ be a solution to the mean-field BSDE (2.6).
-Then the stochastic process $\left(\int_{0}^{t}Z_{s}\!\cdot dW_{s}\right)_{0\leq t\leq T}$ defines a BMO martingale.
-Moreover, its BMO norm satisfies the following bound:
-$\left\|\int_{0}^{\cdot}Z_{s}\!\cdot dW_{s}\right\|_{\mathrm{BMO}}^{2}\leq(1\vee T)\big(4K_{0}+K_{z}+K_{\tilde{z}}\big)\,\frac{e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}}{3K_{z}}\left(\frac{K_{\tilde{z}}}{K_{z}}e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}+1\right).$ (4.10)
+Then the stochastic process $\big(\int_{0}^{t}Z_{s}\cdot dW_{s}\big)_{0\leq t\leq T}$ defines a BMO martingale.
+Moreover, its BMO norm satisfies the bound
+$\left\|\int_{0}^{\cdot}Z_{s}\cdot dW_{s}\right\|_{\mathrm{BMO}}^{2}\leq(1\vee T)\,(4K_{0}+K_{z}+K_{\tilde{z}})\,\frac{e^{(4K_{z}+K_{\tilde{z}})(1+\|Y\|_{\mathbb{S}^{\infty}})}}{4K_{z}+K_{\tilde{z}}}\left(\frac{K_{z}+4K_{\tilde{z}}}{4K_{z}+K_{\tilde{z}}}\,e^{(4K_{z}+K_{\tilde{z}})\|Y\|_{\mathbb{S}^{\infty}}}+1\right).$ (4.10)
 Proof.
-The argument follows from arguments in Proposition 2.1 of briand2008stobsdes with modified $C^{2}$ function
-$\varphi(x):=\frac{e^{3K_{z}|x|}-1-6cK_{z}|x|}{(3K_{z})^{2}},\qquad c\in\mathbb{R},x\in\mathbb{R}.$
-In Step 1 we express the $\mathrm{BMO}$ norm of $Z$ in terms of its $\mathbb{H}^{2}$ norm; in Step 2 we derive an explicit $\mathbb{H}^{2}$ bound in terms of the model coefficients and $\|Y\|_{\mathbb{S}^{\infty}}$.
-Step 1 (Estimate of $\|Z\|_{\mathbb{H}^{2}_{\mathrm{BMO}}}$ in terms of $\|Z\|_{\mathbb{H}^{2}}$). Fix $c=\tfrac{1}{2}$ then $\varphi(0)=0$ and, for all $x\in\mathbb{R}$,
-$\varphi(x)\geq 0,\qquad\varphi^{\prime}(x)=\frac{e^{3K_{z}|x|}-1}{3K_{z}}\operatorname{sgn}(x),\qquad\varphi^{\prime\prime}(x)=e^{3K_{z}|x|},\qquad\varphi^{\prime\prime}(x)-3K_{z}\,|\varphi^{\prime}(x)|\equiv 1.$
+The argument follows Proposition 2.1 of Briand and Confortola (2008), with constants adapted to the growth estimate (4.9). Define
+[…]
+Recall that, under $(\mathrm{H}_{0})$} and $(\mathrm{H}_{q}^{\mathrm{unif}})$}, the estimate (4.9) reads
+$\widetilde{\mathbb{E}}\,|g(t,Z_{t},\tilde{Z}_{t})|\leq K+A|Z_{t}|^{2}+B\,\widetilde{\mathbb{E}}|\tilde{Z}_{t}|^{2},\qquad t\in[0,T].$
+Define the $C^{2}$ function with parameter $c\in\mathbb{R}$
+$\varphi(x):=\frac{e^{\lambda|x|}-1-2c\lambda|x|}{\lambda^{2}},\qquad x\in\mathbb{R},$
+where $\lambda:=2A=4K_{z}+K_{\tilde{z}}$.
+In Step 1 we express the BMO norm of $Z$ in terms of its $\mathbb{H}^{2}$ norm; in Step 2 we derive an explicit $\mathbb{H}^{2}$ bound in terms of the model coefficients and $\|Y\|_{\mathbb{S}^{\infty}}$.
+Step 1 (Estimate of $\|Z\|_{\mathbb{H}^{2}_{\mathrm{BMO}}}$ in terms of $\|Z\|_{\mathbb{H}^{2}}$). Fix $c=\tfrac{1}{2}$. Then $\varphi(0)=0$ and, for all $x\in\mathbb{R}$,
+$\varphi(x)\geq 0,\qquad\varphi^{\prime}(x)=\frac{e^{\lambda|x|}-1}{\lambda}\,\operatorname{sgn}(x),\qquad\varphi^{\prime\prime}(x)=e^{\lambda|x|},\qquad\varphi^{\prime\prime}(x)-\lambda\,|\varphi^{\prime}(x)|\equiv 1.$
 Moreover, for $|x|\leq\|Y\|_{\mathbb{S}^{\infty}}$ we have the deterministic bound
-$0\leq|\varphi^{\prime}(x)|\leq C_{Y}:=\frac{1}{3K_{z}}\,e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}.$ (4.11)
-Let $\tau$ be any stopping time and apply Itô’s formula to $\varphi(Y)$ and using the growth estimate 4.9 for $g$ together with the symmetry $\widetilde{\mathbb{E}}|\tilde{Z}_{s}|^{2}=\mathbb{E}|Z_{s}|^{2}$ yields
+$0\leq|\varphi^{\prime}(x)|\leq C_{Y}:=\frac{1}{\lambda}\,e^{\lambda\|Y\|_{\mathbb{S}^{\infty}}}.$ (4.11)
+Let $\tau$ be any stopping time. Applying Itô’s formula to $\varphi(Y)$ on $\tau,T]$ and using (4.9) together with $\widetilde{\mathbb{E}}|\tilde{Z}_{s}|^{2}=\mathbb{E}|Z_{s}|^{2}$ yields
 $\varphi(Y_{\tau})$ $=\varphi(\xi)+\int_{\tau}^{T}\Big(-\varphi^{\prime}(Y_{s})\,\widetilde{\mathbb{E}}[g(s,Z_{s},\tilde{Z}_{s})]-\tfrac{1}{2}\,\varphi^{\prime\prime}(Y_{s})|Z_{s}|^{2}\Big)\,ds-\int_{\tau}^{T}\varphi^{\prime}(Y_{s})Z_{s}\cdot dW_{s}$
-$\leq\varphi(\xi)+\!\int_{\tau}^{T}\!\Big(\frac{3K_{z}}{2}|\varphi^{\prime}(Y_{s})|-\tfrac{1}{2}\varphi^{\prime\prime}(Y_{s})\Big)|Z_{s}|^{2}ds+\!\int_{\tau}^{T}\!|\varphi^{\prime}(Y_{s})|\Big(K+\tfrac{3}{2}K_{\tilde{z}}\,\mathbb{E}|Z_{s}|^{2}\Big)ds$
-$\leq\varphi(\xi)-\tfrac{1}{2}\!\int_{\tau}^{T}\!|Z_{s}|^{2}ds+\!\int_{\tau}^{T}\!|\varphi^{\prime}(Y_{s})|\Big(K+\tfrac{3}{2}K_{\tilde{z}}\,\mathbb{E}|Z_{s}|^{2}\Big)ds-\!\int_{\tau}^{T}\!\varphi^{\prime}(Y_{s})Z_{s}\cdot dW_{s},$
-where $K:=K_{0}+\frac{1}{2}K_{z}+\frac{1}{2}K_{\tilde{z}}$.
-Taking $\mathbb{E}[\cdot\mid\mathcal{F}_{\tau}]$ and using (4.11) together with $|\varphi(x)|\leq C_{Y}|x|$ (since $\varphi(0)=0$ and $\varphi^{\prime}$ is bounded on the range of $Y$), we get
-$\tfrac{1}{2}\,\mathbb{E}\!\Big[\!\int_{\tau}^{T}\!|Z_{s}|^{2}ds\,\Big|\,\mathcal{F}_{\tau}\Big]$ $\leq\mathbb{E}\!\big[\varphi(\xi)-\varphi(Y_{\tau})\,\big|\,\mathcal{F}_{\tau}\big]+C_{Y}K(T-\tau)+\tfrac{3}{2}K_{\tilde{z}}C_{Y}\!\!\int_{0}^{T}\!\mathbb{E}|Z_{s}|^{2}ds$
-$\leq C_{Y}\,\mathbb{E}\!\big[|\xi|\,\big|\,\mathcal{F}_{\tau}\big]+C_{Y}KT+\tfrac{3}{2}K_{\tilde{z}}C_{Y}\!\!\int_{0}^{T}\!\mathbb{E}|Z_{s}|^{2}ds.$
-we thus obtain
-$\tfrac{1}{2}\,{\mathbb{E}}\!\Big[\!\int_{\tau}^{T}\!|Z_{s}|^{2}ds\,\Big|\,\mathcal{F}_{\tau}\Big]\leq C_{Y}\left(\|\xi\|_{L^{\infty}}+K\,T+\tfrac{3}{2}K_{\tilde{z}}\!\!\int_{0}^{T}\!\!{\mathbb{E}}|Z_{s}|^{2}ds\right).$ (4.12)
+$\leq\varphi(\xi)+\int_{\tau}^{T}\Big(A|\varphi^{\prime}(Y_{s})|-\tfrac{1}{2}\varphi^{\prime\prime}(Y_{s})\Big)|Z_{s}|^{2}\,ds+\int_{\tau}^{T}|\varphi^{\prime}(Y_{s})|\Big(K+B\,\mathbb{E}|Z_{s}|^{2}\Big)\,ds$
+$\hskip 56.9055pt-\int_{\tau}^{T}\varphi^{\prime}(Y_{s})Z_{s}\cdot dW_{s},$
+Since $\lambda=2A$ and $\varphi^{\prime\prime}-\lambda|\varphi^{\prime}|\equiv 1$, we have
+$A|\varphi^{\prime}(Y_{s})|-\tfrac{1}{2}\varphi^{\prime\prime}(Y_{s})=-\tfrac{1}{2}\big(\varphi^{\prime\prime}(Y_{s})-2A|\varphi^{\prime}(Y_{s})|\big)=-\tfrac{1}{2}.$
+Taking $\mathbb{E}[\cdot\,|\,\mathcal{F}_{\tau}]$, using (4.11), and also $|\varphi(x)|\leq C_{Y}|x|$ on the range of $Y$ (since $\varphi(0)=0$ and $\varphi^{\prime}$ is bounded there), we obtain
+$\tfrac{1}{2}\,\mathbb{E}\!\Big[\int_{\tau}^{T}|Z_{s}|^{2}\,ds\ \Big|\ \mathcal{F}_{\tau}\Big]$ $\leq\mathbb{E}\!\big[\varphi(\xi)-\varphi(Y_{\tau})\,\big|\,\mathcal{F}_{\tau}\big]+C_{Y}K(T-\tau)+BC_{Y}\int_{0}^{T}\mathbb{E}|Z_{s}|^{2}\,ds$
+$\leq C_{Y}\,\|\xi\|_{L^{\infty}}+C_{Y}KT+BC_{Y}\int_{0}^{T}\mathbb{E}|Z_{s}|^{2}\,ds.$
+Hence,
+$\tfrac{1}{2}\,\mathbb{E}\!\Big[\int_{\tau}^{T}|Z_{s}|^{2}\,ds\ \Big|\ \mathcal{F}_{\tau}\Big]\leq C_{Y}\left(\|\xi\|_{L^{\infty}}+K\,T+B\int_{0}^{T}\mathbb{E}|Z_{s}|^{2}\,ds\right).$ (4.12)
 Since $\tau$ is arbitrary, (4.12) shows that $\big\|\int_{0}^{\cdot}Z_{s}\cdot dW_{s}\big\|_{\mathrm{BMO}}^{2}$ is controlled by $\|Z\|_{\mathbb{H}^{2}}^{2}$ and model parameters.
-Step 2 (An explicit $\mathbb{H}^{2}$ bound). Write $\alpha:=K_{\tilde{z}}/K_{z}>0$. We may, without loss of generality, assume $\alpha$ satisfies
-$0<\alpha\;<\;\min\!\Big\{\tfrac{1}{2},\ \frac{e^{3K_{z}}-2}{e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}}\Big\}.$
-by enlarging $K_{z}$ or $K_{\tilde{z}}$. Define positive constants
-$c_{1}:=\frac{K_{\tilde{z}}}{2K_{z}}e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}+\frac{1}{2},\qquad c_{2}:=\tfrac{1}{2}e^{3K_{z}}.$
-Note that, by definition of $\alpha$, $c_{1}$ can also be expressed as $c_{1}=\tfrac{1}{2}\big(\alpha e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}+1\big)$.
-By the choice of $\alpha$, it is easy to verify $c_{1}<c_{2}$.
-Pick any $c\in(c_{1},c_{2})$ and consider the same $\varphi$ at in Step 1 with a different $c$ value, so that $\varphi(0)=0$ and, for all $x$,
-$|\varphi^{\prime}(x)|=\frac{e^{3K_{z}|x|}-2c}{3K_{z}}>\frac{e^{3K_{z}|x|}-2c_{2}}{3K_{z}}\geq 0.$
-We also have identity $\varphi^{\prime\prime}(\cdot)-3K_{z}\left|\varphi^{\prime}(\cdot)\right|=2c$. Applying Itô to $\varphi(Y)$ on $[0,T]$ and arguing as in Step 1 with $\tau=0$ gives
-$\varphi(Y_{0})$ $\leq\varphi(\xi)-\int_{0}^{T}c\,|Z_{s}|^{2}ds+\frac{3K_{\tilde{z}}}{2}\int_{0}^{T}|\varphi^{\prime}(Y_{s})|\,\mathbb{E}|Z_{s}|^{2}ds+K\int_{0}^{T}|\varphi^{\prime}(Y_{s})|ds-\int_{0}^{T}\varphi^{\prime}(Y_{s})Z_{s}\cdot dW_{s}.$
-Taking expectations and using $|\varphi^{\prime}(x)|=\frac{e^{3K_{z}x}-2c}{3K_{z}}\leq\frac{e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}}{3K_{z}}=C_{Y}$ for $|x|\leq\|Y\|_{\mathbb{S}^{\infty}}$,
-$\varphi(Y_{0})\leq\mathbb{E}\!\left[\varphi(\xi)\right]+KC_{Y}T-c\,\mathbb{E}\!\left[\int_{0}^{T}|Z_{s}|^{2}ds\right]+\frac{3K_{\tilde{z}}}{2}\,C_{Y}\,\mathbb{E}\!\left[\int_{0}^{T}|Z_{s}|^{2}ds\right].$
-By the definition of $c_{1}$ and the fact that $c>c_{1}=\frac{3K_{\tilde{z}}}{2}C_{Y}+\tfrac{1}{2}$ we have $\frac{3K_{\tilde{z}}}{2}C_{Y}<c-\tfrac{1}{2}$. Hence
-$\frac{1}{2}\,\mathbb{E}\!\left[\int_{0}^{T}|Z_{s}|^{2}ds\right]\ \leq\ \mathbb{E}\!\left[\varphi(\xi)\right]+KC_{Y}T\ \leq\ \frac{e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}}{3K_{z}}\,\big(\left\|\xi\right\|_{L^{\infty}}+KT\big).$ (4.13)
-Finally, plugging (4.13) into (4.12) and taking the essential supremum over $\tau$ yields an explicit $\mathrm{BMO}$ bound for $\int_{0}^{\cdot}Z_{s}\cdot dW_{s}$ in terms of $\|Y\|_{\mathbb{S}^{\infty}}$ and the model parameters.
-This completes the proof.
-∎
-Next, suppose the solution $Z\in\mathbb{H}^{2}_{\mathrm{BMO}}$.
-Then, for any solution pair $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ to the quadratic mean-field BSDE, we derive the following proposition showing that $Y\in\mathbb{S}^{\infty}$ which is independent of the BMO norm $\left\|\int_{0}^{\cdot}Z_{s}\!\cdot dW_{s}\right\|_{\mathrm{BMO}}$.
-We introduce a simple but useful lemma which will be used later on when we work under equivalent probability measures.
-Lemma 4.3.
-Let $\mathbb{P}$ and $\mathbb{Q}$ be equivalent probability measure on $(\Omega,\mathcal{F})$. For any random variable $X$,
-$\|X\|_{L^{\infty}(\mathbb{P})}=\|X\|_{L^{\infty}(\mathbb{Q})}.$
-In particular, $X$ is bounded $\mathbb{P}$-a.s. iff it is bounded $\mathbb{Q}$-a.s., with the same bound.
-Proof.
-Set $K:=\|X\|_{L^{\infty}(\mathbb{P})}$.
-For every $\varepsilon>0$, $\mathbb{P}(|X|>K+\varepsilon)=0$.
-Since $\mathbb{Q}\ll\mathbb{P}$, also $\mathbb{Q}(|X|>K+\varepsilon)=0$, so $\|X\|_{L^{\infty}(\mathbb{Q})}\leq K$.
-By symmetry ($\mathbb{P}\ll\mathbb{Q}$), the reverse inequality holds.
-∎
+Step 2 (An explicit $\mathbb{H}^{2}$ bound). We assume (without loss of generality, by enlarging $K_{z}$ relative to $K_{\tilde{z}}$) that
+$0<\alpha<\min\Big\{\tfrac{1}{2},\ \frac{e^{\lambda}-2}{e^{\lambda\|Y\|_{\mathbb{S}^{\infty}}}}\Big\}.$ (4.13)
+Define
+$\widehat{C}_{Y}:=\frac{e^{\lambda(1+\|Y\|_{\mathbb{S}^{\infty}})}}{\lambda},\qquad c_{1}:=B\,\widehat{C}_{Y}+\frac{1}{2},\qquad c_{2}:=\frac{1}{2}e^{\lambda}.$
+Then (4.13) implies $c_{1}<c_{2}$. Fix any $c\in(c_{1},c_{2})$ and consider the shifted $C^{2}$ function
+$\varphi(x):=\frac{e^{\lambda(1+|x|)}-1-2c\lambda(1+|x|)}{\lambda^{2}},\qquad x\in\mathbb{R}.$ (4.14)
+For $x\neq 0$ we have
+$\varphi^{\prime}(x)=\frac{e^{\lambda(1+|x|)}-2c}{\lambda}\,\operatorname{sgn}(x),\qquad\varphi^{\prime\prime}(x)=e^{\lambda(1+|x|)},\qquad\text{and}\qquad\varphi^{\prime\prime}-\lambda|\varphi^{\prime}|\equiv 2c,$
+and the same identities hold in the generalized Itô–Tanaka sense at $x=0$. Moreover, since $1+|x|\geq 1$, we have for all $x\in\mathbb{R}$,
+$|\varphi^{\prime}(x)|=\frac{e^{\lambda(1+|x|)}-2c}{\lambda}>\frac{e^{\lambda(1+|x|)}-2c_{2}}{\lambda}\geq 0,$
+and for $|x|\leq\|Y\|_{\mathbb{S}^{\infty}}$,
+$|\varphi^{\prime}(x)|\leq\frac{e^{\lambda(1+\|Y\|_{\mathbb{S}^{\infty}})}}{\lambda}=\widehat{C}_{Y}.$
+Applying Itô’s formula to $\varphi(Y)$ on $[0,T]$ (as in Step 1, with $\tau=0$) and using (4.9) yields
+$\varphi(Y_{0})$ $\leq\varphi(\xi)-\int_{0}^{T}c\,|Z_{s}|^{2}\,ds+B\int_{0}^{T}|\varphi^{\prime}(Y_{s})|\,{\mathbb{E}}|Z_{s}|^{2}\,ds+K\int_{0}^{T}|\varphi^{\prime}(Y_{s})|\,ds-\int_{0}^{T}\varphi^{\prime}(Y_{s})Z_{s}\cdot dW_{s}.$
+Taking expectations and using $|\varphi^{\prime}(Y_{s})|\leq\widehat{C}_{Y}$ gives
+$\varphi(Y_{0})\leq{\mathbb{E}}[\varphi(\xi)]+K\widehat{C}_{Y}T-c\,{\mathbb{E}}\!\left[\int_{0}^{T}|Z_{s}|^{2}\,ds\right]+B\widehat{C}_{Y}\,{\mathbb{E}}\!\left[\int_{0}^{T}|Z_{s}|^{2}\,ds\right].$
+Since $c>c_{1}=B\widehat{C}_{Y}+\tfrac{1}{2}$, we can absorb the last term and obtain
+$\frac{1}{2}\,{\mathbb{E}}\!\left[\int_{0}^{T}|Z_{s}|^{2}\,ds\right]\leq{\mathbb{E}}[\varphi(\xi)]+K\widehat{C}_{Y}T.$
+Finally, since $\varphi(0)=0$ and $|\varphi^{\prime}(x)|\leq\widehat{C}_{Y}$ for $|x|\leq\|Y\|_{\mathbb{S}^{\infty}}$, we have $|\varphi(x)|\leq\widehat{C}_{Y}|x|$, hence
+$\frac{1}{2}\,{\mathbb{E}}\!\left[\int_{0}^{T}|Z_{s}|^{2}\,ds\right]\leq\widehat{C}_{Y}\big(\|\xi\|_{L^{\infty}}+KT\big).$ (4.15)
+Finally, plugging (4.15) into (4.12) and taking the essential supremum over $\tau$ completes the proof. ∎
+Next, suppose that the solution satisfies $Z\in\mathbb{H}^{2}_{\mathrm{BMO}}$.
+Then, for any solution pair $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ to the quadratic mean-field BSDE, we derive the following proposition showing that $Y\in\mathbb{S}^{\infty}$, with a bound that is independent of the BMO norm $\left\|\int_{0}^{\cdot}Z_{s}\!\cdot dW_{s}\right\|_{\mathrm{BMO}}$.
 Proposition 4.3.
-Suppose Assumptions $(\rm H_{0})$} and $(\rm H_{q})$} hold.
-Then for any solution $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ of BSDE (2.6), we have $Y\in\mathbb{S}^{\infty}$ and the following estimate holds:
+Suppose Assumptions $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q}^{\mathrm{unif}})$} hold.
+Then for any solution $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ of BSDE (2.6), we have $Y\in\mathbb{S}^{\infty}$ and
 $\left\|Y\right\|_{\mathbb{S}^{\infty}}\leq\left\|\xi\right\|_{L^{\infty}}.$ (4.16)
 Proof.
-The proof relies on a linearization argument combined with the BMO property. Using Remark 2.6, without the loss of generality, we assume
+The proof relies on a linearization argument combined with the BMO property. Using Remark 2.9, without loss of generality we assume
 $g(t,0,\tilde{z})=0,\qquad\forall(t,\tilde{z})\in[0,T]\times\mathbb{R}^{d}.$
-Then, the BSDE (2.6) can be rewritten as
+Then BSDE (2.6) can be rewritten as
 $Y_{t}=\xi+\int_{t}^{T}\Big(\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big)\cdot Z_{s}\,ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},\qquad t\in[0,T],$ (4.17)
 where the process $(L_{s})_{0\leq s\leq T}$ is defined by
-$L_{s}:=\frac{g(s,Z_{s},\tilde{Z}_{s})-g(s,0,\tilde{Z}_{s})}{|\Delta Z_{s}|^{2}}\,\Delta Z_{s}\,\mathbf{1}_{\{\Delta Z_{s}\neq 0\}}\qquad 0\leq s\leq T,$
-Since $g(s,0,\tilde{z})=0$ by construction, the numerator is simply $g(s,Z_{s},\tilde{Z}_{s})$. Moreover, Assumption $(\rm H_{q})$} yields
-$\Big|\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big|\leq\int_{\tilde{\Omega}}\left|L_{s}\right|\,d\tilde{\mathbb{P}}\leq K_{z}\big(1+\left|Z_{s}\right|\big),\qquad 0\leq s\leq T.$
-Since $Z\in\mathbb{H}^{2}_{\mathrm{BMO}}$ by assumption, $\big(\int_{0}^{t}\Big(\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big)\cdot dW_{s}\big)_{t\in[0,T]}$ is a BMO martingale, and its stochastic exponential $\mathcal{E}_{t}$ satisfies Kazamaki’s condition.
-Define the probability measure $\mathbb{P}^{L}$ by $d\mathbb{P}^{L}:=\mathcal{E}_{T}\,d\mathbb{P}$.
-By the above BMO argument, $(\mathcal{E}_{t})$ is a uniformly integrable martingale, so $\mathbb{P}^{L}$ is equivalent to $\mathbb{P}$.
-Girsanov’s theorem then implies that the process
+$L_{s}:=\frac{g(s,Z_{s},\tilde{Z}_{s})-g(s,0,\tilde{Z}_{s})}{|Z_{s}|^{2}}\,Z_{s}\,\mathbf{1}_{\{Z_{s}\neq 0\}},\qquad 0\leq s\leq T.$
+[…]
+$|g(s,z,\tilde{z})-g(s,0,\tilde{z})|\leq K_{z}\big(1+|z|+|\tilde{z}|\big)\,|z|.$
+Therefore,
+$\Big|\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big|\leq K_{z}\big(1+|Z_{s}|+{\mathbb{E}}|Z_{s}|\big),\qquad 0\leq s\leq T.$ (4.18)
+Define the process
+$M_{t}:=\int_{0}^{t}\Big(\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big)\cdot dW_{s},\qquad t\in[0,T].$
+We claim that $M$ is a BMO martingale. Indeed, by (4.18) and $(a+b+c)^{2}\leq 3(a^{2}+b^{2}+c^{2})$,
+$\Big|\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big|^{2}\leq 3K_{z}^{2}\Big(1+|Z_{s}|^{2}+{\mathbb{E}}|Z_{s}|^{2}\Big).$
+Hence, for any stopping time $\tau$,
+${\mathbb{E}}\!\Big[\langle M\rangle_{T}-\langle M\rangle_{\tau}\,\big|\,\mathcal{F}_{\tau}\Big]$ $={\mathbb{E}}\!\Big[\int_{\tau}^{T}\Big|\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big|^{2}ds\,\Big|\,\mathcal{F}_{\tau}\Big]$
+$\leq 3K_{z}^{2}\Big(T+{\mathbb{E}}\!\Big[\int_{\tau}^{T}|Z_{s}|^{2}ds\,\Big|\,\mathcal{F}_{\tau}\Big]+\int_{0}^{T}{\mathbb{E}}|Z_{s}|^{2}ds\Big).$
+Since $Z\in\mathbb{H}^{2}_{\mathrm{BMO}}$, and $\mathbb{H}^{2}\subset\mathbb{H}^{2}_{\mathrm{BMO}}$.
+Therefore $\|M\|_{\mathrm{BMO}}<\infty$, proving the claim.
+By Kazamaki’s condition, the stochastic exponential $\mathcal{E}(M)$ is a uniformly integrable martingale.
+Define the probability measure $\mathbb{P}^{L}$ by $d\mathbb{P}^{L}:=\mathcal{E}(M)_{T}\,d\mathbb{P}$; then $\mathbb{P}^{L}$ is equivalent to $\mathbb{P}$.
+By Girsanov’s theorem, the process
 $W_{t}^{L}:=W_{t}+\int_{0}^{t}\Big(\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big)\,ds,\qquad t\in[0,T],$
-is a dd–dimensional Brownian motion under $\mathbb{P}^{L}$, thus we obtain
-$Y_{t}=\xi-\int_{t}^{T}Z_{s}\!\cdot dW_{s}^{L}.$
-Hence,
-$Y_{t}={\mathbb{E}}^{\mathbb{P}^{L}}\big[\xi\,\big|\,\mathcal{F}_{t}\big],\qquad 0\leq t\leq T.$
-Since $\mathbb{P}^{L}$ is equivalent to $\mathbb{P}$, the same bound holds $\mathbb{P}$–almost surely by Lemma 4.3, and taking the essential supremum over $t\in[0,T]$ completes the proof. ∎
-Then we have a direct result from the previous two propositions.
+is a dd–dimensional Brownian motion under $\mathbb{P}^{L}$, and (4.17) becomes
+$Y_{t}=\xi-\int_{t}^{T}Z_{s}\cdot dW_{s}^{L},\qquad 0\leq t\leq T.$
+Thus
+$Y_{t}={\mathbb{E}}^{\mathbb{P}^{L}}\big[\xi\,\big|\,\mathcal{F}_{t}\big],\qquad 0\leq t\leq T,$
+and consequently $|Y_{t}|\leq\|\xi\|_{L^{\infty}}$ for all $t\in[0,T]$, $\mathbb{P}^{L}$–a.s.
+Since $\mathbb{P}^{L}\sim\mathbb{P}$, the same bound holds $\mathbb{P}$–a.s., and taking the essential supremum over $t\in[0,T]$ completes the proof.
+∎
+We now state a direct consequence of the previous two propositions.
 Corollary 4.4.
-Let $(\rm H_{0})$} and $(\rm H_{q})$} hold. Then, for any solution pair $(Y,Z)$ to the quadratic BSDE (2.6), if either
+Let $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q}^{\mathrm{unif}})$} hold. Then, for any solution pair $(Y,Z)$ to the quadratic BSDE (2.6), if either
 $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}\quad\text{or}\quad(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}_{\mathrm{BMO}},$
@@
 Proof.
-We only sketch how the two estimates interact, as the details follow directly from Propositions 4.2 and 4.4. If $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}$, then Proposition 4.2 yields
+We only sketch how the two estimates interact here. If $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}$, then Proposition 4.2 yields
 $Z\in\mathbb{H}^{2}_{\mathrm{BMO}}\quad\text{and}\quad\Big\|\textstyle\int_{0}^{\cdot}Z_{s}\!\cdot dW_{s}\Big\|_{\mathrm{BMO}}\ \text{is bounded in terms of }\|Y\|_{\mathbb{S}^{\infty}}.$
@@
 ∎
+Corollary 4.5.
+Assume $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ is a solution to the BSDE (2.6) with $\xi$ satisfying:
+  • $({\rm H_{\xi}})$
+    There is a process $H\in\mathbb{H}^{\infty}$ satisfying
+    $\xi=\mathbb{E}\xi+\int_{0}^{T}H_{s}dW_{s}.$
+Then $Z\cdot W$ is sliceable and moreover, $\left\|\int_{0}^{T}|Z_{s}|^{2}ds\right\|_{L^{\infty}}<\infty.$
+Proof.
+From the proof of Proposition 4.3, we have
+$Y_{t}=\xi-\int_{t}^{T}Z_{s}\cdot dW^{L}_{s}.$
+In other words,
+$\int_{0}^{t}Z_{s}\cdot dW^{L}_{s}=Y_{t}-Y_{0}=\mathbb{E}^{L}_{t}[\xi]-\mathbb{E}^{L}[\xi].$
+Since the quadratic variation does not change under an equivalent change of measure, we have
+$\left\|\int_{0}^{T}|Z_{s}|^{2}ds\right\|_{L^{\infty}}=\left\|\int_{0}^{T}|H_{s}|^{2}ds\right\|_{L^{\infty}}<\infty.$
+Since the class of sliceable BMO martingales is the BMO closure of martingales with bounded quadratic variation, this shows that $Z\cdot W$ is sliceable. ∎
 Theorem 4.6.
-For $i=1,2$, let $(Y^{i},Z^{i})\in\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}$ solve
+For $i=1,2$, consider $(Y^{i},Z^{i})\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ solve
 $Y_{t}^{i}=\xi^{i}+\int_{t}^{T}\!\int_{\tilde{\Omega}}g(s,Z_{s}^{i},\tilde{Z}_{s}^{i})\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}^{i}\cdot dW_{s},\qquad t\in[0,T],$
-with $\xi^{1},\xi^{2},g$ satisfying $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q})$. Then there exist $p^{*}>1$ and, for every $p>p^{*}$, a constant $C_{p}>0$ such that
-$\|Y^{1}-Y^{2}\|_{\mathbb{S}^{2p}}+\|Z^{1}-Z^{2}\|_{\mathbb{H}^{p}}\ \leq\ C_{p}\;\|\xi^{1}-\xi^{2}\|_{L^{2p}},$
-where $p^{*}>1$ and $C_{p}$ depend only on $T$, $K_{z}$, $K_{\tilde{z}}$, $K_{0}$, and do not depend on $\|Z^{i}\|_{\mathbb{S}^{\infty}}$.
+with $\xi^{1},\xi^{2},g$ satisfying $(\mathrm{H}_{0}),(\mathrm{H}_{q})$ and $(\mathrm{H}_{\xi})$. Then there exists a constant $C>0$ such that
+$\|Y^{1}-Y^{2}\|_{\mathbb{S}^{\infty}}+\|Z^{1}-Z^{2}\|_{\mathbb{H}^{2}_{\mathrm{BMO}}}\ \leq\ C\;\|\xi^{1}-\xi^{2}\|_{L^{\infty}},$
+where $C$ depends only on $(T,K_{z},K_{\tilde{z}},K_{0})$ and the quadratic variation bounds of $\mathbb{E}_{t}[\xi^{1}]$ and $\mathbb{E}_{t}[\xi^{2}]$ in Corollary 4.5.
 Proof.
@@
 $\Delta g(s,Z_{s},\tilde{Z}_{s}):=g(s,Z_{s}^{1},\tilde{Z}_{s}^{1})-g(s,Z_{s}^{2},\tilde{Z}_{s}^{2}),\qquad s\in[0,T].$
-Then we obtain
-$\Delta Y_{t}=\Delta\xi+\int_{t}^{T}\!\int_{\tilde{\Omega}}\!\Delta g(s,Z_{s},\tilde{Z}_{s})\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}\Delta Z_{s}\cdot dW_{s},\qquad t\in[0,T].$ (4.17)
-Preliminary step (linearization, RHI and sliceability).
-As in the proof of Proposition 4.4, we use a linearization argument: there exist processes $L_{s}$ and $\tilde{L}_{s}$ on the product space such that, for all $s\in[0,T]$,
-$L_{s}:=\frac{g(s,Z^{1}_{s},\tilde{Z}^{1}_{s})-g(s,Z^{2}_{s},\tilde{Z}^{1}_{s})}{|\Delta Z_{s}|^{2}}\,\Delta Z_{s}\,\mathbf{1}_{\{\Delta Z_{s}\neq 0\}},\qquad\tilde{L}_{s}:=\frac{g(s,Z^{2}_{s},\tilde{Z}^{1}_{s})-g(s,Z^{2}_{s},\tilde{Z}^{2}_{s})}{|\Delta\tilde{Z}_{s}|^{2}}\,\Delta\tilde{Z}_{s}\,\mathbf{1}_{\{\Delta\tilde{Z}_{s}\neq 0\}}$
-with estimates
-$|L_{s}|\leq K_{z}\big(1+|Z_{s}^{1}|+|Z_{s}^{2}|\big),\qquad|\tilde{L}_{s}|\leq K_{\tilde{z}}\big(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|\big),$ (4.18)
-and $\Delta g(s,Z_{s},\tilde{Z}_{s})=L_{s}\Delta Z_{s}+\tilde{L}_{s}\,\Delta\tilde{Z}_{s}$.
-According to Proposition 4.2, we have $Z^{1},Z^{2}\in\mathbb{H}^{2}_{\mathrm{BMO}}$, so in particular $\big(\int_{0}^{t}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\big)_{t\in[0,T]}$ is a $\mathrm{BMO}$ martingale by (4.18) and symmetry.
-Moreover, by Corollary 4.5, there exists a constant $C>0$, depending only on $(T,K_{z},K_{\tilde{z}},K_{0})$, such that
-$\Big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\Big\|_{\mathrm{BMO}}+\Big\|\int_{0}^{\cdot}Z^{1}_{s}\cdot dW_{s}\Big\|_{\mathrm{BMO}}+\Big\|\int_{0}^{\cdot}Z^{2}_{s}\cdot dW_{s}\Big\|_{\mathrm{BMO}}\;\leq\;C.$ (4.19)
-Hence $\Lambda:=\mathcal{E}(\int_{0}^{\cdot}\widetilde{\mathbb{E}}[L_{s}]\cdot dW_{s})$ defines an equivalent probability measure $\mathbb{P}^{L}$ by Girsanov’s theorem with Brownian motion $W^{L}_{t}:=W_{t}+\int_{0}^{t}\widetilde{\mathbb{E}}[L_{s}]\,ds$.
-In addition, by Lemma 2.2, the same processes admit $\mathrm{BMO}$ norms under the equivalent measure $\mathbb{P}^{L}$, up to multiplicative constants depending only on $C$.
+We consider a sequence of stopping times $0=\tau_{0}\leq\tau_{1}\leq\cdots\leq\tau_{N}=T$ that slices the BMO martingales $Z^{1}\cdot W$ and $Z^{2}\cdot W$ into pieces with BMO norm of size $\varepsilon$ and such that $\tau_{j+1}-\tau_{j}<\varepsilon$ for all $j=0,1,\dots,N-1$.
+Then, for $\tau_{j}\leq t\leq\tau_{j+1}$, we obtain
+$\Delta Y_{t}=\Delta Y_{\tau_{j+1}}+\int_{t}^{\tau_{j+1}}\!\int_{\tilde{\Omega}}\!\Delta g(s,Z_{s},\tilde{Z}_{s})\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{\tau_{j+1}}\Delta Z_{s}\cdot dW_{s}.$ (4.20)
+We define the sliced BMO norm on $\tau_{j},\tau_{j+1}]$ by
+$\big\|Z^{i}\cdot W\big\|_{\mathrm{BMO}_{j}(\mathbb{P})}:=\sup_{\tau\in\mathcal{T}}\Big\|\mathbb{E}_{\tau}\Big[\int_{\tau}^{\tau_{j+1}}|Z^{i}_{s}|^{2}\,ds\Big]\Big\|_{L^{\infty}}^{1/2},$
+where $\mathcal{T}$ denotes the collection of all stopping times valued in $\tau_{j},\tau_{j+1}]$.
+Similarly, $\Lambda:=\mathcal{E}(\int_{0}^{\cdot}\widetilde{\mathbb{E}}[L_{s}]\cdot dW_{s})$ defines an equivalent probability measure $\mathbb{P}^{L}$ by Girsanov’s theorem with Brownian motion $W^{L}_{t}:=W_{t}+\int_{0}^{t}\widetilde{\mathbb{E}}[L_{s}]\,ds$, we set
+$\big\|Z^{i}\cdot W^{L}\big\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})}:=\sup_{\tau\in\mathcal{T}_{j}}\Big\|\mathbb{E}_{\tau}^{\mathbb{P}^{L}}\Big[\int_{\tau}^{\tau_{j+1}}|Z^{i}_{s}|^{2}\,ds\Big]\Big\|_{L^{\infty}}^{1/2}.$
+According to Proposition 4.2, we have $Z^{1},Z^{2}\in\mathbb{H}^{2}_{\mathrm{BMO}}$, so in particular $\big(\int_{0}^{t}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\big)_{t\in[0,T]}$ is a $\mathrm{BMO}$ martingale.
+Moreover, by Corollary 4.4 and Lemma 2.2, there exists a constant $C>0$, depending only on $(T,K_{z},K_{\tilde{z}},K_{0})$, such that
+$\Big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\Big\|_{\mathrm{BMO}}$ $+\Big\|\int_{0}^{\cdot}Z^{1}_{s}\cdot dW_{s}\Big\|_{\mathrm{BMO}}+\Big\|\int_{0}^{\cdot}Z^{2}_{s}\cdot dW_{s}\Big\|_{\mathrm{BMO}}\;\leq\;C,$ (4.21)
+$\frac{1}{C}\left\|Z^{i}\cdot W^{L}\right\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})}$ $\leq\left\|Z^{i}\cdot W\right\|_{\mathrm{BMO}_{j}(\mathbb{P})}\leq C\left\|Z^{i}\cdot W^{L}\right\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})},\qquad i=1,2,\ j=0,1,\dots,N-1,$
 Under $\mathbb{P}^{L}$ we can rewrite (4.20) as
-$\Delta Y_{u}=\Delta\xi+\int_{u}^{T}\widetilde{\mathbb{E}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds-\int_{u}^{T}\Delta Z_{s}\cdot dW^{L}_{s},\qquad u\in[0,T].$ (4.20)
-The Doléans–Dade exponential $\Lambda=\mathcal{E}\!\left(\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\right)$ satisfies a reverse Hölder inequality by Lemma 2.1: there exists $q^{*}>1$ such that, for all $1\leq q\leq q^{*}$,
-$\bar{{\mathbb{E}}}\!\left[\left(\Lambda_{T}\right)^{q}\,\big|\,\bar{\mathcal{F}}_{t}\right]\ \leq\ C_{q}^{*}\,\left(\Lambda_{t}\right)^{q},\qquad 0\leq t\leq T,$
+$\Delta Y_{t}=\Delta Y_{\tau_{j+1}}+\int_{t}^{\tau_{j+1}}\widetilde{\mathbb{E}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds-\int_{t}^{\tau_{j+1}}\Delta Z_{s}\cdot dW^{L}_{s},\qquad\tau_{j}\leq t\leq\tau_{j+1}.$ (4.22)
+By the Itô formula, for any $t\in[\tau_{j},\tau_{j+1}]$, we have
+$|\Delta Y_{t}|^{2}+\mathbb{E}^{L}_{t}\int_{t}^{\tau_{j+1}}|\Delta Z_{s}|^{2}ds$ $\leq\mathbb{E}^{L}_{t}|\Delta Y_{\tau_{j+1}}|^{2}+2\mathbb{E}^{L}_{t}\int_{t}^{\tau_{j+1}}|\Delta Y_{s}|\left|\widetilde{\mathbb{E}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\right|ds$ (4.23)
+$\leq\mathbb{E}^{L}_{t}|\Delta Y_{\tau_{j+1}}|^{2}+2\left\|\sup_{s\in[t,\tau_{j+1}]}|\Delta Y_{s}|\right\|_{L^{\infty}}\mathbb{E}^{L}_{t}\int_{t}^{\tau_{j+1}}\widetilde{\mathbb{E}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]ds.$
+Note that there exists a constant $A$ such that
+$\left|\widetilde{\mathbb{E}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\right|$ $\leq A\widetilde{\mathbb{E}}[(1+|Z^{1}_{s}|+|Z^{2}_{s}|+|\tilde{Z}^{1}_{s}|+|\tilde{Z}^{2}_{s}|)|\Delta\tilde{Z}_{s}|]$
+$\leq A\left[\left(|Z^{1}_{s}|+|Z^{2}_{s}|\right)\mathbb{E}|\Delta Z_{s}|+\mathbb{E}[(1+|Z^{1}_{s}|+|Z^{2}_{s}|)|\Delta Z_{s}|]\right].$
+The Fefferman’s inequality for BMO martingales yields that
+$\int_{t}^{\tau_{j+1}}\mathbb{E}\left[(1+|Z_{s}^{1}|+|Z_{s}^{2}|)|\Delta Z_{s}|\right]ds$ $\leq\sqrt{2}\mathbb{E}\left[\left(\int_{t}^{\tau_{j+1}}(1+|Z_{s}^{1}|+|Z_{s}^{2}|)^{2}ds\right)^{1/2}\right]\left\|\Delta Z\cdot W\right\|_{\mathrm{BMO}_{j}(\mathbb{P})}$
+$\leq\sqrt{6}C\left(\varepsilon+\left\|Z^{1}\cdot W\right\|_{\mathrm{BMO}_{j}(\mathbb{P})}+\left\|Z^{2}\cdot W\right\|_{\mathrm{BMO}_{j}(\mathbb{P})}\right)\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})}$
+$\leq 3\sqrt{6}C\varepsilon\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})}.$
+On the other hand,
+$\mathbb{E}^{L}_{t}\int_{t}^{\tau_{j+1}}\left(|Z^{1}_{s}|+|Z^{2}_{s}|\right)\mathbb{E}|\Delta Z_{s}|ds$ $\leq\sqrt{2}\mathbb{E}^{L}_{t}\left(\int_{t}^{\tau_{j+1}}\left(|Z^{1}_{s}|^{2}+|Z^{2}_{s}|^{2}\right)ds\right)^{1/2}\left(\int_{t}^{\tau_{j+1}}\mathbb{E}|\Delta Z_{s}|^{2}ds\right)^{1/2}$
+$\leq\sqrt{2}C\left(\left\|Z^{1}\cdot W^{L}\right\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})}+\left\|Z^{2}\cdot W^{L}\right\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})}\right)\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})}$
+$\leq 2\sqrt{2}C^{2}\varepsilon\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})}.$
+As a result, (4.23) becomes
+$|\Delta Y_{t}|^{2}+\mathbb{E}^{L}_{t}\int_{t}^{\tau_{j+1}}|\Delta Z_{s}|^{2}ds$ $\leq\mathbb{E}^{L}_{t}|\Delta Y_{\tau_{j+1}}|^{2}+4A(3\sqrt{6}C+2\sqrt{2}C^{2})\varepsilon\left\|\sup_{s\in[t,{\tau_{j+1}}]}|\Delta Y_{s}|\right\|_{L^{\infty}}\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})}$
+$\leq\left\|\Delta Y_{\tau_{j+1}}\right\|^{2}_{L^{\infty}}+\frac{1}{4}\left(\left\|\sup_{s\in[{\tau_{j}},{\tau_{j+1}}]}|\Delta Y_{s}|\right\|^{2}_{L^{\infty}}+\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})}^{2}\right),$
+by letting $\varepsilon$ small enough.
+Then, we have $\left\|\sup_{s\in[{\tau_{j}},{\tau_{j+1}}]}|\Delta Y_{s}|^{2}\right\|_{L^{\infty}}$ and $\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})}^{2}$ are both bounded by
+$\left\|\Delta Y_{\tau_{j+1}}\right\|^{2}_{L^{\infty}}+\frac{1}{4}\left\|\sup_{s\in[{\tau_{j}},{\tau_{j+1}}]}|\Delta Y_{s}|\right\|^{2}_{L^{\infty}}+\frac{1}{4}\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})}^{2}.$
+This implies that
+$\left\|\sup_{s\in[{\tau_{j}},{\tau_{j+1}}]}|\Delta Y_{s}|^{2}\right\|_{L^{\infty}}+\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}_{j}(\mathbb{P}^{L})}^{2}\leq 4\left\|\Delta Y_{\tau_{j+1}}\right\|^{2}_{L^{\infty}}.$
+Applying the above estimates iteratively from $T$ to 0, we obtain the estimate
+$\left\|\sup_{s\in[0,T]}|\Delta Y_{s}|^{2}\right\|_{L^{\infty}}+\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}(\mathbb{P}^{L})}^{2}\leq N4^{N}\left\|\Delta\xi\right\|_{L^{\infty}}^{2},$
+where $N$ depends only on $(T,K_{z},K_{\tilde{z}},K_{0})$. Another application of Lemma 2.2 proves the claim. ∎
+We can remove $\mathrm{(H_{\xi})}$ under a more restrictive structural assumption on $g$.
+Theorem 4.7.
+For $i=1,2$, consider $(Y^{i},Z^{i})\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ solve
+$Y_{t}^{i}=\xi^{i}+\int_{t}^{T}\!\int_{\tilde{\Omega}}g(s,Z_{s}^{i},\tilde{Z}_{s}^{i})\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}^{i}\cdot dW_{s},\qquad t\in[0,T],$
+with $\xi^{1},\xi^{2},g$ satisfying $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q}^{\mathrm{sep}})$. Then there exists a constant $C>0$ such that
+$\|Y^{1}-Y^{2}\|_{\mathbb{S}^{\infty}}+\|Z^{1}-Z^{2}\|_{\mathbb{H}^{2}_{\mathrm{BMO}}}\ \leq\ C\;\|\xi^{1}-\xi^{2}\|_{L^{\infty}},$
+where $C$ depends only on $(T,K_{z},K_{\tilde{z}},K_{0})$.
+Proof.
+In the previous proof, we can take $\tau_{j}=\frac{jT}{N}$, where $N$ is large enough. Consider $u=\tau_{j}$ and $v=\tau_{j+1}$. By the Itô formula, for any $t\in[u,v]$, we have
+$|\Delta Y_{t}|^{2}+\mathbb{E}^{L}_{t}\int_{t}^{v}|\Delta Z_{s}|^{2}ds$ $\leq\mathbb{E}^{L}_{t}|\Delta Y_{v}|^{2}+2K_{\tilde{z}}\mathbb{E}^{L}_{t}\int_{t}^{v}|\Delta Y_{s}|\mathbb{E}\left[(1+|Z_{s}^{1}|+|Z_{s}^{2}|)|\Delta Z_{s}|\right]ds$
+$\leq\mathbb{E}^{L}_{t}|\Delta Y_{v}|^{2}+2K_{\tilde{z}}\mathbb{E}^{L}_{t}\sup_{s\in[t,v]}|\Delta Y_{s}|\int_{t}^{v}\mathbb{E}\left[(1+|Z_{s}^{1}|+|Z_{s}^{2}|)|\Delta Z_{s}|\right]ds$
+$\leq\mathbb{E}^{L}_{t}|\Delta Y_{v}|^{2}+2K_{\tilde{z}}\left(\mathbb{E}^{L}_{t}\sup_{s\in[u,v]}|\Delta Y_{s}|\right)\mathbb{E}\int_{u}^{v}\left[(1+|Z_{s}^{1}|+|Z_{s}^{2}|)|\Delta Z_{s}|\right]ds.$ (4.24)
+On the other hand, by the Fefferman’s inequality for BMO martingale, we have
+$\mathbb{E}\int_{u}^{v}\left[(1+|Z_{s}^{1}|+|Z_{s}^{2}|)|\Delta Z_{s}|\right]ds\leq\sqrt{2}\mathbb{E}\left[\left(\int_{u}^{v}(1+|Z_{s}^{1}|+|Z_{s}^{2}|)^{2}ds\right)^{1/2}\right]\left\|\Delta Z\cdot W\right\|_{\mathrm{BMO}_{[u,v]}(\mathbb{P})},$
 where
-$q^{*}$ $:=\phi\!\left(\Big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\Big\|_{\mathrm{BMO}}\right),\qquad\phi(x):=\left(1+\frac{1}{x^{2}}\log\frac{2x-1}{2x-2}\right)^{\!1/2}-1,$
-$C_{q}^{*}$ $:=2\left(1-\frac{2q-1}{2q-2}\,\exp\!\left\{q^{2}\Big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2}+2\Big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\Big\|_{\mathrm{BMO}}\right\}\right)^{-1}\!.$
-By boundedness of $\Big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\Big\|_{\mathrm{BMO}}$, the constants $q^{*},C_{q}^{*}$ depend only on $(T,K_{z},K_{\tilde{z}},K_{0})$. Finally, set
-$M_{t}:=\int_{0}^{t}(\tilde{\mathbb{E}}|K_{\tilde{z}}\big(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|\big)|^{2})^{1/2}dW_{s},\qquad t\in[0,T].$
-Then $M\in\mathcal{H}^{\infty}(\mathbb{P})$, and by schachermayer1996institut it is sliceable.
-Furthermore, since $Z^{1},Z^{2}\in\mathbb{S}^{\infty}$, we can (by (jackson2022triangle, Remark 2.8)) construct a deterministic partition $0=t_{0}<t_{1}<\dots<t_{N}=T$ with arbitrarily small mesh size bounded by a prescribed $\varepsilon>0$.
-$\mathbb{S}^{p}$–control of $\Delta Y$.
-Fix $\varepsilon$, $k\in\{0,\dots,N-1\}$ and write $I_{k}=[t_{k},t_{k+1}]$.
-By subtracting (4.20) at $u=t_{k+1}$ from that at $u\in I_{k}$, we obtain the local BSDE
-$\Delta Y_{u}=\Delta Y_{t_{k+1}}+\int_{u}^{t_{k+1}}\widetilde{\mathbb{E}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds-\int_{u}^{t_{k+1}}\Delta Z_{s}\cdot dW^{L}_{s},\qquad u\in I_{k}.$
-Using Cauchy-Schwarz on the copy space, the growth bound (4.18), and then Fefferman’s inequality (Theorem 2.5 in kazamaki2006BMO) on $I_{k}$, we obtain
-$\int_{u}^{t_{k+1}}\tilde{\mathbb{E}}|\tilde{L}_{s}||\Delta\tilde{Z}_{s}|ds$ $\leq\mathbb{E}\int_{0}^{T}(\tilde{\mathbb{E}}\big|K_{\tilde{z}}\big(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|\big)\big|^{2})^{1/2}(\tilde{\mathbb{E}}|\Delta\tilde{Z}_{s}|^{2})^{1/2}ds$ (4.21)
-$\leq\sqrt{2}\left\|\int_{0}^{\cdot}(\tilde{\mathbb{E}}|K_{\tilde{z}}\big(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|\big)|^{2})^{1/2}dW_{s}\right\|_{BMO}\int_{u}^{t_{k+1}}(\mathbb{E}|\Delta Z_{s}|^{2})^{1/2}ds$
-$\leq\varepsilon\int_{u}^{t_{k+1}}(\mathbb{E}|\Delta Z_{s}|^{2})^{1/2}ds\qquad u\in I_{k}.$
-Taking $\mathbb{E}^{L}[\cdot\,|\,\mathcal{F}_{u}]$ in the local BSDE yields
-$|\Delta Y_{u}|$ $\leq\mathbb{E}^{L}[|\Delta Y_{t_{k+1}}|\,|\,\mathcal{F}_{u}]+\varepsilon\int_{u}^{t_{k+1}}\big(\mathbb{E}|\Delta Z_{s}|^{2}\big)^{1/2}ds,\qquad u\in I_{k}.$
-Iterating this estimate over the intervals $I_{k+1},I_{k+2},\dots,I_{N-1}$ and using $\Delta Y_{T}=\Delta\xi$, we obtain
-$|\Delta Y_{u}|\leq\mathbb{E}^{L}[|\Delta\xi|\,|\,\mathcal{F}_{u}]+\varepsilon\sqrt{T}\,\left(\int_{u}^{T}\mathbb{E}|\Delta Z_{s}|^{2}\,ds\right)^{1/2},\qquad u\in[0,T].$
-By BMO energy inequality (kazamaki2006BMO, Sec. 2.1) and Corollary 4.5, the last term is bounded by a constant depending only on $(T,K_{z},K_{\tilde{z}},K_{0})$ since:
-$\int_{u}^{T}\Big({\mathbb{E}}|Z_{s}^{1}|^{2}+{\mathbb{E}}|Z_{s}^{2}|^{2}\Big)\,ds\;$ $\leq\;\Big\|\int_{0}^{\cdot}Z_{s}^{1}\,\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2}+\Big\|\int_{0}^{\cdot}Z_{s}^{2}\,\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2}.$
-Therefore, letting $\varepsilon\to 0$ yields
-$|\Delta Y_{u}|\leq\mathbb{E}^{L}[|\Delta\xi|\,|\,\mathcal{F}_{u}]=(\Lambda_{u})^{-1}\,\mathbb{E}[\Lambda_{T}|\Delta\xi|\,|\,\mathcal{F}_{u}],\qquad u\in[0,T].$ (4.22)
-Now fix $q\in(1,q^{*})$ and let $p$ be its conjugate exponent. By the reverse Hölder inequality and conditional Hölder’s inequality,
-$|\Delta Y_{u}|^{p}$ $\leq(\Lambda_{u})^{-p}\,\mathbb{E}[(\Lambda_{T})^{q}\,|\,\mathcal{F}_{u}]^{p/q}\,\mathbb{E}[|\Delta\xi|^{p}\,|\,\mathcal{F}_{u}]$
-$\leq(C_{q}^{*})^{p/q}\,\mathbb{E}[|\Delta\xi|^{p}\,|\,\mathcal{F}_{u}],\qquad u\in[0,T].$
-By Doob’s maximal inequality,
-$\mathbb{E}\Big(\sup_{u\in[0,T]}|\Delta Y_{u}|^{p}\Big)\ \leq\ C_{Y}(p)\,\mathbb{E}|\Delta\xi|^{p},\qquad p\geq p^{*},$ (4.23)
-where $p^{*}$ is the conjugate exponent of $q^{*}$ and $C_{Y}(p):=\Big(\frac{p}{p-1}\Big)^{p}(C_{q}^{*})^{p/q}$.
-They both depending only on $(T,K_{z},K_{\tilde{z}},K_{0})$ as we discussed in the Preliminary step.
-$\mathbb{H}^{p}$–control of $\Delta Z$. Fix $p>p^{*}$, and applying Itô’s formula to $|\Delta Y|^{2}$ on $[0,T]$. we obtain
-$|\Delta Y_{0}|^{2}+\!\int_{0}^{T}|\Delta Z_{s}|^{2}\,ds=|\Delta\xi|^{2}+2\int_{0}^{T}\Delta Y_{s}\,\widetilde{\mathbb{E}}[\Delta g(s,Z_{s},\tilde{Z}_{s})]\,ds-2\int_{0}^{T}\Delta Y_{s}\,Z_{s}\cdot dW_{s}$
-Using the same Fefferman/Cauchy–Schwarz argument as above and $\|\Delta Y\|_{\mathbb{S}^{\infty}}<\infty$, we obtain
-$2\int_{0}^{T}\Delta Y_{s}\,\widetilde{{\mathbb{E}}}\!\big[\Delta g(s,Z_{s},\tilde{Z}_{s})\big]\,ds$ $\leq 2K_{z}\int_{0}^{T}(1+|Z_{s}^{1}|+|Z_{s}^{2}|)\,|\Delta Y_{s}|\,|\Delta Z_{s}|\,ds$
-$\quad+2K_{\tilde{z}}\int_{0}^{T}|\Delta Y_{s}|\,\widetilde{{\mathbb{E}}}\!\Big[(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|)\,|\Delta\tilde{Z}_{s}|\Big]\,ds$
-$\leq\sup_{t\in[0,T]}|\Delta Y_{t}|^{2}\Big(2K_{z}^{2}\int_{0}^{T}(1+|Z_{s}^{1}|+|Z_{s}^{2}|)^{2}\,ds\Big)+\frac{1}{2}\int_{0}^{T}|\Delta Z_{s}|^{2}\,ds$
-$\quad+2K_{\tilde{z}}\|\Delta Y\|_{\mathbb{S}^{\infty}}\int_{0}^{T}\widetilde{{\mathbb{E}}}\!\Big[(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|)\,|\Delta\tilde{Z}_{s}|\Big]\,ds.$
-The last term vanishes in the limit $\varepsilon\to 0$ by the same slicing argument as in (4.21). Next, let $K:=\int_{0}^{T}(1+|Z_{s}^{1}|+|Z_{s}^{2}|)^{2}\,ds$. Then
-${\mathbb{E}}\Big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{p}\,K^{p/2}\Big]\leq\Big({\mathbb{E}}\big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{2p}\big]\Big)^{1/2}\Big({\mathbb{E}}\big[K^{p}\big]\Big)^{1/2}.$ (4.24)
-If $\int_{0}^{\cdot}Z_{s}^{i}\cdot dW_{s}$ is a BMO martingale ($i=1,2$), the BMO energy inequality (see (kazamaki2006BMO, Sec. 2.1)) gives
-${\mathbb{E}}\Big[\Big(\int_{0}^{T}|Z_{s}^{i}|^{2}ds\Big)^{p}\Big]\leq p!\,\Big\|\int_{0}^{\cdot}Z_{s}^{i}\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2p}.$
-Therefore,
-${\mathbb{E}}[K^{p}]\leq 3^{p-1}\Big((3T)^{p}+3^{p}p!\,\Big\|\int_{0}^{t}Z_{s}^{1}\,\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2p}+3^{p}p!\,\Big\|\int_{0}^{t}Z_{s}^{2}\,\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2p}\Big)=:C_{\mathrm{energy}}(p).$ (4.25)
-Lastly, with the help of BDG inequality and Cauchy–Schwarz, we have
-${\mathbb{E}}\Big|\int_{0}^{T}(\Delta Y_{s})\Delta Z_{s}\cdot dW_{s}\Big|^{p/2}$ $\leq C_{\mathrm{BDG}}\,{\mathbb{E}}\Big[\Big(\int_{0}^{T}|\Delta Y_{s}|^{2}|\Delta Z_{s}|^{2}\,ds\Big)^{p/4}\Big]$
-$\leq C_{\mathrm{BDG}}\,\Big({\mathbb{E}}\big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{p}\big]\Big)^{1/2}\Big({\mathbb{E}}\big[(\int_{0}^{T}|\Delta Z_{s}|^{2}\,ds)^{p/2}\big]\Big)^{1/2}.$ (4.26)
-Putting (4.24), (4.25) and (4) together, we obtain
-${\mathbb{E}}\Big[\Big(\int_{0}^{T}|\Delta Z_{s}|^{2}ds\Big)^{p/2}\Big]$ $\leq 2^{\frac{3p}{2}-1}{\mathbb{E}}[|\Delta\xi|^{p}]+2^{\frac{3p}{2}-1}(2K_{z}^{2})^{p/2}\sqrt{C_{\mathrm{energy}}(p)}\Big({\mathbb{E}}[\sup_{t\in[0,T]}|\Delta Y_{t}|^{2p}]\Big)^{1/2}$
-$\quad+2^{4p-4}(C_{\mathrm{BDG}})^{2}\,{\mathbb{E}}[\sup_{t\in[0,T]}|\Delta Y_{t}|^{p}],$
-From previous step (4.23), there exist constants $C_{Y}(p),C_{Y}(2p)$ such that
-$\mathbb{E}\Big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{p}\Big]\leq C_{Y}(p)\,\mathbb{E}\big[|\Delta\xi|^{p}\big],\qquad\Big(\mathbb{E}\Big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{2p}\Big]\Big)^{1/2}\leq\sqrt{C_{Y}(2p)}\,\Big(\mathbb{E}\big[|\Delta\xi|^{2p}\big]\Big)^{1/2}.$
-Substituting these bounds into the previous inequality, we obtain for all $p\geq p^{*}$,
-$\mathbb{E}\Big[\Big(\int_{0}^{T}|\Delta Z_{s}|^{2}ds\Big)^{p/2}\Big]$ $\leq 2^{\frac{3p}{2}-1}\mathbb{E}\big[|\Delta\xi|^{p}\big]+2^{4p-4}(C_{\mathrm{BDG}})^{2}\,C_{Y}(p)\,\mathbb{E}\big[|\Delta\xi|^{p}\big]$
-$\quad+2^{\frac{3p}{2}-1}(2K_{z}^{2})^{p/2}\sqrt{C_{\mathrm{energy}}(p)}\,\sqrt{C_{Y}(2p)}\,\Big(\mathbb{E}\big[|\Delta\xi|^{2p}\big]\Big)^{1/2}.$
-Finally, since $\mathbb{E}|\Delta\xi|^{p}\leq\big(\mathbb{E}|\Delta\xi|^{2p}\big)^{1/2}$, we may absorb the first two terms into the last one and conclude that
-$\mathbb{E}\Big[\Big(\int_{0}^{T}|\Delta Z_{s}|^{2}ds\Big)^{p/2}\Big]$ $\leq C_{Z}(p)\,\Big(\mathbb{E}\big[|\Delta\xi|^{2p}\big]\Big)^{1/2},$
-where
-$C_{Z}(p):=2^{\frac{3p}{2}-1}+2^{4p-4}(C_{\mathrm{BDG}})^{2}\,C_{Y}(p)+2^{\frac{3p}{2}-1}(2K_{z}^{2})^{p/2}\sqrt{C_{\mathrm{energy}}(p)}\,\sqrt{C_{Y}(2p)}.$
-Combing the stability estimate (4.23) completes the proof. ∎
+$\left\|\Delta Z\cdot W\right\|_{\mathrm{BMO}_{[u,v]}(\mathbb{P})}=\left(\sup_{\tau\in\mathcal{T}}\left\|\mathbb{E}_{\tau}\int_{\tau}^{v}|\Delta Z|^{2}ds\right\|_{L^{\infty}}\right)^{1/2}$
+and $\mathcal{T}$ denotes the collection of all stopping times valued in $[u,v]$.
+By Lemma 2.2 applied on the interval $[u,v]$ and Corollary 4.4, there exists a constant $C$, which only depends on $(T,K_{z},K_{\tilde{z}},K_{0})$, such that
+$\left\|\Delta Z\cdot W\right\|_{\mathrm{BMO}_{[u,v]}(\mathbb{P})}\leq\frac{C}{2\sqrt{2}K_{\tilde{z}}}\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}_{[u,v]}(\tilde{\mathbb{P}})}$.
+In addition, since $t\mapsto\left(\mathbb{E}\int_{0}^{t}(1+|Z_{s}^{1}|+|Z_{s}^{2}|)^{2}ds\right)^{1/2}$ is uniformly continuous, for any $\varepsilon>0$, there exists $\delta>0$ such that $v-u<\delta$ implies
+$\left(\mathbb{E}\int_{u}^{v}(1+|Z_{s}^{1}|+|Z_{s}^{2}|)^{2}ds\right)^{1/2}<\frac{\varepsilon}{C}.$ (4.25)
+As a result, (4) becomes
+$|\Delta Y_{t}|^{2}+\mathbb{E}^{L}_{t}\int_{t}^{v}|\Delta Z_{s}|^{2}ds$ $\leq\mathbb{E}^{L}_{t}|\Delta Y_{v}|^{2}+\varepsilon\mathbb{E}^{L}_{t}\sup_{s\in[u,v]}|\Delta Y_{s}|\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}_{[u,v]}(\tilde{\mathbb{P}})}$
+$\leq\left\|\Delta Y_{v}\right\|^{2}_{L^{\infty}}+\frac{\varepsilon}{2}\left\|\sup_{s\in[u,v]}|\Delta Y_{s}|\right\|^{2}_{L^{\infty}}+\frac{\varepsilon}{2}\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}_{[u,v]}(\tilde{\mathbb{P}})}^{2}.$
+By taking $\varepsilon=1/2$, by the similar argument in the proof of previous theorem, we have
+$\left\|\sup_{s\in[u,v]}|\Delta Y_{s}|^{2}\right\|_{L^{\infty}}+\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}_{[u,v]}(\tilde{\mathbb{P}})}^{2}\leq 4\left\|\Delta Y_{v}\right\|^{2}_{L^{\infty}}.$
+If we partition $[0,T]$ into intervals of size less than $\delta$ and apply the above estimates iteratively from $T$ to 0, we obtain the estimate
+$\left\|\sup_{s\in[0,T]}|\Delta Y_{s}|^{2}\right\|_{L^{\infty}}+\left\|\Delta Z\cdot W^{L}\right\|_{\mathrm{BMO}(\tilde{\mathbb{P}})}^{2}\leq N4^{N}\left\|\Delta\xi\right\|_{L^{\infty}}^{2},$
+where $N$ is the number of subintervals in the partition. Another application of Lemma 2.2 proves the claim. ∎
 5 Proof: existence and uniqueness
-In this section, we establish the main result for the mean-field BSDE (2.6) under Assumptions $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q})$.
-The proof proceeds in two main steps.
-First, we prove the existence of a solution when the terminal condition $\xi$ is Malliavin differentiable with bounded derivative.
-This is achieved in Lemma 5.1 via a truncation and regularization argument.
-With the help of the estimates for Lipschitz BSDEs derived in Proposition 3.1, this allows us to apply the classical result from Buckdahn2009MFBSDE.
-In the second step, we extend this existence result to the case of general bounded terminal conditions by approximating $\xi$ with a sequence of Malliavin differentiable random variables.
-For each approximating BSDE, existence is guaranteed by Lemma 5.1, and we invoke the stability result from Theorem 4.6 to show that the sequence of solutions is Cauchy in suitable norms.
-The limit of this sequence is then shown to solve the original mean-field BSDE (2.6).
-Uniqueness follows directly from the same stability result.
-The full argument is completed in the proof of the main theorem below.
+In this section, we prove the existence and uniqueness results for the mean-field BSDE (2.6) in the two quadratic regimes introduced earlier.
+Under Assumptions $(\mathrm{H}_{0})$} and $(\mathrm{H}_{q}^{\mathrm{unif}})$}, we establish Theorem 5.3 for terminal conditions of the form $\xi=\phi(W_{t_{1}},\ldots,W_{t_{n}})$ with $\phi$ bounded and $\alpha$-Hölder continuous.
+The generalization to Theorem 2.6 can be achieved by measure-change and time-change.
+Under Assumptions $(\mathrm{H}_{0})$} and $(\mathrm{H}_{q}^{\mathrm{sep}})$}, we establish Theorem 5.4 for terminal conditions of the form $\xi=\phi(W_{[0,T]})$ with $\phi$ bounded and uniformly continuous in the sup norm.
+The proofs follow the same general scheme, but with different approximation methods adapted to the corresponding terminal regularity.
+We first prove in Lemma 5.1 that if the terminal condition $\xi$ is Malliavin differentiable with uniformly bounded derivative, then (2.6) admits a solution in $\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}$.
+The argument is based on truncating the quadratic generator to obtain Lipschitz generator.
+Next, we approximate the desired terminal condition by a sequence $(\xi^{n})$ with bounded Malliavin derivatives.
+In the unified case $(\mathrm{H}_{q}^{\mathrm{unif}})$, we use mollification of the Hölder function $\phi$ on $(\mathbb{R}^{d})^{n}$ and verify, via Lemma 5.2, the uniform quadratic-variation bounds required in Theorem 4.6.
+In the separable case $(\mathrm{H}_{q}^{\mathrm{sep}})$, we use the Moreau–Yosida approximation on path space, which yields Lipschitz functionals with bounded Malliavin derivatives.
+For each $n$, Lemma 5.1 provides a solution $(Y^{n},Z^{n})$, and the corresponding stability theorem (Theorem 4.6 under $(\mathrm{H}_{q}^{\mathrm{unif}})$ or Theorem 4.7 under $(\mathrm{H}_{q}^{\mathrm{sep}})$) implies that $(Y^{n},Z^{n})$ is Cauchy in the Banach space $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$.
+Passing to the limit yields a solution to (2.6), and uniqueness follows from the same stability result.
 Lemma 5.1.
-Let $(\rm H_{0})$} and $(\rm H_{q})$} holds and suppose $\xi$ is Malliavin differentiable with bounded derivative $\left\|D\xi\right\|_{\mathbb{S}^{\infty}}<\infty$, then the mean-field BSDE admits a solution in $\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}.$
+Assume $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q}^{\mathrm{unif}})$} hold, and suppose $\xi$ is Malliavin differentiable with bounded derivative $\left\|D\xi\right\|_{\mathbb{S}^{\infty}}<\infty$.
+Then the mean-field BSDE admits a solution in $\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}$.
 Proof.
-We define a sequence of driver functions $(\bar{g}_{k})_{k\in\mathbb{N}}$ as follows:
-$\bar{g}_{k}(t,z,\tilde{z}):=g\left(t,\frac{\left|z\right|\wedge k}{\left|z\right|}z,\frac{\left|\tilde{z}\right|\wedge k}{\left|\tilde{z}\right|}\tilde{z}\right),\quad k\in\mathbb{N}.$
-Since $g$ satisfies Assumption $(\rm H_{q})$}, it is straightforward to verify that the new driver $\bar{g}_{k}$ is Lipschitz, with
+For $k\in\mathbb{N}$, define the truncation map $\pi_{k}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ by
+$\pi_{k}(z):=\frac{|z|\wedge k}{|z|}\,z\,\mathbf{1}_{\{z\neq 0\}},\qquad\pi_{k}(0):=0,$
+and define a sequence of drivers $(\bar{g}_{k})_{k\in\mathbb{N}}$ by
+$\bar{g}_{k}(t,z,\tilde{z}):=g\big(t,\pi_{k}(z),\pi_{k}(\tilde{z})\big),\qquad(t,z,\tilde{z})\in[0,T]\times\mathbb{R}^{d}\times\mathbb{R}^{d}.$
+Note that $|\pi_{k}(z)|\leq k$ for all $z$, and $\pi_{k}$ is 11–Lipschitz:
+$|\pi_{k}(z)-\pi_{k}(z^{\prime})|\leq|z-z^{\prime}|,\qquad\forall z,z^{\prime}\in\mathbb{R}^{d}.$ (5.26)
 […]
-for all $k\in\mathbb{N}$ and $(t,z,z^{\prime},\tilde{z},\tilde{z}^{\prime})\in[0,T]\times[\mathbb{R}^{d}\times\mathbb{R}^{d}]^{2}$.
-Thus, the corresponding BSDE with driver $\bar{g}_{k}$ satisfies $(\rm H_{0})$} and $(H_{\mathrm{lip}})$.
-Then, the mean-field BSDE
+[…]
+[…]
+Since $|\pi_{k}(\cdot)|\leq k$, the common factor is bounded by $1+4k$, and by (5.26),
+[…]
+Thus, $\bar{g}_{k}$ satisfies $(\mathrm{H}_{0})$ and $(\mathrm{H}_{\mathrm{lip}})$. Consequently, the mean-field BSDE with driver $\bar{g}_{k}$,
 $Y_{t}^{k}=\xi+\int_{t}^{T}\int_{\tilde{\Omega}}\bar{g}_{k}(s,Z_{s}^{k},\tilde{Z}_{s}^{k})\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}^{k}\cdot dW_{s},\quad 0\leq t\leq T,$ (5.28)
-admits a unique solution $(Y^{k},Z^{k})\in\mathbb{S}^{2}\times\mathbb{H}^{2}$ (see Theorem 3.1 in Buckdahn2009MFBSDE).
-Moreover, Proposition 3.1 ensures that for each $k\in\mathbb{N}$, the pair $(Y^{k},Z^{k})$ belongs to $\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}$, and the process $Z^{k}$ satisfies a uniform bound independent of the Lipschitz constant of the driver with respect to $z$ and $\tilde{z}$.
-In particular, there exists a constant $C_{Z}>0$ such that $\left\|Z^{k}\right\|_{\mathbb{S}^{\infty}}\leq C_{Z}$ for all $k\in\mathbb{N}$.
-Since $(\tilde{Y}^{k},\tilde{Z}^{k})$ are independent copies of $(Y^{k},Z^{k})$, the same uniform bound holds, we can choose $k^{*}\in\mathbb{N}$ large enough so that
-$k^{*}>\sup_{k\in\mathbb{N}}\left\|Z^{k}\right\|_{\mathbb{S}^{\infty}}.$
-For such $k^{*}$, the truncation is inactive and
-$\bar{g}_{k^{*}}(t,Z_{s}^{k^{*}},\tilde{Z}_{s}^{k^{*}})=g(t,Z_{s}^{k^{*}},\tilde{Z}_{s}^{k^{*}}),\quad 0\leq s\leq T.$
-Hence, the pair $(Y^{k^{*}},Z^{k^{*}})$ indeed solves the mean-field BSDE (2.6), which completes the proof. ∎
+admits a unique solution $(Y^{k},Z^{k})\in\mathbb{S}^{2}\times\mathbb{H}^{2}$ (see Theorem 3.1 in Buckdahn et al.
+(2009b)).
+Moreover, Proposition 3.1 ensures that for each $k\in\mathbb{N}$, $(Y^{k},Z^{k})\in\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}$.
+In particular, there exists a constant $C_{Z}>0$ such that
+$\|Z^{k}\|_{\mathbb{S}^{\infty}}\leq C_{Z},\qquad\forall k\in\mathbb{N}.$
+Since $(\tilde{Y}^{k},\tilde{Z}^{k})$ are independent copies of $(Y^{k},Z^{k})$, the same bound holds for $\tilde{Z}^{k}$.
+Choose $k^{*}\in\mathbb{N}$ such that $k^{*}>C_{Z}$.
+Then for this $k^{*}$ the truncation is inactive along the solution and therefore
+$\bar{g}_{k^{*}}(s,Z_{s}^{k^{*}},\tilde{Z}_{s}^{k^{*}})=g(s,Z_{s}^{k^{*}},\tilde{Z}_{s}^{k^{*}}),\qquad 0\leq s\leq T.$
+Hence, $(Y^{k^{*}},Z^{k^{*}})$ solves the original mean-field BSDE (2.6), which completes the proof. ∎
+For the following results, let us define the time partition $0=t_{0}\leq t_{1}\leq t_{2}\leq\cdots\leq t_{n}=T$.
+Lemma 5.2.
+Consider a bounded function $\phi:(\mathbb{R}^{d})^{n}\to\mathbb{R}$ which is $\alpha$-Hölder continuous, that is, there exists $K>0$ and $\alpha\in(0,1]$ such that
+$|\phi(x)-\phi(x^{\prime})|\leq K\max_{i=1,2,\cdots,n}|x_{i}-x^{\prime}_{i}|^{\alpha},$
+for any $x,x^{\prime}\in(\mathbb{R}^{d})^{n}$. Then, for $H$ satisfying
+$\phi(W_{t_{1}},W_{t_{2}},\cdots,W_{t_{n}})=\mathbb{E}\phi(W_{t_{1}},W_{t_{2}},\cdots,W_{t_{n}})+\int_{0}^{T}H_{s}dW_{s},$
+there exists a constant $C$ such that $\int_{0}^{T}|H_{s}|^{2}ds\leq C$ a.s.
+Proof.
+Let us first prove for the case when $\phi(\omega)=\psi(\omega_{T})$ where $\psi$ is $\alpha$-Hölder continuous function from $\mathbb{R}^{d}$ to $\mathbb{R}$.
+Then, there exists a function $u(t,x)$ such that $u(T,x)=\psi(x)$ and $\partial_{t}u+\frac{1}{2}\partial^{2}_{xx}u=0$.
+Then, we have
+$\psi(W_{T})=u(0,0)+\int_{0}^{T}\partial_{x}u(t,W_{t})dW_{t}\text{ and }u(t,x)=\mathbb{E}\left[\psi(W_{T})|W_{t}=x\right].$
+Since
+$u(t,x)=\mathbb{E}\psi(x+W_{T}-W_{t})=\int\psi(x+y)f_{T-t}(y)dy=\int\psi(u)f_{T-t}(u-x)du,$
+when $f_{T-t}$ is the PDF of N(0,Tt)N(0,T-t), for $Z\sim N(0,T-t)$, we have
+$\partial_{x}u(t,x)$ $=\int\psi(u)\frac{u-x}{T-t}f_{T-t}(u-x)du=\int\psi(x+y)\frac{y}{T-t}f_{T-t}(y)dy$
+$=\mathbb{E}\left[\psi(x+Z)\frac{Z}{T-t}\right]=\mathbb{E}\left[\left(\psi(x+Z)-\psi(x)\right)\frac{Z}{T-t}\right]$
+$\leq\frac{K}{T-t}\mathbb{E}\left[|Z|^{\alpha+1}\right]=KC_{\alpha}(T-t)^{\frac{\alpha+1}{2}-1},$
+[…]
+$\int_{0}^{T}|H_{s}|^{2}ds=\int_{0}^{T}|\partial_{x}u(t,W_{t})|^{2}dt\leq K^{2}C_{\alpha}^{2}\int_{0}^{T}(T-t)^{\alpha-1}dt\leq K^{2}C_{\alpha}^{2}T^{\alpha}.$
+For general case, we can iterate the Step 1 backwards in time to obtain $L^{\infty}$ bound of $\int_{0}^{T}|H_{s}|^{2}ds$. ∎
 We are now in a position to state the main result of this section.
-For the reader’s convenience, we restate the main theorem below.
-To complete the proof of following theorem, it remains to extend the existence result from Malliavin differentiable terminal conditions to the general bounded case.
+For the reader’s convenience, we first restate the main theorem below in a simplified setting and then generalize the statement.
+To complete the proof of the following theorem, it remains to extend the existence result from Malliavin differentiable terminal conditions to the general bounded case.
 Uniqueness then follows directly from the stability result established in Theorem 4.6.
 Theorem 5.3.
-Let Assumptions $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q})$ hold. Then, there exists a unique solution $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ to the BSDE (2.6).
+Let Assumptions $(\mathrm{H}_{0})$} and $(\mathrm{H}_{q}^{\mathrm{unif}})$} hold. In addition, assume that
+$\xi=\phi(W_{t_{1}},W_{t_{2}},\cdots,W_{t_{n}})$
+for a bounded function $\phi:(\mathbb{R}^{d})^{n}\to\mathbb{R}$ which is $\alpha$-Hölder continuous, that is, there exists $K>0$ and $\alpha\in(0,1]$ such that
+$|\phi(x)-\phi(x^{\prime})|\leq K\max_{i=1,2,\cdots,n}|x_{i}-x^{\prime}_{i}|^{\alpha},$
+for any $x,x^{\prime}\in(\mathbb{R}^{d})^{n}$. Then, there exists a unique solution $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ to the BSDE (2.6).
 Proof.
-We begin by approximating the bounded terminal condition $\xi$ with a sequence of Malliavin differentiable random variables of the form $\xi^{n}=\Phi^{n}(W_{t_{1}},\dots,W_{t_{n}})$, where each $\Phi^{n}$ is a bounded continuous function.
-Since $\xi$ is bounded, the sequence $(\xi^{n})_{n\in\mathbb{N}}$ can be chosen such that $\xi^{n}$ converges to $\xi$ in in any $L^{p}$ space for $p\geq 1$.
-To ensure that the approximating BSDEs satisfy Assumption $(\mathrm{H}_{0})$, we truncate each $\Phi^{n}$ by setting $\Phi^{n}:=\Phi^{n}\wedge\left\|\xi\right\|_{L^{\infty}}$.
-Hence, each $\xi^{n}$ remains bounded, and $(\xi^{n},g)$ satisfies $(\mathrm{H}_{0})$ for all $n\in\mathbb{N}$.
+Let $\eta_{m}$ be a mollifier on $(\mathbb{R}^{d})^{n}$ with compact support $([-\frac{1}{m},\frac{1}{m}]^{d})^{n}$.
+Consider $\phi^{m}(x):=\eta_{m}*\phi$, which is Lipschitz.
+Then $\xi^{m}:=\phi^{m}(W_{t_{1}},W_{t_{2}},\cdots,W_{t_{n}})$ has bounded Malliavin derivatives, and Lemma 5.1 yields a corresponding solution $(Y^{m},Z^{m})$.
+Let $H^{m}$ be the martingale representation integrand of $\xi^{m}$, that is,
+$\xi^{m}=\mathbb{E}\xi^{m}+\int_{0}^{T}H^{m}_{s}dW_{s},$
+and show that $\sup_{m}\int_{0}^{T}|H^{m}_{s}|^{2}ds\leq C$ almost surely for some $C$.
+Since $\phi^{m}$ has the same Hölder constant as $\phi$ for any $m$, we can apply Lemma 5.2 to obtain such a constant $C$.
+Finally, Theorem 4.6 provides the convergence of $(Y^{m},Z^{m})$ to $(Y,Z)$ in $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$.
+Uniqueness follows from Theorem 4.6 as well.
+∎
+Theorem 5.4.
+Let Assumptions $(\mathrm{H}_{0})$} and $(\mathrm{H}_{q}^{\mathrm{sep}})$} hold.
+In addition, assume that there exists a function $\phi:C([0,T];\mathbb{R}^{d})\to\mathbb{R}$ such that $\xi=\phi(W_{[0,T]})$ and uniformly continuous in sup norm.
+Then, there exists a unique solution $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ to the BSDE (2.6).
+Proof.
+Let us consider the Moreau-Yosida approximation of $\phi$:
+$\phi^{n}(\omega):=\inf_{y\in C([0,T];\mathbb{R}^{d})}\left\{\phi(y)+n\sup_{t\in[0,T]}\left|y(t)-\omega(t)\right|\right\}.$
+Since $\inf_{y\in C([0,T];\mathbb{R}^{d})}\phi(y)\leq\phi^{n}(\omega)\leq\phi(\omega),$ $\phi^{n}$ is bounded. In addition, note that, for any $\omega,\omega^{\prime}$
+$\phi^{n}(\omega)$ $=\inf\left\{\phi(y)+n\sup_{t\in[0,T]}\left|y(t)-\omega(t)\right|\right\}\leq\inf\left\{\phi(y)+n\sup_{t\in[0,T]}\left|y(t)-\omega^{\prime}(t)\right|\right\}+n\sup_{t\in[0,T]}\left|\omega^{\prime}(t)-\omega(t)\right|$
+$\leq\phi^{n}(\omega^{\prime})+n\sup_{t\in[0,T]}\left|\omega^{\prime}(t)-\omega(t)\right|.$
+Therefore, $\phi^{n}$ is Lipschitz in sup norm.
+Lastly, since $\phi$ is uniformly continuous, for $\varepsilon>0$, there exists $N$ such that $\sup_{t\in[0,T]}\left|y(t)-\omega(t)\right|<\frac{\varepsilon}{2n}$ implies $|\phi(y)-\phi(\omega)|\leq\varepsilon/2$ for all $n\geq N$.
+Then,
+$|\phi^{n}(\omega)-\phi(\omega)|$ $=\left|\inf_{y\in C([0,T];\mathbb{R}^{d})}\left\{\phi(y)-\phi(\omega)+n\sup_{t\in[0,T]}\left|y(t)-\omega(t)\right|\right\}\right|<\varepsilon.$
+Therefore, $\phi^{n}(\omega)$ converges to $\phi(\omega)$ uniformly in $\omega$.By Proposition 3.2 of Cheridito and Nam (2014), $\xi^{n}:=\phi^{n}(W)$ has bounded Malliavin derivatives.
 By Lemma 5.1, each approximating BSDE with terminal condition $\xi^{n}$ admits a solution $(Y^{n},Z^{n})\in\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}$, satisfying
 $Y_{t}^{n}=\xi^{n}+\int_{t}^{T}\int_{\tilde{\Omega}}g(s,Z_{s}^{n},\tilde{Z}_{s}^{n})\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}^{n}\cdot dW_{s}.$ (5.29)
-Furthermore, the stability result in Theorem 4.6 implies that
-$\|Y^{n}-Y^{m}\|_{\mathbb{S}^{2p}}+\|Z^{n}-Z^{m}\|_{\mathbb{H}^{p}}\ \leq\ C_{p}\;\|\xi^{n}-\xi^{m}\|_{L^{2p}}$
-where $C_{p}$ depends only on $(T,K_{z},K_{\tilde{z}},K_{0})$, thus is independent of $n$ and $m$.
-Pick $(\xi^{n})_{n\in\mathbb{N}}$ is Cauchy in $L^{2p}$, it follows that $(Y^{n},Z^{n})_{n\in\mathbb{N}}$ is also Cauchy in $\mathbb{S}^{2p}\times\mathbb{H}^{p}$.
-As each $(Y^{n},Z^{n})$ lies in the Banach space $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$, we conclude the existence of a limit $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$.
-Passing to the limit in (5.28), we verify that the pair $(Y,Z)$ indeed solves the original mean-field BSDE (2.6), establishing existence.
-Uniqueness then follows directly from the stability result in Theorem 4.6, which concludes the proof.
+Moreover, since $\phi^{n}(\omega)$ converges to $\phi(\omega)$ uniformly in $\omega$, we have $\left\|\xi^{n}-\xi\right\|_{L^{\infty}}\to 0$.
+Then, by Theorem 4.7, we obtain the convergence of $(Y^{n},Z^{n})$ to $(Y,Z)$ in $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{BMO}$.
+The uniqueness follows from Theorem 4.7 again.
 ∎
-6 Application: Utility Maximization
-In this section, we provide an extension of the foundational work of hu2005utility to the mean-field case.
+Now let us generalize above theorem to the version we stated in Section 2.
+Theorem 5.5.
+Assume that there is a constant $C$ such that $b:[0,T]\times\mathbb{R}^{k}\to\mathbb{R}^{k}$ and $\sigma:[0,T]\to\mathbb{R}^{k\times n}$ are measurable functions with
+  • $|b(t,x)-b(t,x^{\prime})|\leq C|x-x^{\prime}|,|b(t,0)|\leq C$ for every $(t,x)\in[0,T]\times\mathbb{R}^{k}$, and
+  • $C^{-1}|v|^{2}\leq v^{\intercal}(\sigma\sigma^{\intercal})(t)v\leq C|v|^{2}$ for all $v\in\mathbb{R}^{k}$ and $t\in[0,T]$.
+Then, let $X$ be the strong solution of
+$dX_{t}=b(t,X_{t})dt+\sigma(t)dW_{t};\qquad X_{0}=x_{0}.$
+Then, we can replace $\phi(W_{[0,T]})$ to $\phi(X_{[0,T]})$ Theorem 5.4, and the statement still holds true.
+In addition, if $b(t,x)$ is uniformly bounded, we can replace $\phi((W_{t_{i}})_{i=1,2,\cdots,n}))$ to $\phi((X_{t_{i}})_{i=1,2,\cdots,n}))$ in Theorem 5.3 as well.
+Proof.
+First, let us consider Theorem 5.3.
+Without loss of generality, let us assume $k=1$: otherwise, we can use a dynamic change of coordinate to diagonalize $\sigma\sigma^{\intercal}(t)$ and treat each coordinate separately.
+First of all, note that $D_{t}X_{s}$ is bounded due to Lemma 4.2 of Cheridito and Nam (2014).
+Therefore, in the proof of Theorem 5.3,
+$\xi^{m}:=\phi^{m}(X_{t_{1}},X_{t_{2}},\cdots,X_{t_{n}})$
+has uniformly bounded Malliavin derivatives, which gives the corresponding solution $(Y^{m},Z^{m})$.
+If we show that $\int_{0}^{T}|H_{t}^{m}|^{2}ds<\infty$ for $H^{m}$ defined by the martingale representation of $\xi^{m}$, then Theorem 4.6 gives the existence and uniqueness of solution.
+Note that $H^{m}$ does not change under equivalent measure change.
+Let us define $\mathbb{Q}$ by
+$\left.\frac{d\mathbb{Q}}{d\mathbb{P}}\right|_{t}=\mathcal{E}_{t}\left(\int_{0}^{\cdot}(\sigma^{-1}b)(t,X_{t})^{\intercal}dW_{t}\right).$
+Then we can apply Girsanov’s theorem to conclude that $\mathbb{Q}$ is equivalent to $\mathbb{P}$ and that
+$B_{t}=W_{t}+\int_{0}^{t}\left(\sigma^{-1}b\right)(s,X_{s})ds$
+is $\mathbb{Q}$-Brownian motion.
+We have $dX_{t}=\sigma(t)dB_{t}$ and, under time change, $X_{t}=\tilde{B}_{\int_{0}^{t}|\sigma(s)|^{2}ds}$ for some Brownian motion $\tilde{B}$.
+Since $|\sigma|^{2}>C^{-1}>0$, there exist deterministic times $s_{i}$ such that $X_{t_{i}}=\tilde{B}_{s_{i}}$.
+Since
+$\xi^{m}=\phi^{m}(X_{t_{1}},X_{t_{2}},\cdots,X_{t_{n}})=\phi^{m}(\tilde{B}_{s_{1}},\tilde{B}_{s_{2}},\cdots,\tilde{B}_{s_{n}})$
+with $\int_{0}^{s_{n}}|\sigma(s)|^{2}ds=T$. This implies $s_{n}<\infty$ because $|\sigma|^{2}$ is bounded from below by $C^{-1}$. Using the Lemma 5.2, we obtain the uniform bound on $H^{m}$.
+On the other hand, let us consider Theorem 5.4 where $k$ can be an arbitrary natural number.
+Note that we only need to prove that, for Lipschitz $\phi:C([0,T];\mathbb{R})\to\mathbb{R}$, the Malliavin derivative of $\phi(X_{[0,T]})$ is bounded: if this statement is true, $\phi^{n}$ in the proof of Theorem 5.4 has bounded Malliavin derivatives, so we can use the same argument to prove the existence and uniqueness.
+Note that, if $\phi$ is Lipschitz, for linear interpolator $l^{m}:\mathbb{R}^{m}\to C([0,T];\mathbb{R})$ and $t_{i}=\frac{iT}{m}$, we have
+$\mathbb{E}\left[\left|\phi(X_{[0,T]})-(\phi\circ l^{m})(X_{t_{1}},X_{t_{2}},\cdots,X_{t_{m}})\right|^{p}\right]$ $\leq C^{2}\mathbb{E}\left[\max_{i=1,2,\cdots,m}\sup_{t\in(t_{i},t_{i+1}]}|X_{t}-X_{t_{i}}|^{p}\right],$
+for $p>2$. Note that we have
+$\sup_{t\in(t_{i},t_{i+1}]}|X_{t}-X_{t_{i}}|^{p}$ $\leq\sup_{t\in(t_{i},t_{i+1}]}\left|\int_{t_{i}}^{t}b(s,X_{s})ds+\int_{t_{i}}^{t}\sigma(s)dW_{s}\right|^{p}$
+$\leq 2^{p-1}\sup_{t\in(t_{i},t_{i+1}]}\left(\left|\int_{t_{i}}^{t}b(s,X_{s})ds\right|^{p}+\left|\int_{t_{i}}^{t}\sigma(s)dW_{s}\right|^{p}\right)$
+$\leq 2^{p-1}\left((t_{i+1}-t_{i})^{p-1}\int_{t_{i}}^{t_{i+1}}|b(s,X_{s})|^{p}ds+\sup_{t\in(t_{i},t_{i+1}]}\left|\int_{t_{i}}^{t}\sigma(s)dW_{s}\right|^{p}\right)$
+$\leq 2^{p-1}\left((2T)^{p-1}m^{-(p-1)}C^{p}\int_{0}^{T}(1+|X_{s}|^{2})ds+\sup_{t\in(t_{i},t_{i+1}]}\left|\int_{t_{i}}^{t}\sigma(s)dW_{s}\right|^{p}\right).$
+It is easy to see that, as $m\to\infty$,
+$\mathbb{E}\max_{i=1,2,\cdots,m}m^{-(p-1)}\int_{0}^{T}(1+|X_{s}|^{2})ds\leq\mathbb{E}m^{-(p-2)}\int_{0}^{T}(1+|X_{s}|^{2})ds\to 0$
+because $X\in\mathbb{H}^{2}$ from our assumption. On the other hand
+$\mathbb{E}\max_{i=1,2,\cdots,m}\sup_{t\in(t_{i},t_{i+1}]}\left|\int_{t_{i}}^{t}\sigma(s)dW_{s}\right|^{p}$ $\leq\sum_{i=1}^{m}\mathbb{E}\sup_{t\in(t_{i},t_{i+1}]}\left|\int_{t_{i}}^{t}\sigma(s)dW_{s}\right|^{p}\leq C_{p}\sum_{i=1}^{m}\mathbb{E}\left|\int_{t_{i}}^{t_{i+1}}|\sigma(s)|^{2}ds\right|^{p/2}$
+$\leq C_{p}\sum_{i=1}^{m}(t_{i+1}-t_{i})^{p/2-1}\int_{t_{i}}^{t_{i+1}}|\sigma(s)|^{p}ds$
+$\leq C_{p}\left(\frac{T}{m}\right)^{p/2-1}\int_{0}^{T}|\sigma(s)|^{p}ds\to 0,$
+as $m\to\infty$.
+Here, $C_{p}$ is the constant from the Burkholder-Davis-Gundy inequality.
+As a result, $(\phi\circ l^{m})(X_{t_{1}},X_{t_{2}},\cdots,X_{t_{m}})\to\phi(X_{[0,T]})$ in $L^{p}$.
+In addition, by (Nualart, 2006, Proposition 1.2.4) and the boundedness of $D_{t}X_{s}$, we know that $(\phi\circ l^{m})(X_{t_{1}},X_{t_{2}},\cdots,X_{t_{m}})$ has bounded Malliavin derivative.
+Therefore, by (Nualart, 2006, Lemma 1.2.3), we conclude that $\phi(X_{[0,T]})$ has bounded Malliavin derivative whenever $\phi$ is Lipschitz.
+∎
+6 Multidimensional MFBSDE with Small Terminal Condition and Utility Maximization
+In this section, we provide an extension of the foundational work of Hu et al.
+(2005) to the mean-field case.
 Specifically, we explore how a trader’s decision-making process evolves in a financial market influenced by collective behavior.
 Unlike previous studies that assume fixed constraints, we consider a dynamic setting where a trader’s strategy is shaped not only by individual preferences but also by the actions of others.
-This interaction leads to a more flexible and realistic model, where part of the trading constraints are given by market conditions rather than being imposed arbitrarily.
-This section is structured as follows: In Section 6.1, we introduce the utility maximization problem and the associated mathematical framework.
-We then extend the given constraint set to incorporate a dependence on the stochastic process.
-It is important to note that, at this stage, we do not make any assumptions about this process beyond considering it as a general stochastic process.
-Next, we establish the connection between the utility maximization problem and the BSDE (6.30) with a thoughtfully chosen driver.
-We will subsequently demonstrate that the process defined earlier for the constraint set is precisely the solution $Z$ in the BSDE (6.30).
+This interaction leads to a more flexible and realistic model, in which part of the trading constraints are determined by market conditions rather than imposed arbitrarily.
+This section is structured as follows.
+In Section 6.1, we introduce the utility maximization problem and the associated mathematical framework.
+We then extend the constraint set to incorporate dependence on a stochastic process.
+At this stage, we do not impose any assumptions on this process beyond considering it as a general stochastic process.
+Next, we establish the connection between the utility maximization problem and the BSDE (6.31) with a suitably chosen driver.
+We then show that the process used to define the constraint set is precisely the solution component $Z$ in BSDE (6.31).
 With the complete framework in place, we offer a financial interpretation of the components, which is discussed in Remarks 6.4.
-To connect this application to our earlier analysis about BSDE, we generalize the separately quadratic driver from BSDE (2.6) to a fully coupled quadratic driver in BSDE (6.32) and establish the existence and uniqueness of solutions in Proposition 6.7 when the terminal condition is small enough.
-To do this, we adopt a localization method in 6.6, then we extend it to the global case.
-Then, after we established the existence and uniqueness for BSDE (6.32), we then connect this back to the utility maximization problem by presenting the final result in Theorem 6.9.
-We first derive a key estimate in Lemma 6.3.
-This estimate is then used in Theorem 6.9 to verify that the driver of BSDE (6.30) satisfies conditions $(\rm H_{0})$} and $(\rm H_{q}^{\prime})$}, ensuring the well-posedness of BSDE (6.32).
-Leveraging these lemmas, we conclude with the main result in Theorem 6.9, which is supported by the proof of Theorem 7 in hu2005utility.
+To relate this application to our theoretical results, we move beyond the structural assumptions on the terminal condition used in Sections 45.
+The driver in (6.32) involves the distance to a constraint set that depends on the aggregated signal $\mathbb{E}[Z]$, and its quadratic growth is generally fully coupled in $(Z,\mathbb{E}[Z])$.
+Consequently, we fit (6.31) into the unified frameworks used earlier via a different representation.
+We introduce the fully coupled BSDE (6.33) and prove a dedicated well-posedness result under a small terminal oscillation assumption.
+More precisely, we establish local existence and uniqueness for the centered BSDE (6.34) on short time intervals in Lemma 6.6 via a contraction argument, and then obtain a global solution on $[0,T]$ by a concatenation procedure in Proposition 6.7.
+The smallness requirement is imposed on the centered terminal condition $\eta:=\xi-\mathbb{E}[\xi]$, which is natural in the utility maximization setting where the terminal liability need not be small but may be close to its mean.
+With this well-posedness result in hand, we return to the utility maximization problem.
+We first derive an elementary growth estimate for the (volatility-scaled) constraint set in Lemma 6.3.
+This estimate is used in Theorem 6.9 to verify that the driver $f$ in (6.32) satisfies $(\mathrm{H}_{0})$ and falls within the fully coupled quadratic class covered by Proposition 6.7.
+Finally, we conclude that the BSDE (6.31) admits a unique solution, and we identify the value function and an optimal strategy by adapting the verification argument in Theorem 7 of Hu et al.
+(2005) to the present mean-field, market-adaptive constraint setting.
 6.1 Financial market model setup
@@
 $\frac{\mathrm{d}S_{t}^{i}}{S_{t}^{i}}=b_{t}^{i}\,\mathrm{d}t+\sigma_{t}^{i}\,\mathrm{d}W_{t},\quad i=1,\dots,d,$
-where $W$ is $n$-dimensional Brownian motion, $b^{i}$ is a real-valued predictable, bounded stochastic process, representing the drift, and $\sigma^{i}$ is an $\mathbb{R}^{n}$-valued predictable bounded stochastic process, representing the volatility.
-The volatility matrix $\sigma_{t}=(\sigma_{t}^{1},\dots,\sigma_{t}^{d})^{\top}$ is assumed to have full rank $d$.
-And we assume that $\sigma_{t}\sigma_{t}^{\top}$ is uniformly elliptic.
-Thus market price of risk $\theta_{t}$, also referred to as the risk premium, is defined by:
+where $W$ is an $n$-dimensional Brownian motion, $b^{i}$ is a real-valued predictable, bounded stochastic process representing the drift, and $\sigma^{i}$ is an $\mathbb{R}^{n}$-valued predictable, bounded stochastic process representing the volatility.
+The volatility matrix $\sigma_{t}=(\sigma_{t}^{1},\dots,\sigma_{t}^{d})^{\top}$ is assumed to have full rank $d$, and we assume that $\sigma_{t}\sigma_{t}^{\top}$ is uniformly elliptic.
+Thus, the market price of risk $\theta_{t}$, also referred to as the risk premium, is defined by
 $\theta_{t}:=\sigma_{t}^{\top}(\sigma_{t}\sigma_{t}^{\top})^{-1}b_{t},\quad t\in[0,T],$
@@
 $U(x)=-\exp(-\alpha x),\quad x\in\mathbb{R}.$
-For the investors other than the representative, we assume the risk aversion parameter $\alpha>0$ is identical.
-The model for the risky assets is of the same type but with $b,\sigma,W$ being independent and identically distributed to the representative’s model.
-In addition, we assume the terminal liabilities are independent and have distribution $F$.
+For investors other than the representative, we assume that the risk aversion parameter $\alpha>0$ is identical.
+The model for the risky assets is of the same type, but with $b$, $\sigma$ and $W$ being independent and identically distributed relative to the representative’s model.
+In addition, we assume that the terminal liabilities are independent and have distribution $F$.
 6.2 Constraint set and aggregated market signal
@@
 Below, we outline the constraint set in Definition 6.1, which accounts for market-wide influences.
-To incorporate the influence of market-wide behavior, for $t\in[0,T]$ and $\omega\in\Omega$, we define two stochastic processes that satisfy the necessary integrability conditions such that
+To incorporate the influence of market-wide behavior, for $t\in[0,T]$ and $\omega\in\Omega$, we define two stochastic processes satisfying the necessary integrability conditions such that
 $Z_{t}(\omega):=\hat{Z_{t}}(\omega)\sigma_{t}(\omega),$
 where $Z_{t}(\omega)\in\mathbb{R}^{1\times n}$ and $\hat{Z_{t}}(\omega)\in\mathbb{R}^{1\times d}$.
-These two stochastic processes are essentially equivalent because the matrix $\sigma$ has full rank.
-Later on, in Remark 6.4, we will see $\hat{Z}_{t}\sigma_{t}+\frac{\theta_{t}}{\alpha}=Z_{t}+\frac{\theta_{t}}{\alpha}$ yields the optimal strategy at time $t\in[0,T]$ if no constants are imposed.
-Thus $Z$ can be interpreted as the trading strategy adjusted by $\theta/\alpha$.
-Later, we will show that this $Z$ will precisely correspond to the martingale density process of BSDE (6.30) as well.
-We now present the following expressions for the constraint set under different settings.
+These two processes are essentially equivalent because the matrix $\sigma$ has full rank.
+Later, in Remark 6.4, we will see that $\hat{Z}_{t}\sigma_{t}+\frac{\theta_{t}}{\alpha}=Z_{t}+\frac{\theta_{t}}{\alpha}$ yields the optimal strategy at time $t\in[0,T]$ if no constraints are imposed.
+Thus, $Z$ can be interpreted as the trading strategy adjusted by $\theta/\alpha$.
+Later, we will show that this $Z$ corresponds precisely to the martingale density process of BSDE (6.31).
+We now present expressions for the constraint set under different settings.
 Specifically, we denote the constraint without market influence by $\hat{D}\subseteq\mathbb{R}^{d}$, where $\hat{D}$ is a closed set.
@@
 where $\gamma:=(\gamma_{1},\dots,\gamma_{d})$ is the market sensitivity vector and $\eta\in\mathbb{R}^{1\times d}$ is the market shift vector, with both being constant vectors.
-With the help of the previous definitions, for a fixed $t\in[0,T]$, we start with a set $\hat{D}\subset\mathbb{R}^{d}$, which represents certain constraints, such as risk limits or position constraints.
-Upon incorporating information from the market, we could apply scaling and shifting to this set, which now reflects both the original constraints and the market’s influence.
-This results in a new set that is much more flexible and realistic, capturing the dynamic behavior of the market.
-Therefore, we denote this updated set as $\hat{C}(\mathbb{E}[\hat{Z}_{t}])$, emphasizing its dependence on the aggregated market signal $\mathbb{E}[\hat{Z_{t}}$] at time $t\in[0,T]$.
+With the help of the previous definitions, for a fixed $t\in[0,T]$, we start with a set $\hat{D}\subset\mathbb{R}^{d}$, which represents certain constraints such as risk limits or position constraints.
+After incorporating information from the market, we apply scaling and shifting to this set so that it reflects both the original constraints and the market’s influence.
+This results in a new set that is more flexible and realistic, capturing the dynamic behavior of the market.
+We denote this updated set by $\hat{C}(\mathbb{E}[\hat{Z}_{t}])$, emphasizing its dependence on the aggregated market signal $\mathbb{E}[\hat{Z}_{t}]$ at time $t\in[0,T]$.
 Definition 6.2 (Admissible strategies with constraints $\hat{\mathcal{A}}$).
@@
 $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right):=\hat{C}(\mathbb{E}[\hat{Z}_{t}])\,\sigma_{t}(\omega)\quad\text{and}\quad p_{t}:=\pi_{t}\sigma_{t}.$
-Note $Z_{t}=\hat{Z_{t}}\sigma_{t}$ and $\mathbb{E}[Z_{t}]=\mathbb{E}[\hat{Z_{t}}]\sigma_{t}$.
+Note that $Z_{t}=\hat{Z}_{t}\,\sigma_{t}$ and $\mathbb{E}[Z_{t}]=\mathbb{E}[\hat{Z}_{t}]\,\sigma_{t}$.
 Lemma 6.3.
@@
 With the help of these notations, we can rewrite the problem as
-$\sup_{p\in\mathcal{A}}\mathbb{E}\left[-\exp\left(-\alpha\left(x+\int_{0}^{T}p_{t}(\mathop{}\!dW_{t}+\theta_{t}\mathop{}\!dt)-F\right)\right)\right]=\mathbb{E}\left[-\exp(-\alpha(X_{T}^{p^{*}}-F))\right]$
+$\sup_{p\in\mathcal{A}}\mathbb{E}\left[-\exp\left(-\alpha\left(x+\int_{0}^{T}p_{t}(\mathop{}\!dW_{t}+\theta_{t}\mathop{}\!dt)-F\right)\right)\right]=\mathbb{E}\left[-\exp(-\alpha(X_{T}^{p^{*}}-F))\right],$
 for the optimal control $p^{*}\in\mathcal{A}$. Let us introduce a family of processes $\left\{V^{(p)}(\cdot,x):p\in\mathcal{A}\right\}$ as follows: let
 $V^{(p)}(t,x):=-\exp(-\alpha(X_{t}^{p}-Y_{t})),\quad t\in[0,T],\,p\in\mathcal{A},$
-where $(Y,Z)$ is a solution of the BSDE
-$Y_{t}=F+\int_{t}^{T}f(s,Z_{s},\mathbb{E}[Z_{s}])\,ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},\quad t\in[0,T],.$ (6.30)
-Here, the driver $f$ to be chosen such that $V^{(p)}(\cdot,x)$ is a supermartingale for all $p\in\mathcal{A}$, and there exists $p^{*}\in\mathcal{A}$ such that $V^{(p^{*})}(\cdot,x)$ is a true martingale.
-If we are able to choose such $f$, which will be shown later, we have
-$\mathbb{E}V^{(p)}(T,x)\leq V^{(p)}(0,x)=V^{(p^{*})}(0,x)=\mathbb{E}V^{(p^{*})}(T,x).$
+where $(Y,Z)$ is a solution of the MFBSDE
+$Y_{t}=F+\int_{t}^{T}f(s,Z_{s},\mathbb{E}[Z_{s}])\,ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},\quad t\in[0,T].$ (6.31)
+Here, the driver $f$ will be chosen so that $V^{(p)}(\cdot,x)$ is a supermartingale for all $p\in\mathcal{A}$, and there exists $p^{*}\in\mathcal{A}$ such that $V^{(p^{*})}(\cdot,x)$ is a true martingale.
+If we can choose such an $f$ (as we show below), then
+$\mathbb{E}V^{(p)}(T,x)\leq V^{(p)}(0,x)=V^{(p^{*})}(0,x)=\mathbb{E}V^{(p^{*})}(T,x),$
 for all $p\in\mathcal{A}$ and the optimal control $p^{*}\in\mathcal{A}$.
 Earlier, we described the process $(Z_{t})_{0\leq t\leq T}$ simply as a stochastic process.
-However, in this context, we now define it as the martingale density process $(Z_{t})_{0\leq t\leq T}$ for the BSDE in equation (6.30).
-Consequently, as the constraint set depends on $\mathbb{E}[Z_{t}]$, the driver should has dependence on $\mathbb{E}[Z_{t}]$ as well: see (6.31).
-To verify this, let us specify the driver $f$ for the construction of $\left\{V^{(p)}(\cdot,x):p\in\mathcal{A}\right\}$.
+However, in this context, we now define it as the martingale density process $(Z_{t})_{0\leq t\leq T}$ for the BSDE in (6.31).
+Consequently, since the constraint set depends on $\mathbb{E}[Z_{t}]$, the driver should also depend on $\mathbb{E}[Z_{t}]$; see (6.32).
+To verify this, we now specify the driver $f$ for the construction of $\left\{V^{(p)}(\cdot,x):p\in\mathcal{A}\right\}$.
 By Itô’s formula,
@@
 is a true martingale as $\alpha(Z-p)$ is square integrable.
-We require $\tilde{A}^{p}$ to be a non-increasing process such that $v(t,p_{t},Z_{t})\geq 0$ for all $p\in\mathcal{A}$, and there exists a $p^{*}\in\mathcal{A}$ such that $v(t,p^{*}_{t},Z_{t})=0$.
+We require $\tilde{A}^{p}$ to be non-increasing, which holds if $v(t,p_{t},Z_{t})\geq 0$ for all $p\in\mathcal{A}$, and there exists a $p^{*}\in\mathcal{A}$ such that $v(t,p^{*}_{t},Z_{t})=0$.
 Therefore, if we choose
@@
 $=\frac{\alpha}{2}\left|p_{t}-\left(Z_{t}+\frac{\theta_{t}}{\alpha}\right)\right|^{2}-\frac{\alpha}{2}\text{dist}^{2}\left(Z_{t}+\frac{\theta_{t}}{\alpha},C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)\right)\geq 0.$
-Thus, a family of processes $\left\{V^{(p)}(\cdot,x):p\in\mathcal{A}\right\}$ becomes a colletion of supermartingales where, for $p_{t}^{*}\in\Pi_{C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)}\left(Z_{t}+\frac{\theta_{t}}{\alpha}\right)$, we have $v(\cdot,p^{*},Z)=0$ for every $t\in[0,T]$.
-Therefore, $V^{(p^{*})}(\cdot,x)$ becomes a true martingale.
-We have now completed the construction, and it is clear how the utility maximization problem in our setting connects to a type of mean-field type BSDEs as in hibon2017MFquadratic.
-In hu2005utility, the optimization problem was solved under the assumption that the random endowment $F$ is bounded and that the trading strategy $p$ takes values in the admissible set $\mathcal{A}$.
-However, this set differs from ours because, in their case, the constraint set for trading strategies is simply closed in $\mathbb{R}^{d}$.
-In contrast, we introduce a particular constraint set where the trader’s behavior depends on the market (i.e., the behavior of other traders) in a mean-field sense.
-Specifically, we expand the driver to include the expectation of the control process $Z$.
-Thus, we define a constraint set that depends on time, reflecting the market’s influence at each moment.
-This approach introduces greater flexibility and a more realistic model.
-To establish the connection between the control process $Z$ in the BSDE (6.30) and the trading strategy $p\in\mathcal{A}$, we provide the following remarks, which bridge the mathematical framework with financial interpretations.
+Thus, the family of processes $\left\{V^{(p)}(\cdot,x):p\in\mathcal{A}\right\}$ becomes a collection of supermartingales.
+Moreover, for $p_{t}^{*}\in\Pi_{C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)}\left(Z_{t}+\frac{\theta_{t}}{\alpha}\right)$, we have $v(\cdot,p^{*},Z)=0$ for every $t\in[0,T]$, hence $V^{(p^{*})}(\cdot,x)$ becomes a true martingale.
+We have now completed the construction, and it is clear how the utility maximization problem in our setting connects to a class of mean-field quadratic BSDEs (cf.
+Hibon et al.
+(2023)).
+In Hu et al.
+(2005), the optimization problem is solved under the assumptions that the random endowment $F$ is bounded and that the trading strategy $p$ takes values in an admissible set $\mathcal{A}$.
+However, their admissible set differs from ours: in their case, the constraint set for trading strategies is simply closed in $\mathbb{R}^{d}$.
+In contrast, we introduce a constraint set in which the trader’s behavior depends on the market (i.e., the behavior of other traders) in a mean-field sense.
+In particular, the driver depends on the expectation of the control process $Z$, and the constraint set becomes time-dependent, reflecting the market’s influence at each moment.
+To connect the control process $Z$ in BSDE (6.31) with the trading strategy $p\in\mathcal{A}$, we provide the following remarks bridging the mathematical framework with financial interpretations.
 Remark 6.4.
@@
 Thus, the process $Z_{t}=p_{t}^{*}-\frac{\theta_{t}}{\alpha}$ can be interpreted as the unconstrained trading strategy adjusted by $\frac{\theta_{t}}{\alpha}$ at time $t\in[0,T]$.
-In general situation, the constraint set $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)$ depends on the average of the adjusted trading strategies among the market participants.
+In general, the constraint set $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)$ depends on the average of the adjusted trading strategies among the market participants.
 The convergence is guaranteed by the homogeneity and exchangeability of market participants.
@@
 We are now ready to present the final result.
-We first borrow some ideas from tevzadze2008solvability, under small terminal value and lipschitz coefficient, the mean-field BSDE 6.32 admits a unique adapted solution.
-Next, in Proposition 6.7 we demonstrate that the driver in BSDE 6.30 guarantees the existence and uniqueness of solution.
-Finally, we complete this section by presenting the maximum expected utility value and corresponding trading strategy in in Theorem 6.9.
-For the purpose of the final result, we introduce the BSDE (6.32) with the same formulation and the same $(\rm H_{0})$} condition in the main result but with a fully coupled quadratic growth condition and assume the following BSDE with driver $g$ and terminal condition $\xi$, satisfying the conditions specified below:
+We first adapt ideas from Tevzadze (2008): under a smallness condition on the centred terminal value and a (local) Lipschitz component, the fully coupled mean-field BSDE (6.33) admits a unique solution.
+We then globalize this local result by a localization/concatenation procedure in Lemma 6.6 and Proposition 6.7.
+Finally, we return to the utility maximization problem and identify both the value function and an optimal strategy in Theorem 6.9.
+For the purpose of the final result, we introduce the BSDE (6.33) with the same formulation and the same $(\mathrm{H}_{0})$ condition in the main result but with a different representation and assume the following BSDE with driver $g$ and terminal condition $\xi$, satisfying the conditions specified below:
 $Y_{t}=\xi+\int_{t}^{T}\int_{\tilde{\Omega}}g(s,Z_{s},\tilde{Z}_{s})\mathop{}\!d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}\cdot\mathop{}\!dW_{s},\quad t\in[0,T].$ (6.33)
-  • $(\rm H_{q}^{\prime})$}
-    There exist constants $L_{z}>0$, and $K_{z}>0$ such that:
+  • $(\mathrm{H}_{q}^{\mathrm{LK}})$}
+    There exist constants $L_{z}>0$, and $K_{q}>0$ such that:
 […]
     for any $(t,z,\tilde{z},z^{\prime},\tilde{z}^{\prime})\in[0,T]\times[\mathbb{R}^{d}\times\mathbb{R}^{d}]^{2}$.
-This is the fully coupled quadratic case, in contrast with the separately quadratic $(\rm H_{q})$} of the previous sections. We provide an explicit example here.
-Example 6.5.
-For simplicity, consider the one-dimensional case $d=1$ and define
-$g(t,z,\tilde{z}):=z\,\tilde{z},\qquad(t,z,\tilde{z})\in[0,T]\times\mathbb{R}^{2}.$
-This $g$ satisfies the fully coupled quadratic condition $(\rm H_{q}^{\prime})$}, but does not satisfy the separately quadratic condition $(\rm H_{q})$}.
+This is again the fully coupled quadratic case, that is, $(\mathrm{H}_{q}^{\mathrm{unif}})$} and $(\mathrm{H}_{q}^{\mathrm{LK}})$} are equivalent up to a change of constants.
 We begin by establishing existence and uniqueness of solutions on a sufficiently small time interval in Lemma 6.6, where the required smallness conditions for the terminal value and time horizon are satisfied.
 By concatenating these local solutions, we extend the result to the global interval $[0,T]$ in Proposition 6.7.
-First, we note that the terminal condition in BSDE 6.30 is $F$, representing a random liability.
+First, we note that the terminal condition in BSDE 6.31 is $F$, representing a random liability.
 It is not realistic to assume $F$ is close to zero.
 Instead, we relax this requirement and require that $F$ is close to its mean.
-Accordingly, we rewrite the general BSDE 6.32 in terms of the centered random variable $\eta$:
+Accordingly, we rewrite the general BSDE 6.33 in terms of the centered random variable $\eta$:
 $U_{t}=\eta+\int_{t}^{T}\int_{\tilde{\Omega}}\big[g(s,Z_{s},\tilde{Z}_{s})-g(s,0,0)\big]\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},\qquad 0\leq t\leq T,$ (6.34)
 where
-$U_{t}:=Y_{t}-g(t,0,0)-\mathbb{E}[\xi],\qquad\eta:=\xi-\mathbb{E}[\xi]$
-are both well-defined because of $(\rm H_{0})$}.
+$U_{t}:=Y_{t}-\int_{0}^{t}g(s,0,0)ds-\mathbb{E}[\xi]\text{ and }\eta:=\xi-\mathbb{E}[\xi]$
+are both well-defined because of $(\mathrm{H}_{0})$.
+Remark 6.5.
+When $b$ and $\sigma$ are deterministic, (6.34) has a unique solution if we assume the terminal condition $\xi$ satisfies the condition described in Theorem 2.6.
+In the following, we give the results when $b$ and $\sigma$ are not deterministic. In this case, we need smallness in the spread of the terminal condition from its mean.
 Lemma 6.6.
-Let $T<(64L_{z}^{2})^{-1}$ and assume that the driver $g$ satisfies $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q}^{\prime})$ and terminal condition $\eta$ be a bounded terminal condition such that
+Let $T<(64L_{z}^{2})^{-1}$ and assume that the driver $g$ satisfies $(\mathrm{H}_{0})$} and $(\mathrm{H}_{q}^{\mathrm{LK}})$ and terminal condition $\eta$ be a bounded terminal condition such that
 $\left\|\eta\right\|_{L^{\infty}}\leq\rho,$
 where
-$\rho^{2}:=\frac{1-64L_{z}^{2}T}{65K_{z}}.$
-Then the BSDE 6.33 admits a unique solution $(U,Z)\in\mathcal{B}_{\rho}$, where
+$\rho^{2}:=\frac{1-64L_{z}^{2}T}{65K_{q}}.$
+Then the BSDE 6.34 admits a unique solution $(U,Z)\in\mathcal{B}_{\rho}$, where
 $\mathcal{B}_{\rho}:=\left\{(U,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}:\left\|U\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|Z\right\|_{BMO}^{2}\leq\rho^{2}\right\}.$
@@
 Step 1. We first establish existence and uniqueness for the auxiliary BSDE:
-$U_{t}=\eta+\int_{t}^{T}\int_{\tilde{\Omega}}g(s,Z_{s},\tilde{Z}_{s})-g(s,0,0)\,d\tilde{\mathbb{P}}ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},\qquad 0\leq t\leq T.$
+$U_{t}=\eta+\int_{t}^{T}\int_{\tilde{\Omega}}\big[g(s,Z_{s},\tilde{Z}_{s})-g(s,0,0)\big]\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},\qquad 0\leq t\leq T.$
 For each $(u,z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$, with $\tilde{z}$ a copy of $z$, define $(U,Z)=\Gamma(u,z)$ to be the solution to
-$U_{t}=\eta+\int_{t}^{T}\int_{\tilde{\Omega}}g(s,z_{s},\tilde{z}_{s})-g(s,0,0)\,d\tilde{\mathbb{P}}ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s}.$
+$U_{t}=\eta+\int_{t}^{T}\int_{\tilde{\Omega}}\big[g(s,z_{s},\tilde{z}_{s})-g(s,0,0)\big]\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s}.$
 Applying Itô’s formula to $|U_{t}|^{2}$ on $[t,T]$, we obtain
@@
 Taking the conditional expectation and applying $2ab\leq\frac{1}{2}a^{2}+2b^{2}$, we deduce
-$|U_{t}|^{2}+\mathbb{E}\left[\int_{t}^{T}|Z_{s}|^{2}ds\mid\mathcal{F}_{t}\right]$ $\leq\left\|\eta\right\|_{L^{\infty}}^{2}+\frac{1}{2}\left\|U\right\|_{\mathbb{S}^{\infty}}^{2}+2\mathbb{E}\left[\int_{t}^{T}\int_{\tilde{\Omega}}|g(s,z_{s},\tilde{z}_{s})-g(s,0,0)|d\tilde{\mathbb{P}}ds\Big|\mathcal{F}_{t}\right]^{2}.$ (6.34)
-Together with the $(\rm H_{q}^{\prime})$} condition and Jensen’s inequality, we have
-$\left|\int_{\tilde{\Omega}}g(s,z_{s},\tilde{z}_{s})-g(s,0,0)d\tilde{\mathbb{P}}\right|\leq L_{z}\big(|z_{s}|+\mathbb{E}|z_{s}|\big)+2K_{z}\big(|z_{s}|^{2}+\mathbb{E}|z_{s}|^{2}\big).$
+$|U_{t}|^{2}+\mathbb{E}_{t}\left[\int_{t}^{T}|Z_{s}|^{2}ds\right]$ $\leq\left\|\eta\right\|_{L^{\infty}}^{2}+\frac{1}{2}\left\|U\right\|_{\mathbb{S}^{\infty}}^{2}+2\mathbb{E}_{t}\left[\int_{t}^{T}\int_{\tilde{\Omega}}|g(s,z_{s},\tilde{z}_{s})-g(s,0,0)|d\tilde{\mathbb{P}}ds\right]^{2}.$ (6.35)
+Together with the $(\mathrm{H}_{q}^{\mathrm{LK}})$} condition and Jensen’s inequality, we have
+$\left|\int_{\tilde{\Omega}}\big[g(s,z_{s},\tilde{z}_{s})-g(s,0,0)\big]\,d\tilde{\mathbb{P}}\right|\leq L_{z}\big(|z_{s}|+\mathbb{E}|z_{s}|\big)+2K_{q}\big(|z_{s}|^{2}+\mathbb{E}|z_{s}|^{2}\big).$
 Thus,
-$\mathbb{E}\left[\int_{t}^{T}\left|\int_{\tilde{\Omega}}g(s,z_{s},\tilde{z}_{s})-g(s,0,0)d\tilde{\mathbb{P}}\right|ds\mid\mathcal{F}_{t}\right]$ (6.35)
-$\leq\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{z}|z_{s}|^{2}ds\mid\mathcal{F}_{t}\right]+\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{z}|z_{s}|^{2}ds\right].$
+$\mathbb{E}_{t}\left[\int_{t}^{T}\big|\int_{\tilde{\Omega}}g(s,z_{s},\tilde{z}_{s})-g(s,0,0)d\tilde{\mathbb{P}}\big|ds\right]$ (6.36)
+$\leq\mathbb{E}_{t}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{q}|z_{s}|^{2}ds\right]+\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{q}|z_{s}|^{2}ds\right].$
 Combining (6.35) and (6.36) gives
 $\frac{1}{2}\left\|U\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|Z\right\|_{BMO}^{2}$
-$\leq\left\|\eta\right\|_{L^{\infty}}^{2}+2\operatorname*{ess\,sup}_{(t,\omega)\in[0,T]\times\Omega}\left[\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{z}|z_{s}|^{2}ds\mid\mathcal{F}_{t}\right]+\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{z}|z_{s}|^{2}ds\right]\right]^{2}$
-$\leq\left\|\eta\right\|_{L^{\infty}}^{2}+4\operatorname*{ess\,sup}_{(t,\omega)\in[0,T]\times\Omega}\left(\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{z}|z_{s}|^{2}ds\mid\mathcal{F}_{t}\right]^{2}+\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{z}|z_{s}|^{2}ds\right]^{2}\right)$
-$\leq\left\|\eta\right\|_{L^{\infty}}^{2}+16\left(L_{z}^{2}T\left\|z\right\|_{BMO}^{2}+4K_{z}^{2}\left\|z\right\|_{BMO}^{4}\right).$
-The last inequality uses Jensen’s inequality:
-$\left(\int_{t}^{T}L_{z}|z_{s}|\,ds\right)^{2}\leq(T-t)\left(\int_{t}^{T}L_{z}^{2}|z_{s}|^{2}\,ds\right).$
+$\leq\left\|\eta\right\|_{L^{\infty}}^{2}+2\operatorname*{ess\,sup}_{(t,\omega)\in[0,T]\times\Omega}\left[\mathbb{E}_{t}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{q}|z_{s}|^{2}ds\right]+\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{q}|z_{s}|^{2}ds\right]\right]^{2}$
+$\leq\left\|\eta\right\|_{L^{\infty}}^{2}+4\operatorname*{ess\,sup}_{(t,\omega)\in[0,T]\times\Omega}\left(\mathbb{E}_{t}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{q}|z_{s}|^{2}ds\right]^{2}+\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{q}|z_{s}|^{2}ds\right]^{2}\right)$
+$\leq\left\|\eta\right\|_{L^{\infty}}^{2}+16\left(L_{z}^{2}T\left\|z\right\|_{BMO}^{2}+4K_{q}^{2}\left\|z\right\|_{BMO}^{4}\right).$
+The last inequality uses Jensen’s inequality: $\left(\int_{t}^{T}L_{z}|z_{s}|\,ds\right)^{2}\leq(T-t)\left(\int_{t}^{T}L_{z}^{2}|z_{s}|^{2}\,ds\right)$.
 Now, using $a^{2}+b^{2}\leq(|a|+|b|)^{2}$, it follows that
@@
 $\leq 4\left\|\eta\right\|_{L^{\infty}}^{2}+\alpha^{2}(\left\|u\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|z\right\|_{BMO}^{2})+\beta^{2}(\left\|u\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|z\right\|_{BMO}^{2})^{2},$
-where $\alpha:=8L_{z}\sqrt{T}$, $\beta:=16K_{z}$. To obtain a uniform bound, set $R>0$ such that
+where $\alpha:=8L_{z}\sqrt{T}$, $\beta:=16K_{q}$. To obtain a uniform bound, set $R>0$ such that
 $4\left\|\eta\right\|_{L^{\infty}}^{2}+\alpha^{2}R^{2}+\beta^{2}R^{4}\leq R^{2},$
@@
 $\mathcal{B}_{R}:=\left\{(U,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{BMO}:\left\|U\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|Z\right\|_{BMO}^{2}\leq R^{2}\right\}$
-is then such that $\Gamma(\mathcal{B}_{R})\subset\mathcal{B}_{R}$.
+satisfies $\Gamma(\mathcal{B}_{R})\subset\mathcal{B}_{R}$.
 Step 2. Next, we prove that $\Gamma$ is a contraction on $\mathcal{B}_{R}$. Let $(u^{i},z^{i})\in\mathcal{B}_{R}$ ($i=1,2$), with $(U^{i},Z^{i})=\Gamma(u^{i},z^{i})$, and set
@@
 $\frac{1}{2}\left\|\Delta U\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|\Delta Z\right\|_{BMO}^{2}$
-$\leq 2\,\operatorname*{ess\,sup}_{(t,\omega)\in[0,T]\times\Omega}\left[\mathbb{E}\left(\int_{t}^{T}|g(s,z^{1}_{s},\tilde{z}^{1}_{s})-g(s,z^{2}_{s},\tilde{z}^{2}_{s})|ds\mid\mathcal{F}_{t}\right)\right]^{2}$
-$\leq 4L_{z}^{2}T\,\operatorname*{ess\,sup}_{(t,\omega)}\mathbb{E}\left[\int_{t}^{T}|\Delta z_{s}|^{2}+\mathbb{E}|\Delta z_{s}|^{2}ds\mid\mathcal{F}_{t}\right]$
-$\quad+16K_{z}^{2}\,\operatorname*{ess\,sup}_{(t,\omega)}\mathbb{E}\left[\int_{t}^{T}(|z^{1}_{s}|^{2}+|z^{2}_{s}|^{2}+\mathbb{E}|z^{1}_{s}|^{2}+\mathbb{E}|z^{2}_{s}|^{2})ds\mid\mathcal{F}_{t}\right]\mathbb{E}\left[\int_{t}^{T}(|\Delta z_{s}|+\mathbb{E}|\Delta z_{s}|)^{2}ds\mid\mathcal{F}_{t}\right]$
-$\leq\left(8L_{z}^{2}T+32K_{z}^{2}(\left\|z^{1}\right\|_{BMO}^{2}+\left\|z^{2}\right\|_{BMO}^{2})\right)\left(\left\|\Delta u\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|\Delta z\right\|_{BMO}^{2}\right).$
+$\leq 2\,\operatorname*{ess\,sup}_{(t,\omega)\in[0,T]\times\Omega}\left[\mathbb{E}_{t}\left(\int_{t}^{T}|g(s,z^{1}_{s},\tilde{z}^{1}_{s})-g(s,z^{2}_{s},\tilde{z}^{2}_{s})|ds\right)\right]^{2}$
+$\leq 4L_{z}^{2}T\,\operatorname*{ess\,sup}_{(t,\omega)}\mathbb{E}_{t}\left[\int_{t}^{T}|\Delta z_{s}|^{2}+\mathbb{E}|\Delta z_{s}|^{2}ds\right]$
+$\quad+16K_{q}^{2}\,\operatorname*{ess\,sup}_{(t,\omega)}\mathbb{E}_{t}\left[\int_{t}^{T}(|z^{1}_{s}|^{2}+|z^{2}_{s}|^{2}+\mathbb{E}|z^{1}_{s}|^{2}+\mathbb{E}|z^{2}_{s}|^{2})ds\right]\mathbb{E}_{t}\left[\int_{t}^{T}(|\Delta z_{s}|+\mathbb{E}|\Delta z_{s}|)^{2}ds\right]$
+$\leq\left(8L_{z}^{2}T+32K_{q}^{2}(\left\|z^{1}\right\|_{BMO}^{2}+\left\|z^{2}\right\|_{BMO}^{2})\right)\left(\left\|\Delta u\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|\Delta z\right\|_{BMO}^{2}\right).$
 Since $\left\|z^{i}\right\|_{BMO}^{2}\leq R^{2}$ for $i=1,2$, we obtain
-$\left\|\Delta U\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|\Delta Z\right\|_{BMO}^{2}\leq(16L_{z}^{2}T+128K_{z}^{2}R^{2})\left(\left\|\Delta u\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|\Delta z\right\|_{BMO}^{2}\right).$
+$\left\|\Delta U\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|\Delta Z\right\|_{BMO}^{2}\leq(16L_{z}^{2}T+128K_{q}^{2}R^{2})\left(\left\|\Delta u\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|\Delta z\right\|_{BMO}^{2}\right).$
 Now, recall from the previous step that $R^{2}=\frac{8}{(1-\alpha^{2})}\left\|\eta\right\|_{L^{\infty}}^{2}$.
@@
 then it follows that
-$16L_{z}^{2}T+128K_{z}^{2}R^{2}<1,$
+$16L_{z}^{2}T+128K_{q}^{2}R^{2}<1,$
 and thus the mapping $\Gamma$ is a contraction and admits a unique fixed point in $\mathcal{B}_{R}$.
@@
 Consequently, if we define $\rho>0$ by
-$\rho^{2}:=\frac{1-64L_{z}^{2}T}{65K_{z}},$
+$\rho^{2}:=\frac{1-64L_{z}^{2}T}{65K_{q}},$
 then, provided that $\left\|\eta\right\|_{L^{\infty}}\leq\rho$, The general BSDE admits a unique adapted solution $(U,Z)\in\mathcal{B}_{\rho}$ on this local time interval, under the imposed smallness conditions on both the terminal value and the time horizon.
@@
 Proposition 6.7.
-Let $T>0$, and assume that $g$ satisfies $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q}^{\prime})$. Suppose there exists $\rho>0$ such that
+Let $T>0$, and assume that $g$ satisfies $(\mathrm{H}_{0})$} and $(\mathrm{H}_{q}^{\mathrm{LK}})$. Suppose there exists $\rho>0$ such that
 $\left\|\eta\right\|_{L^{\infty}}\leq\rho,$
 where
-$\rho^{2}=\frac{1}{65^{2}K_{z}}.$
+$\rho^{2}:=\frac{1}{65^{2}K_{q}}.$
 Then the BSDE (6.34) admits a unique solution $(U,Z)\in\mathcal{B}_{\rho}$ on $[0,T]$, where
@@
 Let $\delta:=(65L_{z}^{2})^{-1}<(64L_{z}^{2})^{-1}$, and divide $[0,T]$ into $N$ subintervals of length at most $\delta$: $0=t_{0}<t_{1}<\cdots<t_{N}=T$, where $t_{k+1}-t_{k}\leq\delta$ for all $k$.
-By construction, $\rho^{2}=\frac{1-64L_{z}^{2}\delta}{65K_{z}}=\frac{1}{65^{2}K_{z}}$.
+By construction, $\rho^{2}=\frac{1-64L_{z}^{2}\delta}{65K_{q}}=\frac{1}{65^{2}K_{q}}$.
 We construct the solution recursively, starting from the last interval $t_{N-1},t_{N}]$.
@@
 ∎
-We remark that, given the unique solution $(U,Z)\in\mathcal{B}_{\rho}$ constructed above for the auxiliary BSDE 6.33, the solution to the original BSDE 6.32 is obtained by the explicit transformation
-$Y_{t}=U_{t}+g(t,0,0)+\mathbb{E}[\xi],\qquad 0\leq t\leq T.$
+We remark that, given the unique solution $(U,Z)\in\mathcal{B}_{\rho}$ constructed above for the auxiliary BSDE 6.34, the solution to the original BSDE 6.33 is obtained by the explicit transformation
+$Y_{t}=U_{t}+\int_{0}^{t}g(s,0,0)ds+\mathbb{E}[\xi],\qquad 0\leq t\leq T.$
 Furthermore, we emphasize that the smallness condition $\left\|\eta\right\|_{L^{\infty}}\leq\rho$, with $\eta:=\xi-\mathbb{E}[\xi]$, requires the terminal condition $\xi$ to be close to its mean.
@@
 Suppose the liability $F$, the terminal condition of BSDE (6.31) satisfies the condition such that $\left\|F-\mathbb{E}[F]\right\|_{L^{\infty}}\leq\rho$ where
-$\rho^{2}=\frac{1}{65^{2}K_{z}}.$
+$\rho^{2}:=\frac{1}{65^{2}K_{q}}.$
 Let $(Y,Z)$ denote the unique solution to BSDE (6.31) with driver $f$ (6.32).
@@
 Proof.
-First we will show the driver $f$ (6.31) satisfies the condition $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q}^{\prime})$.
+First we will show the driver $f$ (6.32) satisfies the condition $(\mathrm{H}_{0})$} and $(\mathrm{H}_{q}^{\mathrm{sep}})$}.
 The verification of the first condition is straightforward due to the boundedness of $\theta$.
-To verify $(\mathrm{H}_{q}^{\prime})$, we apply Lemma 6.3.
+To verify $(\mathrm{H}_{q}^{\mathrm{sep}})$}, we apply Lemma 6.3.
 For any $\omega\in\Omega$ and $t\in[0,T]$, since $Z_{t}(\omega)\in\mathbb{R}^{d}$, we have
@@
 $\quad\leq 2|\theta_{t}|\,|Z_{t}^{1}-Z_{t}^{2}|+\frac{\alpha}{2}(1\vee K_{1})^{2}\left(|Z_{t}^{1}|+|Z_{t}^{2}|+\mathbb{E}|Z_{t}^{1}|+\mathbb{E}|Z_{t}^{2}|\right)\left(|Z_{t}^{1}-Z_{t}^{2}|+\mathbb{E}|Z_{t}^{1}-Z_{t}^{2}|\right),\quad\text{a.s.}$
-Together with the boundedness of $\theta$, we obtain a sufficient condition for $(\rm H_{q}^{\prime})$}.
+Together with the boundedness of $\theta$, we obtain a sufficient condition for $(\mathrm{H}_{q}^{\mathrm{LK}})$}.
 This ensures the existence and uniqueness of a solution $(Y,Z)$ to BSDE (6.31) by Proposition 6.7.
@@
 $V^{(p)}(t,x):=-\exp\left(-\alpha(X_{t}^{p}-Y_{t})\right),\quad t\in[0,T],$
-is a supermartingale, and that it becomes a martingale when $p=p^{*}$ as defined above. This verification follows directly from arguments in Theorem 7 of hu2005utility, and we omit the details here. ∎
+is a supermartingale, and that it becomes a martingale when $p=p^{*}$ as defined above.
+This verification follows directly from arguments in Theorem 7 of Hu et al.
+(2005), and we omit the details here.
+∎
 7 Conclusion
-In this paper we have developed a general framework for mean-field BSDEs whose generators have quadratic growth in the control and its law.
-Working under the separately quadratic structure (Hq) with bounded terminal condition, we proved global existence and uniqueness of solutions in the Banach space $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$, without imposing any smallness condition.
-The key ingredients were a Malliavin calculus argument for Lipschitz mean-field BSDEs, which yields uniform bounds on the control process, together with BMO estimates and a global stability result that allows us to pass from Lipschitz to quadratic generators.
-We then extended the analysis to the fully coupled quadratic regime (Hq), in which the generator admit genuine cross-interaction terms.
-In this more general setting we established existence and uniqueness under a natural smallness condition on the centred terminal condition.
-As an application, we showed how the fully coupled theory can be used to formulate and solve a mean-field exponential utility maximization problem, thereby generalizing the classical framework of hu2005utility.
+In this paper, we study mean-field BSDEs whose generators of quadratic growth in the control component and in the mean-field term.
+Our main well-posedness results cover two complementary structural regimes.
+First, under the unified quadratic condition $(\mathrm{H}_{q}^{\mathrm{unif}})$, we establish global existence and uniqueness of solutions in $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ for bounded terminal conditions of the form $\xi=\phi(W_{t_{1}},\ldots,W_{t_{n}})$, where $\phi$ is bounded and Hölder continuous.
+Second, under the separably quadratic condition $(\mathrm{H}_{q}^{\mathrm{sep}})$, we prove existence and uniqueness in the same solution class for bounded terminal conditions generated by uniformly continuous functionals of the Brownian path, i.e.
+$\xi=\phi(W_{[0,T]})$ with $\phi:C([0,T];\mathbb{R}^{d})\to\mathbb{R}$ uniformly continuous in the supremum norm.
+Finally, motivated by a mean-field exponential utility maximization problem with market-adaptive constraints, we require a fully coupled quadratic condition allowing cross-interaction terms.
+For this purpose we work under the quadratic condition $(\mathrm{H}_{q}^{\mathrm{LK}})$ (equivalent to $(\mathrm{H}_{q}^{\mathrm{unif}})$) and develop a dedicated concatenation approach.
+This yields existence and uniqueness for the associated BSDE under a smallness assumption on the centred terminal condition $\eta:=\xi-\mathbb{E}[\xi]$.
+As an application, we show that the resulting BSDE characterizes the value function and an optimal strategy, thereby extending the classical framework of Hu et al.
+(2005) to a mean-field market with constraints driven by an aggregated trading signal.